### v4

# A Comprehensive Library for Benchmarking Multi-class Visual Anomaly Detection

Abstract
Visual anomaly detection aims to identify anomalous regions in images through unsupervised learning paradigms, with increasing application demand and value in fields such as industrial inspection and medical lesion detection. Despite significant progress in recent years, there is a lack of comprehensive benchmarks to adequately evaluate the performance of various mainstream methods across different datasets under the practical multi-class setting. The absence of standardized experimental setups can lead to potential biases in training epochs, resolution, and metric results, resulting in erroneous conclusions. This paper addresses this issue by proposing a comprehensive visual anomaly detection benchmark, ADer, which is a modular framework that is highly extensible for new methods. The benchmark includes multiple datasets from industrial and medical domains, implementing fifteen state-of-the-art methods and nine comprehensive metrics. Additionally, we have proposed the GPU-assisted ADEval package to address the slow evaluation problem of metrics like time-consuming mAU-PRO on large-scale data, significantly reducing evaluation time by more than 1000-fold. Through extensive experimental results, we objectively reveal the strengths and weaknesses of different methods and provide insights into the challenges and future directions of multi-class visual anomaly detection. We hope that ADer will become a valuable resource for researchers and practitioners in the field, promoting the development of more robust and generalizable anomaly detection systems. Full codes have been attached in Appendix and will be open-sourced.
1 Introduction
In recent years, with the rapid advancement in model iteration and computational power, Visual Anomaly Detection (VAD) has made significant progress across various fields [25, 7]. It has become a crucial component in key tasks such as industrial quality inspection and medical lesion detection. Due to its unsupervised experimental setup, VAD demonstrates immense application value in real-world scenarios where the yield rate is high, defect samples are difficult to obtain, and potential defect patterns are diverse. However, the field faces challenges such as small dataset sizes and insufficient evaluation metrics, resulting in potentially unfair comparison outcomes due to differing training recipes among methods. Moreover, most methods have not been compared on the latest large-scale datasets (e.g., Real-IAD  [32] and COCO-AD [42]) and new evaluation metrics (e.g., mAD [40] and mIoU-max [42]). The fundamental issue is the absence of standardized training strategies, akin to those in object detection, to evaluate different algorithms. Training epoch and resolution factors can potentially affect evaluation results, leading to erroneous conclusions.
To address this pressing issue, we believe that establishing a comprehensive and fair benchmark is crucial for the sustained and healthy development of this field. Therefore, we have constructed an integrated ADer library, benchmarking state-of-the-art methods by utilizing a unified evaluation interface under the more practical multi-class setting. This library is designed as a highly extensible modular framework (see Sec. 3), allowing for the easy implementation of new methods. Specifically, the framework integrates multiple datasets from industrial, medical, and general-purpose domains (see Sec. 3.2), and implements fifteen state-of-the-art methods (including augmentation-based, embedding-based, reconstruction-based, and hybrid methods, see Sec. 3.1) and nine comprehensive evaluation metrics (see Sec. 3.3), ensuring thorough and unbiased performance evaluation for each method. Additionally, to address the efficiency issue of evaluating time-consuming metrics like mAU-PRO on large-scale data, we have developed and open-sourced the GPU-assisted ADEval package (see Sec. 3.6), significantly reducing evaluation time by over 1000 times, making previously impractical extensive detailed evaluations feasible on large-scale datasets.
Figure 1: Intuitive benchmarked results comparison on MVTec AD [3] (Left), VisA [47] (Middle), and Real-IAD [32] (Right) datasets among mainstream methods. For each dataset, the horizontal axis represents the training time for different methods, the left vertical axis represents mAD [40] (marked as circles, with radius indicating model parameter count), and the right vertical axis represents mIoU-max [42] (marked as squares, with side length indicating model FLOPs).
Through extensive and fair experiments, we objectively reveal the strengths and weaknesses of different visual anomaly detection methods, comparing their efficiency (i.e., model parameter count and FLOPs) and training resource consumption across different datasets, as shown in Fig. 1. Detailed results and analyses (see Sec. 4 and Appendix) elucidate the challenges of multi-class visual anomaly detection and provide valuable insights for future research directions.
In summary, the contributions of this paper are as follows: 1) Comprehensive benchmark: We introduce a modular and extensible library termed ADer for visual anomaly detection, which implements and evaluates 15 state-of-the-art anomaly detection methods on 11 popular datasets with 9 comprehensive evaluation metrics. 2) GPU-assisted evaluation package: We develop and will open-source the ADEval package for large-scale evaluation, significantly reducing the evaluation time of complex metrics by over 1000 times. 3) Extensive experimental analysis: We conduct extensive experiments to objectively evaluate the performance of different methods, providing insights into their strengths, weaknesses, and potential areas for improvement. 4) Open-source resources: We will open-source the complete ADer code, making it a valuable resource for the research community and promoting further advancements in the field.
2 Background and Related Work
Figure 2: A comparative diagram of different frameworks for the benchmarked methods in Sec. 3.1.
2.1 Problem Definition and Objective
Visual anomaly detection (VAD) is a critical task in computer vision, aiming at identifying patterns or instances in visual data that deviate significantly from the norm. These anomalies can manifest as industrial defects, medical lesion, or rare objects that are not typically present in the training data. The primary objective of VAD is to develop algorithms capable of discerning these irregularities with high accuracy and reliability. This task is particularly challenging due to the inherent variability and complexity of visual data, the scarcity of anomalous examples, and the need for robust generalization across diverse scenarios. In a formal context, multi-class VAD can be defined as follows: Given a training dataset $D_{train}=\{x_{1},x_{2},...,x_{n}\}$ with $C$ categories and each visual image $x_{i}$ belonging to a specific category, the goal is to learn a unified AD model $M$ that can predict an anomaly score $s_{i}=M(x_{i})$ for each image. This score reflects the likelihood of each pixel in $x_{i}$ being an anomaly. The model $M$ is typically trained on $D_{train}$ that predominantly contains normal instances, with the assumption that anomalies are rare and not well-represented in the training set. Inevitably, there are some mislabeled or inaccurately labeled noisy samples, which constitute inherent biases within the dataset. These are typically disregarded under standard settings. The performance of the model is then evaluated based on its ability to correctly identify anomalous images and their defect regions in a unified testset that contains normal and anomalous images.
2.2 Challenges in Multi-class VAD
The complexity of VAD arises from several factors: 1) Data Imbalance: Anomalies are rare, leading to highly imbalanced datasets where normal instances and region areas vastly outnumber anomalous ones. 2) Variability of Anomalies: Anomalies can vary widely in appearance, making it difficult to capture all possible variations during training. 3) Context Sensitivity: The definition of what constitutes an anomaly can be context-dependent, requiring models to understand the broader context in which the visual data is situated. 4) Efficiency Requirements: Many applications of VAD require real-time processing and limited GPU memory, necessitating efficient and scalable algorithms. 5) Comprehensive and Fair Evaluation: Current methods exhibit significant differences in training hyperparameters and insufficient evaluation of performance metrics, so it is necessary to benchmark them using fair and standardized criteria. In this benchmark study, we systematically evaluate a range of state-of-the-art VAD methods (Sec. 3.1) across multiple datasets (Sec. 3.2) and comprehensive metrics (Sec. 3.3). Our goal is to provide a comprehensive assessment of current capabilities, identify key challenges, and suggest directions for future research in visual anomaly detection.
2.3 Visual Anomaly Detection
Visual anomaly detection methods can generally be categorized into three types: 1) Augmentation-based methods generate pseudo-supervised information for anomalies by creating abnormal regions [23, 37], constructing anomalous data [38, 19], or adding feature perturbations [27, 31]. This enables the model to learn the differences between normal and abnormal distributions. 2) Embedding-based methods leverage pretrained models to extract powerful feature representations and judge anomalies in high-dimensional space. Typical approaches include distribution-map-based methods [14, 22], teacher-student frameworks [4, 33], and memory-bank techniques [29, 13]. 3) Reconstruction-based methods use encoder-decoder architectures to locate anomalies by analyzing the reconstruction error. They typically include both image-level [2, 24, 16] and feature-level approaches [10, 40, 42]. There are also some hybrid methods [35, 31, 44] that attempt to integrate multiple techniques to further enhance model performance.
Basic network structures of VAD. Early visual anomaly detection methods typically employ UNet-based autoencoder architectures [1, 24, 37]. With advancements in foundational visual model structures [17, 34, 26, 39, 43, 41] and pretraining techniques [18, 8], more recent methods often utilize models pretrained on ImageNet-1K [11] as feature extractors, such as the ResNet [17] series, Wide ResNet-50 [36], and EfficientNet-b4 [30]. Recently, benefiting from the dynamic modeling capabilities of Vision Transformers (ViT) [12], some studies [9, 35, 40] have attempted to incorporate this architecture into anomaly detection.
3 Methodology: ADer Benchmark
Method Hyper Params. Efficiency Train Mem. (M) Test Mem. (M) MVTec AD VisA Real-IAD
BS Optim. LR Params. FLOPs Backbone Train T. Test T. Train T. Test T. Train T. Test T.
Aug. DRAEM [37] 8 Adam 1e-4 97.4M 198G UNet 12,602 2,858 3.4h 35s 8.0h 36s 39.2 18m41s
SimpleNet [27] 32 AdamW 1e-4 72.8M 17.715G WRN50 2,266 4,946 1.7h 5m50s 3.9h 7m21s 15.9h 4h51m
RealNet [45] 16 Adam 1e-4 591M 115G WRN50 14,004 3,794 2.6h 41s 5.4h 41s - -
Emb. CFA [21] 4 AdamW 1e-3 38.6M 55.3G WRN50 4,364 2,826 1.2h 18s 2.7h 17s 10.9h 14m20s
PatchCore [29] 8 - - - - WRN50 - - 0.6h 9h22m - OOM - OOM
CFLOW-AD [14] 32 Adam 2e-4 237M 28.7G WRN50 3,048 1,892 5.2h 56s 10.4h 1m15s 40.9h 22m49s
PyramidalFlow [22] 2 Adam 2e-4 34.3M 962G RN18 3,904 2,836 7.9h 1m30s 10.5h 2m43s 45h 38m15s
Rec. RD [10] 16 Adam 5e-3 80.6M 28.4G WRN50 3,286 1,464 0.8h 13s 1.1h 18s 4.1h 7m48s
DiAD [16] 12 Adam 1e-5 1331M 451.5G RN50 26,146 20,306 9.1h 16m 21.6h 19m 90h 16h20m
ViTAD [40] 8 AdamW 1e-4 39.0M 9.7G ViT-S 1,470 800 0.8h 15s 1.1h 15s 5.2h 10m2s
InvAD [42] 32 Adam 1e-3 95.6M 45.4G WRN50 5,920 3,398 1.0h 31s 2.3h 33s 9.2h 21m
InvAD-lite [42] 32 Adam 1e-3 17.1M 9.3G RN34 1,846 1,100 0.8h 20s 1.1h 31s 3.4h 9m27s
MambaAD [15] 16 AdamW 5e-3 25.7M 8.3G RN34 6,542 1,484 2.4h 34s 5.6h 23s 23.6h 24m6s
Hybrid UniAD [35] 8 AdamW 1e-4 24.5M 3.4G EN-b4 1,856 968 0.8h 22s 1.0h 18s 4.1h 7m2s
RD++ [31] 16 Adam 1e-3 96.1M 37.5G WRN50 4,772 1,480 3.3h 28s 7.8h 33s 42.4h 15m17s
DesTSeg [44] 32 SGD 0.4 35.2M 30.7G RN18 3,446 1,240 1.0h 19s 1.7h 16s 6.8h 8m13s
Table 1: Attribute comparison for mainstream representative methods. Notations: Augmentation-based (Aug.), Embedding-based (Emb.), Reconstruction-based (Rec.), Parameters (Params), Memory (Mem.), Batch Size (BS), Optimizer (Optim.), Time (T.), ResNet (RN), Wide-ResNet (WRN), EfficientNet (EN), hours (h), minutes (m), seconds (s), unavailable (-), out-of-memory (OOM). Train and test time are evaluated under the standard setting described in Sec. 4.1 in one L40S GPU. Memory is tested under the standard setting with a batch size of 8, and the results for different methods are presented in Sec. 4.2. Bold, underline, and wavy-line represent the best, second-best, and third-best results, respectively.
3.1 Supported VAD Methods
Following the categories of current VAD methods in Sec. 2.3, we choose representative models for each category. The selection criteria are based on the method’s popularity, effectiveness, and ease of use. 1) For Augmentation-based methods, we choose DRAEM [37], SimpleNet [27], and RealNet [45]. 2) For Embedding-based methods, we select CFA [21], PatchCore [29], CFLOW-AD [14], and PyramidalFlow [22]. 3) For Reconstruction-based methods, we include RD [10], DiAD [16], ViTAD [40], InvAD [42], InvAD-lite [42], and MambaAD [15]. Additionally, UniAD [35], RD++ [31], and DesTSeg [44] are categorized as hybrid methods due to their use of multiple techniques. Fig. 2 presents schematic diagrams and comparisons of the frameworks for each method belonging to different types, facilitating a better understanding of the differences among these methods. Tab. 1 provides a direct comparison of the hyperparameters, efficiency, and training time on three mainstream datasets for different methods, using one L40S GPU. Note that different methods may yield varying results when tested on different hardware, but the overall relative trends remain largely unchanged.
3.2 VAD Datasets
Dataset Category Number Image Quantity Epoch Setting in ADer
Train Test Train Test
Normal Anomaly Normal
MVTec AD [3] 15 15 03,629 01,258 00467 100 300
MVTec AD 3D [6] 10 10 02,950 00249 00948 100 300
MVTec LOCO-AD [5] 05 05 01,772 00993 00575 100 300
VisA [47] 12 12 08,659 00962 01,200 100 300
BTAD [28] 03 03 01,799 00580 00451 100 300
MPDD [20] 06 06 00888 00282 00176 100 300
MAD_Real [46] 10 10 00490 00221 00050 100 300
MAD_Sim [46] 20 20 04,200 04,951 00638 100 300
Real-IAD [32] 30 30 36,465 51,329 63,256 100 -
Uni-Medical [40] 03 03 13,339 04,499 02,514 100 300
COCO-AD [42] 61 81 30,438 01,291 03,661 100 -
65,133 02,785 02,167 100 -
79,083 03,328 01,624 100 -
77,580 03,253 01,699 100 -
Table 2: Comparison of representative VAD datasets, i.e., industrial, medical, and general-purpose fields, respectively. Large-scale Real-IAD and COCO-AD only employ 100 epoch setting.
To comprehensively evaluate the effectiveness, stability, and generalization of different methods, we benchmark extensive and fair experiments on three types of datasets: 1) Real and synthetic industrial anomaly detection (AD) datasets, i.e., MVTec AD [3], MVTec AD 3D [6], MVTec LOCO-AD [5], VisA [47], BTAD [28], MPDD [20], MAD_Real [46], MAD_Sim [46], and Real-IAD [32]. 2) The medical Uni-Medical [40] dataset. 3) The general-purpose COCO-AD [42] dataset. Detailed descriptions of the datasets are provided in Tab. 2, including the categories and scales of the datasets. Note that COCO-AD is inherently a multi-class dataset with four splits, and the average is taken when evaluating the comprehensive results.
3.3 Evaluation Metrics
[…]
Figure 3: Core sub-modules of the framework in ADer. The blue area represents standard components, while the red area indicates that a new method requires only three corresponding files.
3.4 Simplify Implementation by Structured ADer
To ensure fair comparison among different methods, we construct a standardized ADer framework. As shown in Fig. 3, it includes shared foundational training/testing components and implements various metric calculations (compatible with our ADEval). The standardized dataset allows for easy comparison, eliminating potential unfair settings from different codebases. Additionally, ADer is highly extensible for new methods, requiring only compliant model, trainer, and configuration files.
3.5 Feature comparison with current benchmarks.
The existing vision anomaly detection benchmark works are primarily open-iad and anomalib. However, their updates for general AD models only extend up to 2022, and they have not yet implemented the latest and practical multi-class anomaly detection methods. We briefly discuss the relationship between the most popular Anomalib and ADer as follows: 1) From the framework perspective: Anomalib is based on PyTorch Lightning that is deeply encapsulated, whereas ADer has a shallower encapsulation, exposing more interfaces to facilitate rapid algorithm iteration. 2) From the methods perspective: Anomalib supports general AD models only up to 2022, while ADer supports more recent models up to 2024. 3) From the data and metrics perspective: Compared to Anomalib, ADer supports large-scale industrial Real-IAD [32], medical Uni-Medical [40], and general-purpose COCO-AD [42] datasets, as well as more application-relevant metrics like mIoU-max [42] and averaged mAD [40]. 4) From the setting perspective: ADer focuses more on the recently popular and future research trend of multi-class settings.
3.6 ADEval: Fast and Memory-efficient Routines for mAU-ROC/mAP/mAU-PRO
The speed of metric evaluation is crucial for the iterative process of model algorithms. When the number of test images increases and the resolution becomes higher, the pixel-level evaluation algorithms implemented naively using sklearn and skimage packages become time-consuming. This is particularly evident with large-scale datasets such as Real-IAD [32] and COCO-AD [42], where evaluation times can exceed one hour. To address this issue, we have released the GPU-assisted ADEval library, which employs an iterative-accumulating algorithm with CUDA acceleration. By designing specialized histogram bins and employing a weighted accumulation process, the computation of metrics such as AUROC, AUPR, and AUPRO, which require sliding confidence thresholds to derive curves and then calculate the area under these curves, is optimized into an iterative-accumulative form. This approach avoids the need to cache the complete set of ground truth and predicted values during the evaluation of large-scale, high-resolution test sets.
Given a test set with inputs $X=\{x_{i}\}$, labels $Y=\{y_{i}\mid i\in I\}$, and inference results $P_{X,Y}=\{\hat{y}_{i}\mid i\in I\}$, the ROC curve $L=\{(tpr_{i},fpr_{i},\hat{y}_{i})\mid i\in I\}$ can be sampled as $\hat{L}=\{(\widehat{tpr}_{k},\widehat{fpr}_{k},\widehat{thr}_{k})\mid 0\leq k<N\}$, where $\widehat{thr}_{k}$ is strictly increasing. The area under the ROC curve (AUROC) is given by $\operatorname{Trapz}(\{(fpr_{i},tpr_{i})\mid i\in I\})\approx\operatorname{Trapz}(\{\widehat{fpr}_{k},\widehat{tpr}_{k}\mid 0\leq k<N\})$, where $N$ is the number of sampling points on the ROC curve. Let $\widehat{thr}_{N}=\max(P_{X,Y})$.
The true positive rate (TPR) and false positive rate (FPR) are defined as follows:
[…]
Let $p_{k}=\left|\{i\mid i\in I\wedge y_{i}=1\wedge\widehat{thr}_{k}\leq\hat{y}_{i}\leq\widehat{thr}_{k+1}\}\right|$ and $q_{k}=\left|\{i\mid i\in I\wedge y_{i}=0\wedge\widehat{thr}_{k}\leq\hat{y}_{i}\leq\widehat{thr}_{k+1}\}\right|$.
Then, the true positive rate and false positive rate can be expressed as:
$\widehat{tpr}_{k}=\frac{\sum_{l=k}^{N-1}p_{l}}{\left|\{i\mid i\in I\wedge y_{i}=1\}\right|},\widehat{fpr}_{k}=\frac{\sum_{l=k}^{N-1}q_{l}}{\left|\{i\mid i\in I\wedge y_{i}=0\}\right|},$
Thus, the sampled ROC curve $\hat{L}$ can be fully derived from $P=\{p_{k}\mid 0\leq k<N\}$ and $Q=\{q_{k}\mid 0\leq k<N\}$. These sets $P$ and $Q$ can be computed iteratively and cumulatively as follows:
For $P$, maintain a histogram $H=\{h_{k}\mid 0\leq k<N\}$ initialized to $h_{k}=0$ for all $0\leq k<N$. For each batch $b$ of inference results $\{\hat{y}_{i}\mid i\in b\}$, update the histogram as:
$h_{k}\leftarrow h_{k}+\left|\{i\mid i\in b\wedge y_{i}=1\wedge\widehat{thr}_{k}\leq\hat{y}_{i}\leq\widehat{thr}_{k+1}\}\right|.$
Continue this process until all samples have been processed.
This method requires a space complexity of $O(N)$, which is significantly less than the traditional method that requires $O(|X|)$ space to store all inference results. Additionally, each iterative update can be efficiently implemented using mature histogram operators, leveraging parallelism and GPU acceleration. This approach is particularly advantageous for computing large-scale, high-resolution evaluation metrics, especially pixel-level metrics.
Table 3: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3 under 100/300 epochs. Bold, underline, and wavy-line represent the best, second-best, and third-best results, respectively. Patchcore requires no training that shares results under different epoch settings.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Aug. DRAEM [37] 54.5/55.2 76.3/77.0 83.6/83.9 47.6/48.7 03.2/03.1 06.7/06.3 14.3/15.8 03.5/03.3 40.9/41.4
SimpleNet [27] 95.4/79.2 98.3/90.8 95.7/87.6 96.8/82.4 48.8/24.0 51.9/29.0 86.9/62.0 36.4/17.8 82.0/65.0
RealNet [45] 84.8/82.9 94.1/93.3 90.9/90.9 72.6/69.8 48.2/50.0 41.4/40.4 56.8/51.2 28.8/28.5 69.8/68.4
Emb. CFA [21] 57.6/55.8 78.3/78.8 84.7/84.5 54.8/43.9 11.9/04.8 14.7/08.9 25.3/19.3 08.9/04.7 46.8/42.3
PatchCore [29] 98.8/0-00 99.5/0-00 98.4/0-00 98.3/0-00 59.9/0-00 61.0/0-00 94.2/0-00 44.9/0-00 87.2/0-00
CFLOW-AD [14] 91.6/92.7 96.7/97.2 93.4/94.0 95.7/95.8 45.9/46.8 48.6/49.6 88.3/89.0 33.2/34.0 80.0/80.7
PyramidalFlow [22] 70.2/66.2 85.5/84.3 85.5/85.1 80.0/74.2 22.3/17.2 22.0/19.6 47.5/40.0 12.8/11.4 59.0/55.2
Rec. RD [10] 93.6/90.5 97.2/95.0 95.6/95.1 95.8/95.9 48.2/47.1 53.6/52.1 91.2/91.2 37.0/35.8 82.2/81.0
DiAD [16] 88.9/92.0 95.8/96.8 93.5/94.4 89.3/89.3 27.0/27.3 32.5/32.7 63.9/64.4 21.1/21.3 70.1/71.0
ViTAD [40] 98.3/98.4 99.3/99.4 97.3/97.5 97.6/97.5 55.2/55.2 58.4/58.1 92.0/91.7 42.3/42.0 85.4/85.4
InvAD [42] 98.1/98.9 99.0/99.6 97.6/98.2 98.0/98.1 56.3/57.1 59.2/59.6 94.4/94.4 42.8/43.1 86.1/86.6
InvAD-lite [42] 97.9/98.1 99.2/99.1 96.8/96.8 97.3/97.3 54.4/55.0 57.8/58.1 93.3/93.1 41.4/41.7 85.2/85.4
MambaAD [15] 97.8/98.5 99.3/99.5 97.3/97.7 97.4/97.6 55.1/56.1 57.6/58.7 93.4/93.6 41.2/42.3 85.4/86.0
Hybrid UniAD [35] 92.5/96.8 97.3/98.9 95.4/97.0 95.8/96.8 42.7/45.0 48.0/50.2 89.3/91.0 32.5/34.2 80.1/82.2
RD++ [31] 97.9/95.8 98.8/98.0 96.4/96.6 97.3/97.3 54.7/53.0 58.0/57.0 93.2/92.9 41.5/40.5 85.2/84.4
DesTSeg [44] 96.4/96.3 98.6/98.8 96.2/96.1 92.0/92.6 71.1/75.8 68.2/71.3 83.4/82.6 52.8/56.6 86.6/87.6
Table 4: Benchmarked results on VisA dataset [47] by the suggested metrics under 100/300 epochs.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Aug. DRAEM [37] 55.1/56.2 62.4/64.6 72.9/74.9 37.5/45.0 00.6/00.7 01.7/01.8 10.0/16.0 00.9/00.9 34.3/37.0
SimpleNet [27] 86.4/80.7 89.1/83.8 82.8/79.3 96.6/94.4 34.0/29.2 37.8/33.1 79.2/74.2 25.7/22.1 72.3/67.8
RealNet [45] 71.4/79.2 79.5/84.8 74.7/78.3 61.0/65.4 25.7/29.2 22.6/27.9 27.4/33.9 13.5/17.4 51.8/57.0
Emb. CFA [21] 66.3/67.1 74.3/73.8 74.2/75.3 81.3/83.0 22.1/13.7 26.2/18.7 50.8/48.7 17.0/11.3 56.5/54.3
CFLOW-AD [14] 86.5/87.2 88.8/89.5 84.9/85.1 97.7/97.8 33.9/34.2 37.2/37.2 86.8/87.3 24.9/24.9 73.7/74.0
PyramidalFlow [22] 58.2/69.0 66.3/72.9 74.4/75.8 77.0/79.1 07.2/07.9 09.6/08.7 42.8/52.6 05.6/04.7 47.9/52.3
Rec. RD [10] 90.6/93.9 90.9/94.8 89.3/90.4 98.0/98.1 35.4/38.4 42.5/43.7 91.9/91.9 27.9/29.0 76.9/78.7
DiAD [16] 84.8/90.5 88.5/91.4 86.9/90.4 82.5/83.4 17.9/19.2 23.2/25.0 44.5/44.3 14.9/16.2 61.2/63.5
ViTAD [40] 90.4/90.3 91.1/91.2 86.0/86.4 98.2/98.2 36.4/36.4 41.0/40.9 85.7/85.8 27.5/27.5 75.5/75.6
InvAD [42] 95.4/95.6 95.7/96.0 91.6/92.3 98.9/99.0 43.3/43.7 46.8/46.9 93.1/93.0 32.5/32.6 80.7/80.9
InvAD-lite [42] 94.9/95.3 95.2/95.8 90.7/91.0 98.6/98.7 40.2/41.2 44.0/44.9 93.1/93.2 29.8/30.6 79.5/80.0
MambaAD [15] 94.5/93.6 94.9/93.9 90.2/89.8 98.4/98.2 39.3/34.0 43.7/39.3 92.1/90.5 29.5/25.9 79.0/77.0
Hybrid UniAD [35] 89.0/91.4 91.0/93.3 85.8/87.5 98.3/98.5 34.5/35.3 39.6/40.2 86.5/89.0 26.4/26.5 75.0/76.5
RD++ [31] 93.9/93.1 94.7/94.1 90.2/90.0 98.4/98.4 42.3/40.4 46.3/44.8 91.9/91.4 31.2/29.9 79.7/78.9
DesTSeg [44] 89.9/89.0 91.4/90.3 86.7/85.9 86.7/84.8 46.6/43.3 47.2/44.4 61.1/57.5 32.7/30.1 72.8/70.7
4 Results and Analysis
4.1 Experimental Setup
Different methods potentially introduce various factors that can impact model performance. To ensure a fair and comprehensive evaluation of the effectiveness and convergence of different methods, we fix the most influential parameters, i.e., resolution (2562) and training epochs (100 and 300). The reason lies in the fact that tasks such as classification, detection, and segmentation typically set specific resolutions and standard training epochs. We observe that for most methods, 100 epochs generally suffice to reach saturation [40, 42], with only a few methods [35] requiring more epochs for training. Therefore, we also establish a setting with 300 epochs. Meanwhile, we maintain consistency with the original papers for batch size, optimizer, learning rate, and data augmentation. We report the evaluation results corresponding to the final epoch at the end of training to ensure fairness. All experiments are conducted on one L40S GPU.
Table 5: Benchmarked results on Real-IAD dataset [32] by the suggested metrics under 100 epoch.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Aug. DRAEM [37] 50.9 45.9 61.3 44.0 00.2 00.4 13.6 00.2 30.9
SimpleNet [27] 54.9 50.6 61.5 76.1 01.9 04.9 42.4 02.5 41.8
Emb. CFA [21] 55.7 50.5 61.9 81.3 01.6 03.8 48.8 02.0 43.4
CFLOW-AD [14] 77.0 75.8 69.9 94.8 17.6 21.7 80.4 12.4 62.5
PyramidalFlow [22] 54.4 48.0 62.0 71.1 01.2 01.1 34.9 00.5 39.0
Rec. RD [10] 82.7 79.3 74.1 97.2 25.2 32.8 90.0 20.0 68.8
DiAD [16] 75.6 66.4 69.9 88.0 02.9 07.1 58.1 03.7 52.6
ViTAD [40] 82.7 80.2 73.7 97.2 24.3 32.3 84.8 19.6 67.9
InvAD [42] 89.4 87.0 80.2 98.4 32.6 38.9 92.7 24.6 74.2
InvAD-lite [42] 87.2 85.2 77.8 98.0 31.7 37.9 92.0 23.8 72.8
MambaAD [15] 87.0 85.3 77.6 98.6 32.4 38.1 91.2 23.9 72.9
Hybrid UniAD [35] 83.1 81.2 74.5 97.4 23.3 30.9 87.1 18.6 68.2
RD++ [31] 83.6 80.6 74.8 97.7 25.9 33.6 90.7 20.5 69.6
DesTSeg [44] 79.3 76.7 70.7 80.3 36.9 40.3 56.1 26.2 62.9
Table 6: Benchmarked results on all other datasets by the mIoU/mAD metrics under 100 epoch.
Method MVTec 3D MVTec LOCO BTAD MPDD MAD_Real MAD_Sim Uni-Medical COCO-AD
Aug. DRAEM [37] 01.0/42.5 05.6/38.5 03.4/43.3 02.5/32.8 00.8/43.8 00.7/48.0 03.0/33.5 08.0/36.0
SimpleNet [27] 13.9/67.8 21.2/65.0 28.6/74.3 24.5/76.2 06.3/54.3 04.2/58.0 23.3/67.7 11.5/41.0
RealNet [45] 0-00/0-00 0-00/0-00 36.6/73.7 28.2/70.1 0-00/0-00 0-00/0-00 0-00/0-00 0-00/0-00
Emb. CFA [21] 09.3/55.2 09.1/49.0 33.6/78.2 16.6/62.3 08.7/56.5 04.6/55.1 14.7/55.0 08.9/38.7
PatchCore [29] 24.5/76.7 20.4/65.4 38.0/81.5 35.0/81.9 16.6/66.6 0-00/0-00 0-00/0-00 0-00/0-00
CFLOW-AD [14] 15.8/69.8 17.3/61.7 33.8/77.7 20.1/68.3 08.8/61.2 02.7/57.1 17.7/68.6 16.0/51.5
PyramidalFlow [22] 06.4/59.7 08.0/44.7 18.3/66.1 10.4/62.9 05.1/56.7 02.5/54.5 09.4/45.7 08.0/36.2
Rec. RD [10] 22.2/73.7 15.8/60.6 42.1/83.2 31.4/79.0 07.2/58.2 04.5/60.0 26.9/70.3 11.5/44.3
DiAD [16] 05.4/62.8 14.9/56.5 15.7/68.5 08.2/58.1 03.6/55.8 04.2/58.9 23.2/69.1 11.6/44.1
ViTAD [40] 20.4/73.5 19.8/62.5 40.1/81.5 27.7/75.6 05.0/55.2 05.0/60.1 33.7/74.7 20.1/52.2
InvAD [42] 27.4/78.6 23.1/67.0 44.3/84.5 34.0/80.2 16.8/64.7 08.6/67.4 32.6/74.6 14.3/49.0
InvAD-lite [42] 26.9/78.0 20.6/64.9 42.6/82.9 30.9/78.2 08.1/60.2 06.0/62.1 26.5/70.7 13.7/47.6
MambaAD [15] 25.9/77.4 20.6/64.3 39.0/80.9 26.8/76.3 07.2/58.2 05.0/60.8 33.5/75.0 12.9/47.1
Hybrid UniAD [35] 16.7/70.3 21.6/62.4 36.9/81.1 12.5/61.7 05.8/56.1 03.5/58.3 27.6/70.7 10.9/42.4
RD++ [31] 25.2/76.4 17.6/62.2 42.8/83.3 33.6/79.3 08.5/59.2 04.4/60.0 29.4/71.6 11.8/45.0
DesTSeg [44] 28.4/70.4 20.3/61.3 29.0/74.6 25.6/69.3 04.5/50.1 04.1/49.7 21.2/58.6 08.5/38.6
4.2 Benchmark Results on Industrial, Medical, and General-purpose UAD Datasets
To thoroughly evaluate the effectiveness of different methods and their adaptability to various data domains, we conduct experiments on multiple datasets across three domains. Due to space constraints, we report the average metrics for the popular MVTec AD (see Tab. 3), VisA (see Tab. 4), and Real-IAD (see Tab. 5) datasets in the main paper. For the remaining datasets, we report the mAD and mIoU-max metrics (see Tab. 6). Full results for each category are provided in Appendix A to  V.
Quantitative results. InvAD [42] consistently shows excellent performance across all datasets. ViTAD [40] and MambaAD [15], specifically designed for multi-class settings, also achieve good results. In contrast, DiAD [16] and UniAD [35] require more epochs to converge and do not perform well under the 100/300 epoch standard we set. DeSTSeg [44] exhibits outstanding performance in pixel-level segmentation. Methods designed for single-class settings, such as RD [10], RD++ [31], CFLOW-AD [14], and RealNet [45], also perform well in multi-class settings. However, single-class methods like DRAEM [37], SimpleNet [27], CFA [21], and PyramidFlow [22] show significant performance gaps in multi-class anomaly detection and are not suitable for such tasks. Considering the training time, model parameters, and FLOPs shown in Fig. 2, InvAD, InvAD-lite, and ViTAD achieve a good balance of effectiveness and efficiency. RD, UniAD, and DeSTSeg also perform well in terms of both efficiency and effectiveness. On the other hand, methods like DiAD, PyramidFlow, CFLOW-AD, RD++, RealNet, MambaAD, and SimpleNet have significantly longer training times compared to others.
Qualitative results. Fig. 4 in Appendix presents intuitive visualization results under the 100 epochs training setting on popular MVTec AD [3] and VisA [47] datasets, as well as the medical Uni-Medical [40] and large-scale Real-IAD [32] datasets.
Convergence analysis. From Tab. 3 and Tab. 4, we analyze the convergence of different methods by comparing the results after training for 100/300 epochs. The methods can be categorized into three groups: 1) Methods that show no significant improvement in performance after 300 epochs compared to 100 epochs, indicating rapid convergence within 100 epochs. These models include DRAEM [37], SimpleNet [27], CFA [21], PyramidFlow [22], ViTAD [40], InvAD-lite [42], RD++ [31], and DeSTSeg [44].2) Methods that show improvement with continued training on the VisA dataset but no improvement or a decline on the MVTec AD dataset, indicating slower convergence on larger datasets. These models include RealNet and RD. 3) Methods that show significant improvement after 300 epochs compared to 100 epochs, indicating slower convergence. These models include CFLOW-AD [14], DiAD [16], InvAD [42], MambaAD [15], and UniAD [35].
Stability analysis. For current anomaly detection algorithms, most authors select the best epoch’s results as the model’s performance. However, this method of epoch selection is unscientific and may indicate significant model instability. Therefore, we further analyze model stability using Tab. 3 and Tab. 4, comparing the results at 100/300 epochs to identify any substantial differences. The results show that SimpleNet [27] and PyramidFlow [22] exhibit considerable differences, indicating poor model stability, while other methods do not show significant fluctuations.
Figure 4: A Pearson correlation coefficient analysis for different methods on several datasets.
Cross-domain dataset correlation. To analyze the adaptability of different methods across various datasets and the relationships and differences between different types of datasets, we employ Pearson correlation analysis to examine the correlations among these datasets. Specifically, we select four distinct datasets for analysis: MVTec AD [3], Uni-Medical [40], Real-IAD [32], and COCO-AD [42]. MVTec AD represents a fundamental industrial dataset, Uni-Medical consists of medical images from CT scans, Real-IAD is a large-scale multi-view industrial dataset from real-world scenarios, and COCO-AD is a large-scale panoptic segmentation dataset from real-life scenes. We evaluate four categories of methods using eight metrics: image and pixel mAU-ROC, mAP, mF1-max, region mAU-PRO, and segmentation mIoU-max. The results, as shown in Fig. 4, indicate that the COCO-AD and Uni-Medical datasets exhibit lower Pearson correlation coefficients due to significant differences in data distribution compared to general industrial datasets. Although the Pearson correlation coefficient between the Uni-Medical and Real-IAD datasets is relatively high, Tab. 5 and Tab. A11 in the Appendix reveal that this is because all methods perform poorly on these two datasets. Additionally, it is observed that the methods SimpleNet and DeSTSeg show considerable instability in their results across different datasets. This instability may be attributed to the inherent instability of the data augmentation algorithms they employ.
Training-free PatchCore. PatchCore [29] does not require model training. It extracts all features from the training data, then selects a core subset and stores it in a Memory Bank. During testing, each test image is compared with the Memory Bank to compute an anomaly score. Because it stores the core subset of all normal features in the Memory Bank, PatchCore is only feasible for multi-class anomaly detection tasks on small-scale datasets. For large-scale datasets, it faces limitations due to insufficient GPU and memory resources. Although it achieves excellent results on the MVTec AD dataset, as shown in Tab. 1, its testing time is nearly a thousand times longer than other methods. In summary, PatchCore performs exceptionally well on small-scale datasets but is constrained by large-scale datasets and testing time.
Dataset Analysis. The experimental results indicate that there is room for improvement in the VisA [47] and Real-IAD [32] datasets due to the very small defect areas, necessitating models with stronger capabilities for detecting minor defects. The MAD_Real and MAD_Sim [46] datasets, due to their small data volume and varying difficulty levels, result in similar performance across all models, particularly in the mF1-max metric. The Uni-Medical [40] dataset, consisting of images converted from CT scans, has a data distribution that significantly differs from other industrial datasets, suggesting the need for specialized detection networks tailored to medical datasets. COCO-AD [42], as a newly proposed large-scale dataset for general scenarios, presents high complexity. Current industrial AD networks are unable to achieve effective results on the COCO-AD.
4.3 Challenges for Current VAD
Immature method. For challenging anomaly detection datasets such as MVTecLOCO, pose-agnostic MAD, and general-purpose COCO-AD, current methods perform poorly in a multi-class setting. Future research should focus on designing more robust methods to address this issue.
Efficiency. Most methods do not consider model complexity during design, resulting in high FLOPs. This issue becomes more pronounced when applied to real-world high-resolution scenarios. Incorporating lightweight characteristics in model design could be a potential solution.
Dataset scale. Mainstream datasets in the VAD field, such as MVTec AD and Real-IAD, are relatively small compared to those in detection and segmentation fields and are tailored to specific industrial scenarios. This limitation could hinder technological development. Collecting larger-scale, general-scene AD datasets is crucial for the advancement of the VAD field.
VAD-specific metric. Metrics like mAU-ROC and mAP are not uniquely designed for the BAD field. Developing more reliable evaluation methods to better meet practical application needs is essential.
Augmentation and tricks. In fields such as classification, detection, and segmentation, data augmentation and tricks are extremely important for model training. However, few studies explore their role in the AD field, potentially limiting model performance.
Model interpretability. In many applications, understanding why a model detects a particular anomaly is crucial.
5 Conclusion and Discussion
This paper addresses the urgent need for a comprehensive and fair benchmark in the field of visual anomaly detection. We introduce a modular and scalable ADer library designed to fairly facilitate the evaluation of fifteen advanced VAD methods across multiple mainstream datasets. ensuring a thorough and unbiased assessment of each method’s performance. Our extensive experiments reveal the strengths and weaknesses of different methods, providing valuable insights into their efficiency and training resource consumption. We also develop and open-source a GPU-assisted ADEval package to reduce the evaluation time, enabling extensive assessments. Experimental results highlight the challenges of various VAD methods and offer valuable insights for future research directions.
Broader Impacts. The open-sourcing ADer can accelerate the development of new VAD technology for the open-source community and become a valuable resource for practitioners in the field.
Appendix B Detailed Quantitative Results on All Dataset Under 100 epochs
Table A1: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 54.5 76.3 83.6 47.6 03.2 06.7 14.3 03.5 40.9
SimpleNet [27] 95.4 98.3 95.7 96.8 48.8 51.9 86.9 36.4 82.0
RealNet [45] 84.8 94.1 90.9 72.6 48.2 41.4 56.8 28.8 69.8
Emb. CFA [21] 57.6 78.3 84.7 54.8 11.9 14.7 25.3 08.9 46.8
PatchCore [29] 98.8 99.5 98.4 98.3 59.9 61.0 94.2 44.9 87.2
CFLOW-AD [14] 91.6 96.7 93.4 95.7 45.9 48.6 88.3 33.2 80.0
PyramidalFlow [22] 70.2 85.5 85.5 80.0 22.3 22.0 47.5 12.8 59.0
Rec. RD [10] 93.6 97.2 95.6 95.8 48.2 53.6 91.2 37.0 82.2
DiAD [16] 88.9 95.8 93.5 89.3 27.0 32.5 63.9 21.1 70.1
ViTAD [40] 98.3 99.3 97.3 97.6 55.2 58.4 92.0 42.3 85.4
InvAD [42] 98.1 99.0 97.6 98.0 56.3 59.2 94.4 42.8 86.1
InvAD-lite [42] 97.9 99.2 96.8 97.3 54.4 57.8 93.3 41.4 85.2
MambaAD [15] 97.8 99.3 97.3 97.4 55.1 57.6 93.4 41.2 85.4
Hybrid UniAD [35] 92.5 97.3 95.4 95.8 42.7 48.0 89.3 32.5 80.1
RD++ [31] 97.9 98.8 96.4 97.3 54.7 58.0 93.2 41.5 85.2
DesTSeg [44] 96.4 98.6 96.2 92.0 71.1 68.2 83.4 52.8 86.6
Table A2: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 55.7 81.8 88.8 46.6 00.8 01.9 21.9 01.0 42.5
SimpleNet [27] 75.8 92.3 90.4 94.7 17.3 23.4 81.0 13.9 67.8
Emb. CFA [21] 67.1 89.3 89.6 71.6 09.7 16.1 43.2 09.3 55.2
PatchCore [29] 84.1 95.1 92.5 98.6 33.7 38.5 94.4 24.5 76.7
CFLOW-AD [14] 73.1 91.0 90.2 96.8 21.6 26.6 89.0 15.8 69.8
PyramidalFlow [22] 58.7 85.6 88.6 90.8 07.2 12.0 74.9 06.4 59.7
Rec. RD [10] 75.8 91.2 91.5 98.4 29.6 35.5 93.9 22.2 73.7
DiAD [16] 75.4 91.0 90.9 91.7 05.3 10.2 74.9 05.4 62.8
ViTAD [40] 79.6 93.4 91.6 98.0 27.2 33.3 91.6 20.4 73.5
InvAD [42] 86.9 96.2 93.3 98.7 37.4 42.4 95.2 27.4 78.6
InvAD-lite [42] 84.7 95.6 92.6 98.6 38.3 41.9 94.2 26.9 78.0
MambaAD [15] 84.1 95.1 92.2 98.6 36.9 40.8 94.2 25.9 77.4
Hybrid UniAD [35] 75.3 92.1 90.3 96.6 21.2 27.9 88.9 16.7 70.3
RD++ [31] 81.8 93.9 92.8 98.5 34.4 39.6 94.1 25.2 76.4
DesTSeg [44] 79.1 93.8 90.7 81.1 41.3 42.7 64.0 28.4 70.4
Table A3: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 52.5 65.4 77.6 46.7 04.5 10.4 12.4 05.6 38.5
SimpleNet [27] 81.6 88.5 82.9 76.5 29.0 32.7 63.8 21.2 65.0
Emb. CFA [21] 66.5 78.6 77.9 58.5 14.2 16.3 31.1 09.1 49.0
PatchCore [29] 80.5 89.0 81.5 75.1 29.9 31.8 69.9 20.4 65.4
CFLOW-AD [14] 75.1 84.2 79.8 70.7 25.0 27.9 69.5 17.3 61.7
PyramidalFlow [22] 53.4 69.3 77.5 57.7 15.0 14.5 25.7 08.0 44.7
Rec. RD [10] 74.6 85.3 78.9 70.7 21.0 26.1 67.7 15.8 60.6
DiAD [16] 71.3 79.5 80.9 71.8 19.0 24.8 48.4 14.9 56.5
ViTAD [40] 76.4 85.2 79.6 73.4 28.7 31.2 63.1 19.8 62.5
InvAD [42] 82.0 89.9 82.0 76.4 31.1 34.8 73.0 23.1 67.0
InvAD-lite [42] 79.4 87.8 80.9 75.8 28.5 32.2 69.4 20.6 64.9
MambaAD [15] 75.8 85.5 81.0 76.5 28.8 32.5 70.2 20.6 64.3
Hybrid UniAD [35] 73.2 82.1 79.6 73.1 30.7 33.4 64.7 21.6 62.4
RD++ [31] 76.4 86.8 79.2 71.5 24.3 28.3 68.8 17.6 62.2
DesTSeg [44] 77.9 86.7 81.3 64.1 35.2 30.5 53.4 20.3 61.3
Table A4: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 55.1 62.4 72.9 37.5 00.6 01.7 10.0 00.9 34.3
SimpleNet [27] 86.4 89.1 82.8 96.6 34.0 37.8 79.2 25.7 72.3
RealNet [45] 71.4 79.5 74.7 61.0 25.7 22.6 27.4 13.5 51.8
Emb. CFA [21] 66.3 74.3 74.2 81.3 22.1 26.2 50.8 17.0 56.5
CFLOW-AD [14] 86.5 88.8 84.9 97.7 33.9 37.2 86.8 24.9 73.7
PyramidalFlow [22] 58.2 66.3 74.4 77.0 07.2 09.6 42.8 05.6 47.9
Rec. RD [10] 90.6 90.9 89.3 98.0 35.4 42.5 91.9 27.9 76.9
DiAD [16] 84.8 88.5 86.9 82.5 17.9 23.2 44.5 14.9 61.2
ViTAD [40] 90.4 91.1 86.0 98.2 36.4 41.0 85.7 27.5 75.5
InvAD [42] 95.4 95.7 91.6 98.9 43.3 46.8 93.1 32.5 80.7
InvAD-lite [42] 94.9 95.2 90.7 98.6 40.2 44.0 93.1 29.8 79.5
MambaAD [15] 94.5 94.9 90.2 98.4 39.3 43.7 92.1 29.5 79.0
Hybrid UniAD [35] 89.0 91.0 85.8 98.3 34.5 39.6 86.5 26.4 75.0
RD++ [31] 93.9 94.7 90.2 98.4 42.3 46.3 91.9 31.2 79.7
DesTSeg [44] 89.9 91.4 86.7 86.7 46.6 47.2 61.1 32.7 72.8
Table A5: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 71.3 78.5 78.0 49.0 03.7 06.5 16.2 03.4 43.3
SimpleNet [27] 93.2 97.3 93.3 96.3 41.5 44.3 69.8 28.6 76.5
RealNet [45] 89.7 95.3 92.8 84.0 48.1 52.7 53.4 36.6 73.7
Emb. CFA [21] 92.7 97.5 93.5 96.3 47.4 50.2 69.5 33.6 78.2
PatchCore [29] 94.4 98.2 94.6 97.5 55.0 54.9 76.0 38.0 81.5
CFLOW-AD [14] 91.2 94.8 88.3 96.8 45.6 50.1 72.7 33.8 77.1
PyramidalFlow [22] 87.0 83.1 81.0 90.9 29.6 26.9 64.1 18.3 66.1
Rec. RD [10] 94.4 96.6 94.0 98.1 59.6 59.2 80.7 42.1 83.2
DiAD [16] 90.1 88.4 92.6 91.7 19.6 26.7 70.4 15.7 68.5
ViTAD [40] 93.6 96.8 93.3 97.6 59.2 56.7 73.2 40.1 81.5
InvAD [42] 96.1 97.8 94.3 98.1 62.6 61.4 80.9 44.3 84.5
InvAD-lite [42] 93.1 97.4 94.6 97.9 59.2 59.7 78.6 42.6 82.9
MambaAD [15] 93.1 96.1 92.5 97.7 52.5 56.0 78.2 39.0 80.9
Hybrid UniAD [35] 94.8 98.3 94.2 97.2 50.3 53.8 78.8 36.9 81.1
RD++ [31] 94.6 97.8 94.1 98.0 59.6 59.8 79.0 42.8 83.3
DesTSeg [44] 92.8 95.9 92.3 92.2 34.8 44.3 70.0 29.0 74.6
Table A6: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 35.6 53.3 72.5 42.1 02.2 04.6 19.0 02.5 32.8
SimpleNet [27] 88.4 92.0 87.9 96.5 32.0 34.6 89.0 24.5 74.3
RealNet [45] 85.1 90.2 88.3 83.3 36.1 39.6 68.1 28.2 70.1
Emb. CFA [21] 81.6 87.7 85.7 84.9 19.6 22.9 53.5 16.6 62.3
PatchCore [29] 94.6 96.9 93.4 98.9 46.1 47.6 95.7 35.0 81.9
CFLOW-AD [14] 75.7 80.1 81.7 96.8 26.3 28.0 89.5 20.1 68.3
PyramidalFlow [22] 73.6 77.0 79.4 94.1 21.1 17.8 77.2 10.4 62.9
Rec. RD [10] 91.3 93.6 91.8 98.3 40.4 41.8 95.5 31.4 79.0
DiAD [16] 68.3 77.9 80.1 90.4 10.9 13.1 66.1 08.2 58.1
ViTAD [40] 87.8 90.5 88.0 97.7 35.2 37.4 92.8 27.7 75.6
InvAD [42] 93.7 93.9 93.0 98.2 42.4 45.7 94.8 34.0 80.2
InvAD-lite [42] 90.9 92.9 89.5 98.0 39.7 42.6 94.0 30.9 78.2
MambaAD [15] 88.7 93.2 90.8 97.5 33.6 38.1 92.3 26.8 76.3
Hybrid UniAD [35] 70.5 76.0 78.0 93.9 13.7 19.9 79.7 12.5 61.7
RD++ [31] 90.2 93.3 90.5 98.5 43.0 44.1 95.5 33.6 79.3
DesTSeg [44] 91.3 90.8 90.2 82.0 32.6 34.6 63.3 25.6 69.3
Table A7: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 59.1 88.1 90.4 49.4 00.6 01.7 22.3 00.8 44.5
SimpleNet [27] 47.8 81.9 90.7 83.1 05.4 11.6 59.8 06.3 54.3
Emb. CFA [21] 49.3 81.7 91.4 83.3 10.5 15.2 64.2 08.7 56.5
PatchCore [29] 59.7 84.8 93.5 95.4 20.2 27.3 85.5 16.6 66.6
CFLOW-AD [14] 59.8 86.0 91.2 90.9 10.7 15.4 74.4 08.8 61.2
PyramidalFlow [22] 52.1 83.6 90.3 88.4 05.0 09.4 68.2 05.1 56.7
Rec. RD [10] 55.5 85.8 90.6 87.0 08.2 12.8 67.8 07.2 58.2
DiAD [16] 49.9 85.1 90.8 87.7 02.9 06.8 67.4 03.6 55.8
ViTAD [40] 47.0 81.3 90.2 87.3 04.5 09.3 66.6 05.0 55.2
InvAD [42] 66.5 87.7 93.8 88.3 22.1 26.2 73.2 16.8 65.4
InvAD-lite [42] 60.4 84.3 92.2 88.2 09.0 14.3 72.9 08.1 60.2
MambaAD [15] 56.2 84.3 90.3 87.8 07.5 12.8 68.5 07.2 58.2
Hybrid UniAD [35] 47.6 81.3 90.4 88.2 05.8 10.6 68.6 05.8 56.1
RD++ [31] 54.7 83.9 92.3 88.5 09.4 14.9 70.4 08.5 59.2
DesTSeg [44] 45.4 80.3 90.7 72.7 05.0 08.2 48.3 04.5 50.1
Table A8: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 49.5 88.3 93.8 48.4 00.8 01.5 24.5 00.7 43.8
SimpleNet [27] 58.9 91.7 93.8 85.5 04.2 08.0 63.9 04.2 58.0
Emb. CFA [21] 56.4 91.1 93.8 76.6 04.2 08.7 54.7 04.6 55.1
CFLOW-AD [14] 55.8 90.6 93.8 86.9 03.4 05.2 63.9 02.7 57.1
PyramidalFlow [22] 50.8 88.9 93.8 82.0 02.3 04.8 58.8 02.5 54.5
Rec. RD [10] 58.6 91.7 93.8 88.3 04.7 08.5 74.2 04.5 60.0
DiAD [16] 65.7 93.3 94.1 87.2 03.9 07.9 60.1 04.2 58.9
ViTAD [40] 58.5 91.5 93.8 89.0 05.0 09.5 73.5 05.0 60.1
InvAD [42] 70.4 94.6 93.9 90.5 08.8 15.7 79.3 08.6 64.7
InvAD-lite [42] 64.7 93.3 93.8 89.4 06.4 11.2 76.1 06.0 62.1
MambaAD [15] 63.0 92.9 93.8 88.3 05.2 09.4 73.3 05.0 60.8
Hybrid UniAD [35] 54.2 90.0 93.8 88.1 03.5 06.8 71.8 03.5 58.3
RD++ [31] 59.3 91.7 93.8 88.3 04.6 08.5 73.8 04.4 60.0
DesTSeg [44] 57.9 91.8 93.8 54.5 06.2 07.7 35.8 04.1 49.7
Table A9: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 51.6 60.6 70.6 33.8 02.0 05.7 10.5 03.0 33.5
SimpleNet [27] 77.5 77.7 76.7 94.3 34.4 36.0 77.0 23.3 67.7
Emb. CFA [21] 71.0 73.3 73.1 78.3 19.5 24.9 44.7 14.7 55.0
CFLOW-AD [14] 79.2 79.6 77.2 95.4 40.4 26.9 81.8 17.7 68.6
PyramidalFlow [22] 53.9 57.3 70.5 78.8 08.4 17.1 34.2 09.4 45.7
Rec. RD [10] 76.1 75.7 78.2 96.5 38.8 39.8 86.8 26.9 70.3
DiAD [16] 78.8 77.2 77.7 95.8 34.2 35.5 84.3 23.2 69.1
ViTAD [40] 81.8 80.7 80.0 97.1 48.3 48.2 86.7 33.7 74.7
InvAD [42] 82.4 80.5 80.5 97.3 46.2 46.1 89.3 32.6 74.6
InvAD-lite [42] 77.9 77.3 79.1 96.2 39.0 39.3 85.8 26.5 70.7
MambaAD [15] 83.9 80.8 81.9 96.8 45.8 47.5 88.2 33.5 75.0
Hybrid UniAD [35] 79.0 76.1 77.1 96.6 39.3 41.1 86.0 27.6 70.7
RD++ [31] 77.2 76.5 79.2 96.7 41.5 42.2 87.8 29.4 71.6
DesTSeg [44] 78.5 77.0 78.2 65.7 41.7 34.0 35.3 21.2 58.6
Table A10: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 50.9 45.9 61.3 44.0 00.2 00.4 13.6 00.2 30.9
SimpleNet [27] 54.9 50.6 61.5 76.1 01.9 04.9 42.4 02.5 41.8
Emb. CFA [21] 55.7 50.5 61.9 81.3 01.6 03.8 48.8 02.0 43.4
CFLOW-AD [14] 77.0 75.8 69.9 94.8 17.6 21.7 80.4 12.4 62.5
PyramidalFlow [22] 54.4 48.0 62.0 71.1 01.2 01.1 34.9 00.5 39.0
Rec. RD [10] 82.7 79.3 74.1 97.2 25.2 32.8 90.0 20.0 68.8
DiAD [16] 75.6 66.4 69.9 88.0 02.9 07.1 58.1 03.7 52.6
ViTAD [40] 82.7 80.2 73.7 97.2 24.3 32.3 84.8 19.6 67.9
InvAD [42] 89.4 87.0 80.2 98.4 32.6 38.9 92.7 24.6 74.2
InvAD-lite [42] 87.2 85.2 77.8 98.0 31.7 37.9 92.0 23.8 72.8
MambaAD [15] 87.0 85.3 77.6 98.6 32.4 38.1 91.2 23.9 72.9
Hybrid UniAD [35] 83.1 81.2 74.5 97.4 23.3 30.9 87.1 18.6 68.2
RD++ [31] 83.6 80.6 74.8 97.7 25.9 33.6 90.7 20.5 69.6
DesTSeg [44] 79.3 76.7 70.7 80.3 36.9 40.3 56.1 26.2 62.9
Table A11: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 53.5 48.5 62.5 49.9 07.6 14.4 15.3 08.0 36.0
SimpleNet [27] 55.6 49.4 62.3 60.2 14.5 19.2 26.1 11.5 41.0
Emb. CFA [21] 56.7 51.2 62.6 56.2 10.3 16.0 17.9 08.9 38.7
CFLOW-AD [14] 67.7 57.9 64.5 76.0 20.3 26.4 47.7 16.0 51.5
PyramidalFlow [22] 51.6 47.4 61.6 50.0 13.4 14.4 15.0 08.0 36.2
Rec. RD [10] 57.6 49.9 62.0 66.5 13.9 20.0 39.8 11.5 44.2
DiAD [16] 59.0 53.0 63.2 68.1 20.5 14.2 30.8 11.6 44.1
ViTAD [40] 66.9 59.3 63.7 76.2 27.6 32.2 39.1 20.1 52.1
InvAD [42] 64.2 56.0 63.4 71.2 18.5 24.1 45.8 14.3 49.0
InvAD-lite [42] 63.3 56.0 62.9 69.9 18.2 23.1 39.7 13.7 47.6
MambaAD [15] 62.8 55.2 62.8 68.9 16.7 22.0 41.6 12.9 47.1
Hybrid UniAD [35] 55.2 49.3 61.7 64.6 12.8 19.0 34.3 10.9 42.4
RD++ [31] 57.5 49.3 62.0 68.2 15.5 20.1 42.2 11.8 45.0
DesTSeg [44] 54.4 49.4 61.8 54.5 10.6 15.4 24.4 08.5 38.6
Appendix C Detailed Quantitative Results on All Dataset Under 300 epochs
Table A12: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 55.2 77.0 83.9 48.7 03.1 06.3 15.8 03.3 41.4
SimpleNet [27] 79.2 90.8 87.6 82.4 24.0 29.0 62.0 17.8 65.0
RealNet [45] 82.9 93.3 90.9 69.8 50.0 40.4 51.2 28.5 68.4
Emb. CFA [21] 55.8 78.8 84.5 43.9 04.8 08.9 19.3 04.7 42.3
PatchCore [29] 98.8 99.5 98.4 98.3 59.9 61.0 94.2 44.9 88.6
CFLOW-AD [14] 92.7 97.2 94.0 95.8 46.8 49.6 89.0 34.0 80.7
PyramidalFlow [22] 66.2 84.3 85.1 74.2 17.2 19.6 40.0 11.4 55.2
Rec. RD [10] 90.5 95.0 95.1 95.9 47.1 52.1 91.2 35.8 81.0
DiAD [16] 92.0 96.8 94.4 89.3 27.3 32.7 64.4 21.3 71.0
ViTAD [40] 98.4 99.4 97.5 97.5 55.2 58.1 91.7 42.0 85.4
InvAD [42] 98.9 99.6 98.2 98.1 57.1 59.6 94.4 43.1 86.6
InvAD-lite [42] 98.1 99.1 96.8 97.3 55.0 58.1 93.1 41.7 85.4
MambaAD [15] 98.5 99.5 97.7 97.6 56.1 58.7 93.6 42.3 86.0
Hybrid UniAD [35] 96.8 98.9 97.0 96.8 45.0 50.2 91.0 34.2 82.2
RD++ [31] 95.8 98.0 96.6 97.3 53.0 57.0 92.9 40.5 84.4
DesTSeg [44] 96.3 98.8 96.1 92.6 75.8 71.3 82.6 56.6 87.6
Table A13: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 49.2 79.4 89.0 50.7 00.7 02.2 23.4 01.1 42.1
SimpleNet [27] 74.9 91.8 90.2 94.0 14.4 21.4 79.2 12.4 66.6
Emb. CFA [21] 60.1 86.3 89.5 52.0 07.0 11.9 28.7 06.9 47.9
PatchCore [29] 84.1 95.1 92.5 98.6 33.7 38.5 94.4 24.5 78.4
CFLOW-AD [14] 74.8 91.4 90.5 96.9 21.6 26.8 89.2 15.9 70.2
PyramidalFlow [22] 52.5 80.8 88.6 87.2 05.9 10.4 68.0 05.6 56.2
Rec. RD [10] 79.3 93.0 92.3 98.3 28.6 35.4 93.5 22.1 74.3
DiAD [16] 78.8 91.3 91.7 91.8 05.2 10.5 75.1 05.6 63.5
ViTAD [40] 78.7 93.3 91.5 98.0 27.0 33.0 91.3 20.2 73.3
InvAD [42] 86.3 95.9 93.4 98.7 36.2 41.1 94.9 26.4 78.1
InvAD-lite [42] 85.1 95.6 92.3 98.4 34.4 39.0 93.4 24.6 76.9
MambaAD [15] 85.8 95.5 92.6 98.6 37.1 40.8 94.1 26.0 77.8
Hybrid UniAD [35] 77.0 92.6 91.0 96.8 23.7 30.6 89.4 18.6 71.6
RD++ [31] 80.1 93.5 92.2 98.4 31.5 37.2 93.6 23.4 75.2
DesTSeg [44] 82.5 95.1 90.9 82.1 43.8 45.2 65.1 30.2 72.1
Table A14: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 51.4 66.9 77.5 48.0 05.2 10.5 13.7 05.6 39.0
SimpleNet [27] 81.8 89.1 82.5 70.9 28.3 32.1 61.2 20.9 63.7
Emb. CFA [21] 60.6 73.7 77.6 64.7 16.5 18.8 36.4 10.8 49.8
PatchCore [29] 80.5 89.0 81.5 75.1 29.9 31.8 69.9 20.4 67.7
CFLOW-AD [14] 76.0 85.0 79.9 70.9 25.6 28.4 70.0 17.7 62.3
PyramidalFlow [22] 58.1 70.4 77.4 55.5 16.8 14.4 20.4 08.1 44.7
Rec. RD [10] 73.7 84.6 79.4 70.7 21.0 26.2 67.5 15.9 60.4
DiAD [16] 77.2 84.6 81.5 72.1 18.6 25.3 54.4 15.2 59.1
ViTAD [40] 76.2 85.0 80.1 73.2 27.9 30.6 62.0 19.3 62.1
InvAD [42] 82.3 89.9 82.4 77.7 30.9 34.1 72.8 22.4 67.2
InvAD-lite [42] 78.0 86.8 81.6 77.0 28.0 32.0 69.4 20.4 64.7
MambaAD [15] 78.2 87.1 81.8 77.5 28.0 32.9 68.6 20.8 64.9
Hybrid UniAD [35] 78.7 87.1 80.9 74.6 32.6 35.8 70.7 23.4 65.8
RD++ [31] 75.7 85.9 79.3 72.1 24.2 28.5 67.9 17.8 61.9
DesTSeg [44] 81.2 88.6 81.8 63.7 31.8 31.9 59.2 21.1 62.6
Table A15: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 56.2 64.6 74.9 45.0 00.7 01.8 16.0 00.9 37.0
SimpleNet [27] 80.7 83.8 79.3 94.4 29.2 33.1 74.2 22.1 67.8
RealNet [45] 79.2 84.8 78.3 65.4 29.2 27.9 33.9 17.4 57.0
Emb. CFA [21] 67.1 73.8 75.3 83.0 13.7 18.7 48.7 11.3 54.3
CFLOW-AD [14] 87.2 89.5 85.1 97.8 34.2 37.2 87.3 24.9 74.0
PyramidalFlow [22] 69.0 72.9 75.8 79.1 07.9 08.7 52.6 04.7 52.3
Rec. RD [10] 93.9 94.8 90.4 98.1 38.4 43.7 91.9 29.0 78.7
DiAD [16] 90.5 91.4 90.4 83.4 19.2 25.0 44.3 16.2 63.5
ViTAD [40] 90.3 91.2 86.4 98.2 36.4 40.9 85.8 27.5 75.6
InvAD [42] 95.6 96.0 92.3 99.0 43.7 46.9 93.0 32.6 80.9
InvAD-lite [42] 95.3 95.8 91.0 98.7 41.2 44.9 93.2 30.6 80.0
MambaAD [15] 93.6 93.9 89.8 98.2 34.0 39.3 90.5 25.9 77.0
Hybrid UniAD [35] 91.4 93.3 87.5 98.5 35.3 40.2 89.0 26.5 76.5
RD++ [31] 93.1 94.1 90.0 98.4 40.4 44.8 91.4 29.9 78.9
DesTSeg [44] 89.0 90.3 85.9 84.8 43.3 44.4 57.5 30.1 70.7
Table A16: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 69.3 76.8 79.5 47.7 03.2 06.1 16.1 03.2 42.7
SimpleNet [27] 94.0 97.9 93.9 96.2 41.0 43.7 69.6 28.1 76.6
RealNet [45] 93.1 96.3 92.5 87.2 48.0 55.5 57.9 38.7 75.8
Emb. CFA [21] 92.7 97.8 94.6 96.2 46.7 49.7 69.0 33.2 78.1
PatchCore [29] 94.4 98.2 94.6 97.5 55.0 54.9 76.0 38.0 83.3
CFLOW-AD [14] 91.6 95.6 89.7 96.9 46.0 49.0 72.7 33.0 77.4
PyramidalFlow [22] 85.3 81.7 79.0 76.1 27.7 23.4 43.1 15.0 59.5
Rec. RD [10] 94.1 96.8 93.8 98.0 57.1 58.0 79.9 41.0 82.5
DiAD [16] 90.2 88.3 92.6 91.9 20.5 27.0 70.3 16.0 68.7
ViTAD [40] 94.0 97.0 93.7 97.6 58.3 56.5 72.8 39.9 81.4
InvAD [42] 95.3 97.3 93.7 97.9 58.7 58.8 78.8 41.7 82.9
InvAD-lite [42] 94.2 97.9 94.4 97.7 55.7 57.0 77.4 40.0 82.0
MambaAD [15] 92.9 96.2 93.0 97.6 51.2 55.1 77.3 38.2 80.5
Hybrid UniAD [35] 94.5 98.4 94.9 97.4 52.4 55.5 78.9 38.4 81.7
RD++ [31] 94.5 97.6 94.3 98.0 57.7 58.2 78.9 41.2 82.7
DesTSeg [44] 93.5 96.7 93.8 94.8 39.1 38.5 72.9 24.6 75.6
Table A17: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 60.2 67.5 75.1 50.4 05.4 04.7 21.8 02.6 40.7
SimpleNet [27] 90.6 94.1 89.7 97.1 33.6 35.7 90.0 25.6 75.8
RealNet [45] 86.0 90.0 87.3 74.7 39.2 39.7 52.3 28.0 67.0
Emb. CFA [21] 92.0 93.5 90.2 95.7 29.0 33.0 83.2 23.8 73.8
PatchCore [29] 94.6 96.9 93.4 98.9 46.1 47.6 95.7 35.0 83.5
CFLOW-AD [14] 80.4 85.1 84.7 97.3 28.4 30.1 90.9 21.5 71.0
PyramidalFlow [22] 72.5 76.2 77.6 89.0 12.4 15.8 68.5 09.1 58.9
Rec. RD [10] 90.3 92.8 90.5 98.3 39.6 40.6 95.2 30.2 78.2
DiAD [16] 85.8 89.2 86.5 91.4 15.3 19.2 66.1 12.0 64.8
ViTAD [40] 87.4 90.8 87.0 97.8 34.6 37.8 92.9 28.0 75.5
InvAD [42] 96.0 97.1 94.5 98.5 44.1 46.4 95.3 34.4 81.7
InvAD-lite [42] 92.8 93.2 91.4 98.3 41.4 43.9 94.6 32.1 79.4
MambaAD [15] 89.2 93.1 90.3 97.7 33.5 38.6 92.8 27.2 76.5
Hybrid UniAD [35] 80.1 83.2 85.1 95.4 19.0 25.6 83.8 16.8 67.5
RD++ [31] 90.6 93.3 90.0 98.4 42.8 43.0 95.3 32.7 79.1
DesTSeg [44] 92.6 91.8 92.8 90.8 30.6 32.9 78.3 24.1 72.8
Table A18: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 46.3 82.5 90.2 42.5 00.5 01.2 14.9 00.6 39.7
SimpleNet [27] 36.1 77.0 89.6 77.2 02.6 06.3 48.9 03.3 48.2
Emb. CFA [21] 48.5 81.2 91.4 83.3 10.4 15.0 64.1 08.6 56.3
PatchCore [29] 59.7 84.8 93.5 95.4 20.2 27.3 85.5 16.6 68.2
CFLOW-AD [14] 57.4 85.9 91.2 90.9 10.8 15.3 74.1 08.8 60.8
PyramidalFlow [22] 50.6 82.9 90.5 88.6 07.5 12.0 68.8 06.7 57.3
Rec. RD [10] 56.0 84.9 90.7 86.9 08.7 13.2 66.9 07.4 58.2
DiAD [16] 58.0 85.7 91.1 87.8 05.1 09.6 69.4 05.2 58.1
ViTAD [40] 44.5 81.3 90.2 86.9 04.5 08.9 65.0 04.8 54.5
InvAD [42] 63.5 86.5 92.7 91.6 18.9 24.2 77.2 14.9 64.9
InvAD-lite [42] 57.3 84.8 91.9 90.8 11.4 16.3 74.7 09.5 61.0
MambaAD [15] 54.7 83.1 92.2 90.8 13.1 19.2 74.7 11.4 61.1
Hybrid UniAD [35] 45.4 80.3 90.8 87.8 06.4 11.6 68.3 06.4 55.8
RD++ [31] 52.3 84.0 91.3 88.7 09.3 14.4 70.4 08.2 58.6
DesTSeg [44] 36.1 77.9 89.8 82.5 03.2 06.0 52.7 03.2 49.7
Table A19: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 48.8 88.1 93.8 47.9 00.7 01.5 20.0 00.8 43.0
SimpleNet [27] 51.5 88.6 93.8 86.1 02.3 04.5 63.9 02.3 55.8
Emb. CFA [21] 55.7 91.1 93.8 71.8 03.7 07.7 45.3 04.1 52.7
CFLOW-AD [14] 56.9 90.9 93.8 86.9 03.4 05.1 63.7 02.6 57.2
PyramidalFlow [22] 49.0 88.1 93.8 84.4 02.4 05.0 62.7 02.6 55.1
Rec. RD [10] 57.9 91.3 93.8 88.3 04.4 08.2 73.7 04.3 59.7
DiAD [16] 65.6 93.4 94.1 87.7 04.4 08.8 61.4 04.7 59.4
ViTAD [40] 59.0 91.4 93.8 89.0 04.8 09.3 73.6 04.9 60.1
InvAD [42] 70.4 94.5 94.0 90.7 08.9 15.3 79.1 08.4 64.7
InvAD-lite [42] 65.3 93.4 93.9 89.7 06.6 11.4 76.4 06.1 62.4
MambaAD [15] 61.4 92.4 93.8 88.3 04.8 09.0 73.0 04.8 60.4
Hybrid UniAD [35] 59.4 91.5 93.8 89.0 04.4 08.4 74.6 04.4 60.2
RD++ [31] 58.6 91.7 93.8 88.3 04.4 08.1 73.2 04.2 59.7
DesTSeg [44] 59.2 91.9 93.8 56.8 05.7 08.6 41.8 04.5 51.1
Table A20: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 58.5 63.2 71.7 32.5 02.0 06.1 09.1 03.2 34.7
SimpleNet [27] 72.5 75.3 76.7 94.9 35.9 37.0 79.2 24.2 67.4
Emb. CFA [21] 56.3 62.8 70.4 38.1 04.8 06.6 08.9 03.4 35.4
CFLOW-AD [14] 79.0 78.5 77.3 95.4 39.6 26.3 81.7 17.0 68.3
PyramidalFlow [22] 60.0 62.8 71.9 82.2 08.8 14.9 46.9 08.3 49.6
Rec. RD [10] 76.4 75.8 77.9 96.4 38.9 39.8 86.5 27.0 70.2
DiAD [16] 82.9 81.9 79.4 96.0 35.9 38.2 85.4 25.2 71.4
ViTAD [40] 81.5 80.6 80.1 97.0 46.8 46.9 86.5 32.7 74.2
InvAD [42] 81.3 80.0 79.9 97.2 45.4 45.5 89.0 32.3 74.0
InvAD-lite [42] 77.6 77.2 79.3 96.2 38.7 39.7 85.7 27.1 70.6
MambaAD [15] 82.3 78.9 81.3 96.7 43.0 46.1 87.9 32.4 73.7
Hybrid UniAD [35] 80.4 76.6 77.5 96.5 39.0 41.3 85.8 27.8 71.0
RD++ [31] 77.2 76.1 78.5 96.7 41.0 41.9 87.4 29.1 71.3
DesTSeg [44] 82.0 85.0 79.7 83.2 50.6 50.5 66.1 34.1 71.0
Appendix D Detailed Quantitative Results on MVTec AD Dataset Under 100 epochs
Table A21: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bottle Aug. DRAEM [37] 49.4 74.0 86.3 57.8 06.5 13.2 17.3 07.1 43.5
SimpleNet [27] 100. 100. 100. 97.4 57.3 65.9 89.1 49.1 87.1
RealNet [45] 95.6 98.8 95.1 69.8 53.9 46.6 60.9 30.4 74.4
Emb. CFA [21] 64.2 87.4 86.3 65.3 11.0 15.1 35.5 08.2 52.1
PatchCore [29] 100. 100. 100. 98.9 79.0 77.1 96.1 62.7 93.0
CFLOW-AD [14] 99.9 100. 99.2 97.4 61.8 63.6 91.9 46.6 87.7
PyramidalFlow [22] 77.1 93.1 86.3 77.7 16.9 23.9 40.9 13.6 59.4
Rec. RD [10] 40.6 77.5 86.3 93.4 29.6 48.5 80.4 32.0 65.2
DiAD [16] 97.7 99.1 99.2 93.5 45.7 49.5 76.0 32.9 80.1
ViTAD [40] 100. 100. 100. 98.8 79.9 75.8 94.9 61.0 92.8
InvAD [42] 100. 100. 100. 98.6 75.2 73.6 95.7 58.2 91.9
InvAD-lite [42] 100. 100. 100. 98.4 75.2 73.5 95.7 58.1 91.8
MambaAD [15] 99.8 100. 99.2 98.7 79.7 75.7 96.0 60.9 92.7
Hybrid UniAD [35] 99.4 99.8 98.4 98.1 73.5 70.7 95.3 54.7 90.7
RD++ [31] 99.8 99.9 98.4 98.2 72.5 71.6 95.0 55.8 90.8
DesTSeg [44] 100. 100. 100. 94.9 87.2 81.1 91.8 68.3 93.6
cable Aug. DRAEM [37] 50.2 61.7 76.0 55.6 03.2 05.7 16.6 02.9 38.4
SimpleNet [27] 97.1 98.3 93.9 96.6 48.5 55.0 86.0 37.9 82.2
RealNet [45] 70.0 83.0 76.0 61.5 23.5 25.9 33.3 14.9 53.3
Emb. CFA [21] 50.3 59.9 76.3 52.8 03.7 07.5 22.3 03.9 39.0
PatchCore [29] 99.8 99.9 99.5 97.7 51.6 55.2 93.6 38.1 85.3
CFLOW-AD [14] 89.6 93.7 86.0 89.8 27.3 33.3 78.0 20.0 71.1
PyramidalFlow [22] 58.2 73.7 76.0 82.8 19.0 16.4 41.6 08.9 52.5
Rec. RD [10] 81.9 88.7 81.4 84.0 26.6 34.4 77.3 20.8 67.8
DiAD [16] 88.5 94.3 88.9 90.7 27.0 33.3 63.9 20.0 69.5
ViTAD [40] 98.8 99.3 96.3 96.3 43.6 46.8 90.6 30.6 81.7
InvAD [42] 97.9 98.7 95.1 97.1 46.7 50.8 92.1 34.0 82.6
InvAD-lite [42] 97.0 98.2 94.0 90.9 39.4 47.5 87.7 31.2 79.2
MambaAD [15] 99.3 99.6 97.4 95.2 42.0 47.9 90.4 31.5 81.7
Hybrid UniAD [35] 93.0 95.7 89.3 96.1 45.0 50.7 84.9 34.0 79.2
RD++ [31] 93.7 96.3 91.3 93.2 37.7 45.1 86.8 29.1 77.7
DesTSeg [44] 97.5 98.7 94.6 95.2 67.2 64.0 85.1 47.1 86.0
capsule Aug. DRAEM [37] 58.0 86.3 90.5 48.3 00.8 01.8 14.8 00.9 42.9
SimpleNet [27] 85.4 96.7 92.7 98.1 36.9 46.8 87.1 30.6 77.7
RealNet [45] 64.4 90.3 90.5 54.6 23.7 12.1 23.4 06.4 51.3
Emb. CFA [21] 40.8 82.5 90.5 27.0 00.6 01.8 07.0 00.9 35.7
PatchCore [29] 99.2 99.8 98.6 99.0 49.0 52.9 95.7 35.9 84.9
CFLOW-AD [14] 86.1 96.7 93.4 98.5 41.0 44.2 92.8 28.4 79.0
PyramidalFlow [22] 55.6 87.0 90.5 90.3 13.5 19.6 57.3 10.9 59.1
Rec. RD [10] 98.4 99.7 96.9 98.8 45.4 51.8 96.1 34.9 83.9
DiAD [16] 88.6 97.7 91.8 90.9 16.0 23.5 54.5 13.3 66.1
ViTAD [40] 95.3 99.0 95.9 98.2 42.0 47.1 92.4 30.8 81.4
InvAD [42] 94.3 98.7 96.0 98.9 46.7 50.8 95.5 34.0 83.0
InvAD-lite [42] 93.9 98.6 94.8 98.4 42.2 47.4 94.4 31.1 81.4
MambaAD [15] 91.8 98.2 93.2 98.3 43.5 47.8 93.0 31.4 80.8
Hybrid UniAD [35] 69.8 90.4 91.6 97.5 32.6 36.9 86.9 22.6 72.2
RD++ [31] 98.3 99.7 97.3 99.0 47.1 52.1 96.3 35.3 84.3
DesTSeg [44] 90.3 97.6 94.6 89.3 56.4 58.1 60.6 41.0 78.1
carpet Aug. DRAEM [37] 66.7 87.4 87.0 51.9 01.7 03.2 16.6 01.6 44.9
SimpleNet [27] 97.1 99.1 94.6 97.7 42.8 46.9 90.0 30.6 81.2
RealNet [45] 96.5 99.1 96.0 89.2 69.0 64.3 84.0 47.4 85.4
Emb. CFA [21] 47.5 78.0 86.4 50.9 07.0 11.4 31.9 06.1 44.7
PatchCore [29] 98.4 99.5 97.8 99.1 66.4 64.6 95.2 47.7 88.7
CFLOW-AD [14] 99.2 99.8 97.2 98.8 55.9 59.4 94.3 42.3 86.4
PyramidalFlow [22] 44.6 73.5 86.4 79.2 08.5 15.1 52.3 08.2 51.4
Rec. RD [10] 99.0 99.7 98.3 99.1 59.8 61.1 95.8 44.0 87.5
DiAD [16] 64.5 87.8 86.4 87.9 17.0 26.5 67.0 15.3 62.4
ViTAD [40] 99.5 99.9 99.4 99.0 59.8 64.1 94.7 47.2 88.1
InvAD [42] 98.6 99.6 97.2 99.1 60.8 62.1 95.8 45.1 87.6
InvAD-lite [42] 99.8 99.9 99.4 99.1 58.0 61.4 96.4 44.3 87.7
MambaAD [15] 99.8 99.9 99.4 99.2 64.0 63.8 97.3 46.8 89.1
Hybrid UniAD [35] 99.8 99.9 99.4 98.5 53.6 55.6 95.9 38.5 86.1
RD++ [31] 97.6 99.3 96.0 98.9 57.2 58.5 94.8 41.3 86.0
DesTSeg [44] 95.9 98.7 95.0 95.7 68.7 64.9 93.4 48.0 87.5
Table A22: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
grid Aug. DRAEM [37] 47.3 72.2 84.4 46.4 00.6 01.4 13.5 00.7 38.0
SimpleNet [27] 98.2 99.5 98.2 96.9 26.4 33.0 88.3 19.7 77.2
RealNet [45] 97.2 99.2 96.4 82.6 41.2 45.7 77.7 29.7 77.1
Emb. CFA [21] 51.1 76.4 84.4 39.9 00.5 01.4 05.2 00.7 37.0
PatchCore [29] 99.4 99.8 98.2 98.9 40.8 44.3 94.0 28.4 82.2
CFLOW-AD [14] 86.7 95.5 88.0 92.9 18.2 25.1 81.0 14.4 69.6
PyramidalFlow [22] 75.5 89.9 85.1 85.7 09.8 16.6 66.9 09.0 61.4
Rec. RD [10] 98.7 99.6 97.4 99.2 47.6 48.8 97.5 32.3 84.1
DiAD [16] 99.0 99.7 98.2 85.7 04.0 09.5 56.7 05.0 64.7
ViTAD [40] 99.6 99.9 99.1 98.6 31.3 36.5 95.8 22.3 80.1
InvAD [42] 99.5 99.8 97.4 99.1 45.1 46.6 97.1 30.4 83.5
InvAD-lite [42] 100. 100. 100. 99.2 49.4 49.5 97.7 32.9 85.1
MambaAD [15] 99.8 99.9 99.1 98.9 48.4 48.5 96.9 32.0 84.5
Hybrid UniAD [35] 99.0 99.7 98.2 94.8 26.6 34.0 89.3 20.5 77.4
RD++ [31] 99.2 99.7 97.3 99.2 46.8 50.2 97.0 33.5 84.2
DesTSeg [44] 99.6 99.9 98.2 96.6 47.7 52.0 91.9 35.1 83.7
hazelnut Aug. DRAEM [37] 45.7 68.1 77.8 41.1 03.6 04.2 12.4 02.1 36.1
SimpleNet [27] 100. 100. 100. 98.1 49.0 52.4 93.9 35.5 84.8
RealNet [45] 100. 100. 100. 77.5 44.2 48.9 75.4 32.4 78.0
Emb. CFA [21] 51.8 67.7 78.0 45.7 04.5 08.3 17.3 04.3 39.0
PatchCore [29] 100. 100. 100. 99.0 61.2 63.6 96.2 46.6 88.6
CFLOW-AD [14] 99.9 100. 99.3 98.5 59.1 57.9 95.6 40.7 87.2
PyramidalFlow [22] 92.1 95.0 91.2 92.7 33.2 36.9 84.2 22.6 75.0
Rec. RD [10] 100. 100. 100. 98.7 58.6 59.6 96.5 42.5 87.6
DiAD [16] 96.9 98.6 96.3 95.3 27.2 36.9 81.3 22.6 76.1
ViTAD [40] 100. 100. 100. 99.0 64.8 64.3 95.6 47.4 89.1
InvAD [42] 100. 100. 100. 99.0 61.8 62.6 96.6 45.5 88.6
InvAD-lite [42] 99.7 99.8 99.3 98.9 58.4 63.0 95.7 46.0 87.8
MambaAD [15] 100. 100. 100. 99.1 67.0 66.1 95.3 49.4 89.6
Hybrid UniAD [35] 99.6 99.8 98.6 98.0 53.0 53.9 93.7 36.9 85.2
RD++ [31] 94.8 96.2 92.0 98.8 59.1 58.7 96.4 41.6 85.1
DesTSeg [44] 99.6 99.8 98.6 96.8 84.7 80.2 92.9 66.9 93.2
leather Aug. DRAEM [37] 53.2 81.4 85.2 45.4 01.1 02.9 16.3 01.5 40.8
SimpleNet [27] 100. 100. 100. 98.5 27.8 33.9 95.5 20.4 79.4
RealNet [45] 100. 100. 100. 97.9 70.4 68.0 98.0 51.5 90.6
Emb. CFA [21] 62.2 82.7 87.8 65.4 07.7 14.8 40.7 08.0 51.6
PatchCore [29] 100. 100. 100. 99.4 52.1 51.2 97.8 34.5 85.8
CFLOW-AD [14] 100. 100. 100. 99.2 45.0 46.1 98.1 30.0 84.1
PyramidalFlow [22] 67.2 77.3 89.9 87.7 06.4 15.2 74.0 08.2 59.7
Rec. RD [10] 100. 100. 100. 99.3 39.1 46.6 97.9 30.4 83.3
DiAD [16] 97.6 99.1 95.8 90.6 05.9 11.9 68.3 06.3 67.0
ViTAD [40] 100. 100. 100. 99.5 51.1 55.8 98.0 38.7 86.3
InvAD [42] 100. 100. 100. 99.5 54.3 54.3 98.1 37.3 86.6
InvAD-lite [42] 100. 100. 100. 99.3 45.6 47.4 98.5 31.0 84.4
MambaAD [15] 100. 100. 100. 99.3 50.6 50.4 98.7 33.7 85.6
Hybrid UniAD [35] 100. 100. 100. 99.1 41.1 43.5 98.3 27.8 83.1
RD++ [31] 100. 100. 100. 99.3 40.4 47.1 97.5 30.8 83.5
DesTSeg [44] 100. 100. 100. 99.3 79.6 71.9 98.4 56.2 92.7
metal_nut Aug. DRAEM [37] 58.1 82.5 89.4 38.8 08.9 21.3 11.0 11.9 44.3
SimpleNet [27] 99.0 99.8 97.8 97.7 82.2 79.2 87.6 65.6 91.9
RealNet [45] 78.6 94.7 89.4 52.5 32.3 21.0 39.6 11.7 58.3
Emb. CFA [21] 78.3 94.4 89.9 71.4 37.4 36.5 30.1 22.4 62.6
PatchCore [29] 100. 100. 100. 98.8 88.9 85.8 95.6 75.1 95.6
CFLOW-AD [14] 96.9 99.3 96.2 96.0 71.2 71.7 88.5 55.8 88.5
PyramidalFlow [22] 63.0 89.9 90.2 81.6 41.9 45.8 37.3 29.7 64.2
Rec. RD [10] 99.9 100. 99.5 93.9 62.8 64.7 92.0 47.9 87.5
DiAD [16] 88.1 97.4 89.9 94.5 74.0 72.7 53.5 57.2 81.4
ViTAD [40] 99.7 99.9 98.9 96.1 74.0 75.7 93.0 60.9 91.0
InvAD [42] 100. 100. 100. 97.4 79.0 81.2 93.1 68.4 93.0
InvAD-lite [42] 99.8 100. 98.9 97.2 78.3 79.5 93.8 66.0 92.5
MambaAD [15] 99.7 99.9 98.4 96.7 74.2 78.3 92.9 64.4 91.4
Hybrid UniAD [35] 98.0 99.5 98.4 93.4 55.0 67.2 80.8 50.6 84.6
RD++ [31] 100. 100. 100. 96.7 74.7 78.4 94.1 64.5 92.0
DesTSeg [44] 99.8 100. 98.9 95.8 88.2 82.5 94.8 70.2 94.3
Table A23: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pill Aug. DRAEM [37] 57.1 87.5 92.1 34.8 02.0 06.6 04.4 03.4 40.6
SimpleNet [27] 90.6 98.1 94.2 96.7 74.7 70.6 85.3 54.6 87.2
RealNet [45] 68.5 92.8 91.6 54.4 47.8 08.9 35.1 04.6 57.0
Emb. CFA [21] 71.3 93.4 92.2 88.7 66.4 63.7 42.1 46.8 74.0
PatchCore [29] 97.0 99.4 96.5 98.1 76.8 74.0 96.1 58.7 91.1
CFLOW-AD [14] 84.3 96.5 93.2 96.7 59.5 56.3 90.5 39.2 82.4
PyramidalFlow [22] 67.0 92.8 91.6 83.3 18.3 26.0 65.2 14.9 63.5
Rec. RD [10] 97.2 99.5 96.8 97.4 63.0 64.8 96.1 47.9 87.8
DiAD [16] 94.7 98.9 97.2 92.8 49.1 49.7 65.2 33.1 78.2
ViTAD [40] 96.9 99.5 96.1 98.7 77.7 75.0 95.5 59.9 91.3
InvAD [42] 98.6 99.8 97.5 97.9 67.7 68.7 96.6 52.3 89.5
InvAD-lite [42] 96.3 99.4 95.4 97.6 65.4 67.9 96.4 51.4 88.3
MambaAD [15] 94.7 99.0 96.5 96.2 55.0 58.9 95.2 41.7 85.1
Hybrid UniAD [35] 75.1 94.3 91.6 90.0 21.4 28.6 89.2 16.7 70.0
RD++ [31] 98.0 99.6 97.5 98.2 72.7 70.3 96.4 54.2 90.4
DesTSeg [44] 92.7 98.6 94.8 90.1 78.7 73.9 70.2 58.6 85.6
screw Aug. DRAEM [37] 48.4 71.5 85.3 41.4 00.2 00.5 09.3 00.3 36.7
SimpleNet [27] 77.5 90.6 86.7 95.8 15.8 23.8 83.1 13.5 67.6
RealNet [45] 71.3 86.6 87.0 51.8 15.4 04.5 18.5 02.3 47.9
Emb. CFA [21] 62.4 78.6 88.5 26.5 00.2 00.7 09.9 00.3 38.1
PatchCore [29] 98.3 99.4 97.1 99.5 41.5 44.6 97.2 28.7 82.5
CFLOW-AD [14] 63.4 85.1 85.3 96.5 13.6 18.5 87.7 10.2 64.3
PyramidalFlow [22] 63.7 78.7 87.0 71.4 00.9 02.1 21.5 01.0 46.5
Rec. RD [10] 97.8 99.3 96.3 99.4 42.3 47.3 97.1 31.0 82.8
DiAD [16] 63.2 82.8 86.9 86.3 03.3 09.2 57.4 04.8 55.6
ViTAD [40] 91.6 97.2 91.4 98.9 32.3 39.9 93.6 24.9 77.8
InvAD [42] 91.7 95.1 96.0 99.5 44.0 46.4 97.6 30.2 81.5
InvAD-lite [42] 93.0 97.4 93.6 99.4 47.2 48.9 97.5 32.4 82.4
MambaAD [15] 92.7 96.8 94.0 99.3 45.3 45.1 97.0 29.1 81.5
Hybrid UniAD [35] 88.4 94.4 92.3 98.2 19.2 28.4 90.9 16.5 73.1
RD++ [31] 97.5 99.2 95.8 99.5 45.5 48.9 97.7 32.3 83.4
DesTSeg [44] 86.1 95.2 88.6 76.2 52.1 50.0 53.3 33.3 71.6
tile Aug. DRAEM [37] 68.0 84.7 84.0 48.2 06.3 13.3 14.0 07.1 45.5
SimpleNet [27] 100. 100. 99.4 95.4 59.1 60.4 82.5 43.3 85.3
RealNet [45] 97.5 99.3 97.6 93.9 84.1 76.8 90.5 62.3 91.4
Emb. CFA [21] 46.0 74.5 83.6 40.3 08.5 13.2 16.6 07.0 40.4
PatchCore [29] 100. 100. 99.4 96.4 58.7 67.1 86.6 50.5 86.9
CFLOW-AD [14] 99.8 99.9 99.4 96.0 56.0 62.1 86.5 45.0 85.7
PyramidalFlow [22] 78.6 92.2 83.6 75.7 55.1 28.1 34.4 16.3 64.0
Rec. RD [10] 99.0 99.6 97.0 95.3 48.7 61.0 86.3 43.9 83.8
DiAD [16] 93.4 96.6 95.1 76.1 24.1 32.9 49.7 19.7 66.8
ViTAD [40] 100. 100. 100. 96.7 56.6 69.0 88.7 52.7 87.3
InvAD [42] 100. 100. 100. 95.8 49.8 62.4 88.3 45.4 85.2
InvAD-lite [42] 98.0 99.2 96.5 93.9 44.9 54.5 80.1 37.4 81.0
MambaAD [15] 96.8 98.8 93.8 93.0 43.9 52.6 79.5 35.7 79.8
Hybrid UniAD [35] 96.7 98.8 95.8 89.2 38.9 47.5 77.6 31.1 77.8
RD++ [31] 100. 100. 99.4 96.7 56.8 66.4 89.8 49.7 87.0
DesTSeg [44] 99.7 99.9 98.2 97.5 89.4 81.9 95.7 69.4 94.6
toothbrush Aug. DRAEM [37] 44.2 72.8 83.3 46.8 01.4 03.1 14.5 01.6 38.0
SimpleNet [27] 89.4 95.8 90.9 98.0 53.6 55.9 80.6 38.8 80.6
RealNet [45] 76.4 91.5 83.3 84.8 50.1 56.1 34.1 39.0 68.0
Emb. CFA [21] 61.4 72.0 87.9 69.6 03.3 08.5 43.8 04.4 49.5
PatchCore [29] 91.9 96.2 95.2 99.0 54.8 58.5 91.0 41.3 83.8
CFLOW-AD [14] 84.4 94.5 86.2 98.2 45.7 47.1 84.5 30.8 77.2
PyramidalFlow [22] 89.4 95.7 90.6 73.0 42.3 31.3 23.2 18.6 63.6
Rec. RD [10] 98.3 99.4 96.6 98.9 51.8 57.7 92.6 40.5 85.0
DiAD [16] 95.3 98.1 95.2 89.6 11.0 17.5 65.3 09.6 67.4
ViTAD [40] 98.6 99.5 96.8 99.1 51.8 62.2 91.3 45.1 85.6
InvAD [42] 93.9 97.4 95.2 99.0 50.6 59.7 91.6 42.6 83.9
InvAD-lite [42] 97.8 99.1 95.2 98.9 48.1 59.1 91.9 41.9 84.3
MambaAD [15] 96.4 98.5 96.8 98.9 47.5 59.7 92.0 42.5 84.3
Hybrid UniAD [35] 83.9 93.1 90.9 98.4 39.7 49.9 85.8 33.3 77.4
RD++ [31] 95.8 98.3 95.1 99.1 57.3 61.8 92.2 44.8 85.7
DesTSeg [44] 91.7 96.5 95.2 95.6 60.6 72.1 65.5 56.4 82.5
Table A24: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
transistor Aug. DRAEM [37] 52.8 44.1 57.1 51.7 04.6 09.3 17.8 04.9 33.9
SimpleNet [27] 97.8 96.6 91.6 95.4 60.4 57.6 82.5 40.4 83.1
RealNet [45] 79.7 83.0 73.3 60.9 40.2 28.3 44.6 16.5 58.6
Emb. CFA [21] 68.9 64.8 63.4 60.5 17.6 20.4 22.0 11.3 45.4
PatchCore [29] 99.8 99.7 97.5 96.0 62.3 59.2 90.8 42.0 86.5
CFLOW-AD [14] 90.5 91.4 84.2 84.8 38.3 39.0 73.0 24.2 71.6
PyramidalFlow [22] 50.8 53.8 57.1 75.9 13.2 19.5 26.1 10.8 42.3
Rec. RD [10] 94.2 94.8 89.5 85.1 41.8 44.7 74.5 28.8 74.9
DiAD [16] 99.0 98.1 97.5 97.8 71.8 67.6 85.7 51.1 88.2
ViTAD [40] 97.7 96.7 91.6 93.6 57.5 54.3 76.6 37.3 81.1
InvAD [42] 97.8 96.8 92.7 95.7 62.0 59.7 91.9 42.6 85.2
InvAD-lite [42] 98.2 97.4 92.3 95.0 61.0 58.9 88.3 41.8 84.4
MambaAD [15] 99.9 99.8 98.7 96.0 63.8 61.6 90.4 44.5 87.2
Hybrid UniAD [35] 99.3 98.9 97.6 97.3 66.8 66.4 92.2 49.7 88.4
RD++ [31] 95.2 95.0 89.5 89.3 47.7 49.0 77.9 32.4 77.7
DesTSeg [44] 96.9 96.2 89.5 73.9 54.1 50.5 76.6 33.8 76.8
wood Aug. DRAEM [37] 74.3 90.2 87.2 49.7 04.4 07.5 14.6 03.9 46.8
SimpleNet [27] 99.5 99.8 98.3 92.5 40.3 42.4 80.0 26.9 79.0
RealNet [45] 99.6 99.9 99.2 90.4 76.1 71.9 88.8 56.1 89.4
Emb. CFA [21] 72.5 90.1 87.0 54.4 06.8 09.3 31.6 04.9 50.2
PatchCore [29] 98.7 99.6 96.7 95.2 52.1 52.3 91.3 35.4 83.7
CFLOW-AD [14] 98.4 99.5 96.6 94.2 45.7 49.8 90.2 33.1 82.1
PyramidalFlow [22] 86.7 95.5 88.5 62.6 39.5 17.9 32.3 09.8 60.4
Rec. RD [10] 99.6 99.9 98.3 95.5 49.7 51.6 91.7 34.8 83.8
DiAD [16] 94.3 97.8 94.3 82.8 15.1 22.2 57.5 12.5 66.3
ViTAD [40] 98.8 99.6 96.7 96.3 61.3 59.2 89.3 42.0 85.9
InvAD [42] 99.2 99.8 98.3 95.3 51.3 52.1 91.7 35.2 84.0
InvAD-lite [42] 96.9 99.0 95.9 94.4 44.8 47.6 89.1 31.2 81.1
MambaAD [15] 98.5 99.5 96.0 94.0 46.9 48.4 92.0 32.0 82.2
Hybrid UniAD [35] 98.2 99.5 96.6 93.1 40.6 45.1 89.5 29.1 80.4
RD++ [31] 98.9 99.7 97.5 95.6 52.2 52.4 91.0 35.5 83.9
DesTSeg [44] 98.6 99.5 97.6 93.2 79.5 72.5 95.0 56.8 90.8
zipper Aug. DRAEM [37] 44.4 79.9 88.1 55.4 02.7 06.1 20.7 03.2 42.5
SimpleNet [27] 99.1 99.8 97.5 97.9 57.8 55.4 91.9 38.3 85.6
RealNet [45] 77.3 94.1 88.1 67.6 51.9 42.6 47.7 27.1 67.0
Emb. CFA [21] 34.8 72.3 88.1 64.4 03.2 07.5 23.4 03.9 42.0
PatchCore [29] 99.6 99.9 98.7 98.9 63.4 64.6 96.4 47.8 88.8
CFLOW-AD [14] 95.6 98.5 96.7 97.9 50.2 54.8 92.4 37.7 83.7
PyramidalFlow [22] 83.3 95.1 89.1 81.0 15.4 16.3 55.7 08.9 62.3
Rec. RD [10] 99.5 99.9 99.2 98.5 56.0 61.0 95.5 43.8 87.1
DiAD [16] 73.5 90.3 90.0 84.4 14.3 24.0 56.7 13.7 61.9
ViTAD [40] 97.5 99.2 97.0 95.9 44.1 50.1 89.5 33.4 81.9
InvAD [42] 99.4 99.8 98.3 98.2 50.2 57.7 94.5 40.5 85.4
InvAD-lite [42] 97.5 99.3 96.7 98.6 58.4 61.4 95.8 44.3 86.8
MambaAD [15] 98.2 99.5 96.7 98.1 55.0 58.9 94.5 41.7 85.8
Hybrid UniAD [35] 86.9 95.5 92.8 95.8 33.5 41.0 88.7 25.8 76.3
RD++ [31] 99.3 99.8 98.7 98.4 53.5 59.2 94.8 42.1 86.2
DesTSeg [44] 97.6 99.0 98.3 90.6 71.9 66.9 86.2 50.3 87.2
Avg Aug. DRAEM [37] 54.5 76.3 83.6 47.6 03.2 06.7 14.3 03.5 40.9
SimpleNet [27] 95.4 98.3 95.7 96.8 48.8 51.9 86.9 36.4 82.0
RealNet [45] 84.8 94.1 90.9 72.6 48.2 41.4 56.8 28.8 69.8
Emb. CFA [21] 57.6 78.3 84.7 54.8 11.9 14.7 25.3 08.9 46.8
PatchCore [29] 98.8 99.5 98.4 98.3 59.9 61.0 94.2 44.9 87.2
CFLOW-AD [14] 91.6 96.7 93.4 95.7 45.9 48.6 88.3 33.2 80.0
PyramidalFlow [22] 70.2 85.5 85.5 80.0 22.3 22.0 47.5 12.8 59.0
Rec. RD [10] 93.6 97.2 95.6 95.8 48.2 53.6 91.2 37.0 82.2
DiAD [16] 88.9 95.8 93.5 89.3 27.0 32.5 63.9 21.1 70.1
ViTAD [40] 98.3 99.3 97.3 97.6 55.2 58.4 92.0 42.3 85.4
InvAD [42] 98.1 99.0 97.6 98.0 56.3 59.2 94.4 42.8 86.1
InvAD-lite [42] 97.9 99.2 96.8 97.3 54.4 57.8 93.3 41.4 85.2
MambaAD [15] 97.8 99.3 97.3 97.4 55.1 57.6 93.4 41.2 85.4
Hybrid UniAD [35] 92.5 97.3 95.4 95.8 42.7 48.0 89.3 32.5 80.1
RD++ [31] 97.9 98.8 96.4 97.3 54.7 58.0 93.2 41.5 85.2
DesTSeg [44] 96.4 98.6 96.2 92.0 71.1 68.2 83.4 52.8 86.6
Appendix E Detailed Quantitative Results on MVTec 3D Dataset Under 100 epochs
Table A25: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bagel Aug. DRAEM [37] 54.3 84.7 88.9 53.5 02.5 06.3 32.2 03.3 46.1
SimpleNet [27] 79.5 94.8 88.9 95.2 20.0 29.3 71.1 17.1 68.4
Emb. CFA [21] 82.6 95.6 89.3 84.3 19.1 30.6 48.8 18.1 64.3
PatchCore [29] 91.8 98.0 92.9 98.8 43.7 46.8 93.3 30.5 80.8
CFLOW-AD [14] 78.8 94.2 89.3 98.3 31.3 38.9 89.7 24.2 74.4
PyramidalFlow [22] 50.5 84.4 88.9 94.3 05.3 11.5 79.3 06.1 59.2
Rec. RD [10] 86.2 96.5 90.5 98.7 40.2 44.5 92.2 28.6 78.4
DiAD [16] 75.8 93.5 90.5 89.7 04.3 10.1 66.4 05.3 61.5
ViTAD [40] 92.1 98.0 93.4 98.7 40.3 44.2 92.4 28.4 79.9
InvAD [42] 95.3 98.8 94.6 99.1 45.4 49.6 95.0 33.0 82.5
InvAD-lite [42] 87.8 97.0 90.9 98.6 41.9 45.6 92.1 29.5 79.1
MambaAD [15] 93.4 98.3 93.3 98.7 41.0 45.3 94.6 29.3 80.7
Hybrid UniAD [35] 88.3 97.2 90.3 98.1 35.5 40.3 91.3 25.2 77.3
RD++ [31] 87.2 96.6 91.5 98.7 39.5 46.9 92.5 30.7 79.0
DesTSeg [44] 95.4 98.8 94.5 94.6 63.2 60.2 86.3 43.1 84.7
cable_gland Aug. DRAEM [37] 61.1 85.0 89.7 40.1 00.3 00.5 15.7 00.3 41.8
SimpleNet [27] 75.8 92.5 91.1 96.3 15.6 23.9 87.9 13.6 69.0
Emb. CFA [21] 61.2 87.5 89.2 71.0 01.8 04.6 38.7 02.4 50.6
PatchCore [29] 96.4 99.0 96.1 99.5 37.3 43.6 98.3 27.9 81.5
CFLOW-AD [14] 69.7 90.2 90.1 96.7 09.4 16.2 90.0 08.8 66.0
PyramidalFlow [22] 64.4 88.9 89.7 87.6 01.3 02.9 62.9 01.5 56.8
Rec. RD [10] 91.2 97.5 95.0 99.5 39.7 44.8 98.5 28.9 80.9
DiAD [16] 88.0 97.1 92.6 95.0 04.6 10.1 84.2 05.3 67.4
ViTAD [40] 84.5 94.2 92.1 98.7 22.6 31.4 95.7 18.6 74.2
InvAD [42] 96.6 99.1 96.6 99.6 40.6 47.1 98.8 30.8 82.6
InvAD-lite [42] 93.7 98.5 96.0 99.4 41.1 46.9 98.5 30.6 82.0
MambaAD [15] 91.1 97.8 93.6 99.3 38.8 44.6 98.2 28.7 80.5
Hybrid UniAD [35] 73.3 91.0 89.8 97.1 14.2 22.0 91.7 12.4 68.4
RD++ [31] 96.6 99.1 96.0 99.6 46.1 49.3 98.7 32.7 83.6
DesTSeg [44] 89.9 97.5 92.9 89.9 47.8 52.3 85.3 35.4 79.4
carrot Aug. DRAEM [37] 50.8 83.6 90.7 56.2 00.6 00.9 35.4 00.5 45.5
SimpleNet [27] 82.4 96.1 92.5 98.1 13.3 20.6 92.6 11.5 70.8
Emb. CFA [21] 58.6 88.6 91.0 83.1 03.0 09.4 54.0 05.0 55.4
PatchCore [29] 95.2 99.0 95.6 99.4 25.9 30.8 97.7 18.2 77.7
CFLOW-AD [14] 85.2 96.8 92.3 98.9 20.0 27.2 96.0 15.7 73.8
PyramidalFlow [22] 59.9 87.4 90.7 97.8 13.0 18.8 92.5 10.4 65.7
Rec. RD [10] 85.4 96.2 92.5 99.4 27.9 33.8 97.7 20.4 76.1
DiAD [16] 93.4 98.3 98.1 97.6 06.2 11.2 91.1 05.9 70.8
ViTAD [40] 77.4 93.0 92.2 98.8 19.3 25.9 95.9 14.8 71.8
InvAD [42] 90.5 97.6 94.5 99.3 24.9 30.1 97.6 17.7 76.4
InvAD-lite [42] 87.7 96.9 94.6 99.4 27.6 32.7 98.0 19.5 76.7
MambaAD [15] 90.1 97.7 93.9 99.3 29.0 32.7 97.9 19.6 77.2
Hybrid UniAD [35] 75.0 93.8 91.4 98.2 15.6 22.8 94.2 12.9 70.1
RD++ [31] 89.3 97.2 93.9 99.4 28.2 34.0 97.7 20.5 77.1
DesTSeg [44] 78.2 93.7 91.3 67.6 30.4 33.4 46.1 20.0 63.0
cookie Aug. DRAEM [37] 80.6 93.4 91.1 47.3 01.7 05.2 20.5 02.7 48.5
SimpleNet [27] 66.9 89.0 88.0 92.5 27.2 34.6 71.5 21.0 67.1
Emb. CFA [21] 61.2 86.8 88.4 72.4 19.4 27.6 38.3 16.0 56.3
PatchCore [29] 71.0 91.6 88.0 98.0 52.4 50.2 92.5 33.5 77.7
CFLOW-AD [14] 46.3 78.0 88.4 96.9 30.3 32.6 89.2 19.5 66.0
PyramidalFlow [22] 23.4 67.8 88.0 88.8 07.1 11.9 70.8 06.3 51.1
Rec. RD [10] 50.2 80.1 88.0 96.7 28.1 33.2 87.4 19.9 66.2
DiAD [16] 52.6 84.7 88.0 87.2 05.6 10.5 67.2 05.5 56.6
ViTAD [40] 70.7 90.9 88.2 97.7 47.6 48.8 87.6 32.3 75.9
InvAD [42] 63.5 88.5 88.0 98.4 48.8 49.3 92.5 32.7 75.6
InvAD-lite [42] 62.0 88.4 88.0 97.3 39.9 42.8 87.0 27.2 72.2
MambaAD [15] 58.3 85.6 88.4 96.8 38.8 42.0 84.3 26.6 70.6
Hybrid UniAD [35] 62.3 86.9 88.0 96.0 24.2 29.6 88.1 17.4 67.9
RD++ [31] 56.7 82.8 88.4 97.6 39.6 42.6 89.2 27.1 71.0
DesTSeg [44] 61.7 87.9 88.4 83.7 43.9 45.2 57.8 29.2 66.9
Table A26: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
dowel Aug. DRAEM [37] 38.5 75.0 88.9 42.1 00.7 02.2 24.6 01.1 38.9
SimpleNet [27] 90.3 97.5 92.5 97.8 16.2 21.5 90.2 12.0 72.3
Emb. CFA [21] 88.8 96.8 92.9 85.0 07.1 11.1 56.9 05.9 62.7
PatchCore [29] 98.2 99.6 97.5 99.3 44.5 46.7 96.3 30.5 83.2
CFLOW-AD [14] 92.1 98.1 92.9 98.8 28.3 32.7 94.4 19.5 76.8
PyramidalFlow [22] 66.5 89.3 89.7 93.9 06.3 10.5 74.9 05.5 61.6
Rec. RD [10] 97.3 99.3 96.7 99.7 51.1 50.6 98.8 33.9 84.8
DiAD [16] 65.3 89.1 91.9 79.7 04.6 09.2 47.9 04.8 55.4
ViTAD [40] 94.1 98.5 94.8 99.2 37.8 41.7 95.9 26.3 80.3
InvAD [42] 99.3 99.8 98.6 99.7 49.5 48.6 98.6 32.1 84.9
InvAD-lite [42] 96.8 99.3 96.0 99.6 51.4 51.5 97.7 34.7 84.6
MambaAD [15] 96.5 99.2 95.3 99.5 50.1 48.9 97.4 32.3 83.8
Hybrid UniAD [35] 87.1 96.9 90.1 98.9 34.0 39.0 95.3 24.2 77.3
RD++ [31] 99.6 99.9 98.6 99.7 53.9 51.3 98.8 34.5 86.0
DesTSeg [44] 86.0 96.5 90.2 82.8 49.4 50.1 60.8 33.4 73.7
foam Aug. DRAEM [37] 50.3 81.6 88.9 46.7 00.2 00.4 14.4 00.2 40.4
SimpleNet [27] 78.7 94.5 89.4 86.4 12.1 22.1 67.9 12.4 64.4
Emb. CFA [21] 66.1 90.6 89.4 45.1 08.0 21.0 32.6 11.8 50.4
PatchCore [29] 78.9 94.4 89.9 94.9 18.8 29.6 83.4 17.4 70.0
CFLOW-AD [14] 75.0 93.5 89.3 85.5 15.8 27.2 57.5 15.7 63.4
PyramidalFlow [22] 65.6 90.2 88.9 71.4 11.0 18.2 49.9 10.0 56.5
Rec. RD [10] 70.2 91.7 89.4 93.9 14.4 27.1 81.0 15.6 66.8
DiAD [16] 78.3 94.4 89.2 87.5 01.4 03.9 59.7 02.0 59.2
ViTAD [40] 82.9 95.4 90.8 94.0 14.3 26.0 81.3 14.9 69.2
InvAD [42] 83.5 95.8 91.2 93.8 18.8 31.0 81.1 18.3 70.7
InvAD-lite [42] 82.6 95.4 90.1 94.4 24.1 32.9 82.8 19.7 71.8
MambaAD [15] 79.9 94.7 89.9 94.9 23.8 33.4 83.8 20.1 71.5
Hybrid UniAD [35] 76.8 94.0 88.9 85.0 10.3 23.0 60.1 13.0 62.6
RD++ [31] 72.9 92.4 91.9 92.7 16.3 28.7 77.3 16.7 67.5
DesTSeg [44] 80.2 95.0 89.1 84.5 44.3 49.5 72.7 32.9 73.6
peach Aug. DRAEM [37] 61.6 86.9 89.5 46.7 00.5 01.0 22.5 00.5 44.1
SimpleNet [27] 70.0 90.5 90.1 92.2 05.5 11.0 73.3 05.8 61.8
Emb. CFA [21] 49.9 81.6 89.1 67.4 01.2 05.7 39.0 02.9 47.7
PatchCore [29] 75.2 92.7 91.2 98.2 26.4 32.4 93.3 19.3 72.8
CFLOW-AD [14] 72.2 91.3 90.4 97.9 17.1 18.9 92.1 10.4 68.6
PyramidalFlow [22] 62.0 86.7 89.8 96.4 06.6 11.0 87.3 05.8 62.8
Rec. RD [10] 64.0 85.0 91.6 98.4 15.6 21.7 93.7 12.2 67.1
DiAD [16] 92.4 93.5 99.1 95.5 04.4 08.0 84.9 04.2 68.3
ViTAD [40] 66.6 88.8 89.1 98.1 19.5 25.2 92.5 14.4 68.5
InvAD [42] 90.4 97.3 94.1 99.3 41.3 46.5 97.1 30.3 80.9
InvAD-lite [42] 85.4 95.3 94.1 99.3 42.8 44.0 97.2 28.2 79.7
MambaAD [15] 89.3 97.1 93.4 99.4 42.9 44.0 97.5 28.2 80.5
Hybrid UniAD [35] 73.7 91.8 90.4 98.0 19.6 25.5 92.5 14.6 70.2
RD++ [31] 74.2 89.8 92.0 98.8 26.2 31.7 94.8 18.8 72.5
DesTSeg [44] 70.6 91.5 89.4 76.7 46.7 45.5 56.1 29.5 68.1
potato Aug. DRAEM [37] 50.1 80.9 89.8 49.6 00.5 01.3 30.0 00.6 43.2
SimpleNet [27] 61.9 85.3 89.9 95.2 02.6 05.6 83.7 02.9 60.6
Emb. CFA [21] 39.1 77.4 89.3 64.9 00.3 01.0 30.7 00.5 43.2
PatchCore [29] 56.8 83.9 90.2 98.9 09.8 17.0 96.0 09.3 64.7
CFLOW-AD [14] 62.0 86.8 89.3 98.5 11.4 13.9 95.3 07.5 65.3
PyramidalFlow [22] 70.7 91.2 89.8 97.8 09.9 15.9 92.5 08.6 66.8
Rec. RD [10] 63.1 87.6 91.1 99.2 15.3 22.7 96.6 12.8 67.9
DiAD [16] 82.2 95.8 90.2 98.5 08.6 14.7 94.8 07.9 69.3
ViTAD [40] 58.2 86.7 89.8 98.7 11.6 17.9 95.1 09.8 65.4
InvAD [42] 66.9 90.4 90.2 99.1 17.8 24.8 96.4 14.1 69.4
InvAD-lite [42] 67.8 90.5 90.5 99.1 17.8 22.4 95.9 12.6 69.1
MambaAD [15] 59.1 86.0 90.2 99.0 17.7 22.8 95.4 12.8 67.2
Hybrid UniAD [35] 51.6 83.8 89.3 97.9 08.5 13.9 92.3 07.5 62.5
RD++ [31] 67.0 90.0 90.6 99.2 17.5 23.7 96.7 13.5 69.2
DesTSeg [44] 58.2 86.8 90.2 54.8 08.5 09.2 22.5 04.8 47.2
Table A27: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
rope Aug. DRAEM [37] 56.5 69.4 83.6 43.0 00.5 01.0 13.1 00.5 38.2
SimpleNet [27] 93.3 97.7 93.8 99.3 53.8 52.4 93.1 35.5 83.3
Emb. CFA [21] 92.5 96.8 90.1 90.4 35.7 44.1 74.5 28.3 74.9
PatchCore [29] 96.7 98.7 95.5 99.6 55.6 56.9 97.4 39.7 85.8
CFLOW-AD [14] 94.1 97.7 92.2 99.3 45.1 47.3 95.3 31.0 81.6
PyramidalFlow [22] 79.7 90.9 83.2 86.4 08.0 11.9 61.3 06.3 60.2
Rec. RD [10] 95.9 98.4 93.1 99.6 50.6 55.2 97.9 38.1 84.4
DiAD [16] 72.4 87.2 81.7 94.3 09.5 16.1 74.6 08.8 62.3
ViTAD [40] 98.5 99.4 96.4 98.5 41.7 43.3 89.5 27.6 81.0
InvAD [42] 97.0 98.8 94.9 99.6 57.8 58.9 97.1 41.7 86.3
InvAD-lite [42] 94.5 98.0 94.7 99.4 54.2 53.0 96.3 36.1 84.3
MambaAD [15] 96.1 98.4 93.2 99.3 45.1 48.1 96.1 31.7 82.3
Hybrid UniAD [35] 95.9 98.5 95.6 99.3 41.5 47.8 96.4 31.4 82.1
RD++ [31] 95.9 98.5 96.3 99.6 52.7 56.2 97.8 39.0 85.3
DesTSeg [44] 92.7 97.2 92.2 98.8 64.7 62.7 97.2 45.7 86.5
tire Aug. DRAEM [37] 52.7 76.9 87.4 40.6 00.1 00.3 10.2 00.2 38.3
SimpleNet [27] 59.5 85.7 87.4 93.9 06.9 13.1 78.3 07.0 60.7
Emb. CFA [21] 70.7 90.8 87.4 52.5 01.2 05.3 18.9 02.7 46.7
PatchCore [29] 80.1 93.6 87.9 99.1 22.6 31.1 95.8 18.4 72.9
CFLOW-AD [14] 56.1 83.9 87.4 97.5 07.0 11.2 90.7 05.9 62.0
PyramidalFlow [22] 44.2 79.2 87.4 93.7 03.7 07.3 78.0 03.8 56.2
Rec. RD [10] 54.2 79.2 87.4 99.0 13.5 21.5 95.5 12.1 64.3
DiAD [16] 54.0 76.1 87.4 92.5 03.4 08.1 78.1 04.2 57.1
ViTAD [40] 70.6 89.3 89.6 97.9 17.4 28.3 90.6 16.5 69.1
InvAD [42] 86.2 95.8 89.9 99.5 29.3 38.2 97.5 23.6 76.6
InvAD-lite [42] 89.0 96.7 91.3 99.3 42.3 47.4 96.4 31.1 80.3
MambaAD [15] 87.3 95.9 91.1 99.3 42.1 46.3 96.7 30.1 79.8
Hybrid UniAD [35] 68.8 87.5 88.7 97.0 08.4 15.4 87.5 08.4 64.8
RD++ [31] 78.4 92.6 88.7 99.4 23.6 31.8 97.0 18.9 73.1
DesTSeg [44] 78.1 93.0 88.4 77.9 13.7 18.7 55.7 10.3 60.8
Avg Aug. DRAEM [37] 55.7 81.8 88.8 46.6 00.8 01.9 21.9 01.0 42.5
SimpleNet [27] 75.8 92.3 90.4 94.7 17.3 23.4 81.0 13.9 67.8
Emb. CFA [21] 67.1 89.3 89.6 71.6 09.7 16.1 43.2 09.3 55.2
PatchCore [29] 84.1 95.1 92.5 98.6 33.7 38.5 94.4 24.5 76.7
CFLOW-AD [14] 73.1 91.0 90.2 96.8 21.6 26.6 89.0 15.8 69.8
PyramidalFlow [22] 58.7 85.6 88.6 90.8 07.2 12.0 74.9 06.4 59.7
Rec. RD [10] 75.8 91.2 91.5 98.4 29.6 35.5 93.9 22.2 73.7
DiAD [16] 75.4 91.0 90.9 91.7 05.3 10.2 74.9 05.4 62.8
ViTAD [40] 79.6 93.4 91.6 98.0 27.2 33.3 91.6 20.4 73.5
InvAD [42] 86.9 96.2 93.3 98.7 37.4 42.4 95.2 27.4 78.6
InvAD-lite [42] 84.7 95.6 92.6 98.6 38.3 41.9 94.2 26.9 78.0
MambaAD [15] 84.1 95.1 92.2 98.6 36.9 40.8 94.2 25.9 77.4
Hybrid UniAD [35] 75.3 92.1 90.3 96.6 21.2 27.9 88.9 16.7 70.3
RD++ [31] 81.8 93.9 92.8 98.5 34.4 39.6 94.1 25.2 76.4
DesTSeg [44] 79.1 93.8 90.7 81.1 41.3 42.7 64.0 28.4 70.4
Appendix F Detailed Quantitative Results on MVTec LOCO Dataset Under 100 epochs
Table A28: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
breakfast_box Aug. DRAEM [37] 62.6 76.9 77.2 50.0 04.4 12.5 14.9 06.7 42.6
SimpleNet [27] 88.9 94.1 85.5 87.8 56.9 55.3 73.0 38.3 77.4
Emb. CFA [21] 64.3 79.1 77.6 68.0 29.7 28.5 33.9 16.6 54.4
PatchCore [29] 77.1 88.4 78.0 89.4 55.4 49.3 76.1 32.8 73.4
CFLOW-AD [14] 74.9 86.4 77.8 88.6 48.0 46.5 71.1 30.3 70.5
PyramidalFlow [22] 58.3 75.3 77.4 59.7 44.2 28.1 20.6 16.3 51.9
Rec. RD [10] 66.2 80.7 77.2 83.7 27.1 35.6 65.0 21.6 62.2
DiAD [16] 68.5 82.0 77.2 84.6 26.3 37.6 39.7 23.2 59.4
ViTAD [40] 77.7 88.0 78.8 89.1 56.1 53.1 70.3 36.1 73.3
InvAD [42] 82.7 91.1 80.5 90.2 63.7 60.8 79.7 43.6 78.4
InvAD-lite [42] 77.1 87.9 77.6 88.6 50.5 49.0 77.2 32.5 72.6
MambaAD [15] 69.7 83.6 77.5 89.1 54.5 50.1 74.5 33.5 71.3
Hybrid UniAD [35] 73.1 85.5 77.2 89.0 48.8 46.2 71.0 30.1 70.1
RD++ [31] 69.8 84.3 77.2 86.8 41.4 45.1 69.6 29.1 67.7
DesTSeg [44] 85.5 92.7 82.1 82.6 64.1 59.4 72.6 42.2 77.0
juice_bottle Aug. DRAEM [37] 52.1 74.3 83.4 31.3 03.9 10.1 07.3 05.3 37.5
SimpleNet [27] 96.7 98.8 94.8 91.6 48.9 54.6 86.1 37.5 81.6
Emb. CFA [21] 85.1 94.3 85.4 73.6 20.3 22.0 51.7 12.4 61.8
PatchCore [29] 95.2 98.3 94.1 89.9 56.2 57.8 87.1 40.6 82.7
CFLOW-AD [14] 95.9 98.4 93.6 91.7 44.1 47.3 84.1 31.0 79.3
PyramidalFlow [22] 48.6 74.9 83.5 54.2 09.7 10.1 24.9 05.3 43.7
Rec. RD [10] 90.1 96.6 89.3 89.2 46.9 47.7 84.4 31.4 77.7
DiAD [16] 97.3 98.8 96.1 83.2 40.0 42.4 74.9 26.9 76.1
ViTAD [40] 90.3 96.4 88.7 88.6 48.4 49.1 83.4 32.5 77.8
InvAD [42] 95.8 98.5 94.2 90.6 53.9 58.3 88.4 41.2 82.8
InvAD-lite [42] 93.0 97.3 92.7 90.6 53.4 57.2 87.5 40.0 81.7
MambaAD [15] 92.0 97.0 92.3 89.7 47.8 53.2 85.3 36.3 79.6
Hybrid UniAD [35] 92.2 97.0 90.5 90.6 52.5 54.6 87.3 37.6 80.7
RD++ [31] 90.6 96.5 88.8 90.1 49.2 50.9 85.2 34.2 78.8
DesTSeg [44] 98.5 99.5 97.2 78.8 57.0 57.2 79.7 40.1 81.1
pushpins Aug. DRAEM [37] 54.5 55.1 72.8 52.4 00.8 01.8 14.8 00.9 36.0
SimpleNet [27] 74.8 81.8 74.0 68.9 03.7 08.8 62.0 04.6 53.4
Emb. CFA [21] 53.7 63.2 71.4 43.6 00.8 01.7 22.5 00.9 36.7
PatchCore [29] 74.3 82.8 72.7 56.1 03.1 07.5 63.6 03.9 51.4
CFLOW-AD [14] 60.8 69.1 71.4 39.4 01.2 03.0 52.4 01.5 42.5
PyramidalFlow [22] 51.7 60.9 71.4 58.8 01.3 03.2 29.3 01.6 39.5
Rec. RD [10] 71.4 80.4 71.7 51.2 02.2 07.6 68.1 04.0 50.4
DiAD [16] 57.6 65.7 73.0 62.8 01.4 03.2 36.4 01.6 42.9
ViTAD [40] 66.8 73.6 72.2 46.1 02.2 06.0 57.1 03.1 46.3
InvAD [42] 81.0 87.4 76.7 67.2 03.8 09.1 75.1 04.8 57.2
InvAD-lite [42] 77.9 84.0 76.1 60.0 03.0 08.4 67.8 04.4 53.9
MambaAD [15] 79.8 85.9 76.3 58.8 02.6 09.4 72.7 05.0 55.1
Hybrid UniAD [35] 63.5 70.1 71.4 41.6 01.0 03.9 53.8 02.0 43.6
RD++ [31] 77.0 84.0 74.6 52.3 02.5 07.5 67.5 03.9 52.2
DesTSeg [44] 59.9 71.1 71.7 54.0 03.9 08.2 42.3 04.3 44.4
screw_bag Aug. DRAEM [37] 53.1 65.9 78.4 49.3 06.7 13.7 10.1 07.4 39.6
SimpleNet [27] 63.5 76.5 78.3 61.2 10.3 16.5 46.8 09.0 50.4
Emb. CFA [21] 63.2 77.2 78.4 50.2 07.5 13.6 30.2 07.3 45.8
PatchCore [29] 70.1 82.8 78.2 72.1 13.1 22.0 62.1 12.4 57.2
CFLOW-AD [14] 64.4 77.7 78.4 68.7 11.2 20.4 67.0 11.3 55.4
PyramidalFlow [22] 52.7 68.5 78.5 59.0 09.9 15.6 35.5 08.5 45.7
Rec. RD [10] 66.3 79.8 78.2 71.5 12.6 21.8 52.1 12.2 54.6
DiAD [16] 57.6 68.1 79.3 70.2 11.3 21.2 58.5 11.9 52.3
ViTAD [40] 65.2 77.7 79.0 72.5 13.2 21.9 42.2 12.3 53.1
InvAD [42] 68.5 82.1 78.7 71.9 13.2 21.9 53.0 12.3 55.6
InvAD-lite [42] 68.2 79.7 78.7 72.2 13.1 21.9 49.1 12.3 54.7
MambaAD [15] 55.2 70.5 78.2 73.1 13.2 22.7 53.6 12.8 52.4
Hybrid UniAD [35] 56.9 69.1 78.8 66.8 10.8 19.2 64.8 10.6 52.3
RD++ [31] 66.8 80.8 78.2 69.4 12.1 20.2 53.0 11.3 54.4
DesTSeg [44] 68.4 81.9 78.4 51.2 09.7 13.6 44.0 07.3 49.6
Table A29: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
splicing_connectors Aug. DRAEM [37] 40.1 54.7 76.4 50.4 06.5 14.1 15.0 07.6 36.7
SimpleNet [27] 83.8 91.4 81.9 73.2 25.1 28.5 51.1 16.6 62.1
Emb. CFA [21] 66.2 79.3 76.6 57.2 12.9 15.7 17.4 08.5 46.5
PatchCore [29] 85.6 92.9 84.4 68.1 21.5 22.3 60.8 12.6 62.2
CFLOW-AD [14] 79.3 89.3 77.8 65.2 20.3 22.3 73.1 12.5 61.0
PyramidalFlow [22] 55.9 67.0 76.4 56.7 09.8 15.7 18.2 08.5 42.8
Rec. RD [10] 79.2 89.1 78.0 58.0 16.3 17.9 68.8 09.9 58.2
DiAD [16] 75.7 83.0 78.8 58.1 15.9 19.6 32.4 10.9 51.9
ViTAD [40] 82.1 90.6 79.6 70.5 23.7 25.8 62.4 14.8 62.1
InvAD [42] 81.9 90.5 79.7 62.2 20.8 23.8 68.7 13.5 61.1
InvAD-lite [42] 80.8 89.9 79.3 67.9 22.6 24.5 65.5 14.0 61.5
MambaAD [15] 82.4 90.7 80.4 71.8 26.0 26.8 64.9 15.5 63.3
Hybrid UniAD [35] 80.4 88.7 80.1 77.7 40.5 43.2 46.5 27.5 65.3
RD++ [31] 77.7 88.5 77.3 59.2 16.2 17.5 68.9 09.6 57.9
DesTSeg [44] 77.5 88.1 77.0 53.8 41.3 14.0 28.4 07.5 54.3
Avg Aug. DRAEM [37] 52.5 65.4 77.6 46.7 04.5 10.4 12.4 05.6 38.5
SimpleNet [27] 81.6 88.5 82.9 76.5 29.0 32.7 63.8 21.2 65.0
Emb. CFA [21] 66.5 78.6 77.9 58.5 14.2 16.3 31.1 09.1 49.0
PatchCore [29] 80.5 89.0 81.5 75.1 29.9 31.8 69.9 20.4 65.4
CFLOW-AD [14] 75.1 84.2 79.8 70.7 25.0 27.9 69.5 17.3 61.7
PyramidalFlow [22] 53.4 69.3 77.5 57.7 15.0 14.5 25.7 08.0 44.7
Rec. RD [10] 74.6 85.3 78.9 70.7 21.0 26.1 67.7 15.8 60.6
DiAD [16] 71.3 79.5 80.9 71.8 19.0 24.8 48.4 14.9 56.5
ViTAD [40] 76.4 85.2 79.6 73.4 28.7 31.2 63.1 19.8 62.5
InvAD [42] 82.0 89.9 82.0 76.4 31.1 34.8 73.0 23.1 67.0
InvAD-lite [42] 79.4 87.8 80.9 75.8 28.5 32.2 69.4 20.6 64.9
MambaAD [15] 75.8 85.5 81.0 76.5 28.8 32.5 70.2 20.6 64.3
Hybrid UniAD [35] 73.2 82.1 79.6 73.1 30.7 33.4 64.7 21.6 62.4
RD++ [31] 76.4 86.8 79.2 71.5 24.3 28.3 68.8 17.6 62.2
DesTSeg [44] 77.9 86.7 81.3 64.1 35.2 30.5 53.4 20.3 61.3
Appendix G Detailed Quantitative Results on VisA Dataset Under 100 epochs
Table A30: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
candle Aug. DRAEM [37] 36.0 45.5 66.7 32.0 00.1 00.3 05.5 00.2 26.6
SimpleNet [27] 93.6 94.5 84.9 96.4 09.9 19.9 88.4 11.1 69.7
RealNet [45] 52.7 57.8 66.7 51.8 09.8 05.2 28.5 02.7 38.9
Emb. CFA [21] 71.8 73.7 70.6 76.2 03.7 12.2 59.7 06.5 52.6
CFLOW-AD [14] 93.0 93.3 85.3 98.8 13.9 23.0 93.8 13.0 71.6
PyramidalFlow [22] 43.4 44.9 66.7 77.9 00.5 02.3 55.8 01.2 41.6
Rec. RD [10] 93.1 93.4 86.9 99.1 24.2 35.4 95.5 21.5 75.4
DiAD [16] 92.3 93.6 84.5 80.2 01.0 04.1 43.1 02.1 57.0
ViTAD [40] 90.0 91.3 81.6 96.1 16.2 27.9 86.6 16.2 70.0
InvAD [42] 95.3 94.8 90.6 99.2 23.3 34.1 95.8 20.5 76.2
InvAD-lite [42] 97.1 97.2 92.0 99.0 21.7 30.6 96.0 18.1 76.2
MambaAD [15] 95.8 96.4 89.6 99.0 21.3 31.5 95.8 18.7 75.6
Hybrid UniAD [35] 95.7 95.6 89.2 99.0 24.1 32.8 93.7 19.6 75.7
RD++ [31] 92.5 92.9 87.8 99.3 26.2 37.1 95.5 22.7 75.9
DesTSeg [44] 89.7 90.5 86.1 79.6 41.1 43.9 61.7 28.1 70.4
capsules Aug. DRAEM [37] 49.0 68.2 76.9 54.8 00.4 00.7 20.3 00.3 38.6
SimpleNet [27] 76.8 87.0 78.3 95.5 42.0 45.5 66.1 29.4 70.2
RealNet [45] 72.8 85.0 77.0 65.9 31.8 34.8 31.8 21.1 57.0
Emb. CFA [21] 56.5 71.2 77.2 55.2 00.4 00.9 05.9 00.4 38.2
CFLOW-AD [14] 54.8 72.8 76.9 94.2 26.4 30.4 64.3 17.9 60.0
PyramidalFlow [22] 55.8 69.4 76.9 86.9 03.3 08.7 62.4 04.5 51.9
Rec. RD [10] 81.8 90.5 81.3 99.3 60.2 60.7 93.0 43.5 81.0
DiAD [16] 51.5 71.2 78.1 90.6 02.9 06.5 49.7 03.4 50.1
ViTAD [40] 80.4 88.2 78.9 98.2 31.0 41.6 75.5 26.2 70.5
InvAD [42] 90.1 94.2 86.8 99.6 68.6 64.9 95.0 48.1 85.6
InvAD-lite [42] 91.7 94.7 90.1 99.3 60.5 58.5 95.2 41.3 84.3
MambaAD [15] 91.1 94.4 89.6 99.0 58.4 56.7 93.2 39.5 83.2
Hybrid UniAD [35] 71.2 83.9 78.4 97.5 42.2 45.5 73.6 29.4 70.3
RD++ [31] 84.3 91.4 83.5 99.4 64.1 63.6 93.2 46.6 82.8
DesTSeg [44] 85.4 91.8 83.3 95.9 39.0 43.6 82.6 27.9 74.5
cashew Aug. DRAEM [37] 84.4 92.9 84.0 32.3 00.7 02.2 08.7 01.1 43.6
SimpleNet [27] 93.2 96.8 90.2 98.6 67.8 65.0 82.0 48.1 84.8
RealNet [45] 79.8 89.6 84.5 51.0 22.4 03.1 19.7 01.6 50.0
Emb. CFA [21] 74.9 88.1 81.0 97.6 57.3 56.1 76.2 39.0 75.9
CFLOW-AD [14] 95.8 98.2 93.3 99.0 53.8 54.9 94.6 37.9 84.2
PyramidalFlow [22] 84.1 90.7 87.7 49.7 00.6 02.2 15.7 01.1 47.2
Rec. RD [10] 92.1 95.7 91.3 91.3 43.8 49.4 85.1 32.8 78.4
DiAD [16] 77.0 88.3 83.3 85.5 42.4 46.6 19.4 30.3 63.2
ViTAD [40] 89.1 94.9 87.5 98.2 62.5 61.5 79.9 44.4 81.9
InvAD [42] 96.1 98.3 93.2 96.5 58.4 61.1 90.6 44.0 84.9
InvAD-lite [42] 94.1 97.0 90.8 94.4 51.6 55.5 90.5 38.4 82.0
MambaAD [15] 93.8 97.0 91.7 94.0 46.7 51.7 88.2 34.9 80.4
Hybrid UniAD [35] 94.8 97.4 91.5 99.2 61.3 63.9 91.0 46.9 85.6
RD++ [31] 93.4 96.3 92.2 95.1 56.4 58.9 87.2 41.7 82.8
DesTSeg [44] 88.0 94.2 87.5 85.9 65.9 62.6 58.8 45.5 77.6
chewinggum Aug. DRAEM [37] 56.4 72.6 79.0 34.4 00.3 01.1 01.1 00.6 35.0
SimpleNet [27] 97.4 98.8 93.8 98.3 31.5 36.9 75.1 22.6 76.0
RealNet [45] 87.7 94.6 86.0 72.4 40.0 42.0 36.4 26.5 65.6
Emb. CFA [21] 76.3 88.7 80.7 95.6 33.1 36.5 70.6 22.3 68.8
CFLOW-AD [14] 97.3 98.9 95.3 99.2 60.3 60.3 87.9 43.2 85.6
PyramidalFlow [22] 36.3 63.6 80.0 68.5 01.3 01.1 39.2 00.6 41.4
Rec. RD [10] 95.8 98.0 92.6 98.7 60.4 62.6 78.7 45.5 83.8
DiAD [16] 82.5 90.9 83.5 64.0 14.7 29.3 15.7 17.2 54.4
ViTAD [40] 95.8 98.0 93.2 97.8 61.1 58.5 72.2 41.3 82.4
InvAD [42] 98.0 99.0 94.9 98.7 60.4 62.8 81.3 45.8 85.0
InvAD-lite [42] 97.5 98.9 94.8 98.3 56.0 57.6 80.6 40.4 83.4
MambaAD [15] 97.4 98.8 94.7 98.0 54.6 60.0 80.0 42.8 83.4
Hybrid UniAD [35] 97.4 98.9 96.3 99.1 60.3 58.5 85.9 41.4 85.2
RD++ [31] 97.1 98.7 94.2 98.7 58.8 61.8 79.8 44.7 84.2
DesTSeg [44] 96.3 98.6 95.8 98.0 79.2 71.9 68.9 56.1 87.0
Table A31: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
fryum Aug. DRAEM [37] 46.5 68.9 80.0 59.6 04.3 10.3 32.0 05.4 43.1
SimpleNet [27] 84.0 92.8 83.6 94.9 46.2 48.5 81.7 32.0 76.0
RealNet [45] 66.2 83.4 80.6 53.8 35.7 10.4 19.2 05.5 49.9
Emb. CFA [21] 68.7 84.0 81.0 79.4 26.1 32.5 49.8 19.4 60.2
CFLOW-AD [14] 90.1 95.5 88.3 97.5 53.6 53.3 87.4 36.3 80.8
PyramidalFlow [22] 68.7 83.5 81.0 81.7 12.7 19.0 62.7 10.5 58.5
Rec. RD [10] 95.4 97.9 92.5 96.9 47.5 51.6 93.2 34.8 82.1
DiAD [16] 83.7 92.6 84.3 97.1 55.7 58.1 69.0 40.9 77.2
ViTAD [40] 93.9 97.2 90.6 97.4 46.4 50.3 88.7 33.6 80.6
InvAD [42] 96.0 98.2 91.9 97.0 48.5 52.5 92.1 35.6 82.3
InvAD-lite [42] 95.6 98.0 90.9 96.8 47.2 51.6 92.8 34.7 81.8
MambaAD [15] 95.5 97.8 91.5 97.0 48.3 51.7 92.3 34.9 82.0
Hybrid UniAD [35] 86.0 93.4 85.7 97.2 46.3 53.1 81.9 36.2 77.7
RD++ [31] 94.8 97.8 91.9 96.9 47.3 52.7 92.5 35.8 82.0
DesTSeg [44] 92.2 96.6 88.8 68.9 48.9 43.8 49.2 28.0 69.8
macaroni1 Aug. DRAEM [37] 53.4 47.5 64.6 22.3 00.0 00.1 02.6 00.0 27.2
SimpleNet [27] 77.4 75.7 71.4 97.0 03.8 10.6 87.0 05.6 60.4
RealNet [45] 76.2 81.5 72.0 60.1 13.2 19.7 27.5 10.9 50.0
Emb. CFA [21] 57.9 61.6 68.1 73.5 00.4 02.7 58.8 01.4 46.1
CFLOW-AD [14] 80.6 77.5 75.3 99.1 07.6 13.3 94.9 07.1 64.0
PyramidalFlow [22] 41.9 43.4 66.0 81.2 00.1 00.2 39.9 00.1 39.0
Rec. RD [10] 94.8 92.8 88.9 99.8 23.8 31.1 97.4 18.4 75.5
DiAD [16] 97.9 97.2 92.1 54.1 00.0 00.1 03.3 00.0 49.2
ViTAD [40] 84.8 82.5 75.4 98.5 07.2 16.0 90.2 08.7 64.9
InvAD [42] 94.3 92.8 87.1 99.7 21.7 30.1 97.4 17.7 74.7
InvAD-lite [42] 92.9 90.4 86.0 99.7 18.6 25.6 97.2 14.7 72.9
MambaAD [15] 91.5 89.6 82.3 99.5 17.6 26.1 96.0 15.0 71.8
Hybrid UniAD [35] 84.2 81.4 77.0 99.0 08.0 16.4 94.0 08.9 65.7
RD++ [31] 93.4 91.6 86.0 99.7 24.1 34.2 96.9 20.7 75.1
DesTSeg [44] 95.2 93.3 92.7 79.7 15.7 24.4 46.2 13.9 63.9
macaroni2 Aug. DRAEM [37] 52.3 51.5 62.8 48.7 00.0 00.1 12.2 00.0 32.5
SimpleNet [27] 66.7 61.2 66.9 90.9 00.7 04.2 79.1 02.2 52.8
RealNet [45] 57.6 57.5 69.4 51.4 05.3 03.6 16.0 01.8 37.3
Emb. CFA [21] 50.9 51.1 67.6 79.0 00.2 01.6 52.5 00.8 43.3
CFLOW-AD [14] 64.3 62.7 66.2 97.2 01.5 05.5 88.2 02.8 55.1
PyramidalFlow [22] 67.8 71.0 69.2 63.9 00.0 00.1 08.3 00.0 40.0
Rec. RD [10] 87.4 83.0 82.8 99.6 12.8 22.4 97.2 12.6 69.3
DiAD [16] 50.6 42.5 63.4 56.2 00.0 00.1 12.3 00.0 32.2
ViTAD [40] 77.1 69.9 71.4 98.0 03.4 09.1 87.5 04.8 59.5
InvAD [42] 85.9 81.9 81.2 99.6 10.8 18.1 97.4 10.0 67.8
InvAD-lite [42] 86.5 83.0 80.4 99.6 10.9 17.9 97.9 09.8 68.0
MambaAD [15] 84.5 80.6 77.1 99.5 08.8 16.9 96.5 09.2 66.3
Hybrid UniAD [35] 76.5 76.6 69.4 97.3 05.0 12.4 88.0 06.6 60.7
RD++ [31] 83.7 78.6 77.8 99.5 10.3 18.7 96.3 10.3 66.4
DesTSeg [44] 70.6 70.0 71.7 69.2 11.3 19.1 44.7 10.6 50.9
pcb1 Aug. DRAEM [37] 56.1 52.2 68.4 33.6 00.3 00.9 11.4 00.5 31.8
SimpleNet [27] 92.8 93.9 87.9 98.7 80.0 74.9 78.9 59.8 86.7
RealNet [45] 76.6 79.6 71.2 70.5 36.9 41.9 25.1 26.5 57.4
Emb. CFA [21] 80.9 81.5 75.6 93.0 62.4 65.3 46.0 48.5 72.1
CFLOW-AD [14] 93.5 93.5 86.5 99.1 72.4 68.7 87.8 52.3 85.9
PyramidalFlow [22] 83.3 80.9 82.1 92.8 49.3 45.5 40.5 29.5 67.8
Rec. RD [10] 54.8 47.6 74.9 98.6 18.8 38.3 92.5 23.7 60.8
DiAD [16] 92.8 95.3 94.5 96.6 27.9 38.4 60.8 23.8 72.3
ViTAD [40] 95.3 93.7 90.7 99.4 63.2 61.1 89.0 43.9 84.6
InvAD [42] 96.5 96.1 93.1 99.8 81.3 75.9 96.1 61.2 91.3
InvAD-lite [42] 96.2 94.4 93.7 99.8 79.3 73.2 96.1 57.8 90.4
MambaAD [15] 95.7 94.5 93.3 99.8 79.0 72.9 94.3 57.4 89.9
Hybrid UniAD [35] 94.4 93.1 91.1 99.3 62.9 59.9 89.1 42.8 84.3
RD++ [31] 96.6 96.2 91.7 99.3 67.7 61.0 95.2 43.9 86.8
DesTSeg [44] 82.7 81.2 77.0 96.3 70.0 66.4 71.9 49.7 77.9
Table A32: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pcb2 Aug. DRAEM [37] 59.6 70.8 66.7 38.0 00.1 00.4 07.3 00.2 34.7
SimpleNet [27] 90.1 91.8 83.4 96.6 13.9 23.5 80.6 13.3 68.6
RealNet [45] 71.6 78.8 67.4 64.1 12.9 19.8 32.7 11.0 49.6
Emb. CFA [21] 74.3 75.9 71.5 74.7 06.5 14.9 46.5 08.1 52.0
CFLOW-AD [14] 92.4 93.0 85.8 96.6 12.5 18.7 82.3 10.3 68.8
PyramidalFlow [22] 72.8 76.3 70.8 93.9 08.8 15.9 71.9 08.7 58.6
Rec. RD [10] 97.3 96.9 94.1 97.8 22.1 30.0 91.6 17.6 75.7
DiAD [16] 97.7 98.0 92.8 83.8 03.8 09.9 50.9 05.2 62.4
ViTAD [40] 90.2 89.4 85.4 97.9 12.9 21.0 82.7 11.7 68.5
InvAD [42] 97.0 97.1 93.1 99.1 16.1 23.8 92.5 13.5 74.1
InvAD-lite [42] 94.7 94.7 89.9 98.9 12.9 23.0 91.4 13.0 72.2
MambaAD [15] 95.2 95.0 90.2 98.8 13.9 24.1 90.0 13.7 72.5
Hybrid UniAD [35] 90.8 91.5 85.4 97.9 09.0 17.3 83.6 09.4 67.9
RD++ [31] 98.3 98.5 94.9 97.9 28.6 34.6 91.1 20.9 77.7
DesTSeg [44] 90.1 89.1 83.0 93.6 20.9 29.6 66.3 17.4 67.5
pcb3 Aug. DRAEM [37] 36.1 46.1 66.4 38.5 00.2 00.5 07.9 00.3 28.0
SimpleNet [27] 86.4 88.3 78.0 97.6 19.7 27.7 80.8 16.1 68.4
RealNet [45] 78.8 84.0 72.5 71.5 29.8 36.6 35.7 22.4 58.4
Emb. CFA [21] 63.0 68.7 66.9 85.5 11.5 20.3 39.6 11.3 50.8
CFLOW-AD [14] 80.5 83.6 73.6 96.3 22.6 26.8 80.3 15.4 66.2
PyramidalFlow [22] 53.8 56.4 66.4 56.4 00.5 02.7 07.6 01.3 34.8
Rec. RD [10] 96.2 96.1 90.7 98.0 25.3 34.8 93.6 21.1 76.4
DiAD [16] 94.2 93.8 89.1 88.5 03.8 10.2 48.0 05.4 61.1
ViTAD [40] 90.4 90.6 84.6 98.2 23.5 29.6 87.9 17.4 72.1
InvAD [42] 96.6 97.0 92.1 99.2 17.3 28.7 93.1 16.8 74.9
InvAD-lite [42] 93.8 94.3 86.1 99.1 17.6 28.2 92.6 16.4 73.1
MambaAD [15] 95.0 95.2 88.8 99.1 21.7 27.6 92.0 16.0 74.2
Hybrid UniAD [35] 82.8 83.8 76.9 98.2 13.6 24.1 80.9 13.7 65.8
RD++ [31] 95.7 95.8 89.9 98.0 33.2 35.6 93.7 21.6 77.4
DesTSeg [44] 93.4 94.1 87.3 89.9 31.9 33.6 56.9 20.2 69.6
pcb4 Aug. DRAEM [37] 68.0 58.0 76.4 32.8 00.5 01.3 06.6 00.7 34.8
SimpleNet [27] 97.6 97.8 92.0 95.2 22.1 30.9 79.0 18.2 73.5
RealNet [45] 72.5 79.2 68.3 58.3 28.0 23.1 24.0 13.1 50.5
Emb. CFA [21] 72.5 75.7 69.8 75.2 10.0 19.1 44.2 10.6 52.4
CFLOW-AD [14] 98.5 98.4 96.1 96.7 21.9 30.9 85.4 18.3 75.4
PyramidalFlow [22] 48.9 52.7 66.4 89.5 05.0 08.9 66.3 04.7 48.2
Rec. RD [10] 100. 100. 99.5 97.8 30.2 36.0 89.2 22.0 79.0
DiAD [16] 99.7 99.7 99.0 95.4 09.9 18.2 83.5 10.0 72.2
ViTAD [40] 99.1 98.7 97.0 99.0 42.8 47.9 93.4 31.5 82.6
InvAD [42] 99.8 99.8 98.0 98.6 46.6 45.1 91.1 29.2 82.7
InvAD-lite [42] 99.8 99.8 98.0 98.7 50.6 49.4 90.8 32.8 83.9
MambaAD [15] 99.6 99.7 97.0 98.7 47.4 47.6 90.8 31.2 83.0
Hybrid UniAD [35] 99.2 99.2 95.5 97.1 29.1 33.7 83.1 20.3 76.7
RD++ [31] 100. 100. 99.0 97.4 32.8 38.7 86.8 24.0 79.2
DesTSeg [44] 98.8 98.7 95.1 94.3 53.4 52.2 74.3 35.3 81.0
pipe_fryum Aug. DRAEM [37] 63.5 74.5 83.1 22.9 00.7 02.5 04.9 01.2 36.0
SimpleNet [27] 80.7 90.9 83.2 98.8 70.8 66.2 71.3 49.5 80.3
RealNet [45] 64.0 83.5 80.5 61.3 43.2 31.4 32.3 18.6 56.6
Emb. CFA [21] 48.2 71.5 80.0 90.9 53.7 52.7 59.2 35.8 65.2
CFLOW-AD [14] 97.4 98.8 96.0 99.2 60.8 61.3 94.6 44.2 86.9
PyramidalFlow [22] 41.6 63.3 80.0 81.3 04.7 09.1 43.0 04.8 46.1
Rec. RD [10] 98.1 99.0 96.5 99.1 55.8 58.2 96.3 41.1 86.1
DiAD [16] 97.4 98.5 98.0 98.6 52.5 57.1 78.4 40.0 82.9
ViTAD [40] 98.1 99.1 95.5 99.5 67.0 67.2 95.0 50.6 88.8
InvAD [42] 99.2 99.6 97.5 99.4 66.3 64.7 94.8 47.9 88.8
InvAD-lite [42] 98.6 99.3 96.0 99.0 55.3 57.4 95.7 40.2 85.9
MambaAD [15] 99.0 99.5 97.0 99.0 53.9 57.5 95.6 40.3 85.9
Hybrid UniAD [35] 95.5 97.7 92.9 99.0 52.1 58.2 93.3 41.1 84.1
RD++ [31] 96.8 98.3 94.1 99.0 57.6 58.8 95.4 41.6 85.7
DesTSeg [44] 96.3 98.4 92.1 88.8 81.4 75.2 51.5 60.2 83.4
Table A33: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 55.1 62.4 72.9 37.5 00.6 01.7 10.0 00.9 34.3
SimpleNet [27] 86.4 89.1 82.8 96.6 34.0 37.8 79.2 25.7 72.3
RealNet [45] 71.4 79.5 74.7 61.0 25.7 22.6 27.4 13.5 51.8
Emb. CFA [21] 66.3 74.3 74.2 81.3 22.1 26.2 50.8 17.0 56.5
CFLOW-AD [14] 86.5 88.8 84.9 97.7 33.9 37.2 86.8 24.9 73.7
PyramidalFlow [22] 58.2 66.3 74.4 77.0 07.2 09.6 42.8 05.6 47.9
Rec. RD [10] 90.6 90.9 89.3 98.0 35.4 42.5 91.9 27.9 76.9
DiAD [16] 84.8 88.5 86.9 82.5 17.9 23.2 44.5 14.9 61.2
ViTAD [40] 90.4 91.1 86.0 98.2 36.4 41.0 85.7 27.5 75.5
InvAD [42] 95.4 95.7 91.6 98.9 43.3 46.8 93.1 32.5 80.7
InvAD-lite [42] 94.9 95.2 90.7 98.6 40.2 44.0 93.1 29.8 79.5
MambaAD [15] 94.5 94.9 90.2 98.4 39.3 43.7 92.1 29.5 79.0
Hybrid UniAD [35] 89.0 91.0 85.8 98.3 34.5 39.6 86.5 26.4 75.0
RD++ [31] 93.9 94.7 90.2 98.4 42.3 46.3 91.9 31.2 79.7
DesTSeg [44] 89.9 91.4 86.7 86.7 46.6 47.2 61.1 32.7 72.8
Appendix H Detailed Quantitative Results on BTAD Dataset Under 100 epochs
Table A34: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
1.0 Aug. DRAEM [37] 82.6 93.6 84.7 47.0 03.0 06.3 14.3 03.3 47.4
SimpleNet [27] 98.3 99.4 97.9 95.0 41.6 45.1 64.7 29.1 77.4
RealNet [45] 100. 100. 100. 95.7 69.5 67.6 81.3 51.1 87.7
Emb. CFA [21] 98.2 99.4 99.0 94.7 37.8 44.2 64.6 28.4 76.8
PatchCore [29] 99.6 99.8 99.0 97.1 55.5 56.6 75.8 39.5 83.3
CFLOW-AD [14] 94.8 98.1 93.5 94.7 36.6 43.4 66.2 27.7 75.3
PyramidalFlow [22] 94.5 96.7 98.0 89.2 18.9 06.3 66.6 03.3 67.2
Rec. RD [10] 99.8 99.9 99.0 97.8 57.5 58.5 84.4 41.4 85.3
DiAD [16] 99.4 99.8 99.0 90.5 24.3 31.4 73.7 18.6 74.0
ViTAD [40] 97.9 99.2 95.9 96.5 55.0 55.1 69.8 38.1 81.3
InvAD [42] 99.9 100. 99.0 98.0 63.7 62.9 84.1 45.9 86.8
InvAD-lite [42] 97.8 99.2 96.9 97.3 59.3 61.4 77.3 44.3 84.2
MambaAD [15] 97.4 99.0 94.7 97.1 55.3 58.1 76.5 41.0 82.6
Hybrid UniAD [35] 99.7 99.9 98.0 97.4 57.2 58.4 79.2 41.3 84.3
RD++ [31] 98.6 99.5 97.9 97.5 53.6 57.8 80.9 40.6 83.7
DesTSeg [44] 98.1 99.3 96.9 85.8 25.4 39.3 68.2 24.4 73.3
2.0 Aug. DRAEM [37] 70.6 94.7 92.8 51.7 07.9 11.5 18.2 06.1 49.6
SimpleNet [27] 81.7 96.8 93.6 94.5 46.3 49.4 47.4 32.8 72.8
RealNet [45] 72.9 95.3 92.8 83.9 38.8 50.3 38.6 33.6 67.5
Emb. CFA [21] 80.3 96.6 92.8 94.7 51.2 53.9 48.8 36.9 74.0
PatchCore [29] 83.8 97.3 92.8 96.0 59.3 59.8 54.1 42.7 77.6
CFLOW-AD [14] 79.6 96.7 92.8 96.4 61.4 61.0 54.5 43.9 77.5
PyramidalFlow [22] 73.7 95.6 92.8 86.8 64.1 62.5 38.0 45.4 73.4
Rec. RD [10] 84.0 97.3 93.4 96.7 64.5 64.3 58.8 47.4 79.9
DiAD [16] 72.6 95.4 92.8 86.9 27.3 35.5 44.8 21.6 65.0
ViTAD [40] 83.5 97.2 93.0 96.9 74.1 67.8 51.9 51.2 80.6
InvAD [42] 88.7 98.1 93.5 96.6 62.0 61.9 59.6 44.9 80.1
InvAD-lite [42] 81.8 97.0 93.2 96.6 61.8 62.4 60.2 45.4 79.0
MambaAD [15] 82.5 97.2 93.0 96.3 55.1 60.0 59.6 42.8 77.7
Hybrid UniAD [35] 85.0 97.6 92.8 94.8 41.8 51.8 59.4 35.0 74.7
RD++ [31] 85.5 97.6 93.2 96.7 65.2 64.2 57.2 47.2 79.9
DesTSeg [44] 81.2 96.9 92.8 94.6 45.9 58.4 50.9 41.2 74.4
3.0 Aug. DRAEM [37] 60.7 47.1 56.5 48.4 00.3 01.5 16.0 00.8 32.9
SimpleNet [27] 99.7 95.7 88.6 99.4 36.5 38.5 97.5 23.8 79.4
RealNet [45] 96.3 90.5 85.7 72.2 36.0 40.1 40.3 25.1 65.9
Emb. CFA [21] 99.7 96.6 88.9 99.4 53.3 52.5 95.2 35.6 83.7
PatchCore [29] 99.8 97.3 92.1 99.5 50.1 48.2 98.0 31.8 83.6
CFLOW-AD [14] 99.0 89.5 78.6 99.4 38.8 46.0 97.3 29.9 78.4
PyramidalFlow [22] 92.9 57.1 52.2 96.7 05.9 12.0 87.7 06.4 57.8
Rec. RD [10] 99.4 92.5 89.6 99.8 56.8 54.6 98.9 37.6 84.5
DiAD [16] 98.3 69.9 86.1 97.6 07.2 13.1 92.7 07.0 66.4
ViTAD [40] 99.6 94.2 90.9 99.6 48.5 47.2 98.0 30.9 82.6
InvAD [42] 99.7 95.2 90.3 99.8 62.0 59.4 98.9 42.2 86.5
InvAD-lite [42] 99.7 96.0 93.8 99.6 56.5 55.3 98.4 38.2 85.6
MambaAD [15] 99.5 92.2 89.9 99.6 47.1 49.9 98.4 33.2 82.4
Hybrid UniAD [35] 99.8 97.5 91.8 99.5 51.9 51.3 97.9 34.5 84.2
RD++ [31] 99.7 96.4 91.2 99.8 60.1 57.6 99.0 40.4 86.3
DesTSeg [44] 99.2 91.4 87.1 96.3 33.1 35.4 90.8 21.5 76.2
Avg Aug. DRAEM [37] 71.3 78.5 78.0 49.0 03.7 06.5 16.2 03.4 43.3
SimpleNet [27] 93.2 97.3 93.3 96.3 41.5 44.3 69.8 28.6 76.5
RealNet [45] 89.7 95.3 92.8 84.0 48.1 52.7 53.4 36.6 73.7
Emb. CFA [21] 92.7 97.5 93.5 96.3 47.4 50.2 69.5 33.6 78.2
PatchCore [29] 94.4 98.2 94.6 97.5 55.0 54.9 76.0 38.0 81.5
CFLOW-AD [14] 91.2 94.8 88.3 96.8 45.6 50.1 72.7 33.8 77.1
PyramidalFlow [22] 87.0 83.1 81.0 90.9 29.6 26.9 64.1 18.3 66.1
Rec. RD [10] 94.4 96.6 94.0 98.1 59.6 59.2 80.7 42.1 83.2
DiAD [16] 90.1 88.4 92.6 91.7 19.6 26.7 70.4 15.7 68.5
ViTAD [40] 93.6 96.8 93.3 97.6 59.2 56.7 73.2 40.1 81.5
InvAD [42] 96.1 97.8 94.3 98.1 62.6 61.4 80.9 44.3 84.5
InvAD-lite [42] 93.1 97.4 94.6 97.9 59.2 59.7 78.6 42.6 82.9
MambaAD [15] 93.1 96.1 92.5 97.7 52.5 56.0 78.2 39.0 80.9
Hybrid UniAD [35] 94.8 98.3 94.2 97.2 50.3 53.8 78.8 36.9 81.1
RD++ [31] 94.6 97.8 94.1 98.0 59.6 59.8 79.0 42.8 83.3
DesTSeg [44] 92.8 95.9 92.3 92.2 34.8 44.3 70.0 29.0 74.6
Appendix I Detailed Quantitative Results on MPDD Dataset Under 100 epochs
Table A35: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bracket_black Aug. DRAEM [37] 39.0 51.7 74.6 45.1 00.1 00.4 24.0 00.2 33.6
SimpleNet [27] 74.0 83.3 78.5 92.8 03.1 09.7 86.1 05.1 61.1
RealNet [45] 63.0 72.7 75.2 65.9 02.3 06.8 50.7 03.5 48.1
Emb. CFA [21] 58.4 72.8 74.6 74.6 00.2 00.5 31.9 00.2 44.7
PatchCore [29] 88.4 92.7 83.9 98.5 11.0 21.0 96.7 11.8 70.3
CFLOW-AD [14] 60.8 70.3 75.8 95.5 01.7 04.1 85.2 02.1 56.2
PyramidalFlow [22] 65.8 72.7 77.0 95.3 01.8 05.0 81.4 02.6 57.0
Rec. RD [10] 84.2 88.3 83.1 97.7 06.6 12.7 94.1 06.8 66.7
DiAD [16] 70.9 81.7 79.0 89.8 00.4 01.0 65.3 00.5 55.5
ViTAD [40] 78.3 86.2 83.6 95.5 02.5 08.0 89.1 04.2 63.3
InvAD [42] 83.4 90.9 81.0 97.4 07.1 15.2 94.0 08.2 67.0
InvAD-lite [42] 75.4 84.6 78.6 95.2 07.9 16.8 89.6 09.1 64.0
MambaAD [15] 81.9 88.3 81.7 94.5 05.3 11.2 88.4 05.9 64.5
Hybrid UniAD [35] 78.1 84.9 79.2 93.0 00.6 01.6 79.2 00.8 59.5
RD++ [31] 82.7 86.6 82.9 98.0 04.2 10.1 94.1 05.3 65.5
DesTSeg [44] 83.4 87.3 85.7 72.4 05.7 13.5 59.0 07.3 58.1
bracket_brown Aug. DRAEM [37] 42.5 61.1 79.7 46.8 00.8 01.9 27.5 01.0 37.2
SimpleNet [27] 90.4 93.8 93.6 94.4 07.9 16.9 86.7 09.2 69.1
RealNet [45] 91.0 94.0 92.6 69.9 12.1 19.1 50.5 10.6 61.3
Emb. CFA [21] 88.7 94.9 88.9 85.1 03.2 06.9 52.6 03.6 60.0
PatchCore [29] 95.6 97.2 95.3 98.6 26.2 32.0 95.4 19.0 77.2
CFLOW-AD [14] 74.4 77.6 85.2 95.9 05.9 12.0 91.1 06.4 63.2
PyramidalFlow [22] 63.6 76.4 81.0 94.0 05.7 12.5 69.6 06.7 57.5
Rec. RD [10] 86.2 89.6 92.7 96.0 13.1 23.1 93.9 13.1 70.7
DiAD [16] 53.2 69.1 82.9 90.0 02.8 07.0 66.6 03.6 53.1
ViTAD [40] 86.5 86.0 92.6 98.0 16.3 24.9 95.0 14.2 71.3
InvAD [42] 93.7 95.5 95.2 95.5 15.2 22.8 90.8 12.8 72.7
InvAD-lite [42] 89.5 93.0 90.1 97.9 14.9 25.0 94.3 14.3 72.1
MambaAD [15] 95.7 97.5 95.3 97.5 26.3 30.8 92.8 18.2 76.6
Hybrid UniAD [35] 91.4 95.0 92.7 98.0 23.9 32.0 90.7 19.0 74.8
RD++ [31] 84.4 89.9 91.9 96.6 17.2 24.4 94.9 13.9 71.3
DesTSeg [44] 95.0 92.5 96.2 61.6 04.8 10.5 30.1 05.5 55.8
bracket_white Aug. DRAEM [37] 34.0 44.3 66.7 42.3 00.0 00.1 09.0 00.0 28.1
SimpleNet [27] 85.6 89.3 80.0 97.7 02.1 06.5 85.6 03.4 63.8
RealNet [45] 77.3 86.2 80.8 85.2 18.4 30.6 47.4 18.0 60.8
Emb. CFA [21] 80.2 86.0 77.8 89.8 03.9 14.9 47.3 08.0 57.1
PatchCore [29] 93.3 94.8 88.9 98.8 10.0 19.3 93.9 10.7 71.3
CFLOW-AD [14] 62.9 64.1 69.8 98.2 01.2 03.5 93.0 01.8 56.1
PyramidalFlow [22] 76.8 84.8 78.4 98.9 05.6 14.5 91.1 07.8 64.3
Rec. RD [10] 85.3 86.5 81.4 99.0 02.3 06.4 95.2 03.3 65.2
DiAD [16] 84.0 90.7 88.9 96.5 01.0 03.6 82.5 01.8 63.9
ViTAD [40] 78.9 84.0 78.7 95.7 00.9 02.9 86.7 01.5 61.1
InvAD [42] 92.1 94.7 90.0 98.9 09.8 21.9 94.1 12.3 71.6
InvAD-lite [42] 90.2 92.2 84.7 99.2 05.2 10.6 94.8 05.6 68.1
MambaAD [15] 97.1 97.5 90.9 99.3 12.2 20.4 95.0 11.3 73.2
Hybrid UniAD [35] 63.6 67.1 69.6 95.0 00.3 00.8 79.6 00.4 53.7
RD++ [31] 81.4 86.0 76.4 98.7 02.9 10.0 93.4 05.2 64.1
DesTSeg [44] 91.1 89.8 88.5 90.4 01.0 02.7 51.2 01.4 59.2
connector Aug. DRAEM [37] 21.7 22.2 48.3 35.1 00.6 02.1 16.8 01.1 21.0
SimpleNet [27] 96.2 92.0 90.3 98.3 46.4 47.0 94.5 30.7 80.7
RealNet [45] 99.0 97.9 96.6 85.8 38.9 44.1 75.8 28.3 76.9
Emb. CFA [21] 96.9 89.0 93.3 83.5 11.2 16.3 59.2 08.9 64.2
PatchCore [29] 100. 100. 100. 99.5 71.2 63.3 98.4 46.3 90.3
CFLOW-AD [14] 91.2 85.3 78.6 96.7 27.9 29.3 89.4 17.2 71.2
PyramidalFlow [22] 69.0 42.7 63.4 90.7 03.7 07.2 70.0 03.7 49.5
Rec. RD [10] 100. 100. 100. 99.2 60.6 55.5 97.5 38.4 87.5
DiAD [16] 57.1 48.0 56.3 90.1 02.9 06.5 68.4 03.4 47.1
ViTAD [40] 96.2 91.9 86.7 99.3 44.3 47.6 97.7 31.2 80.5
InvAD [42] 96.7 83.4 96.6 98.9 57.9 58.1 96.4 41.0 84.0
InvAD-lite [42] 97.1 90.8 93.3 99.2 62.2 58.1 97.3 41.0 85.4
MambaAD [15] 99.5 99.1 96.3 99.2 54.5 56.5 97.3 39.3 86.1
Hybrid UniAD [35] 69.8 49.9 58.5 95.2 07.6 14.9 83.9 08.1 54.3
RD++ [31] 100. 100. 100. 99.2 67.0 61.8 97.4 44.8 89.3
DesTSeg [44] 92.1 81.4 82.4 76.6 33.8 37.6 61.1 23.1 66.4
Table A36: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
metal_plate Aug. DRAEM [37] 41.9 76.3 84.5 36.3 10.8 21.2 13.3 11.9 40.6
SimpleNet [27] 100. 100. 100. 98.4 89.8 81.7 89.2 69.1 94.2
RealNet [45] 99.9 100. 99.3 97.7 90.5 84.1 93.6 72.6 95.0
Emb. CFA [21] 99.7 99.9 98.6 97.6 89.8 81.8 83.1 69.2 92.9
PatchCore [29] 100. 100. 100. 99.0 93.5 86.1 94.1 75.7 96.1
CFLOW-AD [14] 99.9 100. 99.3 98.5 89.0 83.3 92.4 71.4 94.6
PyramidalFlow [22] 96.7 98.7 95.2 88.7 61.3 21.1 60.8 11.8 74.6
Rec. RD [10] 100. 100. 100. 99.1 92.9 87.7 96.7 78.1 96.6
DiAD [16] 89.3 96.4 89.9 88.4 54.6 52.8 55.5 35.9 75.3
ViTAD [40] 100. 100. 100. 99.3 95.5 88.0 94.9 78.6 96.8
InvAD [42] 100. 100. 100. 99.0 92.2 87.2 96.0 77.2 96.3
InvAD-lite [42] 99.8 99.9 98.6 98.4 86.4 82.9 94.2 70.8 94.3
MambaAD [15] 99.8 99.9 98.6 98.2 83.8 82.2 94.3 69.8 93.8
Hybrid UniAD [35] 47.5 74.3 85.0 90.7 41.2 56.9 73.6 39.8 67.0
RD++ [31] 99.8 99.9 98.6 99.1 92.1 87.2 95.7 77.3 96.1
DesTSeg [44] 100. 100. 100. 95.6 91.4 87.0 88.8 77.0 94.7
tubes Aug. DRAEM [37] 34.8 64.0 81.2 47.0 01.1 02.0 23.2 01.0 36.2
SimpleNet [27] 84.0 93.9 85.0 97.6 42.8 45.8 91.7 29.7 77.3
RealNet [45] 80.1 90.1 85.2 95.3 54.3 53.0 90.4 36.1 78.3
Emb. CFA [21] 65.4 83.7 81.2 78.8 09.2 17.4 47.1 09.5 54.7
PatchCore [29] 90.4 96.6 92.4 98.8 64.5 63.8 95.5 46.9 86.0
CFLOW-AD [14] 65.0 83.3 81.7 96.2 32.0 35.6 86.0 21.7 68.5
PyramidalFlow [22] 69.8 86.6 81.2 97.1 48.2 46.4 90.3 30.2 74.2
Rec. RD [10] 91.9 97.1 93.3 99.0 67.1 65.6 96.0 48.8 87.1
DiAD [16] 55.4 81.1 83.6 87.7 03.9 07.6 58.2 04.0 53.9
ViTAD [40] 86.6 94.8 86.6 98.5 51.5 53.3 93.5 36.4 80.7
InvAD [42] 96.4 98.6 95.5 99.3 72.3 69.0 97.4 52.7 89.8
InvAD-lite [42] 93.0 96.9 91.3 98.3 61.4 62.0 94.0 44.9 85.3
MambaAD [15] 58.4 77.0 81.7 96.1 19.6 27.8 85.8 16.1 63.8
Hybrid UniAD [35] 72.5 84.8 83.1 91.4 08.6 13.5 71.1 07.2 60.7
RD++ [31] 93.0 97.5 93.0 99.3 74.6 70.8 97.2 54.8 89.3
DesTSeg [44] 86.4 93.8 88.4 95.6 59.2 56.1 89.4 39.0 81.3
Avg Aug. DRAEM [37] 35.6 53.3 72.5 42.1 02.2 04.6 19.0 02.5 32.8
SimpleNet [27] 88.4 92.0 87.9 96.5 32.0 34.6 89.0 24.5 74.3
RealNet [45] 85.1 90.2 88.3 83.3 36.1 39.6 68.1 28.2 70.1
Emb. CFA [21] 81.6 87.7 85.7 84.9 19.6 22.9 53.5 16.6 62.3
PatchCore [29] 94.6 96.9 93.4 98.9 46.1 47.6 95.7 35.0 81.9
CFLOW-AD [14] 75.7 80.1 81.7 96.8 26.3 28.0 89.5 20.1 68.3
PyramidalFlow [22] 73.6 77.0 79.4 94.1 21.1 17.8 77.2 10.4 62.9
Rec. RD [10] 91.3 93.6 91.8 98.3 40.4 41.8 95.5 31.4 79.0
DiAD [16] 68.3 77.9 80.1 90.4 10.9 13.1 66.1 08.2 58.1
ViTAD [40] 87.8 90.5 88.0 97.7 35.2 37.4 92.8 27.7 75.6
InvAD [42] 93.7 93.9 93.0 98.2 42.4 45.7 94.8 34.0 80.2
InvAD-lite [42] 90.9 92.9 89.5 98.0 39.7 42.6 94.0 30.9 78.2
MambaAD [15] 88.7 93.2 90.8 97.5 33.6 38.1 92.3 26.8 76.3
Hybrid UniAD [35] 70.5 76.0 78.0 93.9 13.7 19.9 79.7 12.5 61.7
RD++ [31] 90.2 93.3 90.5 98.5 43.0 44.1 95.5 33.6 79.3
DesTSeg [44] 91.3 90.8 90.2 82.0 32.6 34.6 63.3 25.6 69.3
Appendix J Detailed Quantitative Results on MAD_Real Dataset Under 100 epochs
Table A37: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Bear Aug. DRAEM [37] 43.3 83.5 92.3 54.0 00.2 01.1 34.1 00.5 44.1
SimpleNet [27] 24.2 72.7 90.6 94.9 03.0 10.4 87.2 05.5 54.7
Emb. CFA [21] 15.0 69.1 90.6 99.1 04.2 10.3 96.5 05.5 55.0
PatchCore [29] 84.2 96.8 91.7 99.6 17.6 29.5 98.6 17.3 74.0
CFLOW-AD [14] 76.7 94.4 92.3 99.4 06.0 12.3 97.6 06.6 68.4
PyramidalFlow [22] 46.7 78.6 90.6 98.2 02.3 04.9 93.9 02.5 59.3
Rec. RD [10] 76.7 94.6 94.1 97.7 05.2 07.9 93.2 04.1 67.1
DiAD [16] 52.5 88.7 90.6 97.4 05.1 13.5 90.3 07.2 62.6
ViTAD [40] 35.0 76.8 90.6 97.3 01.9 06.9 91.0 03.6 57.1
InvAD [42] 95.8 99.1 98.0 99.7 39.9 44.1 98.9 28.3 82.2
InvAD-lite [42] 82.5 96.1 93.9 98.4 05.0 11.6 95.1 06.2 68.9
MambaAD [15] 32.5 75.1 90.6 93.3 01.0 04.0 80.3 02.0 53.8
Hybrid UniAD [35] 83.3 96.5 90.6 99.0 04.0 09.6 96.3 05.0 68.5
RD++ [31] 77.5 94.4 93.9 99.0 08.5 15.3 96.8 08.3 69.3
DesTSeg [44] 26.7 76.9 90.6 80.3 00.4 01.3 56.6 00.6 47.5
Bird Aug. DRAEM [37] 60.9 89.6 89.8 57.7 00.7 01.0 22.1 00.5 46.0
SimpleNet [27] 58.2 87.8 91.7 82.7 02.4 06.2 53.7 03.2 54.7
Emb. CFA [21] 55.5 83.1 89.8 86.6 04.6 10.9 56.9 05.7 55.3
PatchCore [29] 71.8 92.6 93.6 96.4 32.2 39.7 84.2 24.8 72.9
CFLOW-AD [14] 35.5 74.4 89.8 89.3 02.7 06.3 64.4 03.3 51.8
PyramidalFlow [22] 17.3 69.0 89.8 91.2 02.9 06.4 66.7 03.3 49.0
Rec. RD [10] 51.8 85.9 89.8 87.1 07.7 16.6 57.2 09.1 56.6
DiAD [16] 09.1 70.5 89.8 93.1 04.3 10.4 71.1 05.5 49.8
ViTAD [40] 67.3 89.7 93.6 89.6 09.8 17.6 62.6 09.7 61.5
InvAD [42] 84.5 96.4 93.3 89.0 10.1 20.0 66.1 11.1 65.6
InvAD-lite [42] 53.6 78.6 91.7 92.8 06.2 13.4 73.8 07.2 58.6
MambaAD [15] 86.4 97.1 91.3 95.1 17.8 28.5 77.9 16.6 70.6
Hybrid UniAD [35] 38.2 74.3 89.8 87.1 02.4 06.5 63.3 03.4 51.7
RD++ [31] 45.5 82.2 89.8 89.6 06.8 15.4 63.2 08.4 56.1
DesTSeg [44] 60.0 82.7 93.6 66.2 05.4 13.8 35.4 07.4 51.0
Elephant Aug. DRAEM [37] 77.8 93.9 90.0 50.1 00.4 01.2 18.7 00.6 47.4
SimpleNet [27] 20.0 68.9 87.8 76.4 06.1 14.5 42.9 07.8 45.2
Emb. CFA [21] 35.6 71.5 87.8 86.8 05.7 10.6 58.4 05.6 50.9
PatchCore [29] 27.8 69.5 87.8 93.7 11.1 18.3 78.8 10.1 55.3
CFLOW-AD [14] 23.3 71.6 87.8 90.6 08.6 11.5 66.4 06.1 51.4
PyramidalFlow [22] 51.1 84.2 87.8 93.0 02.8 06.0 76.5 03.1 57.3
Rec. RD [10] 43.3 82.7 87.8 88.2 02.1 05.4 65.6 02.8 53.6
DiAD [16] 36.7 76.9 90.0 91.8 02.0 04.3 74.1 02.2 53.7
ViTAD [40] 46.7 77.4 87.8 96.1 09.1 18.0 85.6 09.9 60.1
InvAD [42] 32.2 72.6 87.8 91.3 06.5 13.2 71.4 07.0 53.6
InvAD-lite [42] 36.7 72.7 87.8 91.5 04.7 10.1 75.2 05.3 54.1
MambaAD [15] 47.8 79.2 87.8 90.7 04.7 09.2 72.5 04.8 56.0
Hybrid UniAD [35] 11.1 63.1 87.8 79.1 01.3 04.0 38.7 02.0 40.7
RD++ [31] 41.1 80.3 87.8 85.2 01.7 04.1 59.6 02.1 51.4
DesTSeg [44] 58.9 83.6 90.0 66.6 01.2 05.0 32.6 02.6 48.3
Parrot Aug. DRAEM [37] 64.3 87.0 92.0 49.2 00.6 01.3 14.6 00.6 44.1
SimpleNet [27] 16.5 71.6 90.2 61.9 00.9 01.8 32.7 00.9 39.4
Emb. CFA [21] 04.3 65.9 90.2 42.3 00.5 01.3 17.0 00.7 31.6
PatchCore [29] 05.2 66.2 90.2 81.7 02.1 05.1 64.0 02.6 44.9
CFLOW-AD [14] 16.5 68.0 90.2 66.7 01.0 02.5 41.9 01.3 41.0
PyramidalFlow [22] 25.2 77.3 90.2 64.2 00.9 02.2 31.8 01.1 41.7
Rec. RD [10] 01.7 64.6 90.2 64.3 00.9 02.0 34.3 01.0 36.9
DiAD [16] 60.9 91.6 90.2 63.6 00.9 02.0 41.7 01.0 50.1
ViTAD [40] 16.5 69.8 90.2 69.8 01.1 02.6 40.8 01.3 41.5
InvAD [42] 11.3 68.6 90.2 52.1 00.7 01.4 32.7 00.7 36.7
InvAD-lite [42] 04.3 65.4 90.2 60.4 00.8 01.8 38.4 00.9 37.3
MambaAD [15] 06.1 66.6 90.2 72.4 01.1 02.6 45.6 01.3 40.7
Hybrid UniAD [35] 33.9 76.4 90.2 61.2 00.8 01.8 37.7 00.9 43.1
RD++ [31] 07.0 66.2 90.2 63.1 00.9 02.0 33.5 01.0 37.6
DesTSeg [44] 03.5 65.1 90.2 48.7 00.6 01.3 18.5 00.6 32.6
Table A38: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Pig Aug. DRAEM [37] 67.1 89.5 87.2 52.2 00.5 01.2 23.8 00.6 45.9
SimpleNet [27] 49.4 76.5 87.2 93.9 10.1 21.9 72.9 12.3 58.8
Emb. CFA [21] 55.3 77.6 89.5 93.0 05.0 11.2 73.0 05.9 57.8
PatchCore [29] 56.5 78.8 89.5 98.5 19.9 28.7 89.7 16.7 65.9
CFLOW-AD [14] 57.6 78.8 87.2 96.5 05.8 12.0 85.2 06.4 60.4
PyramidalFlow [22] 57.6 84.2 87.2 95.7 05.2 10.0 81.9 05.3 60.3
Rec. RD [10] 60.0 80.2 89.5 98.4 21.1 31.0 92.5 18.3 67.5
DiAD [16] 56.5 87.9 87.2 92.9 02.7 05.9 71.7 03.0 57.8
ViTAD [40] 37.6 71.6 87.2 97.7 08.6 17.2 89.8 09.4 58.5
InvAD [42] 58.8 77.1 89.5 98.1 11.7 22.1 91.1 12.4 64.1
InvAD-lite [42] 47.1 73.7 89.5 98.4 13.0 24.1 92.9 13.7 62.7
MambaAD [15] 61.2 80.3 87.2 97.7 12.2 22.8 87.1 12.8 64.1
Hybrid UniAD [35] 50.6 83.9 87.2 88.8 01.6 04.2 62.2 02.1 54.1
RD++ [31] 61.2 80.3 91.4 98.0 14.3 23.0 91.0 13.0 65.6
DesTSeg [44] 48.2 74.8 89.5 71.1 02.3 06.0 43.6 03.1 47.9
Puppy Aug. DRAEM [37] 53.0 82.9 88.9 46.3 00.4 01.9 23.9 01.0 42.5
SimpleNet [27] 68.0 90.6 90.9 94.5 04.1 10.3 72.4 05.4 61.5
Emb. CFA [21] 83.0 95.3 93.0 98.6 07.1 14.5 92.3 07.8 69.1
PatchCore [29] 100. 100. 100. 99.2 13.4 25.1 92.1 14.3 75.7
CFLOW-AD [14] 79.0 94.3 93.0 99.2 27.9 33.3 94.2 20.0 74.4
PyramidalFlow [22] 56.0 85.5 88.9 95.1 07.6 16.9 75.0 09.3 60.7
Rec. RD [10] 78.0 94.4 90.0 95.8 04.4 09.1 87.2 04.7 65.6
DiAD [16] 76.0 94.7 88.9 94.6 01.8 05.0 78.3 02.6 62.8
ViTAD [40] 58.0 87.6 88.9 92.5 01.2 02.9 75.8 01.5 58.1
InvAD [42] 100. 100. 100. 99.8 45.6 44.9 98.1 28.9 84.1
InvAD-lite [42] 96.0 99.0 97.6 96.9 05.9 11.9 90.3 06.3 71.1
MambaAD [15] 82.0 94.7 92.3 88.1 01.8 05.3 71.1 02.7 62.2
Hybrid UniAD [35] 74.0 92.9 90.9 98.6 14.1 21.9 91.0 12.3 69.1
RD++ [31] 66.0 87.6 93.0 98.1 04.8 11.2 92.4 05.9 64.7
DesTSeg [44] 49.0 81.8 88.9 97.6 11.2 11.6 87.7 06.2 61.1
Scorpion Aug. DRAEM [37] 33.0 80.7 90.2 46.8 01.4 02.9 14.6 01.5 38.5
SimpleNet [27] 43.5 84.0 90.2 70.9 06.0 13.0 33.4 06.9 48.7
Emb. CFA [21] 30.4 81.6 90.2 60.9 05.6 13.4 26.4 07.2 44.1
PatchCore [29] 13.0 70.7 90.2 88.2 08.8 17.0 61.6 09.3 49.9
CFLOW-AD [14] 64.3 90.1 90.2 79.1 04.0 07.8 35.9 04.1 53.1
PyramidalFlow [22] 47.8 85.3 90.2 68.0 02.7 05.9 22.1 03.0 46.0
Rec. RD [10] 64.3 91.8 90.2 64.7 02.0 04.8 20.3 02.5 48.3
DiAD [16] 56.5 79.2 93.9 73.2 02.9 07.0 30.8 03.6 49.1
ViTAD [40] 46.1 84.4 90.2 62.9 01.9 04.2 17.3 02.2 43.9
InvAD [42] 45.2 86.6 90.2 68.4 02.3 05.0 22.0 02.6 45.7
InvAD-lite [42] 50.4 83.8 90.2 67.4 02.2 04.9 18.6 02.5 45.4
MambaAD [15] 53.0 84.0 90.2 66.6 02.1 04.8 17.0 02.4 45.4
Hybrid UniAD [35] 32.2 75.9 90.2 81.2 04.8 10.1 44.9 05.3 48.5
RD++ [31] 22.6 73.5 90.2 70.9 02.5 05.8 26.5 03.0 41.7
DesTSeg [44] 31.3 81.2 90.2 48.6 01.3 02.9 11.1 01.5 38.1
Turtle Aug. DRAEM [37] 73.3 93.3 89.4 47.9 01.3 04.0 31.0 02.0 48.6
SimpleNet [27] 95.2 99.0 95.2 98.2 15.2 24.7 93.1 14.1 74.4
Emb. CFA [21] 100. 100. 100. 99.3 41.9 43.0 96.8 27.4 83.0
PatchCore [29] 100. 100. 100. 99.5 43.1 48.4 97.9 31.9 84.1
CFLOW-AD [14] 99.0 99.8 97.7 99.1 30.0 34.8 96.7 21.1 79.6
PyramidalFlow [22] 81.9 95.6 90.5 97.7 15.1 24.5 92.8 13.9 71.2
Rec. RD [10] 90.5 98.0 92.3 98.4 22.8 29.0 94.4 17.0 75.1
DiAD [16] 96.2 99.1 95.5 93.2 07.0 13.8 82.6 07.4 69.6
ViTAD [40] 83.8 96.0 91.3 95.6 06.3 12.8 85.6 06.9 67.3
InvAD [42] 100. 100. 100. 99.7 53.8 55.4 98.7 38.3 86.8
InvAD-lite [42] 96.2 99.2 95.2 97.9 17.7 25.3 94.8 14.5 75.2
MambaAD [15] 87.6 97.1 91.3 96.3 13.7 20.1 90.2 11.2 70.9
Hybrid UniAD [35] 94.3 98.7 95.5 97.8 18.7 28.5 89.6 16.6 74.7
RD++ [31] 100. 100. 100. 99.2 24.3 38.2 96.7 23.6 79.8
DesTSeg [44] 67.6 90.7 89.4 95.1 20.6 26.3 86.2 15.2 68.0
Table A39: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Unicorn Aug. DRAEM [37] 72.4 91.3 91.3 47.5 00.2 00.5 14.5 00.2 45.4
SimpleNet [27] 17.1 70.1 89.4 61.4 00.4 01.5 23.8 00.7 37.7
Emb. CFA [21] 34.3 76.8 89.4 68.2 00.8 03.5 33.8 01.8 43.8
PatchCore [29] 43.8 74.1 93.3 97.3 06.8 14.1 90.3 07.6 60.0
CFLOW-AD [14] 72.4 93.2 91.3 90.3 01.4 03.0 66.6 01.5 59.7
PyramidalFlow [22] 46.7 77.8 89.4 83.6 00.9 02.3 52.2 01.2 50.4
Rec. RD [10] 40.0 76.4 89.4 77.0 00.6 01.4 39.3 00.7 46.3
DiAD [16] 42.9 85.6 89.4 83.4 00.8 02.0 56.4 01.0 51.5
ViTAD [40] 20.0 68.6 89.4 74.7 00.5 01.1 32.1 00.6 40.9
InvAD [42] 37.1 76.5 89.4 85.3 01.0 02.5 54.8 01.3 49.5
InvAD-lite [42] 40.0 75.4 89.4 79.0 00.9 02.3 51.5 01.2 48.4
MambaAD [15] 18.1 70.9 89.4 79.1 00.9 03.1 47.0 01.6 44.1
Hybrid UniAD [35] 21.9 68.2 89.4 91.8 01.7 04.1 72.6 02.1 50.0
RD++ [31] 37.1 77.2 89.4 83.1 00.8 01.8 48.4 00.9 48.3
DesTSeg [44] 33.3 74.0 89.4 59.0 00.4 01.3 29.3 00.7 41.0
Whale Aug. DRAEM [37] 46.2 89.5 92.8 42.6 00.3 01.5 25.6 00.7 42.6
SimpleNet [27] 86.2 97.4 94.1 96.4 05.6 11.7 86.4 06.2 68.3
Emb. CFA [21] 80.0 96.2 93.5 98.6 29.5 32.8 90.5 19.6 74.4
PatchCore [29] 95.0 99.1 98.5 99.7 46.8 47.4 97.8 31.0 83.5
CFLOW-AD [14] 73.1 95.3 92.8 99.1 19.6 30.1 95.0 17.7 72.1
PyramidalFlow [22] 90.6 98.1 98.5 97.4 09.3 15.4 88.7 08.3 71.1
Rec. RD [10] 48.8 89.5 92.8 98.7 15.6 21.0 93.7 11.7 65.7
DiAD [16] 11.9 76.2 92.8 94.0 01.7 04.4 76.4 02.2 51.0
ViTAD [40] 58.8 91.4 92.8 96.8 04.1 09.8 85.2 05.2 62.7
InvAD [42] 100. 100. 100. 99.8 49.4 53.9 98.6 36.9 86.0
InvAD-lite [42] 96.9 99.5 97.0 99.6 34.2 37.8 98.5 23.3 80.5
MambaAD [15] 86.9 97.9 92.8 99.1 19.3 28.1 96.5 16.4 74.4
Hybrid UniAD [35] 36.2 83.4 92.8 97.8 08.2 14.9 89.7 08.0 60.4
RD++ [31] 88.8 97.8 97.0 99.3 29.9 32.3 96.0 19.2 77.3
DesTSeg [44] 75.6 92.1 95.4 94.1 06.3 12.9 81.9 06.9 65.5
Avg Aug. DRAEM [37] 59.1 88.1 90.4 49.4 00.6 01.7 22.3 00.8 44.5
SimpleNet [27] 47.8 81.9 90.7 83.1 05.4 11.6 59.8 06.3 54.3
Emb. CFA [21] 49.3 81.7 91.4 83.3 10.5 15.2 64.2 08.7 56.5
PatchCore [29] 59.7 84.8 93.5 95.4 20.2 27.3 85.5 16.6 66.6
CFLOW-AD [14] 59.8 86.0 91.2 90.9 10.7 15.4 74.4 08.8 61.2
PyramidalFlow [22] 52.1 83.6 90.3 88.4 05.0 09.4 68.2 05.1 56.7
Rec. RD [10] 55.5 85.8 90.6 87.0 08.2 12.8 67.8 07.2 58.2
DiAD [16] 49.9 85.1 90.8 87.7 02.9 06.8 67.4 03.6 55.8
ViTAD [40] 47.0 81.3 90.2 87.3 04.5 09.3 66.6 05.0 55.2
InvAD [42] 66.5 87.7 93.8 88.3 22.1 26.2 73.2 16.8 65.4
InvAD-lite [42] 60.4 84.3 92.2 88.2 09.0 14.3 72.9 08.1 60.2
MambaAD [15] 56.2 84.3 90.3 87.8 07.5 12.8 68.5 07.2 58.2
Hybrid UniAD [35] 47.6 81.3 90.4 88.2 05.8 10.6 68.6 05.8 56.1
RD++ [31] 54.7 83.9 92.3 88.5 09.4 14.9 70.4 08.5 59.2
DesTSeg [44] 45.4 80.3 90.7 72.7 05.0 08.2 48.3 04.5 50.1
Appendix K Detailed Quantitative Results on MAD_Sim Dataset Under 100 epochs
Table A40: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
SimpleNet [27] 62.7 94.7 94.9 91.1 08.0 15.0 69.5 08.1 62.3
Emb. CFA [21] 56.4 93.3 94.9 80.3 09.9 19.5 63.7 10.8 59.7
CFLOW-AD [14] 51.4 92.1 95.1 90.8 04.6 05.2 65.1 02.7 57.8
PyramidalFlow [22] 39.2 89.3 94.9 87.3 03.3 05.0 70.3 02.6 55.6
Rec. RD [10] 57.6 92.0 94.9 93.3 04.9 09.6 77.2 05.0 61.4
DiAD [16] 78.2 97.3 94.9 91.3 03.4 08.1 63.2 04.2 62.3
ViTAD [40] 62.4 92.9 95.1 93.5 05.3 10.0 76.1 05.3 62.2
InvAD [42] 66.9 95.3 94.9 95.1 15.8 25.2 82.8 14.4 68.0
InvAD-lite [42] 68.9 95.0 95.1 93.6 06.7 12.9 78.5 06.9 64.4
MambaAD [15] 72.0 95.5 94.9 92.8 05.5 10.4 75.6 05.5 63.8
Hybrid UniAD [35] 49.6 91.1 94.9 91.9 04.1 07.9 76.1 04.1 59.4
RD++ [31] 54.9 91.1 94.9 92.9 04.2 08.5 75.8 04.5 60.3
DesTSeg [44] 55.4 92.5 94.9 54.4 04.2 06.3 32.1 03.2 48.5
02Unicorn Aug. DRAEM [37] 51.6 91.4 95.2 57.1 00.7 00.9 23.6 00.5 45.8
SimpleNet [27] 76.0 96.7 95.2 83.7 02.0 05.0 59.4 02.6 59.7
Emb. CFA [21] 66.3 95.7 95.5 72.7 01.3 02.9 43.6 01.5 54.0
CFLOW-AD [14] 53.7 93.0 95.2 84.5 02.6 03.7 55.2 01.9 55.4
PyramidalFlow [22] 42.9 89.5 95.5 73.4 00.8 01.8 46.3 00.9 50.0
Rec. RD [10] 60.1 93.9 95.2 85.3 02.0 04.2 71.4 02.1 58.9
DiAD [16] 71.0 96.6 95.2 78.9 01.4 03.9 45.8 02.0 56.1
ViTAD [40] 63.8 95.2 95.2 86.4 02.4 05.6 67.6 02.9 59.5
InvAD [42] 79.4 97.5 95.4 87.5 04.1 11.1 75.6 05.9 64.4
InvAD-lite [42] 72.4 96.6 95.2 86.2 02.5 05.4 73.2 02.8 61.6
MambaAD [15] 66.8 95.7 95.2 84.6 02.7 05.4 69.6 02.8 60.0
Hybrid UniAD [35] 64.7 95.1 95.2 86.2 02.1 05.3 61.9 02.7 58.6
RD++ [31] 64.4 94.5 95.2 85.3 02.2 04.8 69.5 02.5 59.4
DesTSeg [44] 68.3 95.5 95.5 53.2 04.7 06.4 29.3 03.3 50.4
03Mallard Aug. DRAEM [37] 38.8 88.1 95.4 49.8 01.3 01.9 18.2 01.0 41.9
SimpleNet [27] 36.2 87.6 95.4 84.9 04.8 12.1 55.4 06.4 53.8
Emb. CFA [21] 32.2 87.5 95.4 75.4 04.6 12.8 46.7 06.9 50.7
CFLOW-AD [14] 41.3 91.4 95.4 84.9 03.7 05.1 53.1 02.6 53.6
PyramidalFlow [22] 46.0 88.7 95.4 82.1 02.4 04.9 55.0 02.5 53.5
Rec. RD [10] 47.9 91.2 95.4 84.3 04.5 09.1 57.9 04.8 55.8
DiAD [16] 74.2 96.4 95.4 82.4 02.3 04.5 47.5 02.3 57.5
ViTAD [40] 44.9 91.3 95.4 85.5 05.8 11.9 60.4 06.3 56.5
InvAD [42] 44.7 90.0 95.4 85.9 07.2 14.5 61.6 07.8 57.0
InvAD-lite [42] 39.9 89.3 95.4 85.1 05.2 11.7 59.6 06.2 55.2
MambaAD [15] 57.4 93.6 95.4 84.1 05.1 10.5 57.6 05.5 57.7
Hybrid UniAD [35] 29.6 86.4 95.4 85.3 03.4 08.6 59.5 04.5 52.6
RD++ [31] 41.2 88.7 95.4 84.2 03.5 07.8 57.7 04.1 54.1
DesTSeg [44] 32.1 88.0 95.4 56.5 03.7 07.8 32.2 04.0 45.1
04Turtle Aug. DRAEM [37] 40.6 87.6 94.9 53.6 01.0 01.3 27.2 00.7 43.7
SimpleNet [27] 41.9 90.6 94.9 82.8 02.9 06.0 51.7 03.1 53.0
Emb. CFA [21] 27.8 87.2 94.9 71.7 02.6 03.9 45.7 02.0 47.7
CFLOW-AD [14] 64.5 94.5 94.9 86.5 04.2 06.4 58.0 03.3 58.4
PyramidalFlow [22] 33.7 86.7 94.9 80.1 01.6 04.0 48.3 02.0 49.9
Rec. RD [10] 62.7 93.7 94.9 91.3 09.4 16.9 72.5 09.2 63.1
DiAD [16] 86.2 98.4 95.8 91.2 06.1 12.6 60.9 06.7 64.4
ViTAD [40] 59.0 93.6 94.9 90.8 07.8 15.6 70.0 08.5 61.7
InvAD [42] 75.8 96.0 95.1 93.1 13.5 24.6 78.5 14.0 68.1
InvAD-lite [42] 61.8 93.7 94.9 91.8 10.8 19.7 74.2 10.9 63.8
MambaAD [15] 68.9 95.5 94.9 91.0 09.3 18.0 71.2 09.9 64.1
Hybrid UniAD [35] 45.1 88.6 94.9 87.2 03.2 05.8 66.0 03.0 55.8
RD++ [31] 61.9 94.4 94.9 90.5 06.9 12.8 71.2 06.8 61.8
DesTSeg [44] 35.2 89.3 94.9 52.3 02.6 04.0 36.6 02.0 45.0
Table A41: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
05Whale Aug. DRAEM [37] 54.6 91.4 94.4 47.3 00.7 01.4 25.9 00.7 45.1
SimpleNet [27] 63.1 93.9 94.4 82.6 02.3 04.9 60.5 02.5 57.4
Emb. CFA [21] 65.2 94.4 94.4 73.4 02.3 03.8 54.3 01.9 55.4
CFLOW-AD [14] 54.5 91.0 94.4 85.4 02.7 04.4 63.3 02.2 56.5
PyramidalFlow [22] 37.9 84.7 94.4 73.6 01.0 02.4 49.4 01.2 49.1
Rec. RD [10] 56.5 92.6 94.5 87.0 03.6 07.1 74.8 03.7 59.4
DiAD [16] 48.7 91.0 94.4 80.3 02.3 06.6 50.3 03.4 53.4
ViTAD [40] 58.0 92.5 94.4 88.3 04.0 08.4 76.0 04.4 60.2
InvAD [42] 75.7 96.6 94.4 89.7 08.6 17.6 81.5 09.6 66.3
InvAD-lite [42] 64.9 94.5 94.4 89.0 06.1 11.7 78.5 06.2 62.7
MambaAD [15] 60.2 93.5 94.4 87.5 04.9 09.8 74.4 05.1 60.7
Hybrid UniAD [35] 72.4 95.7 94.5 89.5 03.8 08.2 76.9 04.3 63.0
RD++ [31] 52.4 91.7 94.4 86.7 03.7 08.0 74.4 04.1 58.8
DesTSeg [44] 52.0 91.5 94.4 58.1 04.2 09.0 45.7 04.7 50.7
06Bird Aug. DRAEM [37] 44.4 88.1 94.4 44.1 01.0 02.0 23.7 01.0 42.5
SimpleNet [27] 63.6 93.5 94.4 90.3 05.5 09.1 71.1 04.8 61.1
Emb. CFA [21] 65.7 92.3 94.7 83.8 04.6 10.8 62.3 05.7 59.2
CFLOW-AD [14] 63.8 91.5 94.4 90.4 03.9 07.1 70.9 03.7 60.3
PyramidalFlow [22] 75.7 96.1 94.4 92.6 07.0 13.5 76.6 07.2 65.1
Rec. RD [10] 58.5 91.8 94.4 90.6 04.7 08.3 75.4 04.3 60.5
DiAD [16] 79.5 97.3 95.4 93.3 06.0 12.4 77.0 06.6 65.9
ViTAD [40] 64.4 93.3 94.5 91.2 04.8 09.8 75.9 05.1 62.0
InvAD [42] 83.3 97.6 94.7 91.9 06.7 12.5 78.5 06.7 66.5
InvAD-lite [42] 74.8 95.6 94.7 91.5 05.6 09.9 77.2 05.2 64.2
MambaAD [15] 64.5 93.7 94.5 91.2 05.0 09.1 76.4 04.7 62.1
Hybrid UniAD [35] 51.3 89.6 94.4 90.2 03.9 07.5 73.0 03.9 58.6
RD++ [31] 62.9 92.1 94.5 90.9 04.7 08.3 75.7 04.4 61.3
DesTSeg [44] 57.1 91.6 94.4 52.2 04.0 04.6 25.6 02.4 47.1
07Owl Aug. DRAEM [37] 59.2 88.7 92.8 47.6 00.7 01.8 27.1 00.9 45.4
SimpleNet [27] 67.0 92.0 93.4 88.8 03.5 07.5 66.8 03.9 59.9
Emb. CFA [21] 58.3 90.8 93.0 73.4 03.2 07.4 53.8 03.8 54.3
CFLOW-AD [14] 64.1 91.9 92.8 90.0 05.7 05.0 66.4 02.6 59.4
PyramidalFlow [22] 44.7 85.0 93.0 85.3 02.2 05.2 64.2 02.7 54.2
Rec. RD [10] 63.4 92.2 93.0 91.9 05.0 09.2 77.9 04.8 61.8
DiAD [16] 62.6 92.5 92.8 93.8 04.9 10.2 71.6 05.3 61.2
ViTAD [40] 66.8 92.3 92.8 92.5 05.0 10.2 77.6 05.4 62.5
InvAD [42] 71.0 93.8 93.4 94.0 10.3 19.4 82.5 10.8 66.3
InvAD-lite [42] 67.4 93.7 92.8 92.4 06.5 11.9 78.6 06.3 63.3
MambaAD [15] 67.8 93.6 92.8 91.2 04.6 08.2 76.0 04.3 62.0
Hybrid UniAD [35] 61.1 91.5 93.0 91.5 04.1 07.9 74.3 04.1 60.5
RD++ [31] 69.4 93.4 92.8 92.0 05.1 09.2 78.2 04.8 62.9
DesTSeg [44] 65.2 92.6 93.2 53.2 10.7 07.9 35.5 04.1 51.2
08Sabertooth Aug. DRAEM [37] 60.4 90.6 93.2 45.8 00.5 00.9 24.3 00.4 45.1
SimpleNet [27] 54.4 90.2 93.0 83.8 01.7 02.8 57.6 01.4 54.8
Emb. CFA [21] 55.6 89.9 93.0 76.3 02.5 05.1 54.6 02.6 53.9
CFLOW-AD [14] 61.5 90.9 93.2 86.7 01.8 03.4 58.7 01.7 56.6
PyramidalFlow [22] 65.2 93.3 93.0 77.6 01.1 02.4 49.2 01.2 54.5
Rec. RD [10] 67.4 93.2 93.2 87.9 02.1 03.6 67.2 01.8 59.2
DiAD [16] 89.2 98.4 93.8 85.1 01.5 03.0 48.5 01.5 59.9
ViTAD [40] 61.8 91.9 93.2 88.8 02.1 03.8 67.5 02.0 58.4
InvAD [42] 73.7 94.9 93.0 90.5 03.5 05.8 74.0 03.0 62.2
InvAD-lite [42] 69.3 93.7 93.2 88.3 02.8 04.0 68.0 02.1 59.9
MambaAD [15] 63.3 92.5 93.2 86.9 02.5 03.7 63.9 01.9 58.0
Hybrid UniAD [35] 55.5 89.1 93.0 88.2 02.1 03.7 71.2 01.9 57.5
RD++ [31] 64.0 92.2 93.0 87.7 02.2 03.7 66.6 01.9 58.5
DesTSeg [44] 61.1 92.1 93.0 51.6 04.5 04.5 31.0 02.3 48.3
Table A42: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
09Swan Aug. DRAEM [37] 41.6 83.8 93.1 44.7 00.6 01.0 23.4 00.5 41.2
SimpleNet [27] 52.6 89.0 93.1 83.7 02.0 04.8 58.1 02.4 54.8
Emb. CFA [21] 56.6 89.3 93.3 75.1 02.5 07.4 48.4 03.8 53.2
CFLOW-AD [14] 47.4 86.7 93.1 85.8 02.2 04.0 60.6 02.0 54.3
PyramidalFlow [22] 43.9 85.2 93.1 75.7 01.1 02.6 42.6 01.3 49.2
Rec. RD [10] 55.6 89.6 93.1 87.7 02.5 04.5 71.1 02.3 57.7
DiAD [16] 59.6 90.5 93.8 86.8 05.9 11.8 54.7 06.3 57.6
ViTAD [40] 52.8 88.1 93.2 90.0 03.0 06.0 73.1 03.1 58.0
InvAD [42] 68.9 93.7 93.1 91.6 05.9 11.5 80.1 06.1 63.5
InvAD-lite [42] 59.9 90.7 93.1 89.8 04.0 07.8 76.1 04.1 60.2
MambaAD [15] 54.7 88.7 93.1 87.2 02.4 04.3 69.0 02.2 57.1
Hybrid UniAD [35] 52.5 87.5 93.1 88.4 02.7 05.0 71.3 02.5 57.2
RD++ [31] 53.9 88.9 93.1 87.6 02.4 04.5 70.6 02.3 57.3
DesTSeg [44] 54.7 89.3 93.1 50.9 02.5 02.3 29.0 01.2 46.0
10Sheep Aug. DRAEM [37] 44.9 86.4 92.7 44.1 00.8 01.7 27.5 00.9 42.6
SimpleNet [27] 60.4 91.3 92.7 90.5 03.8 06.8 76.8 03.5 60.3
Emb. CFA [21] 65.9 92.2 93.1 85.1 03.9 08.6 67.7 04.5 59.5
CFLOW-AD [14] 61.4 89.8 92.9 92.0 03.8 06.1 79.3 03.1 60.8
PyramidalFlow [22] 54.4 89.6 92.7 90.6 02.3 05.0 71.8 02.6 58.1
Rec. RD [10] 64.4 92.6 92.7 92.6 05.3 09.2 84.5 04.8 63.0
DiAD [16] 85.0 96.8 94.1 93.3 04.1 08.5 76.0 04.4 65.4
ViTAD [40] 64.8 92.0 92.7 93.4 04.8 09.3 84.8 04.9 63.1
InvAD [42] 84.6 97.1 93.6 94.0 08.2 13.6 87.1 07.3 68.3
InvAD-lite [42] 78.7 95.9 93.4 93.5 06.7 10.1 85.2 05.3 66.2
MambaAD [15] 70.1 93.8 92.9 92.8 05.5 09.4 84.0 04.9 64.1
Hybrid UniAD [35] 57.4 90.3 92.7 93.3 04.8 08.2 83.3 04.3 61.4
RD++ [31] 71.4 94.2 92.9 92.7 06.2 10.6 84.3 05.6 64.6
DesTSeg [44] 67.5 93.7 92.7 53.1 08.5 08.1 39.0 04.2 51.8
11Pig Aug. DRAEM [37] 51.2 90.2 94.0 54.8 01.0 01.6 28.9 00.8 46.0
SimpleNet [27] 54.8 90.9 94.0 94.0 06.4 12.8 75.7 06.8 61.2
Emb. CFA [21] 54.7 90.6 94.0 87.4 07.5 14.9 61.9 08.1 58.7
CFLOW-AD [14] 51.4 89.3 94.0 92.3 03.5 05.5 74.5 02.8 58.6
PyramidalFlow [22] 56.4 91.0 94.0 89.8 02.7 05.5 66.9 02.8 58.0
Rec. RD [10] 58.2 91.8 94.0 94.3 06.5 11.8 83.5 06.3 62.9
DiAD [16] 44.0 90.1 94.0 93.7 09.4 16.8 70.9 09.2 59.9
ViTAD [40] 58.4 91.1 94.0 95.9 08.8 17.0 83.3 09.3 64.1
InvAD [42] 72.3 95.1 94.4 96.9 18.9 27.6 87.9 16.0 70.4
InvAD-lite [42] 71.8 95.1 94.2 96.3 16.4 24.4 86.4 13.9 69.2
MambaAD [15] 67.4 94.2 94.2 95.2 10.6 18.5 84.5 10.2 66.4
Hybrid UniAD [35] 46.7 86.8 94.0 93.6 03.8 07.3 78.8 03.8 58.7
RD++ [31] 58.7 92.1 94.0 94.7 07.2 14.1 83.9 07.6 63.5
DesTSeg [44] 63.0 92.9 94.0 57.5 07.1 11.7 37.9 06.2 52.0
12Zalika Aug. DRAEM [37] 43.6 84.1 92.9 43.4 01.1 02.3 25.0 01.2 41.8
SimpleNet [27] 57.1 88.2 93.1 84.2 04.0 08.4 59.5 04.4 56.4
Emb. CFA [21] 54.8 90.0 92.9 73.9 04.2 09.2 53.5 04.8 54.1
CFLOW-AD [14] 51.4 88.0 92.9 86.0 04.4 06.9 55.3 03.6 55.0
PyramidalFlow [22] 43.8 85.3 92.9 79.1 02.2 04.1 57.5 02.1 52.1
Rec. RD [10] 59.3 92.0 92.9 87.2 04.9 08.6 69.9 04.5 59.3
DiAD [16] 46.6 86.6 92.9 85.2 04.4 06.9 42.0 03.6 52.1
ViTAD [40] 54.9 91.0 92.9 87.3 04.7 08.7 67.5 04.6 58.1
InvAD [42] 64.0 92.3 92.9 90.2 10.0 18.1 76.5 10.0 63.4
InvAD-lite [42] 60.9 92.1 92.9 88.6 07.0 11.5 72.3 06.1 60.8
MambaAD [15] 58.3 91.6 92.9 87.2 05.2 08.7 68.9 04.5 59.0
Hybrid UniAD [35] 52.1 88.9 92.9 85.9 03.4 06.0 64.8 03.1 56.3
RD++ [31] 55.5 90.5 92.9 87.0 05.1 08.2 69.3 04.3 58.4
DesTSeg [44] 52.3 88.9 92.9 52.2 04.0 05.7 38.5 02.9 47.8
Table A43: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
13Pheonix Aug. DRAEM [37] 51.9 89.6 94.3 45.4 00.6 01.0 22.8 00.5 43.7
SimpleNet [27] 55.7 90.2 94.4 80.7 01.5 03.3 56.9 01.7 54.7
Emb. CFA [21] 54.1 90.4 94.3 77.4 02.0 05.5 58.0 02.8 54.5
CFLOW-AD [14] 52.1 89.7 94.3 82.4 01.4 02.7 57.2 01.4 54.3
PyramidalFlow [22] 48.0 87.6 94.3 72.8 00.9 02.1 54.1 01.0 51.4
Rec. RD [10] 53.5 90.5 94.3 83.6 02.5 03.5 70.0 01.8 56.8
DiAD [16] 41.5 86.8 94.3 83.2 01.4 03.0 49.1 01.5 51.3
ViTAD [40] 52.6 89.9 94.3 84.8 01.8 03.5 68.4 01.8 56.5
InvAD [42] 62.7 92.9 94.6 87.0 03.6 07.2 77.5 03.7 60.8
InvAD-lite [42] 56.0 91.6 94.3 85.3 03.2 05.3 74.0 02.7 58.5
MambaAD [15] 55.7 91.7 94.4 83.7 02.8 03.7 70.7 01.9 57.5
Hybrid UniAD [35] 55.1 90.2 94.3 83.8 01.5 03.1 64.5 01.6 56.1
RD++ [31] 54.7 90.6 94.4 84.1 01.9 03.6 71.0 01.8 57.2
DesTSeg [44] 53.3 91.0 94.3 52.5 02.2 04.2 35.5 02.1 47.6
14Elephant Aug. DRAEM [37] 50.4 89.5 93.9 52.4 01.1 01.9 25.9 00.9 45.0
SimpleNet [27] 58.7 91.9 93.9 64.9 02.2 04.9 54.8 02.5 53.0
Emb. CFA [21] 54.2 90.9 93.9 61.8 01.9 03.8 45.8 01.9 50.3
CFLOW-AD [14] 55.7 91.4 93.9 68.2 02.2 03.8 49.7 01.9 52.1
PyramidalFlow [22] 44.5 87.3 93.9 63.7 01.4 03.1 45.2 01.6 48.4
Rec. RD [10] 58.3 91.7 93.9 69.9 02.9 06.6 64.6 03.4 55.4
DiAD [16] 68.7 95.1 93.9 74.2 01.9 03.9 44.7 02.0 54.6
ViTAD [40] 57.4 91.9 93.9 70.8 03.1 07.4 64.0 03.8 55.5
InvAD [42] 63.1 93.3 93.9 72.3 04.4 10.8 68.4 05.7 58.0
InvAD-lite [42] 66.2 94.2 93.9 70.5 03.6 08.2 66.0 04.3 57.5
MambaAD [15] 66.2 94.4 93.9 69.6 03.1 07.2 63.5 03.7 56.8
Hybrid UniAD [35] 55.7 90.5 93.9 68.5 01.9 04.1 59.2 02.1 53.4
RD++ [31] 60.7 92.5 93.9 70.2 03.0 06.9 63.5 03.6 55.8
DesTSeg [44] 60.0 91.9 93.9 53.5 05.3 08.1 46.9 04.2 51.4
15Parrot Aug. DRAEM [37] 52.8 86.4 91.8 48.6 00.8 01.4 24.1 00.7 43.7
SimpleNet [27] 62.2 90.3 91.8 80.3 02.4 04.8 53.2 02.4 55.0
Emb. CFA [21] 55.6 88.6 91.8 69.4 02.2 05.0 44.5 02.5 51.0
CFLOW-AD [14] 52.9 86.3 92.0 83.5 02.9 05.2 60.5 02.7 54.8
PyramidalFlow [22] 62.3 90.4 92.0 81.5 03.0 06.6 66.1 03.4 57.4
Rec. RD [10] 54.6 87.9 92.0 83.9 03.2 06.7 67.2 03.5 56.5
DiAD [16] 52.4 88.4 91.8 83.6 02.5 05.6 63.4 02.9 55.4
ViTAD [40] 50.5 87.3 91.8 83.7 03.2 07.1 66.6 03.7 55.7
InvAD [42] 60.8 90.9 92.0 86.3 06.5 13.4 73.9 07.2 60.5
InvAD-lite [42] 57.0 89.6 91.8 85.1 04.2 08.4 69.1 04.4 57.9
MambaAD [15] 55.1 88.9 91.8 84.1 03.5 06.5 65.9 03.4 56.5
Hybrid UniAD [35] 42.9 83.7 91.8 84.4 02.9 05.4 65.8 02.8 53.8
RD++ [31] 49.2 86.2 92.0 84.1 03.2 06.6 66.7 03.4 55.4
DesTSeg [44] 52.4 88.7 91.8 52.0 09.9 05.6 24.8 02.9 46.5
16Cat Aug. DRAEM [37] 50.5 89.9 94.3 43.0 00.6 01.5 24.4 00.7 43.5
SimpleNet [27] 58.3 91.8 94.3 92.1 02.5 05.7 79.3 02.9 60.6
Emb. CFA [21] 54.6 90.9 94.4 82.6 02.7 08.1 59.8 04.2 56.2
CFLOW-AD [14] 52.0 90.1 94.3 93.1 02.3 04.8 80.5 02.5 59.6
PyramidalFlow [22] 49.2 88.6 94.3 91.1 01.7 03.8 70.2 01.9 57.0
Rec. RD [10] 52.4 90.1 94.3 93.6 03.1 06.3 84.8 03.3 60.7
DiAD [16] 45.8 88.0 94.3 93.7 03.0 05.3 81.8 02.7 58.8
ViTAD [40] 57.2 92.0 94.3 93.8 03.2 06.3 84.6 03.3 61.6
InvAD [42] 58.9 92.1 94.3 94.5 05.3 11.0 86.9 05.8 63.3
InvAD-lite [42] 58.5 92.5 94.3 94.2 03.7 07.4 85.6 03.9 62.3
MambaAD [15] 54.5 91.0 94.4 94.0 03.5 07.0 84.9 03.6 61.3
Hybrid UniAD [35] 51.7 89.5 94.3 93.8 02.7 05.6 82.4 02.9 60.0
RD++ [31] 53.9 90.8 94.3 93.4 03.4 06.3 84.5 03.3 60.9
DesTSeg [44] 56.2 90.5 94.3 52.8 01.4 03.1 30.6 01.6 47.0
Table A44: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
17Scorpion Aug. DRAEM [37] 48.4 86.4 92.9 51.5 00.7 01.4 25.6 00.7 43.8
SimpleNet [27] 72.7 94.5 93.1 88.4 04.0 06.5 67.9 03.4 61.0
Emb. CFA [21] 68.3 92.4 93.0 78.9 04.1 07.0 55.9 03.6 57.1
CFLOW-AD [14] 59.7 91.0 92.9 89.6 04.6 04.9 68.1 02.5 58.7
PyramidalFlow [22] 57.6 87.3 93.1 83.4 01.9 05.1 53.1 02.6 54.5
Rec. RD [10] 60.1 91.2 92.9 91.3 04.8 07.4 82.3 03.8 61.4
DiAD [16] 84.8 96.9 93.8 87.0 01.9 03.5 56.5 01.8 60.6
ViTAD [40] 60.1 90.5 93.1 91.2 03.3 07.7 77.6 04.0 60.5
InvAD [42] 81.5 96.7 93.3 93.6 07.2 10.5 86.1 05.6 67.0
InvAD-lite [42] 73.5 94.7 92.9 93.3 06.1 08.6 84.8 04.5 64.8
MambaAD [15] 73.0 94.6 93.0 93.0 05.5 08.2 83.6 04.3 64.4
Hybrid UniAD [35] 63.5 90.9 92.9 91.7 04.4 07.2 79.1 03.8 61.4
RD++ [31] 64.1 92.3 92.9 91.7 05.4 07.9 82.4 04.1 62.4
DesTSeg [44] 77.8 95.9 93.1 60.6 05.4 09.7 44.3 05.1 55.3
18Obesobeso Aug. DRAEM [37] 43.9 87.1 94.1 47.5 01.0 01.8 26.9 00.9 43.2
SimpleNet [27] 63.3 93.6 94.1 92.4 15.8 21.7 81.6 12.2 66.1
Emb. CFA [21] 62.9 93.2 94.1 81.8 12.6 20.0 64.6 11.1 61.3
CFLOW-AD [14] 59.6 91.6 94.1 92.5 05.8 08.8 79.4 04.6 61.7
PyramidalFlow [22] 61.7 93.2 94.1 91.9 04.5 08.8 74.7 04.6 61.3
Rec. RD [10] 64.6 94.0 94.1 93.5 11.6 20.7 87.8 11.5 66.6
DiAD [16] 87.5 98.3 94.4 95.1 10.2 17.7 82.9 09.7 69.4
ViTAD [40] 67.5 94.4 94.1 94.3 18.2 24.7 88.3 14.1 68.8
InvAD [42] 83.5 97.6 94.2 94.8 16.6 26.3 90.8 15.1 72.0
InvAD-lite [42] 70.3 95.3 94.2 93.7 12.9 21.2 88.1 11.8 68.0
MambaAD [15] 69.6 95.0 94.2 93.0 10.9 20.2 86.4 11.3 67.0
Hybrid UniAD [35] 64.9 93.4 94.1 92.2 06.9 12.9 82.4 06.9 63.8
RD++ [31] 67.7 94.0 94.4 93.2 09.1 17.4 86.5 09.5 66.0
DesTSeg [44] 63.1 93.1 94.2 61.1 16.2 22.4 45.0 12.6 56.4
19Bear Aug. DRAEM [37] 54.1 89.3 93.3 49.3 00.8 01.3 23.0 00.7 44.4
SimpleNet [27] 63.0 92.8 93.5 88.4 07.0 12.8 68.6 06.8 60.9
Emb. CFA [21] 63.0 92.4 93.5 80.7 05.5 12.1 57.5 06.4 57.8
CFLOW-AD [14] 60.9 91.3 93.3 88.9 03.7 06.4 64.6 03.3 58.4
PyramidalFlow [22] 52.6 89.6 93.6 87.2 03.6 06.3 61.0 03.3 56.3
Rec. RD [10] 62.9 92.2 93.6 90.8 06.2 10.8 76.8 05.7 61.9
DiAD [16] 59.1 92.9 93.3 89.4 04.2 10.5 62.6 05.5 58.8
ViTAD [40] 60.9 91.2 93.5 91.1 05.5 11.5 75.1 06.1 61.3
InvAD [42] 73.9 95.4 93.3 92.5 13.8 21.7 82.0 12.2 67.5
InvAD-lite [42] 67.7 93.2 93.3 92.2 11.3 18.1 78.9 10.0 65.0
MambaAD [15] 64.2 92.9 93.5 91.2 07.9 13.4 76.1 07.2 62.7
Hybrid UniAD [35] 56.8 90.6 93.3 90.0 04.2 07.9 77.2 04.1 60.0
RD++ [31] 65.3 93.0 93.5 90.8 07.7 11.8 76.9 06.3 62.7
DesTSeg [44] 66.8 93.5 93.7 58.5 09.4 13.7 43.7 07.3 54.2
20Puppy Aug. DRAEM [37] 44.7 84.1 92.7 47.9 00.5 00.9 20.7 00.4 41.6
SimpleNet [27] 54.6 89.8 92.9 83.1 02.6 04.9 54.4 02.5 54.6
Emb. CFA [21] 56.3 90.2 92.9 71.2 03.0 06.9 51.5 03.6 53.1
CFLOW-AD [14] 57.5 90.1 92.7 85.2 01.9 04.5 57.1 02.3 55.6
PyramidalFlow [22] 56.0 89.4 92.9 81.1 01.5 03.2 53.1 01.6 53.9
Rec. RD [10] 54.0 89.6 93.0 86.8 03.6 06.3 66.5 03.3 57.1
DiAD [16] 48.7 88.8 92.7 83.4 01.4 02.7 51.9 01.4 52.8
ViTAD [40] 52.0 87.4 93.0 87.6 02.4 04.9 66.3 02.5 56.2
InvAD [42] 62.9 92.5 92.9 89.3 05.9 11.0 73.3 05.8 61.1
InvAD-lite [42] 54.0 90.0 93.0 87.5 03.5 06.2 68.2 03.2 57.5
MambaAD [15] 49.2 87.7 92.7 86.6 02.6 05.3 64.8 02.7 55.6
Hybrid UniAD [35] 55.6 89.8 92.7 86.5 03.4 07.9 68.4 04.1 57.8
RD++ [31] 59.0 91.5 92.9 87.1 04.1 08.1 66.8 04.2 58.5
DesTSeg [44] 64.6 93.0 92.9 53.7 13.3 09.5 32.9 05.0 51.4
Appendix L Detailed Quantitative Results on Uni-Medical Dataset Under 100 epochs
Table A45: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
brain Aug. DRAEM [37] 47.1 82.8 90.6 33.8 02.2 06.1 07.9 03.1 38.6
SimpleNet [27] 84.8 96.3 91.0 94.5 43.4 43.9 71.4 28.1 75.0
Emb. CFA [21] 82.5 95.9 90.6 75.6 33.0 39.0 36.5 24.2 64.7
CFLOW-AD [14] 85.7 96.2 91.9 94.5 56.6 17.2 73.8 09.4 73.7
PyramidalFlow [22] 62.4 86.1 90.9 85.7 10.4 18.0 50.5 09.9 57.7
Rec. RD [10] 83.9 94.6 91.9 96.5 45.6 48.9 82.8 32.4 77.7
DiAD [16] 93.2 97.7 95.2 94.9 35.2 39.8 79.5 24.8 76.5
ViTAD [40] 89.4 97.3 93.0 97.7 61.8 58.8 84.0 41.7 83.1
InvAD [42] 90.8 97.7 93.2 97.7 59.1 58.3 86.2 41.2 83.3
InvAD-lite [42] 84.9 95.9 91.8 96.5 46.2 49.2 82.5 32.6 78.1
MambaAD [15] 93.7 98.5 94.2 97.9 60.8 61.5 87.5 44.4 84.9
Hybrid UniAD [35] 90.2 97.7 92.6 97.4 56.2 55.6 82.9 38.5 81.8
RD++ [31] 84.8 95.5 91.8 96.9 48.5 51.5 84.2 34.6 79.0
DesTSeg [44] 87.1 95.9 92.2 70.2 52.2 42.7 33.9 27.2 67.7
liver Aug. DRAEM [37] 53.1 49.7 61.6 20.0 00.4 01.5 10.2 00.7 28.1
SimpleNet [27] 60.5 50.6 62.2 94.2 06.0 11.6 81.5 06.2 52.4
Emb. CFA [21] 52.9 48.1 61.3 88.0 04.7 11.1 60.5 05.9 46.7
CFLOW-AD [14] 63.9 56.1 63.0 95.9 02.9 05.6 87.4 02.9 53.5
PyramidalFlow [22] 47.2 43.2 60.9 75.0 03.3 11.5 09.0 06.1 35.7
Rec. RD [10] 55.5 45.8 64.4 96.6 05.5 10.4 91.1 05.5 52.8
DiAD [16] 56.5 51.7 62.0 96.4 05.5 10.6 89.6 05.6 53.2
ViTAD [40] 63.9 54.9 65.1 97.9 13.2 21.4 92.6 12.0 58.4
InvAD [42] 64.2 52.8 65.9 97.0 07.8 14.5 92.9 07.8 56.4
InvAD-lite [42] 57.0 46.6 63.8 96.4 05.3 09.8 90.6 05.1 52.8
MambaAD [15] 64.2 54.6 64.5 97.0 10.0 17.4 92.6 09.5 57.2
Hybrid UniAD [35] 60.3 48.3 63.1 97.1 07.9 14.0 92.7 07.5 54.8
RD++ [31] 55.6 44.8 64.4 96.6 05.5 10.3 91.5 05.4 52.7
DesTSeg [44] 58.5 47.8 61.4 57.6 13.8 15.6 31.4 08.5 40.9
retinal Aug. DRAEM [37] 54.7 49.4 59.7 47.8 03.5 09.6 13.3 05.1 34.0
SimpleNet [27] 87.0 86.2 76.8 94.2 53.9 52.5 78.2 35.6 75.5
Emb. CFA [21] 77.7 75.8 67.4 71.4 20.7 24.5 37.3 14.0 53.5
CFLOW-AD [14] 88.0 86.5 76.8 95.8 61.6 57.9 84.1 40.7 78.7
PyramidalFlow [22] 52.3 42.7 59.7 75.9 11.4 21.8 43.2 12.2 43.9
Rec. RD [10] 89.0 86.8 78.2 96.4 65.1 60.1 86.5 43.0 80.3
DiAD [16] 86.6 82.0 76.0 96.1 61.8 56.2 83.8 39.0 77.5
ViTAD [40] 92.1 90.0 81.9 95.6 69.8 64.2 83.5 47.3 82.4
InvAD [42] 92.2 91.0 82.4 97.1 71.9 65.5 88.7 48.7 84.1
InvAD-lite [42] 91.7 89.2 81.8 95.7 65.3 59.0 84.4 41.9 81.0
MambaAD [15] 93.8 89.1 86.9 95.5 66.4 63.5 84.6 46.6 82.8
Hybrid UniAD [35] 86.3 82.4 75.8 95.3 53.8 53.7 82.5 36.7 75.7
RD++ [31] 91.2 89.2 81.3 96.7 70.4 64.9 87.6 48.1 83.0
DesTSeg [44] 90.0 87.4 81.0 69.2 59.0 43.7 40.4 27.9 67.2
Avg Aug. DRAEM [37] 51.6 60.6 70.6 33.8 02.0 05.7 10.5 03.0 33.5
SimpleNet [27] 77.5 77.7 76.7 94.3 34.4 36.0 77.0 23.3 67.7
Emb. CFA [21] 71.0 73.3 73.1 78.3 19.5 24.9 44.7 14.7 55.0
CFLOW-AD [14] 79.2 79.6 77.2 95.4 40.4 26.9 81.8 17.7 68.6
PyramidalFlow [22] 53.9 57.3 70.5 78.8 08.4 17.1 34.2 09.4 45.7
Rec. RD [10] 76.1 75.7 78.2 96.5 38.8 39.8 86.8 26.9 70.3
DiAD [16] 78.8 77.2 77.7 95.8 34.2 35.5 84.3 23.2 69.1
ViTAD [40] 81.8 80.7 80.0 97.1 48.3 48.2 86.7 33.7 74.7
InvAD [42] 82.4 80.5 80.5 97.3 46.2 46.1 89.3 32.6 74.6
InvAD-lite [42] 77.9 77.3 79.1 96.2 39.0 39.3 85.8 26.5 70.7
MambaAD [15] 83.9 80.8 81.9 96.8 45.8 47.5 88.2 33.5 75.0
Hybrid UniAD [35] 79.0 76.1 77.1 96.6 39.3 41.1 86.0 27.6 70.7
RD++ [31] 77.2 76.5 79.2 96.7 41.5 42.2 87.8 29.4 71.6
DesTSeg [44] 78.5 77.0 78.2 65.7 41.7 34.0 35.3 21.2 58.6
Appendix M Detailed Quantitative Results on Real-IAD Dataset Under 100 epochs
Table A46: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
audiojack Aug. DRAEM [37] 53.0 36.9 51.4 42.8 00.1 00.3 13.7 00.2 28.3
SimpleNet [27] 51.4 36.3 50.9 66.0 00.2 00.9 37.6 00.5 34.8
Emb. CFA [21] 57.3 41.9 52.1 86.5 00.5 01.6 57.1 00.8 42.4
CFLOW-AD [14] 77.4 70.8 60.9 95.4 16.4 15.4 73.2 08.3 58.5
PyramidalFlow [22] 51.7 34.9 50.7 81.7 00.2 00.4 40.7 00.2 37.2
Rec. RD [10] 76.0 62.8 60.6 96.7 13.1 22.1 80.6 12.4 58.8
DiAD [16] 76.5 54.3 65.7 91.6 01.0 03.9 63.3 02.0 50.9
ViTAD [40] 80.6 72.8 63.5 96.8 21.3 31.5 79.5 18.7 63.7
InvAD [42] 85.2 80.7 69.3 98.5 33.9 43.8 89.6 28.0 71.6
InvAD-lite [42] 84.3 78.7 67.8 98.1 32.4 40.8 87.0 25.6 69.9
MambaAD [15] 84.9 77.7 68.2 98.0 22.2 30.4 86.2 17.9 66.8
Hybrid UniAD [35] 81.7 76.2 65.4 97.6 20.7 31.1 84.1 18.4 65.3
RD++ [31] 76.0 60.0 61.5 96.6 10.6 20.1 80.4 11.2 57.9
DesTSeg [44] 81.6 74.7 66.5 88.3 44.9 50.7 64.5 34.0 67.3
bottle_cap Aug. DRAEM [37] 46.9 40.9 60.2 35.2 00.0 00.1 09.7 00.0 27.6
SimpleNet [27] 47.4 41.0 60.2 80.1 00.5 03.7 35.8 01.9 38.4
Emb. CFA [21] 55.7 49.7 60.2 85.2 00.5 02.4 57.0 01.2 44.4
CFLOW-AD [14] 80.0 78.3 71.7 98.8 13.6 22.6 91.9 12.7 65.3
PyramidalFlow [22] 55.4 46.7 60.3 93.2 00.7 02.6 72.0 01.3 47.3
Rec. RD [10] 89.6 86.3 81.2 99.4 18.6 28.6 96.3 16.7 71.4
DiAD [16] 91.6 94.0 87.9 94.6 04.9 11.4 73.0 06.0 65.3
ViTAD [40] 82.2 78.8 74.8 98.7 16.7 27.5 91.5 16.0 67.2
InvAD [42] 95.5 94.9 85.9 99.6 24.9 33.9 97.2 20.4 76.0
InvAD-lite [42] 92.6 91.2 81.9 99.6 21.2 28.5 97.5 16.6 73.2
MambaAD [15] 93.2 92.6 82.7 99.7 29.9 33.8 97.4 20.3 75.6
Hybrid UniAD [35] 91.6 90.6 80.4 99.4 21.0 30.5 95.8 18.0 72.8
RD++ [31] 89.7 86.7 81.5 99.3 18.9 30.4 95.6 17.9 71.7
DesTSeg [44] 75.5 74.5 65.9 71.1 27.3 31.7 43.2 18.9 55.6
button_battery Aug. DRAEM [37] 34.4 46.9 72.4 51.8 00.5 00.9 17.8 00.4 32.1
SimpleNet [27] 60.4 68.6 72.4 78.0 09.4 15.4 38.3 08.4 48.9
Emb. CFA [21] 53.2 57.9 72.5 84.3 01.6 03.2 54.8 01.6 46.8
CFLOW-AD [14] 66.0 75.1 72.7 95.6 29.2 25.8 74.8 14.8 62.7
PyramidalFlow [22] 52.5 56.4 72.5 50.2 17.5 00.8 15.0 00.4 37.8
Rec. RD [10] 74.7 79.7 75.9 97.7 34.5 39.4 86.4 24.5 69.8
DiAD [16] 80.5 71.3 70.6 84.1 01.4 05.3 66.9 02.7 54.3
ViTAD [40] 77.0 82.2 77.6 98.3 49.1 49.6 82.7 32.9 73.8
InvAD [42] 90.4 92.5 84.6 99.1 57.2 56.1 92.6 39.0 81.8
InvAD-lite [42] 82.0 86.9 78.4 98.5 51.1 52.4 89.0 35.5 76.9
MambaAD [15] 82.8 87.4 79.2 98.3 48.6 49.6 88.1 33.0 76.3
Hybrid UniAD [35] 72.8 79.8 74.6 96.5 29.5 33.7 78.7 20.3 66.5
RD++ [31] 76.5 82.2 75.8 97.9 38.4 42.7 87.1 27.2 71.5
DesTSeg [44] 84.6 87.8 81.2 92.2 55.7 53.7 60.4 36.7 73.7
end_cap Aug. DRAEM [37] 48.1 57.5 72.9 54.9 00.1 00.3 21.0 00.1 36.4
SimpleNet [27] 52.4 59.6 72.9 65.5 00.3 01.8 39.0 00.9 41.6
Emb. CFA [21] 55.3 62.5 73.3 72.8 00.4 02.7 40.0 01.4 43.9
CFLOW-AD [14] 63.7 72.7 72.9 87.6 03.5 05.9 60.3 03.0 52.4
PyramidalFlow [22] 55.7 63.8 73.0 77.2 00.4 01.9 33.2 01.0 43.6
Rec. RD [10] 80.0 84.0 78.2 96.5 12.3 21.9 89.4 12.3 66.0
DiAD [16] 85.1 83.4 84.8 81.3 02.0 06.9 38.2 03.6 54.5
ViTAD [40] 72.6 77.1 75.2 93.8 05.7 12.2 75.9 06.5 58.9
InvAD [42] 84.1 87.7 80.3 98.0 14.7 22.7 93.5 12.8 68.7
InvAD-lite [42] 79.8 84.2 78.3 97.4 13.5 21.0 91.5 11.8 66.5
MambaAD [15] 78.6 83.1 77.4 97.2 12.2 19.5 90.3 10.8 65.5
Hybrid UniAD [35] 80.9 86.3 77.8 95.5 09.5 17.7 86.4 09.7 64.9
RD++ [31] 79.7 84.3 77.8 96.8 12.6 22.6 90.4 12.7 66.3
DesTSeg [44] 72.7 76.8 74.9 75.3 10.3 18.7 53.2 10.3 54.6
Table A47: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
eraser Aug. DRAEM [37] 53.8 42.8 55.9 42.6 00.1 00.2 11.2 00.1 29.5
SimpleNet [27] 53.4 46.3 55.8 87.6 03.8 08.7 59.8 04.6 45.1
Emb. CFA [21] 51.2 43.0 55.9 78.5 01.8 05.8 57.6 03.0 42.0
CFLOW-AD [14] 88.6 87.5 77.6 98.9 23.2 26.9 93.4 15.6 70.9
PyramidalFlow [22] 57.8 42.5 57.3 88.8 00.7 02.6 60.3 01.3 44.3
Rec. RD [10] 90.0 88.5 79.4 99.5 30.6 35.9 96.3 21.9 74.3
DiAD [16] 80.0 80.0 77.3 91.1 07.7 15.4 67.5 08.4 59.9
ViTAD [40] 86.0 85.1 74.5 98.5 24.8 32.4 89.3 19.3 70.1
InvAD [42] 92.1 90.3 81.7 99.4 30.4 36.1 95.1 22.0 75.0
InvAD-lite [42] 88.7 87.2 77.4 99.3 28.9 33.6 94.8 20.2 72.8
MambaAD [15] 88.4 86.9 76.7 99.2 27.4 35.7 93.8 21.7 72.6
Hybrid UniAD [35] 90.2 88.8 80.2 99.3 25.4 32.0 94.4 19.1 72.9
RD++ [31] 90.6 89.4 79.7 99.5 29.2 35.9 96.8 21.9 74.4
DesTSeg [44] 80.5 79.5 68.4 83.3 53.8 53.0 56.0 36.1 67.8
fire_hood Aug. DRAEM [37] 56.3 39.5 53.4 42.7 00.1 00.2 09.9 00.1 28.9
SimpleNet [27] 51.3 39.9 52.9 76.8 01.1 04.8 39.9 02.5 38.1
Emb. CFA [21] 56.2 40.6 52.9 79.6 00.4 01.4 42.0 00.7 39.0
CFLOW-AD [14] 80.3 73.8 68.2 98.1 21.6 22.7 86.0 12.8 64.4
PyramidalFlow [22] 56.5 39.1 54.1 50.0 00.2 00.1 15.4 00.1 30.8
Rec. RD [10] 78.4 70.8 64.5 98.9 28.3 34.6 88.4 20.9 66.3
DiAD [16] 83.3 81.7 80.5 91.8 03.2 09.2 66.7 04.8 59.5
ViTAD [40] 75.6 67.8 60.6 98.0 18.0 27.8 82.5 16.1 61.5
InvAD [42] 82.1 75.1 68.3 99.0 28.2 35.3 89.4 21.5 68.2
InvAD-lite [42] 80.0 73.6 65.6 98.8 24.4 31.6 87.5 18.7 65.9
MambaAD [15] 79.9 73.5 65.4 98.7 25.9 33.0 86.8 19.7 66.2
Hybrid UniAD [35] 80.3 74.5 66.8 98.6 25.6 33.5 84.8 20.1 66.3
RD++ [31] 80.5 73.7 66.7 98.9 29.2 36.6 89.4 22.4 67.9
DesTSeg [44] 75.6 65.6 61.5 78.4 26.7 34.1 45.9 20.6 55.4
mint Aug. DRAEM [37] 51.8 46.9 64.0 45.3 00.1 00.1 13.9 00.1 31.7
SimpleNet [27] 55.7 51.9 63.8 73.1 00.5 03.0 26.8 01.5 39.3
Emb. CFA [21] 51.2 49.0 63.8 77.1 00.1 00.3 39.2 00.2 40.1
CFLOW-AD [14] 63.6 64.1 63.7 92.8 08.7 12.8 61.4 06.8 52.4
PyramidalFlow [22] 57.0 51.2 63.8 50.0 00.0 00.1 15.0 00.1 33.9
Rec. RD [10] 66.3 64.6 65.1 94.9 12.8 23.8 73.5 13.5 57.3
DiAD [16] 76.7 76.7 76.0 91.1 05.7 11.6 64.2 06.2 57.4
ViTAD [40] 69.2 70.1 64.9 94.7 13.9 26.6 70.0 15.3 58.5
InvAD [42] 76.8 77.1 68.9 98.2 18.5 29.1 82.8 17.0 64.5
InvAD-lite [42] 72.0 73.1 66.2 97.3 17.9 28.6 80.0 16.7 62.2
MambaAD [15] 72.6 73.8 66.2 97.1 16.7 27.2 76.9 15.8 61.5
Hybrid UniAD [35] 67.6 68.6 64.4 94.6 09.3 22.2 63.2 12.5 55.7
RD++ [31] 67.6 66.6 65.2 96.2 13.7 25.2 76.9 14.4 58.8
DesTSeg [44] 62.8 60.6 64.2 66.5 21.1 27.5 27.5 15.9 47.2
mounts Aug. DRAEM [37] 49.8 34.4 52.1 37.5 00.1 00.2 10.3 00.1 26.3
SimpleNet [27] 49.9 39.2 52.1 87.2 01.4 04.2 63.0 02.2 42.4
Emb. CFA [21] 62.8 49.5 53.9 89.9 03.3 09.3 68.6 04.9 48.2
CFLOW-AD [14] 82.9 74.5 70.3 96.1 16.9 25.5 86.4 14.6 64.7
PyramidalFlow [22] 56.1 41.1 52.1 82.7 01.0 04.0 54.3 02.0 41.6
Rec. RD [10] 88.6 79.8 75.0 99.3 30.7 37.5 95.1 23.0 72.3
DiAD [16] 75.3 74.5 82.5 84.3 00.4 01.1 48.8 00.5 52.4
ViTAD [40] 84.9 74.8 71.4 98.8 25.1 32.2 91.7 19.2 68.4
InvAD [42] 88.9 79.3 77.2 99.4 31.2 35.7 94.2 21.7 72.3
InvAD-lite [42] 87.4 76.6 75.7 99.3 27.4 32.7 94.2 19.5 70.5
MambaAD [15] 87.4 78.3 74.2 99.1 30.9 35.2 92.6 21.3 71.1
Hybrid UniAD [35] 87.2 75.6 76.2 99.4 28.1 33.6 95.3 20.2 70.8
RD++ [31] 88.5 80.6 74.5 99.3 28.8 36.2 95.3 22.1 71.9
DesTSeg [44] 72.8 53.9 62.6 79.6 27.5 37.6 58.0 23.1 56.0
Table A48: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pcb Aug. DRAEM [37] 56.8 64.0 76.2 45.3 00.1 00.3 13.9 00.2 36.7
SimpleNet [27] 54.2 63.4 75.5 76.8 00.4 01.0 45.8 00.5 45.3
Emb. CFA [21] 51.5 62.4 75.5 88.6 01.2 03.5 58.6 01.8 48.8
CFLOW-AD [14] 74.3 83.2 76.6 95.1 15.0 23.8 75.6 13.5 63.4
PyramidalFlow [22] 54.8 64.5 75.7 81.6 00.3 00.7 45.0 00.4 46.1
Rec. RD [10] 79.2 85.4 79.7 97.5 16.1 24.3 88.6 13.8 67.3
DiAD [16] 86.0 85.1 85.4 92.0 03.7 07.4 66.5 03.8 60.9
ViTAD [40] 83.3 89.2 81.2 98.2 25.7 33.6 88.2 20.2 71.3
InvAD [42] 91.9 95.2 86.9 99.4 50.7 53.2 94.9 36.2 81.7
InvAD-lite [42] 90.3 94.3 85.4 99.3 49.1 51.9 94.5 35.1 80.7
MambaAD [15] 90.3 94.3 85.1 99.2 48.4 51.7 93.8 34.9 80.4
Hybrid UniAD [35] 81.0 88.2 78.9 97.2 22.0 32.1 81.7 19.1 68.7
RD++ [31] 82.2 88.3 80.6 97.9 22.6 31.0 89.0 18.3 70.2
DesTSeg [44] 82.0 89.2 79.0 90.0 50.6 52.6 66.1 35.7 72.8
phone_battery Aug. DRAEM [37] 51.7 46.1 58.0 45.8 00.1 00.2 14.6 00.1 30.9
SimpleNet [27] 55.2 48.7 58.1 75.4 01.5 06.0 46.8 03.1 41.7
Emb. CFA [21] 49.9 43.5 58.0 64.0 00.9 03.1 40.6 01.6 37.1
CFLOW-AD [14] 74.9 73.4 64.2 75.2 13.6 22.7 83.9 12.8 58.3
PyramidalFlow [22] 45.5 36.1 58.4 71.6 00.4 01.2 27.2 00.6 34.3
Rec. RD [10] 87.5 82.5 77.4 77.3 22.3 32.1 94.6 19.1 67.7
DiAD [16] 82.3 77.7 75.9 96.8 05.3 11.4 85.4 06.0 62.1
ViTAD [40] 90.7 89.3 79.4 98.8 24.1 32.3 91.0 19.3 72.2
InvAD [42] 91.2 87.7 80.8 86.2 22.2 30.4 95.1 17.9 70.5
InvAD-lite [42] 91.2 89.6 81.8 80.3 25.5 33.8 95.7 20.3 71.1
MambaAD [15] 90.2 88.6 80.6 99.4 35.2 39.9 95.5 24.9 75.6
Hybrid UniAD [35] 84.9 81.2 73.5 89.1 10.5 18.4 88.7 10.1 63.8
RD++ [31] 88.6 85.1 78.4 83.4 22.5 31.9 95.1 19.0 69.3
DesTSeg [44] 83.4 78.5 71.9 75.6 10.5 20.6 62.3 11.5 57.5
plastic_nut Aug. DRAEM [37] 53.7 34.0 51.7 51.4 00.1 00.1 16.7 00.0 29.7
SimpleNet [27] 61.9 43.0 52.6 74.0 00.5 02.3 40.9 01.2 39.3
Emb. CFA [21] 60.2 39.8 52.9 75.3 00.2 00.8 39.1 00.4 38.3
CFLOW-AD [14] 69.7 60.5 53.6 95.6 13.4 14.9 79.1 08.0 55.3
PyramidalFlow [22] 45.0 29.3 49.7 76.3 00.1 00.4 29.3 00.2 32.9
Rec. RD [10] 81.3 70.0 64.5 98.8 22.4 30.4 92.6 17.9 65.7
DiAD [16] 71.9 58.2 65.6 81.1 00.4 03.4 38.6 01.7 45.6
ViTAD [40] 82.3 73.8 64.8 97.6 20.8 28.3 87.1 16.5 65.0
InvAD [42] 89.7 82.7 74.6 99.5 29.0 32.7 96.9 19.6 72.2
InvAD-lite [42] 88.7 81.6 73.3 99.5 30.3 34.6 97.0 20.9 72.1
MambaAD [15] 87.8 82.0 71.9 99.5 34.3 38.0 96.8 23.4 72.9
Hybrid UniAD [35] 80.1 70.5 63.0 98.4 23.0 28.4 89.9 16.5 64.8
RD++ [31] 79.8 69.1 63.7 98.9 21.9 28.8 92.5 16.8 65.0
DesTSeg [44] 84.8 78.4 68.9 85.2 49.2 49.5 62.7 32.9 68.4
plastic_plug Aug. DRAEM [37] 64.1 52.9 54.8 41.6 00.0 00.1 09.4 00.0 31.8
SimpleNet [27] 51.3 40.0 54.6 75.9 00.2 00.8 38.6 00.4 37.3
Emb. CFA [21] 42.0 33.4 54.6 85.6 00.2 00.9 53.7 00.5 38.6
CFLOW-AD [14] 78.8 75.2 64.9 97.2 11.8 20.7 85.7 11.5 62.0
PyramidalFlow [22] 46.2 35.2 54.6 79.7 00.1 00.3 40.5 00.2 36.7
Rec. RD [10] 82.6 75.5 68.8 99.1 21.3 30.2 95.5 17.8 67.6
DiAD [16] 88.7 89.2 90.9 92.9 08.7 15.0 66.1 08.1 64.5
ViTAD [40] 80.8 74.7 66.7 97.5 13.3 22.6 84.6 12.7 62.9
InvAD [42] 89.0 82.9 76.2 99.0 22.9 30.7 94.3 18.1 70.7
InvAD-lite [42] 88.0 83.9 74.6 99.1 26.2 33.7 94.3 20.2 71.4
MambaAD [15] 86.5 83.4 72.8 99.0 25.2 32.3 93.0 19.3 70.3
Hybrid UniAD [35] 80.9 76.6 66.8 98.6 18.4 27.1 89.9 15.7 65.5
RD++ [31] 85.5 77.9 73.2 99.2 21.3 28.8 95.8 16.8 68.8
DesTSeg [44] 73.6 64.7 61.6 71.9 04.4 08.7 41.7 04.6 46.7
Table A49: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
porcelain_doll Aug. DRAEM [37] 45.1 29.1 49.8 46.6 00.1 00.1 14.6 00.1 26.5
SimpleNet [27] 63.5 50.7 51.2 81.4 02.4 07.4 48.9 03.8 43.6
Emb. CFA [21] 69.4 55.1 55.2 80.1 00.1 00.3 41.7 00.2 43.1
CFLOW-AD [14] 83.6 78.1 68.2 97.1 14.0 20.9 86.6 11.7 64.1
PyramidalFlow [22] 48.7 34.0 49.8 84.5 00.2 00.3 48.7 00.2 38.0
Rec. RD [10] 87.3 78.2 72.3 99.2 25.1 33.8 96.1 20.3 70.3
DiAD [16] 72.6 66.8 65.2 93.1 01.4 04.8 70.4 02.5 53.5
ViTAD [40] 86.4 77.5 70.8 99.0 21.6 29.7 94.7 17.4 68.5
InvAD [42] 87.1 78.2 70.5 98.5 23.3 33.0 92.7 19.7 69.0
InvAD-lite [42] 88.6 82.8 74.4 99.3 32.0 36.6 96.4 22.4 72.9
MambaAD [15] 88.4 82.7 74.4 99.2 32.2 36.9 95.9 22.6 72.8
Hybrid UniAD [35] 84.9 76.0 70.3 98.6 14.9 24.8 93.3 14.1 66.1
RD++ [31] 88.4 80.7 72.6 99.3 26.3 34.7 96.7 21.0 71.2
DesTSeg [44] 75.0 61.4 59.4 76.3 35.4 38.7 44.5 24.0 55.8
regulator Aug. DRAEM [37] 46.1 26.1 43.8 38.8 00.0 00.1 11.4 00.0 23.8
SimpleNet [27] 47.5 26.2 43.8 79.9 00.1 00.7 40.8 00.4 34.1
Emb. CFA [21] 55.7 32.2 45.1 81.4 00.4 01.4 41.5 00.7 36.8
CFLOW-AD [14] 50.5 29.5 43.9 88.2 01.2 02.1 58.0 01.0 39.1
PyramidalFlow [22] 55.1 31.3 44.9 76.9 00.1 00.3 42.9 00.1 35.9
Rec. RD [10] 67.9 49.9 48.9 98.1 09.0 17.7 89.6 09.7 54.4
DiAD [16] 72.1 71.4 78.2 84.2 00.4 01.5 44.4 00.7 50.3
ViTAD [40] 61.3 45.6 44.7 95.6 07.6 17.8 78.3 09.8 50.1
InvAD [42] 84.5 78.6 68.8 99.5 25.9 33.7 96.6 20.3 69.7
InvAD-lite [42] 75.7 67.1 57.3 98.7 18.1 29.2 92.6 17.1 62.7
MambaAD [15] 72.1 62.7 53.4 97.9 21.7 30.1 88.5 17.7 60.9
Hybrid UniAD [35] 51.9 34.4 44.3 94.2 07.3 14.9 72.1 08.1 45.6
RD++ [31] 70.3 54.4 51.2 98.3 11.1 20.1 90.8 11.2 56.6
DesTSeg [44] 68.9 56.3 48.6 67.9 29.1 32.6 35.5 19.5 48.4
rolled_strip_base Aug. DRAEM [37] 53.3 68.5 79.8 37.1 00.1 00.3 09.4 00.1 35.5
SimpleNet [27] 66.7 80.1 79.9 77.7 01.5 04.9 50.3 02.5 51.6
Emb. CFA [21] 54.9 68.8 80.1 85.0 00.7 01.6 51.5 00.8 48.9
CFLOW-AD [14] 92.6 96.6 88.9 97.8 10.6 15.0 93.2 08.1 70.7
PyramidalFlow [22] 59.9 74.9 79.8 88.1 01.3 03.5 69.1 01.8 53.8
Rec. RD [10] 97.8 98.8 95.1 99.7 30.5 39.3 98.7 24.4 80.0
DiAD [16] 68.4 55.9 56.8 87.7 00.6 03.2 63.4 01.6 48.0
ViTAD [40] 98.3 99.1 95.5 99.5 27.1 35.5 97.8 21.6 79.0
InvAD [42] 99.3 99.6 97.7 99.8 32.8 42.5 99.0 27.0 81.5
InvAD-lite [42] 98.0 99.0 95.3 99.8 31.6 41.1 99.0 25.9 80.5
MambaAD [15] 98.3 99.2 95.6 99.6 27.4 32.3 98.2 19.2 78.7
Hybrid UniAD [35] 98.7 99.3 96.1 99.5 21.6 32.9 98.2 19.7 78.0
RD++ [31] 96.3 97.9 93.5 99.7 26.8 36.4 98.5 22.2 78.4
DesTSeg [44] 96.5 98.4 93.2 91.2 55.9 55.5 81.4 38.5 81.7
sim_card_set Aug. DRAEM [37] 56.7 62.2 70.7 47.5 00.2 00.3 18.6 00.1 36.6
SimpleNet [27] 50.3 57.9 70.7 73.9 02.8 07.1 30.3 03.7 41.9
Emb. CFA [21] 50.0 57.1 70.7 79.5 00.9 03.0 39.3 01.5 42.9
CFLOW-AD [14] 91.5 92.9 85.3 98.2 30.3 35.2 87.9 21.3 74.5
PyramidalFlow [22] 77.9 75.3 77.1 66.0 01.0 00.6 31.0 00.3 47.0
Rec. RD [10] 91.7 92.0 85.2 98.4 39.3 43.6 89.0 27.9 77.0
DiAD [16] 72.6 53.7 61.5 89.9 01.7 05.8 60.4 03.0 49.4
ViTAD [40] 92.9 92.4 89.5 96.2 23.3 31.1 75.4 18.4 71.5
InvAD [42] 94.1 94.2 87.6 98.0 37.7 43.0 84.7 27.4 77.0
InvAD-lite [42] 94.5 95.0 88.2 98.6 46.7 46.9 88.9 30.6 79.8
MambaAD [15] 94.7 95.4 87.9 98.7 51.0 50.4 89.1 33.7 81.0
Hybrid UniAD [35] 91.4 91.9 84.3 97.8 34.7 41.2 85.8 25.9 75.3
RD++ [31] 93.7 94.1 87.2 98.6 40.4 43.2 89.9 27.6 78.2
DesTSeg [44] 92.9 94.4 86.3 95.7 69.4 65.7 81.0 49.0 83.6
Table A50: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
switch Aug. DRAEM [37] 46.5 52.1 68.6 44.5 00.3 00.6 12.9 00.3 32.2
SimpleNet [27] 47.8 53.1 68.6 69.9 01.2 03.1 49.3 01.6 41.9
Emb. CFA [21] 52.2 56.1 68.6 76.0 01.1 02.9 44.1 01.5 43.0
CFLOW-AD [14] 75.3 79.9 72.5 90.9 13.9 17.7 78.8 09.7 61.3
PyramidalFlow [22] 60.8 62.0 69.9 50.0 00.3 00.6 15.0 00.3 36.9
Rec. RD [10] 84.6 87.5 78.2 94.4 19.4 26.6 91.0 15.3 68.8
DiAD [16] 73.4 49.4 61.2 90.5 01.4 05.3 64.2 02.7 49.3
ViTAD [40] 87.8 90.0 80.6 96.4 28.8 39.2 88.5 24.4 73.0
InvAD [42] 95.8 96.8 91.1 98.9 43.9 51.5 96.0 34.7 82.0
InvAD-lite [42] 92.8 94.5 86.7 99.0 44.3 48.6 95.1 32.1 80.1
MambaAD [15] 92.4 94.3 86.3 98.4 34.9 42.2 93.8 26.7 77.5
Hybrid UniAD [35] 85.0 88.4 77.8 98.8 51.4 55.4 91.1 38.3 78.3
RD++ [31] 82.4 86.1 75.8 95.9 22.0 29.7 90.9 17.5 69.0
DesTSeg [44] 89.4 92.4 82.1 87.9 61.3 57.9 68.3 40.8 77.0
tape Aug. DRAEM [37] 50.0 35.4 54.6 38.0 00.1 00.2 11.4 00.1 27.1
SimpleNet [27] 46.3 37.4 54.6 82.8 01.3 04.0 42.8 02.0 38.5
Emb. CFA [21] 54.5 42.1 55.1 80.2 00.5 01.9 51.1 01.0 40.8
CFLOW-AD [14] 93.5 92.6 84.3 99.2 24.6 24.8 95.6 14.2 73.5
PyramidalFlow [22] 60.4 43.3 58.3 66.9 00.2 00.4 10.6 00.2 34.3
Rec. RD [10] 96.2 95.4 88.4 99.7 42.0 47.4 98.4 31.1 81.1
DiAD [16] 73.9 57.8 66.1 81.7 00.4 02.7 47.3 01.3 47.1
ViTAD [40] 93.6 92.6 84.6 99.4 30.7 38.7 95.9 24.0 76.5
InvAD [42] 97.8 96.9 91.9 99.7 38.9 44.0 98.3 28.2 81.1
InvAD-lite [42] 97.3 96.4 90.1 99.7 41.1 45.8 98.4 29.7 81.3
MambaAD [15] 97.1 96.2 89.6 99.8 45.9 48.4 98.4 31.9 82.2
Hybrid UniAD [35] 96.9 95.9 88.7 99.6 31.6 38.4 97.5 23.8 78.4
RD++ [31] 96.8 95.8 89.8 99.8 40.7 46.1 98.5 30.0 81.1
DesTSeg [44] 90.1 88.7 79.2 87.1 58.3 55.9 67.7 38.8 75.3
terminalblock Aug. DRAEM [37] 44.5 49.0 68.7 46.6 00.1 00.1 13.6 00.1 31.8
SimpleNet [27] 53.6 56.1 68.7 85.0 00.7 02.0 57.7 01.0 46.3
Emb. CFA [21] 55.9 59.9 68.7 89.9 00.7 02.8 64.0 01.4 48.8
CFLOW-AD [14] 81.1 84.3 76.0 97.0 12.1 17.7 86.5 09.7 65.0
PyramidalFlow [22] 57.8 57.5 70.0 91.0 00.5 01.7 67.3 00.8 49.4
Rec. RD [10] 89.5 90.0 83.2 99.5 27.9 35.6 98.0 21.7 74.8
DiAD [16] 62.1 36.4 47.8 75.5 00.1 01.1 38.5 00.5 37.4
ViTAD [40] 89.6 89.3 85.1 98.9 20.1 30.1 94.8 17.7 72.6
InvAD [42] 97.4 98.0 92.5 99.8 35.5 39.1 98.8 24.3 80.2
InvAD-lite [42] 96.9 97.6 91.4 99.8 34.6 38.2 98.8 23.6 79.6
MambaAD [15] 95.3 95.5 89.8 99.6 26.8 32.8 97.6 19.6 76.8
Hybrid UniAD [35] 87.0 88.6 80.6 99.1 22.7 30.8 94.6 18.2 71.9
RD++ [31] 89.3 90.2 82.3 99.5 27.7 35.0 97.8 21.2 74.5
DesTSeg [44] 68.4 73.3 69.4 69.3 22.2 27.2 40.3 15.7 52.9
toothbrush Aug. DRAEM [37] 47.1 50.4 69.8 39.4 00.6 01.6 14.2 00.8 31.9
SimpleNet [27] 61.7 63.5 70.2 79.1 03.0 07.3 46.3 03.8 47.3
Emb. CFA [21] 58.4 60.3 70.4 84.6 03.8 06.7 48.3 03.5 47.5
CFLOW-AD [14] 70.3 74.9 71.7 94.7 18.8 14.0 78.0 07.6 60.3
PyramidalFlow [22] 48.1 50.9 70.1 50.8 08.4 02.6 15.6 01.3 35.2
Rec. RD [10] 81.8 83.9 77.4 96.9 26.3 33.6 89.0 20.2 69.8
DiAD [16] 91.2 93.7 90.9 82.0 01.9 06.6 54.5 03.4 60.1
ViTAD [40] 81.8 84.8 76.5 96.7 25.3 34.9 87.1 21.1 69.6
InvAD [42] 88.2 90.0 81.6 97.6 29.5 37.8 91.0 23.3 73.7
InvAD-lite [42] 87.0 88.5 80.9 97.4 29.3 37.0 90.8 22.7 73.0
MambaAD [15] 86.2 87.5 80.7 97.6 30.1 37.9 91.9 23.4 73.1
Hybrid UniAD [35] 81.9 83.9 77.9 96.1 20.5 30.2 85.9 17.8 68.1
RD++ [31] 83.6 85.5 78.7 97.2 27.5 35.8 90.0 21.8 71.2
DesTSeg [44] 72.4 76.2 71.5 67.3 21.8 27.9 42.0 16.2 54.2
Table A51: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
toy Aug. DRAEM [37] 55.3 60.6 73.6 41.6 00.1 00.2 10.6 00.1 34.6
SimpleNet [27] 53.7 61.8 73.4 75.2 00.2 00.9 36.9 00.5 43.2
Emb. CFA [21] 47.8 56.6 73.4 78.3 00.3 01.0 40.6 00.5 42.6
CFLOW-AD [14] 60.6 68.8 73.7 87.5 02.1 08.2 56.9 04.3 51.1
PyramidalFlow [22] 56.2 65.2 73.4 79.8 00.4 01.6 47.4 00.8 46.3
Rec. RD [10] 70.4 74.8 75.7 95.1 05.1 13.0 82.6 07.0 59.5
DiAD [16] 66.2 57.3 59.8 82.1 01.1 04.2 50.3 02.1 45.9
ViTAD [40] 76.5 82.9 76.0 92.8 09.2 19.8 79.2 11.0 62.3
InvAD [42] 86.8 89.8 83.5 98.0 16.6 25.7 91.8 14.8 70.3
InvAD-lite [42] 84.4 88.9 80.3 96.6 18.1 29.0 89.2 17.0 69.5
MambaAD [15] 83.7 88.4 79.7 96.2 16.8 26.1 88.0 15.0 68.4
Hybrid UniAD [35] 68.5 75.1 74.4 93.6 04.8 11.8 73.0 06.3 57.3
RD++ [31] 73.4 78.1 76.4 96.0 07.8 17.7 85.6 09.7 62.1
DesTSeg [44] 71.5 77.5 75.8 67.5 12.9 19.6 47.5 10.8 53.2
toy_brick Aug. DRAEM [37] 51.5 42.9 57.9 48.5 00.1 00.3 11.3 00.1 30.4
SimpleNet [27] 57.4 48.8 57.9 85.9 04.5 10.6 47.0 05.6 44.6
Emb. CFA [21] 56.7 45.9 58.3 78.7 01.3 03.9 40.6 02.0 40.8
CFLOW-AD [14] 74.2 69.7 64.0 96.1 24.2 26.7 79.2 15.4 62.0
PyramidalFlow [22] 54.6 43.4 58.2 50.1 00.1 00.3 15.3 00.2 31.7
Rec. RD [10] 63.4 56.2 58.9 96.2 16.6 24.4 75.8 13.9 55.9
DiAD [16] 68.4 45.3 55.9 93.5 03.1 08.1 66.4 04.2 48.7
ViTAD [40] 69.1 64.3 60.2 94.6 20.6 29.4 72.9 17.2 58.7
InvAD [42] 74.1 67.4 64.4 97.6 25.4 31.4 82.4 18.6 63.2
InvAD-lite [42] 69.1 62.4 61.1 96.4 17.5 24.2 78.0 13.8 58.4
MambaAD [15] 70.6 64.0 61.8 96.6 18.9 26.8 75.9 15.5 59.2
Hybrid UniAD [35] 78.9 74.6 68.1 97.7 21.8 30.4 82.7 17.9 64.9
RD++ [31] 66.7 58.3 60.6 96.7 17.1 25.3 77.3 14.5 57.4
DesTSeg [44] 72.7 68.5 62.8 82.5 30.1 35.7 55.5 21.8 58.3
transistor1 Aug. DRAEM [37] 57.2 59.6 72.2 43.3 00.2 00.6 13.0 00.3 35.2
SimpleNet [27] 61.7 66.8 72.6 84.5 05.4 10.1 58.5 05.3 51.4
Emb. CFA [21] 53.9 63.0 72.1 87.7 03.5 08.7 56.8 04.5 49.4
CFLOW-AD [14] 91.8 94.7 86.2 98.1 26.2 28.3 90.7 16.5 73.7
PyramidalFlow [22] 49.1 56.2 72.4 77.1 00.4 00.9 35.3 00.4 41.6
Rec. RD [10] 90.9 93.7 85.2 99.1 30.5 36.5 95.5 22.3 75.9
DiAD [16] 73.1 63.1 62.7 88.6 07.2 15.3 58.1 08.3 52.6
ViTAD [40] 91.7 93.9 86.3 98.8 32.0 36.4 93.5 22.2 76.1
InvAD [42] 97.6 98.2 94.2 99.5 38.7 39.9 97.5 24.9 80.8
InvAD-lite [42] 96.4 97.5 92.1 99.5 40.5 42.4 97.4 26.9 80.8
MambaAD [15] 94.9 96.4 89.2 99.4 38.1 39.4 96.8 24.5 79.2
Hybrid UniAD [35] 93.7 95.9 88.8 98.9 27.3 33.0 94.2 19.8 76.0
RD++ [31] 93.4 95.5 89.0 99.3 34.6 38.5 96.2 23.8 78.1
DesTSeg [44] 88.0 91.6 82.5 87.9 44.0 44.1 67.0 28.3 72.2
u_block Aug. DRAEM [37] 55.2 35.2 49.4 43.5 00.1 00.2 10.0 00.1 27.7
SimpleNet [27] 58.4 40.0 51.2 72.8 00.6 02.5 40.2 01.3 38.0
Emb. CFA [21] 54.9 38.6 48.9 74.7 00.3 01.8 36.4 00.9 36.5
CFLOW-AD [14] 80.3 73.0 63.8 98.4 19.9 24.7 89.3 14.1 64.2
PyramidalFlow [22] 44.9 28.6 48.8 81.9 00.5 01.2 39.5 00.6 35.1
Rec. RD [10] 90.1 85.5 75.4 99.6 40.3 45.8 97.2 29.7 76.3
DiAD [16] 75.2 68.4 67.9 88.8 01.6 05.4 54.2 02.8 51.6
ViTAD [40] 84.3 78.2 67.8 99.0 30.3 38.4 90.8 23.8 69.8
InvAD [42] 92.9 89.5 80.5 99.6 33.4 41.3 96.7 26.0 76.3
InvAD-lite [42] 90.4 86.4 76.2 99.6 34.5 43.0 96.8 27.4 75.3
MambaAD [15] 90.0 85.8 74.8 99.5 33.2 42.8 96.1 27.2 74.6
Hybrid UniAD [35] 88.2 83.7 74.4 99.3 24.2 32.7 94.9 19.6 71.1
RD++ [31] 90.2 85.9 75.7 99.6 34.1 41.2 97.3 26.0 74.9
DesTSeg [44] 76.7 70.4 61.0 83.6 57.6 55.0 53.3 37.9 65.4
Table A52: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
usb Aug. DRAEM [37] 45.8 43.8 63.0 48.2 00.1 00.5 19.9 00.3 31.6
SimpleNet [27] 52.2 49.0 63.0 80.4 00.8 02.8 50.8 01.4 42.7
Emb. CFA [21] 56.6 48.6 63.2 85.9 00.7 02.4 51.6 01.2 44.1
CFLOW-AD [14] 68.0 69.8 63.4 94.8 13.0 15.8 75.0 08.6 57.1
PyramidalFlow [22] 48.3 44.5 63.0 79.2 00.2 00.2 39.8 00.1 39.3
Rec. RD [10] 85.2 84.6 75.6 98.1 26.8 35.2 91.2 21.3 71.0
DiAD [16] 58.9 37.4 45.7 78.0 01.0 03.1 28.0 01.6 36.0
ViTAD [40] 86.9 87.3 77.7 98.3 26.1 34.8 90.6 21.1 71.7
InvAD [42] 94.1 93.3 86.8 99.5 38.5 44.0 97.2 28.2 79.1
InvAD-lite [42] 93.1 92.7 85.8 99.5 37.0 43.7 96.8 27.9 78.4
MambaAD [15] 92.7 92.7 85.3 99.3 39.7 44.8 96.0 28.9 78.6
Hybrid UniAD [35] 80.8 81.8 70.9 98.0 21.0 30.9 85.8 18.3 67.0
RD++ [31] 85.5 85.8 75.6 98.0 27.2 36.3 90.8 22.2 71.3
DesTSeg [44] 87.6 88.3 78.5 91.5 45.0 50.9 70.5 34.2 73.2
usb_adaptor Aug. DRAEM [37] 54.0 45.6 56.5 49.5 00.1 00.1 20.0 00.0 32.3
SimpleNet [27] 49.9 40.6 56.7 52.3 00.1 00.4 19.2 00.2 31.3
Emb. CFA [21] 52.1 41.8 56.6 78.4 00.1 00.8 39.5 00.4 38.5
CFLOW-AD [14] 69.9 64.1 59.5 96.2 08.1 16.6 80.8 09.1 56.5
PyramidalFlow [22] 53.4 40.5 56.8 84.2 00.2 00.6 48.4 00.3 40.6
Rec. RD [10] 70.9 61.0 61.9 94.4 09.7 18.7 74.4 10.3 55.9
DiAD [16] 76.9 60.2 67.2 94.0 02.3 06.6 75.5 03.4 54.7
ViTAD [40] 76.1 66.1 66.2 90.8 05.9 13.9 56.2 07.5 53.6
InvAD [42] 80.0 74.4 66.8 96.5 15.8 24.3 79.3 13.8 62.4
InvAD-lite [42] 79.8 74.2 67.3 97.0 16.6 24.8 82.8 14.2 63.2
MambaAD [15] 79.1 75.7 66.0 97.0 15.8 24.9 81.6 14.2 62.9
Hybrid UniAD [35] 76.9 71.5 64.3 96.5 10.9 19.8 80.1 11.0 60.0
RD++ [31] 69.9 62.7 61.1 94.7 10.2 20.1 76.2 11.2 56.4
DesTSeg [44] 78.0 73.3 66.1 80.8 27.6 36.7 51.2 22.4 59.1
vcpill Aug. DRAEM [37] 47.3 37.7 56.2 31.2 00.3 00.8 07.2 00.4 25.8
SimpleNet [27] 62.8 52.2 58.6 82.0 06.5 12.0 47.6 06.4 46.0
Emb. CFA [21] 58.4 47.1 57.3 85.5 04.3 09.3 54.5 04.9 45.2
CFLOW-AD [14] 81.4 79.1 67.5 97.3 35.6 41.3 84.6 26.0 69.5
PyramidalFlow [22] 55.7 43.9 57.8 72.3 00.6 00.8 28.4 00.4 37.1
Rec. RD [10] 85.1 80.3 72.5 98.3 43.7 48.2 89.1 31.8 73.9
DiAD [16] 64.1 40.4 56.2 90.2 01.3 05.2 60.8 02.7 45.5
ViTAD [40] 82.2 80.5 70.4 97.6 44.4 50.0 81.3 33.3 72.3
InvAD [42] 90.1 88.0 78.2 98.9 50.4 55.2 91.1 38.1 78.8
InvAD-lite [42] 84.6 82.2 71.3 98.4 41.4 48.1 87.9 31.7 73.4
MambaAD [15] 88.5 87.6 77.6 98.7 47.9 52.0 89.4 35.1 77.4
Hybrid UniAD [35] 88.6 86.3 75.8 99.2 47.3 47.7 92.7 31.3 76.8
RD++ [31] 85.9 82.6 73.0 98.5 46.0 50.7 90.2 34.0 75.3
DesTSeg [44] 69.5 65.1 58.6 75.3 44.1 45.1 43.5 29.1 57.3
wooden_beads Aug. DRAEM [37] 52.8 44.9 60.2 43.9 00.1 00.3 12.5 00.1 30.7
SimpleNet [27] 56.2 52.2 60.2 75.5 01.2 04.4 36.2 02.2 40.8
Emb. CFA [21] 55.3 46.2 60.2 83.6 00.8 03.1 48.4 01.6 42.5
CFLOW-AD [14] 79.7 79.3 68.6 96.5 20.2 27.3 79.5 15.8 64.4
PyramidalFlow [22] 66.4 56.3 64.8 50.1 00.6 00.3 15.1 00.1 36.2
Rec. RD [10] 81.1 79.1 70.1 97.9 27.2 35.5 85.8 21.6 68.1
DiAD [16] 62.1 56.4 65.9 85.0 01.1 04.7 45.6 02.4 45.8
ViTAD [40] 82.8 82.3 72.9 97.1 31.0 37.4 79.4 23.0 69.0
InvAD [42] 86.6 85.6 76.2 98.4 35.5 40.3 88.0 25.3 72.9
InvAD-lite [42] 85.6 84.6 75.1 98.2 30.5 37.3 87.9 22.9 71.3
MambaAD [15] 83.3 82.3 72.7 98.1 32.4 39.7 85.3 24.8 70.5
Hybrid UniAD [35] 80.9 80.1 70.3 97.7 20.3 27.9 85.0 16.2 66.0
RD++ [31] 83.4 81.4 72.5 98.2 28.9 36.5 87.8 22.3 69.8
DesTSeg [44] 78.2 73.6 69.8 77.9 30.8 34.9 50.5 21.1 59.4
Table A53: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
woodstick Aug. DRAEM [37] 52.9 28.3 45.2 57.3 00.1 00.2 22.9 00.1 29.6
SimpleNet [27] 55.7 35.7 44.6 74.0 02.8 09.1 32.1 04.8 36.3
Emb. CFA [21] 57.8 33.2 46.5 70.2 00.9 04.1 26.5 02.1 34.2
CFLOW-AD [14] 71.3 60.5 52.8 92.6 32.1 38.3 67.9 23.7 59.4
PyramidalFlow [22] 50.1 28.1 44.4 50.0 00.0 00.2 15.0 00.1 26.8
Rec. RD [10] 77.5 61.9 58.5 97.8 31.2 38.8 85.0 24.0 64.4
DiAD [16] 74.1 66.0 62.1 90.9 02.6 08.0 60.7 04.2 52.1
ViTAD [40] 75.6 64.0 56.3 96.3 38.4 43.7 77.8 28.0 64.6
InvAD [42] 80.8 67.6 62.7 98.0 39.0 45.1 85.9 29.1 68.4
InvAD-lite [42] 78.7 65.9 59.3 97.6 36.3 41.0 83.8 25.8 66.1
MambaAD [15] 81.8 71.0 64.4 97.9 42.8 47.0 85.5 30.7 70.1
Hybrid UniAD [35] 81.1 73.7 64.6 94.8 41.5 46.0 79.2 29.9 68.7
RD++ [31] 77.2 61.2 58.2 97.9 30.6 38.0 85.3 23.4 64.1
DesTSeg [44] 80.2 70.6 60.5 91.9 61.5 60.5 69.9 43.4 70.7
zipper Aug. DRAEM [37] 46.5 61.6 76.7 36.4 00.7 01.5 11.3 00.8 33.5
SimpleNet [27] 56.8 67.7 76.7 52.7 01.2 03.8 24.3 01.9 40.5
Emb. CFA [21] 80.6 88.0 79.9 92.5 15.6 22.6 78.3 12.7 65.4
CFLOW-AD [14] 94.3 97.1 90.3 98.2 34.9 35.6 91.0 21.6 77.3
PyramidalFlow [22] 49.2 61.7 77.5 49.9 00.7 01.4 15.6 00.7 36.6
Rec. RD [10] 95.4 97.2 91.6 99.0 43.7 49.9 96.5 33.2 81.9
DiAD [16] 86.0 87.0 84.0 90.2 12.5 18.8 53.5 10.4 61.7
ViTAD [40] 98.8 99.3 95.9 99.1 47.8 50.3 96.4 33.6 83.9
InvAD [42] 98.8 99.2 96.0 99.2 54.3 56.8 97.0 39.7 85.9
InvAD-lite [42] 98.9 99.4 95.6 99.2 54.0 57.4 97.0 40.3 85.9
MambaAD [15] 99.4 99.6 97.1 99.3 58.4 60.9 97.6 43.8 87.5
Hybrid UniAD [35] 98.2 98.9 95.3 98.4 33.1 36.9 95.1 22.6 79.4
RD++ [31] 97.0 98.3 93.0 99.1 47.1 51.6 96.4 34.7 83.2
DesTSeg [44] 93.4 96.0 89.6 70.5 18.9 25.2 71.2 14.4 66.4
Avg Aug. DRAEM [37] 50.9 45.9 61.3 44.0 00.2 00.4 13.6 00.2 30.9
SimpleNet [27] 54.9 50.6 61.5 76.1 01.9 04.9 42.4 02.5 41.8
Emb. CFA [21] 55.7 50.5 61.9 81.3 01.6 03.8 48.8 02.0 43.4
CFLOW-AD [14] 77.0 75.8 69.9 94.8 17.6 21.7 80.4 12.4 62.5
PyramidalFlow [22] 54.4 48.0 62.0 71.1 01.2 01.1 34.9 00.5 39.0
Rec. RD [10] 82.7 79.3 74.1 97.2 25.2 32.8 90.0 20.0 68.8
DiAD [16] 75.6 66.4 69.9 88.0 02.9 07.1 58.1 03.7 52.6
ViTAD [40] 82.7 80.2 73.7 97.2 24.3 32.3 84.8 19.6 67.9
InvAD [42] 89.4 87.0 80.2 98.4 32.6 38.9 92.7 24.6 74.2
InvAD-lite [42] 87.2 85.2 77.8 98.0 31.7 37.9 92.0 23.8 72.8
MambaAD [15] 87.0 85.3 77.6 98.6 32.4 38.1 91.2 23.9 72.9
Hybrid UniAD [35] 83.1 81.2 74.5 97.4 23.3 30.9 87.1 18.6 68.2
RD++ [31] 83.6 80.6 74.8 97.7 25.9 33.6 90.7 20.5 69.6
DesTSeg [44] 79.3 76.7 70.7 80.3 36.9 40.3 56.1 26.2 62.9
Appendix N Detailed Quantitative Results on COCO-AD Dataset Under 100 epochs
Table A54: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
0 Aug. DRAEM [37] 48.8 74.5 85.0 51.9 16.6 27.7 18.1 16.1 46.1
SimpleNet [27] 63.4 82.1 85.0 78.2 39.1 44.6 44.4 28.7 62.4
Emb. CFA [21] 52.1 76.1 85.0 59.3 22.2 29.6 19.4 17.4 49.1
CFLOW-AD [14] 80.2 91.2 87.1 81.9 46.4 48.9 51.8 32.4 69.6
PyramidalFlow [22] 48.5 73.3 85.0 50.0 08.1 27.7 15.0 16.1 43.9
Rec. RD [10] 65.7 82.0 85.0 72.0 30.8 38.3 47.9 23.7 60.2
DiAD [16] 57.5 77.5 85.3 67.0 33.4 26.2 28.8 20.1 53.7
ViTAD [40] 84.8 93.9 88.0 83.1 55.3 53.2 51.0 36.2 72.8
InvAD [42] 74.3 88.1 85.1 78.0 41.6 44.6 52.3 28.7 66.3
InvAD-lite [42] 75.7 89.7 85.3 76.6 40.8 43.0 47.8 27.4 65.6
MambaAD [15] 74.0 88.9 85.2 75.4 38.2 41.5 48.9 26.2 64.6
Hybrid UniAD [35] 66.2 84.5 85.1 70.4 29.3 36.7 39.8 22.4 58.9
RD++ [31] 70.3 85.4 85.1 74.9 34.9 41.1 50.3 25.9 63.1
DesTSeg [44] 57.1 77.4 85.0 54.2 21.2 27.7 28.2 16.1 50.1
1 Aug. DRAEM [37] 47.4 42.7 60.9 48.1 02.4 05.8 13.6 03.0 31.6
SimpleNet [27] 45.0 39.8 60.9 53.3 03.5 06.5 19.3 03.4 32.6
Emb. CFA [21] 52.4 44.8 61.0 56.6 03.4 06.9 13.1 03.6 34.0
CFLOW-AD [14] 59.5 50.0 61.7 76.1 08.2 14.2 46.4 07.6 45.2
PyramidalFlow [22] 52.4 45.7 60.9 50.1 19.5 05.8 15.0 03.0 35.6
Rec. RD [10] 54.6 46.3 61.0 70.1 05.9 10.7 40.9 05.6 41.4
DiAD [16] 54.4 49.8 62.2 71.3 11.8 07.8 28.8 06.2 40.9
ViTAD [40] 58.5 51.2 61.8 78.9 16.3 23.4 38.1 13.3 46.9
InvAD [42] 55.7 47.9 61.0 73.4 07.9 13.4 43.0 07.2 43.2
InvAD-lite [42] 54.0 47.0 61.1 71.9 07.3 12.2 36.6 06.5 41.4
MambaAD [15] 54.9 47.7 61.0 70.8 06.2 11.0 38.4 05.8 41.4
Hybrid UniAD [35] 55.9 47.8 61.0 69.1 05.9 10.4 37.3 05.5 41.1
RD++ [31] 55.8 46.6 61.1 69.2 06.2 08.8 41.1 04.6 41.3
DesTSeg [44] 54.2 47.0 61.1 59.1 05.3 09.6 29.5 05.0 38.0
2 Aug. DRAEM [37] 58.1 37.0 51.0 50.4 05.6 11.9 15.2 06.3 32.7
SimpleNet [27] 55.3 36.8 49.9 55.7 07.8 12.9 22.7 06.9 34.4
Emb. CFA [21] 62.2 42.3 51.6 60.4 09.2 15.1 26.8 08.2 38.2
CFLOW-AD [14] 65.5 44.8 53.7 75.8 14.0 22.8 46.1 12.9 46.1
PyramidalFlow [22] 50.2 32.4 49.4 50.0 09.4 11.9 15.0 06.3 31.2
Rec. RD [10] 58.6 37.3 50.9 67.3 11.3 17.3 40.7 09.5 40.5
DiAD [16] 63.8 43.4 52.5 68.0 19.2 12.2 33.2 10.6 41.8
ViTAD [40] 67.5 53.9 52.6 80.5 28.9 35.9 40.2 21.9 51.4
InvAD [42] 66.5 45.8 54.4 73.5 15.5 23.5 49.5 13.3 47.0
InvAD-lite [42] 67.4 48.9 53.7 72.1 16.2 23.0 43.1 13.0 46.3
MambaAD [15] 65.2 45.3 53.3 70.9 13.9 21.4 45.0 12.0 45.0
Hybrid UniAD [35] 48.8 31.6 49.5 60.0 07.6 14.6 27.9 07.9 34.3
RD++ [31] 51.5 31.0 50.3 70.3 12.7 16.9 43.5 09.2 39.5
DesTSeg [44] 56.6 37.8 50.0 53.8 08.8 11.9 22.3 06.3 34.5
3 Aug. DRAEM [37] 59.6 39.8 53.3 49.1 05.8 12.4 14.3 06.6 33.5
SimpleNet [27] 58.6 38.7 53.4 53.7 07.5 12.9 18.0 06.9 34.7
Emb. CFA [21] 60.0 41.5 52.9 48.4 06.4 12.4 12.2 06.6 33.4
CFLOW-AD [14] 65.5 45.7 55.5 70.1 12.6 19.7 46.3 10.9 45.1
PyramidalFlow [22] 55.2 38.1 51.2 50.0 16.7 12.4 15.0 06.6 34.1
Rec. RD [10] 51.4 34.1 51.3 56.7 07.8 13.7 29.7 07.4 35.0
DiAD [16] 60.1 41.4 52.9 65.9 17.5 10.6 32.3 09.6 40.1
ViTAD [40] 56.6 38.2 52.3 62.3 10.0 16.4 27.1 09.0 37.6
InvAD [42] 60.2 42.0 53.1 59.7 09.0 14.9 38.5 08.0 39.6
InvAD-lite [42] 56.2 38.6 51.6 59.1 08.6 14.3 31.3 07.7 37.1
MambaAD [15] 57.1 38.9 51.5 58.6 08.5 14.2 34.0 07.6 37.5
Hybrid UniAD [35] 49.9 33.2 51.3 59.0 08.2 14.4 32.1 07.8 35.4
RD++ [31] 52.5 34.3 51.6 58.5 08.3 13.5 33.7 07.3 36.1
DesTSeg [44] 49.8 35.3 51.1 50.8 07.2 12.4 17.6 06.6 32.0
Table A55: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 53.5 48.5 62.5 49.9 07.6 14.4 15.3 08.0 36.0
SimpleNet [27] 55.6 49.4 62.3 60.2 14.5 19.2 26.1 11.5 41.0
Emb. CFA [21] 56.7 51.2 62.6 56.2 10.3 16.0 17.9 08.9 38.7
CFLOW-AD [14] 67.7 57.9 64.5 76.0 20.3 26.4 47.7 16.0 51.5
PyramidalFlow [22] 51.6 47.4 61.6 50.0 13.4 14.4 15.0 08.0 36.2
Rec. RD [10] 57.6 49.9 62.0 66.5 13.9 20.0 39.8 11.5 44.2
DiAD [16] 59.0 53.0 63.2 68.1 20.5 14.2 30.8 11.6 44.1
ViTAD [40] 66.9 59.3 63.7 76.2 27.6 32.2 39.1 20.1 52.1
InvAD [42] 64.2 56.0 63.4 71.2 18.5 24.1 45.8 14.3 49.0
InvAD-lite [42] 63.3 56.0 62.9 69.9 18.2 23.1 39.7 13.7 47.6
MambaAD [15] 62.8 55.2 62.8 68.9 16.7 22.0 41.6 12.9 47.1
Hybrid UniAD [35] 55.2 49.3 61.7 64.6 12.8 19.0 34.3 10.9 42.4
RD++ [31] 57.5 49.3 62.0 68.2 15.5 20.1 42.2 11.8 45.0
DesTSeg [44] 54.4 49.4 61.8 54.5 10.6 15.4 24.4 08.5 38.6
Appendix O Detailed Quantitative Results on MVTec AD Dataset Under 300 epochs
Table A56: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bottle Aug. DRAEM [37] 42.8 73.8 86.3 55.0 06.2 11.3 15.8 06.0 41.6
SimpleNet [27] 85.4 95.6 87.6 80.8 21.9 26.6 50.7 15.3 64.1
RealNet [45] 98.7 99.6 96.9 72.5 60.5 53.4 61.5 36.4 77.6
Emb. CFA [21] 36.7 71.5 86.3 23.6 03.5 10.9 02.4 05.8 33.6
PatchCore [29] 100. 100. 100. 98.9 79.0 77.1 96.1 62.7 93.9
CFLOW-AD [14] 99.9 100. 99.2 97.3 62.3 63.9 92.1 47.0 87.8
PyramidalFlow [22] 75.1 91.3 86.9 54.8 07.4 11.8 17.9 06.3 49.3
Rec. RD [10] 99.6 99.9 98.4 97.8 68.0 67.9 94.6 51.5 89.5
DiAD [16] 97.8 99.1 99.2 93.5 45.9 49.4 75.9 32.8 80.1
ViTAD [40] 100. 100. 100. 98.8 80.0 75.3 94.7 60.4 92.7
InvAD [42] 100. 100. 100. 98.4 74.8 72.2 95.2 56.6 91.5
InvAD-lite [42] 100. 100. 100. 98.4 74.8 73.2 95.5 57.7 91.7
MambaAD [15] 100. 100. 100. 98.9 80.3 77.1 96.2 62.7 93.2
Hybrid UniAD [35] 99.8 100. 99.2 98.0 68.9 69.8 94.5 53.6 90.0
RD++ [31] 99.8 100. 99.2 98.2 71.0 71.3 94.8 55.4 90.6
DesTSeg [44] 100. 100. 100. 95.7 88.0 80.8 92.6 67.8 93.9
cable Aug. DRAEM [37] 53.6 66.2 76.0 52.4 02.7 05.6 14.7 02.9 38.7
SimpleNet [27] 71.3 81.6 77.1 76.8 13.0 18.2 51.5 10.0 55.6
RealNet [45] 63.3 80.1 76.0 59.8 29.4 26.6 29.1 15.3 52.0
Emb. CFA [21] 56.4 71.8 76.0 57.1 07.4 08.1 16.3 04.2 41.9
PatchCore [29] 99.8 99.9 99.5 97.7 51.6 55.2 93.6 38.1 87.1
CFLOW-AD [14] 90.3 94.1 88.2 89.9 27.7 33.8 79.4 20.3 71.9
PyramidalFlow [22] 59.6 75.3 76.3 74.7 17.9 17.8 37.5 09.8 51.3
Rec. RD [10] 81.9 88.4 80.6 84.9 25.9 33.0 76.8 19.8 67.4
DiAD [16] 85.9 92.4 85.7 90.8 27.2 33.0 64.4 19.8 68.5
ViTAD [40] 98.8 99.2 96.3 95.4 41.0 45.0 88.9 29.1 80.7
InvAD [42] 99.2 99.5 96.8 97.3 47.0 51.1 92.1 34.3 83.3
InvAD-lite [42] 96.3 97.7 92.6 90.2 36.9 45.7 86.5 29.6 78.0
MambaAD [15] 98.9 99.3 95.6 95.9 42.6 48.0 90.9 31.6 81.6
Hybrid UniAD [35] 96.6 98.0 91.8 97.0 49.3 55.2 86.6 38.2 82.1
RD++ [31] 94.5 96.3 91.6 93.7 37.2 44.5 86.4 28.6 77.7
DesTSeg [44] 93.0 96.2 88.3 92.2 62.8 62.3 70.9 45.3 80.8
capsule Aug. DRAEM [37] 41.4 81.7 90.5 47.9 01.1 01.8 17.6 00.9 40.3
SimpleNet [27] 73.6 92.1 91.1 93.4 19.8 26.6 77.1 15.3 67.7
RealNet [45] 51.0 85.4 90.5 51.8 23.3 05.7 18.8 02.9 46.6
Emb. CFA [21] 40.2 79.1 90.5 81.8 04.0 08.4 56.0 04.4 51.4
PatchCore [29] 99.2 99.8 98.6 99.0 49.0 52.9 95.7 35.9 86.7
CFLOW-AD [14] 87.8 97.2 93.4 98.5 39.9 43.7 93.3 28.0 79.1
PyramidalFlow [22] 54.6 87.9 90.5 89.5 09.1 15.1 55.7 08.2 57.5
Rec. RD [10] 98.2 99.6 96.9 98.8 44.2 50.4 96.0 33.7 83.4
DiAD [16] 97.9 99.6 97.3 90.3 16.2 23.1 54.2 13.0 68.4
ViTAD [40] 95.8 99.1 95.4 98.1 41.6 47.5 92.2 31.2 81.4
InvAD [42] 98.2 99.6 96.9 99.0 47.5 51.1 95.8 34.3 84.0
InvAD-lite [42] 96.5 99.2 96.8 98.6 43.9 49.0 95.1 32.5 82.7
MambaAD [15] 94.8 98.9 95.4 98.4 43.8 47.1 93.9 30.8 81.8
Hybrid UniAD [35] 87.5 96.9 93.5 98.7 45.6 48.5 92.5 32.0 80.5
RD++ [31] 97.6 99.5 96.8 98.8 45.4 51.2 95.7 34.4 83.6
DesTSeg [44] 91.3 98.2 93.4 86.0 57.5 58.4 59.0 41.3 77.7
carpet Aug. DRAEM [37] 34.4 72.8 86.4 50.2 01.6 03.2 15.2 01.6 37.7
SimpleNet [27] 88.8 96.7 89.6 95.4 37.3 44.0 80.9 28.2 76.1
RealNet [45] 97.3 99.2 96.1 87.8 73.0 68.3 82.3 51.8 86.3
Emb. CFA [21] 54.3 81.4 86.4 41.0 03.1 08.0 25.7 04.2 42.8
PatchCore [29] 98.4 99.5 97.8 99.1 66.4 64.6 95.2 47.7 89.9
CFLOW-AD [14] 99.2 99.8 98.3 98.9 59.0 61.4 94.7 44.3 87.3
PyramidalFlow [22] 56.7 80.3 86.4 81.9 07.3 14.9 49.0 08.0 53.8
Rec. RD [10] 98.4 99.5 97.2 99.0 57.7 60.1 95.6 42.9 86.8
DiAD [16] 91.3 97.6 90.5 87.7 16.9 26.6 65.5 15.4 68.0
ViTAD [40] 99.4 99.8 99.4 98.9 61.1 64.2 94.3 47.3 88.2
InvAD [42] 98.8 99.7 97.2 99.1 60.9 62.2 95.6 45.2 87.6
InvAD-lite [42] 99.5 99.8 98.9 99.0 55.6 60.1 96.0 42.9 87.0
MambaAD [15] 99.9 100. 99.4 99.2 62.7 63.3 97.2 46.3 88.8
Hybrid UniAD [35] 99.8 99.9 99.4 98.5 52.9 52.6 95.2 35.7 85.5
RD++ [31] 97.4 99.3 96.6 98.8 52.1 56.3 94.5 39.2 85.0
DesTSeg [44] 93.6 97.9 94.4 96.1 72.3 67.5 94.2 50.9 88.0
Table A57: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
grid Aug. DRAEM [37] 64.2 84.1 85.5 49.3 00.6 01.4 14.3 00.7 42.8
SimpleNet [27] 40.5 69.4 85.7 38.2 00.5 01.4 03.6 00.7 34.2
RealNet [45] 97.4 99.3 98.2 81.0 47.8 49.2 70.3 32.6 77.6
Emb. CFA [21] 51.2 79.2 85.1 27.5 00.4 01.4 01.7 00.7 35.2
PatchCore [29] 99.4 99.8 98.2 98.9 40.8 44.3 94.0 28.4 84.3
CFLOW-AD [14] 89.4 96.5 88.9 93.6 21.3 26.9 82.3 15.5 71.3
PyramidalFlow [22] 78.4 90.2 88.7 90.0 13.9 22.4 75.5 12.6 65.6
Rec. RD [10] 97.8 99.3 96.6 99.2 45.9 47.6 97.3 31.2 83.4
DiAD [16] 99.7 99.9 98.3 85.7 03.9 09.7 57.4 05.1 65.0
ViTAD [40] 99.6 99.9 99.1 98.6 30.6 36.6 95.8 22.4 80.0
InvAD [42] 100. 100. 100. 99.2 45.5 47.3 97.3 31.0 84.2
InvAD-lite [42] 100. 100. 100. 99.3 50.0 50.2 97.7 33.5 85.3
MambaAD [15] 100. 100. 100. 99.2 48.6 48.3 97.6 31.8 84.8
Hybrid UniAD [35] 98.8 99.6 96.6 97.0 23.8 30.0 92.1 17.6 76.8
RD++ [31] 99.1 99.7 98.2 99.1 43.7 48.1 96.7 31.6 83.5
DesTSeg [44] 99.8 99.9 99.1 99.3 65.9 64.9 97.8 48.0 89.5
hazelnut Aug. DRAEM [37] 51.0 69.7 77.8 32.8 02.2 04.2 11.2 02.1 35.6
SimpleNet [27] 93.0 96.1 91.5 92.0 24.7 33.0 81.2 19.8 73.1
RealNet [45] 99.4 99.6 98.6 70.2 47.1 44.9 64.3 29.0 74.9
Emb. CFA [21] 66.1 79.8 78.9 38.0 04.7 08.1 26.7 04.2 43.2
PatchCore [29] 100. 100. 100. 99.0 61.2 63.6 96.2 46.6 90.0
CFLOW-AD [14] 99.9 99.9 99.3 98.5 59.9 56.9 95.7 39.8 87.2
PyramidalFlow [22] 88.5 93.6 87.7 88.1 31.8 39.3 75.5 24.5 72.1
Rec. RD [10] 100. 100. 100. 98.6 58.3 59.3 96.4 42.1 87.5
DiAD [16] 96.6 98.5 96.3 95.3 27.9 37.2 81.0 22.9 76.1
ViTAD [40] 99.8 99.9 98.6 98.9 64.1 62.6 95.2 45.5 88.4
InvAD [42] 100. 100. 100. 98.9 58.7 60.3 96.6 43.2 87.8
InvAD-lite [42] 100. 100. 100. 98.8 58.1 60.9 95.6 43.8 87.6
MambaAD [15] 100. 100. 100. 99.0 64.8 65.6 95.8 48.8 89.3
Hybrid UniAD [35] 99.9 100. 99.3 98.1 53.5 55.2 94.6 38.1 85.8
RD++ [31] 93.4 94.0 94.6 98.7 57.1 58.0 96.4 40.9 84.6
DesTSeg [44] 99.7 99.8 99.3 97.9 90.2 84.2 90.5 72.7 94.5
leather Aug. DRAEM [37] 80.3 91.5 87.8 48.8 00.8 01.3 16.7 00.7 46.7
SimpleNet [27] 87.3 95.2 90.1 82.2 05.3 11.7 74.6 06.2 63.8
RealNet [45] 100. 100. 100. 95.8 73.8 70.4 93.9 54.3 90.6
Emb. CFA [21] 58.1 80.3 86.7 23.7 00.4 01.3 01.7 00.6 36.0
PatchCore [29] 100. 100. 100. 99.4 52.1 51.2 97.8 34.5 87.6
CFLOW-AD [14] 100. 100. 100. 99.3 48.8 49.0 98.3 32.4 85.1
PyramidalFlow [22] 39.1 65.1 87.2 47.9 00.3 01.3 11.6 00.6 36.1
Rec. RD [10] 100. 100. 100. 99.3 38.2 45.4 97.8 29.4 83.0
DiAD [16] 97.4 99.0 96.8 91.5 06.4 13.3 77.9 07.1 68.9
ViTAD [40] 100. 100. 100. 99.5 51.3 55.4 98.0 38.3 86.3
InvAD [42] 100. 100. 100. 99.4 49.9 52.1 97.9 35.2 85.6
InvAD-lite [42] 100. 100. 100. 99.2 41.2 46.4 98.4 30.2 83.6
MambaAD [15] 100. 100. 100. 99.3 47.7 50.1 98.6 33.4 85.1
Hybrid UniAD [35] 100. 100. 100. 98.9 36.3 38.1 97.6 23.5 81.6
RD++ [31] 100. 100. 100. 99.2 39.7 44.4 97.4 28.5 83.0
DesTSeg [44] 100. 100. 99.5 99.5 80.6 73.4 98.8 58.0 93.1
metal_nut Aug. DRAEM [37] 61.2 86.4 89.4 39.8 08.4 21.2 10.8 11.9 45.3
SimpleNet [27] 74.3 92.3 90.1 79.9 33.0 36.8 52.7 22.5 65.6
RealNet [45] 63.4 90.4 89.4 51.6 37.6 21.0 31.2 11.7 54.9
Emb. CFA [21] 55.7 87.0 89.4 58.4 16.2 24.9 10.0 14.2 48.8
PatchCore [29] 100. 100. 100. 98.8 88.9 85.8 95.6 75.1 96.1
CFLOW-AD [14] 97.8 99.5 97.3 96.0 71.4 68.3 88.8 51.9 88.4
PyramidalFlow [22] 68.0 91.1 89.9 80.1 32.8 43.5 31.7 27.8 62.4
Rec. RD [10] 59.9 83.6 93.9 92.1 50.0 52.2 87.5 35.3 74.2
DiAD [16] 94.6 98.8 94.2 94.4 73.7 72.5 53.9 56.9 83.2
ViTAD [40] 99.9 100. 99.5 96.1 74.2 75.9 92.9 61.2 91.2
InvAD [42] 100. 100. 100. 97.7 81.3 82.0 93.7 69.5 93.5
InvAD-lite [42] 99.7 99.9 98.9 97.4 79.6 80.4 93.8 67.2 92.8
MambaAD [15] 99.8 100. 98.9 97.1 77.3 80.2 94.3 66.9 92.5
Hybrid UniAD [35] 98.3 99.5 98.9 92.6 48.8 63.0 79.4 46.0 82.9
RD++ [31] 100. 100. 99.5 96.6 74.0 77.4 93.4 63.1 91.6
DesTSeg [44] 100. 100. 100. 96.9 91.2 86.1 93.2 75.6 95.3
Table A58: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pill Aug. DRAEM [37] 68.0 91.5 92.5 46.0 03.3 06.6 18.4 03.4 46.6
SimpleNet [27] 71.4 93.5 91.6 94.1 53.0 50.7 77.7 34.0 76.0
RealNet [45] 63.3 91.6 91.6 51.1 41.1 06.5 28.3 03.4 53.4
Emb. CFA [21] 62.4 89.5 91.6 48.5 04.8 09.5 27.3 05.0 47.7
PatchCore [29] 97.0 99.4 96.5 98.1 76.8 74.0 96.1 58.7 91.9
CFLOW-AD [14] 85.0 96.8 92.3 96.5 56.2 55.8 90.9 38.7 81.9
PyramidalFlow [22] 68.7 93.0 91.8 77.6 13.9 24.2 61.2 13.8 61.5
Rec. RD [10] 97.0 99.5 97.2 97.5 63.1 64.5 96.0 47.6 87.8
DiAD [16] 94.7 99.0 97.2 93.0 50.4 50.6 65.7 33.9 78.6
ViTAD [40] 96.2 99.4 96.4 98.7 77.5 75.4 95.5 60.6 91.3
InvAD [42] 97.6 99.6 97.2 98.0 69.9 68.3 96.4 51.9 89.6
InvAD-lite [42] 96.5 99.4 96.1 98.0 71.7 70.6 96.6 54.6 89.8
MambaAD [15] 95.5 99.2 96.5 97.1 61.3 64.1 95.6 47.2 87.0
Hybrid UniAD [35] 94.1 98.9 95.1 95.1 40.7 45.9 94.7 29.8 80.6
RD++ [31] 97.6 99.5 97.9 98.2 73.4 70.9 96.3 54.9 90.5
DesTSeg [44] 88.9 97.8 93.0 89.5 83.5 77.4 48.9 63.1 82.7
screw Aug. DRAEM [37] 53.0 75.7 85.6 51.6 00.4 00.5 17.5 00.3 40.6
SimpleNet [27] 63.0 84.6 85.3 91.3 01.7 03.8 70.5 01.9 57.2
RealNet [45] 72.7 89.0 85.6 50.6 09.2 01.9 17.0 01.0 46.6
Emb. CFA [21] 64.2 82.6 85.9 32.2 00.2 00.8 13.1 00.4 39.9
PatchCore [29] 98.3 99.4 97.1 99.5 41.5 44.6 97.2 28.7 84.5
CFLOW-AD [14] 71.4 88.4 85.8 97.0 17.2 24.2 89.3 13.7 67.6
PyramidalFlow [22] 46.7 72.0 85.3 58.5 00.3 01.3 10.7 00.6 39.3
Rec. RD [10] 98.3 99.5 96.1 99.4 41.3 44.9 97.1 29.0 82.4
DiAD [16] 72.4 87.1 87.5 86.4 03.1 08.2 55.4 04.3 57.1
ViTAD [40] 91.2 97.0 92.0 98.9 33.2 39.8 93.6 24.9 78.0
InvAD [42] 97.2 99.0 95.9 99.6 49.5 51.6 97.5 34.8 84.3
InvAD-lite [42] 93.9 97.8 93.5 99.5 46.9 50.4 97.7 33.7 82.8
MambaAD [15] 93.9 97.6 94.0 99.4 50.1 50.2 97.4 33.5 83.2
Hybrid UniAD [35] 92.4 96.6 93.3 98.9 31.5 37.8 94.2 23.3 77.8
RD++ [31] 98.7 99.6 96.7 99.5 41.7 47.5 97.7 31.1 83.1
DesTSeg [44] 86.4 95.3 89.8 74.4 49.2 46.7 50.1 30.5 70.3
tile Aug. DRAEM [37] 63.4 83.5 83.8 48.8 06.4 13.2 14.2 07.0 44.8
SimpleNet [27] 86.5 94.9 88.1 75.0 27.3 35.7 49.0 21.7 65.2
RealNet [45] 98.7 99.6 98.2 91.9 84.8 78.4 87.3 64.5 91.3
Emb. CFA [21] 52.5 76.5 83.6 22.5 04.2 13.2 05.3 07.0 36.8
PatchCore [29] 100. 100. 99.4 96.4 58.7 67.1 86.6 50.5 88.5
CFLOW-AD [14] 99.7 99.9 99.4 95.9 55.9 62.1 86.7 45.0 85.7
PyramidalFlow [22] 82.9 93.8 85.4 79.5 46.6 41.2 49.2 25.9 68.4
Rec. RD [10] 97.9 99.2 95.2 95.2 48.4 60.4 85.7 43.2 83.1
DiAD [16] 93.0 96.6 96.3 76.2 25.0 33.8 50.5 20.3 67.4
ViTAD [40] 99.9 100. 98.8 96.5 56.5 68.7 88.0 52.3 86.9
InvAD [42] 100. 100. 100. 95.5 50.0 61.3 87.6 44.2 84.9
InvAD-lite [42] 98.2 99.3 96.5 94.4 47.9 55.8 82.4 38.7 82.1
MambaAD [15] 98.2 99.3 96.0 93.7 44.7 54.0 80.8 37.0 81.0
Hybrid UniAD [35] 100. 100. 100. 92.6 43.2 51.4 81.5 34.6 81.2
RD++ [31] 99.9 100. 99.4 96.3 52.8 64.3 88.7 47.4 85.9
DesTSeg [44] 100. 100. 100. 98.9 95.9 89.5 98.0 81.0 97.5
toothbrush Aug. DRAEM [37] 43.6 68.4 83.3 49.9 01.7 03.8 16.1 02.0 38.1
SimpleNet [27] 91.4 96.5 90.9 94.4 36.7 40.9 72.1 25.7 74.7
RealNet [45] 85.0 94.6 86.6 72.3 41.5 43.2 25.6 27.6 64.1
Emb. CFA [21] 43.6 72.2 83.3 52.5 05.6 09.5 39.9 05.0 43.8
PatchCore [29] 91.9 96.2 95.2 99.0 54.8 58.5 91.0 41.3 85.1
CFLOW-AD [14] 83.3 94.3 84.6 98.2 46.0 52.3 85.7 35.5 77.8
PyramidalFlow [22] 60.3 82.9 83.3 75.0 17.5 18.1 16.5 10.0 50.5
Rec. RD [10] 35.8 62.6 85.7 97.3 21.3 40.0 86.1 25.0 61.3
DiAD [16] 95.0 98.0 95.2 89.7 11.1 18.4 64.9 10.1 67.5
ViTAD [40] 100. 100. 100. 99.1 54.7 62.5 91.0 45.4 86.8
InvAD [42] 94.2 97.6 95.2 99.0 52.4 59.7 91.5 42.5 84.2
InvAD-lite [42] 98.1 99.2 96.8 99.0 50.3 59.7 91.2 42.6 84.9
MambaAD [15] 98.9 99.6 96.8 99.0 47.8 60.3 92.5 43.2 85.0
Hybrid UniAD [35] 90.3 95.6 95.2 98.5 39.1 49.9 87.9 33.2 79.5
RD++ [31] 96.7 98.7 95.1 99.1 57.0 61.5 92.2 44.4 85.8
DesTSeg [44] 93.9 97.7 92.3 97.5 74.5 70.7 81.1 54.6 86.8
Table A59: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
transistor Aug. DRAEM [37] 42.0 36.8 57.6 50.4 04.2 09.2 14.9 04.8 30.7
SimpleNet [27] 81.8 81.9 70.1 75.8 31.9 38.0 57.0 23.5 62.4
RealNet [45] 73.2 76.8 68.8 57.5 48.6 22.7 31.6 12.8 54.2
Emb. CFA [21] 70.0 59.8 66.0 53.8 08.0 12.0 24.0 06.4 41.9
PatchCore [29] 99.8 99.7 97.5 96.0 62.3 59.2 90.8 42.0 88.0
CFLOW-AD [14] 92.4 93.4 88.9 84.6 38.8 40.3 74.1 25.3 73.2
PyramidalFlow [22] 66.9 64.9 59.8 73.4 17.9 20.4 25.2 11.3 46.9
Rec. RD [10] 93.7 94.4 90.9 85.6 41.7 44.7 74.8 28.8 75.1
DiAD [16] 99.0 98.7 97.5 98.1 72.8 68.5 85.3 52.1 88.6
ViTAD [40] 98.5 97.9 92.1 93.7 57.6 54.5 77.0 37.5 81.6
InvAD [42] 99.7 99.6 96.4 97.1 69.4 66.3 93.1 49.6 88.8
InvAD-lite [42] 96.6 95.6 87.8 95.0 62.4 59.7 85.7 42.5 83.3
MambaAD [15] 100. 99.9 98.8 95.7 64.4 63.5 86.4 46.5 87.0
Hybrid UniAD [35] 100. 100. 100. 97.6 65.2 65.8 94.4 49.0 89.0
RD++ [31] 95.4 95.0 92.3 89.5 48.8 50.2 78.3 33.5 78.5
DesTSeg [44] 99.5 99.3 96.3 76.6 60.4 56.6 78.2 39.5 81.0
wood Aug. DRAEM [37] 52.4 82.0 86.3 47.8 03.7 07.5 13.7 03.9 41.9
SimpleNet [27] 84.4 93.4 91.5 74.6 15.1 23.3 53.6 13.2 62.3
RealNet [45] 99.6 99.9 99.2 87.4 77.0 73.7 85.4 58.3 88.9
Emb. CFA [21] 81.7 92.4 90.3 64.8 07.7 12.9 31.0 06.9 54.4
PatchCore [29] 98.7 99.6 96.7 95.2 52.1 52.3 91.3 35.4 85.5
CFLOW-AD [14] 98.9 99.7 97.5 94.4 47.0 50.4 91.0 33.7 82.7
PyramidalFlow [22] 86.8 95.2 89.9 71.2 35.9 15.2 50.3 08.2 63.5
Rec. RD [10] 99.4 99.8 98.3 95.5 49.4 51.7 91.6 34.9 83.7
DiAD [16] 93.5 97.5 94.4 83.1 15.0 22.8 58.0 12.8 66.3
ViTAD [40] 98.9 99.7 96.8 96.1 60.8 57.9 88.6 40.8 85.5
InvAD [42] 98.9 99.6 98.3 94.8 48.3 50.0 89.9 33.3 82.8
InvAD-lite [42] 97.3 99.1 95.9 94.9 47.3 48.4 88.7 31.9 81.7
MambaAD [15] 98.7 99.6 96.8 94.2 46.7 47.7 92.1 31.3 82.3
Hybrid UniAD [35] 98.6 99.6 96.7 93.3 38.4 43.5 88.0 27.8 79.7
RD++ [31] 99.3 99.8 98.3 95.3 50.2 51.1 90.3 34.3 83.5
DesTSeg [44] 99.2 99.7 98.4 94.1 83.6 76.8 94.8 62.3 92.4
zipper Aug. DRAEM [37] 77.0 90.6 89.8 60.1 03.7 04.1 26.4 02.1 50.2
SimpleNet [27] 95.0 98.7 93.5 92.6 39.2 44.6 78.3 28.7 77.4
RealNet [45] 80.6 94.8 88.1 65.5 55.0 40.5 42.0 25.4 66.6
Emb. CFA [21] 43.4 78.6 88.1 32.7 01.6 04.2 09.1 02.1 36.8
PatchCore [29] 99.6 99.9 98.7 98.9 63.4 64.6 96.4 47.8 90.1
CFLOW-AD [14] 96.0 98.7 97.1 98.0 50.1 55.7 92.8 38.6 84.1
PyramidalFlow [22] 60.9 87.5 88.1 70.0 06.0 08.1 32.7 04.2 50.5
Rec. RD [10] 99.5 99.9 99.2 98.4 52.4 59.9 95.2 42.8 86.4
DiAD [16] 71.8 89.5 90.2 84.0 13.5 23.8 55.5 13.5 61.2
ViTAD [40] 97.7 99.3 97.5 95.9 43.6 50.2 89.4 33.6 81.9
InvAD [42] 99.6 99.9 99.2 98.4 50.8 58.0 95.0 40.9 85.8
InvAD-lite [42] 99.3 99.8 97.5 98.5 57.7 61.0 95.5 43.9 87.0
MambaAD [15] 99.1 99.8 97.4 98.2 58.2 60.5 95.0 43.3 86.9
Hybrid UniAD [35] 95.3 98.5 96.2 97.0 37.7 46.6 91.4 30.4 80.4
RD++ [31] 68.2 89.3 92.2 98.3 50.8 57.9 94.5 40.7 78.7
DesTSeg [44] 99.4 99.8 98.3 94.7 80.8 74.1 90.9 58.8 91.1
Avg Aug. DRAEM [37] 55.2 77.0 83.9 48.7 03.1 06.3 15.8 03.3 41.4
SimpleNet [27] 79.2 90.8 87.6 82.4 24.0 29.0 62.0 17.8 65.0
RealNet [45] 82.9 93.3 90.9 69.8 50.0 40.4 51.2 28.5 68.4
Emb. CFA [21] 55.8 78.8 84.5 43.9 04.8 08.9 19.3 04.7 42.3
PatchCore [29] 98.8 99.5 98.4 98.3 59.9 61.0 94.2 44.9 88.6
CFLOW-AD [14] 92.7 97.2 94.0 95.8 46.8 49.6 89.0 34.0 80.7
PyramidalFlow [22] 66.2 84.3 85.1 74.2 17.2 19.6 40.0 11.4 55.2
Rec. RD [10] 90.5 95.0 95.1 95.9 47.1 52.1 91.2 35.8 81.0
DiAD [16] 92.0 96.8 94.4 89.3 27.3 32.7 64.4 21.3 71.0
ViTAD [40] 98.4 99.4 97.5 97.5 55.2 58.1 91.7 42.0 85.4
InvAD [42] 98.9 99.6 98.2 98.1 57.1 59.6 94.4 43.1 86.6
InvAD-lite [42] 98.1 99.1 96.8 97.3 55.0 58.1 93.1 41.7 85.4
MambaAD [15] 98.5 99.5 97.7 97.6 56.1 58.7 93.6 42.3 86.0
Hybrid UniAD [35] 96.8 98.9 97.0 96.8 45.0 50.2 91.0 34.2 82.2
RD++ [31] 95.8 98.0 96.6 97.3 53.0 57.0 92.9 40.5 84.4
DesTSeg [44] 96.3 98.8 96.1 92.6 75.8 71.3 82.6 56.6 87.6
Appendix P Detailed Quantitative Results on MVTec 3D Dataset Under 300 epochs
Table A60: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bagel Aug. DRAEM [37] 59.8 86.3 88.9 53.2 01.5 05.3 28.4 02.7 46.2
SimpleNet [27] 80.8 94.3 91.3 95.7 19.2 28.0 78.1 16.3 69.6
Emb. CFA [21] 68.2 90.2 88.9 71.5 12.8 23.4 34.8 13.3 55.7
PatchCore [29] 91.8 98.0 92.9 98.8 43.7 46.8 93.3 30.5 82.4
CFLOW-AD [14] 79.6 94.4 89.8 98.3 31.0 39.1 89.6 24.3 74.5
PyramidalFlow [22] 59.9 87.0 88.9 93.9 11.9 17.7 76.8 09.7 62.3
Rec. RD [10] 91.8 97.9 94.3 98.6 39.2 44.8 91.2 28.8 79.7
DiAD [16] 89.6 97.6 91.5 89.6 04.5 10.3 66.2 05.4 64.2
ViTAD [40] 89.3 97.1 93.4 98.6 38.6 43.6 91.8 27.9 78.9
InvAD [42] 91.3 97.8 92.5 98.9 44.7 48.5 93.8 32.0 81.1
InvAD-lite [42] 81.1 94.8 91.1 98.3 40.1 44.5 89.6 28.7 77.1
MambaAD [15] 91.9 98.0 92.0 98.6 38.3 43.6 93.1 27.9 79.4
Hybrid UniAD [35] 84.9 96.1 89.9 98.0 32.9 37.6 89.2 23.2 75.5
RD++ [31] 85.6 96.2 91.8 98.7 39.3 45.9 91.8 29.8 78.5
DesTSeg [44] 96.8 99.3 95.2 96.6 66.6 64.1 89.2 47.1 86.8
cable_gland Aug. DRAEM [37] 26.1 68.5 89.2 56.9 00.4 00.9 26.6 00.5 38.4
SimpleNet [27] 87.1 96.6 92.9 96.9 13.8 20.3 90.3 11.3 71.1
Emb. CFA [21] 42.9 80.5 89.2 43.1 01.3 05.5 23.3 02.8 40.8
PatchCore [29] 96.4 99.0 96.1 99.5 37.3 43.6 98.3 27.9 83.4
CFLOW-AD [14] 75.4 92.1 89.7 97.2 11.4 16.9 91.5 09.2 67.7
PyramidalFlow [22] 55.4 83.9 89.2 75.4 00.8 02.2 47.2 01.1 50.6
Rec. RD [10] 91.5 97.7 93.6 99.4 36.8 41.9 98.2 26.5 79.9
DiAD [16] 88.1 96.5 94.6 94.5 04.1 10.0 81.3 05.3 67.0
ViTAD [40] 85.9 95.0 92.7 98.8 24.2 32.4 96.2 19.4 75.0
InvAD [42] 97.3 99.3 97.7 99.5 40.6 46.9 98.7 30.6 82.9
InvAD-lite [42] 95.1 98.8 94.8 99.4 40.9 47.7 98.4 31.3 82.2
MambaAD [15] 94.6 98.7 94.9 99.4 40.5 46.1 98.4 29.9 81.8
Hybrid UniAD [35] 78.8 91.7 91.6 98.5 23.5 32.4 95.3 19.3 73.1
RD++ [31] 94.6 98.6 95.6 99.5 41.0 45.3 98.4 29.3 81.9
DesTSeg [44] 88.2 97.1 91.3 86.8 45.8 52.1 80.2 35.2 77.4
carrot Aug. DRAEM [37] 32.5 76.1 90.7 57.7 00.7 01.4 31.5 00.7 41.5
SimpleNet [27] 68.2 91.5 91.0 96.5 08.8 15.7 86.3 08.5 65.4
Emb. CFA [21] 48.3 85.8 90.7 56.3 02.6 05.8 32.4 03.0 46.0
PatchCore [29] 95.2 99.0 95.6 99.4 25.9 30.8 97.7 18.2 80.0
CFLOW-AD [14] 86.8 97.2 92.6 98.9 18.6 25.6 96.0 14.7 73.7
PyramidalFlow [22] 64.8 88.4 91.3 96.0 11.6 18.0 86.4 09.9 65.2
Rec. RD [10] 85.4 96.3 93.1 99.3 25.8 32.3 97.6 19.2 75.7
DiAD [16] 92.7 98.1 98.1 97.6 06.3 12.0 91.0 06.4 70.8
ViTAD [40] 79.4 94.0 92.5 98.8 19.5 26.2 95.9 15.1 72.3
InvAD [42] 89.9 97.5 94.9 99.2 24.8 29.6 97.4 17.4 76.2
InvAD-lite [42] 89.0 97.4 93.5 99.3 27.1 32.5 98.0 19.4 76.7
MambaAD [15] 90.7 97.6 95.2 99.4 27.1 32.0 98.0 19.0 77.1
Hybrid UniAD [35] 73.1 93.0 92.0 98.2 13.6 20.5 94.0 11.4 69.2
RD++ [31] 88.8 97.3 93.5 99.3 24.4 29.8 97.2 17.5 75.8
DesTSeg [44] 77.4 94.0 90.7 66.8 34.7 33.5 42.8 20.1 62.8
cookie Aug. DRAEM [37] 51.8 80.7 89.2 44.9 01.2 03.5 14.5 01.8 40.8
SimpleNet [27] 62.0 85.7 88.0 93.1 21.9 29.6 73.1 17.4 64.8
Emb. CFA [21] 48.2 81.2 88.0 49.8 05.2 11.2 22.9 05.9 43.8
PatchCore [29] 71.0 91.6 88.0 98.0 52.4 50.2 92.5 33.5 78.4
CFLOW-AD [14] 46.9 78.7 88.4 96.9 30.8 31.6 89.1 18.8 66.1
PyramidalFlow [22] 24.9 69.8 88.0 85.7 06.3 12.3 62.7 06.6 50.0
Rec. RD [10] 49.9 81.2 88.0 96.3 24.4 31.4 86.4 18.6 65.4
DiAD [16] 53.6 79.3 88.0 87.4 06.1 10.8 67.5 05.7 56.1
ViTAD [40] 71.9 91.7 88.5 97.5 46.3 47.5 87.1 31.2 75.8
InvAD [42] 62.8 87.8 88.0 98.0 46.3 46.6 91.5 30.4 74.4
InvAD-lite [42] 67.1 89.2 88.3 96.2 29.1 34.2 83.9 20.7 69.7
MambaAD [15] 67.2 89.2 88.0 96.5 36.7 39.2 83.4 24.4 71.5
Hybrid UniAD [35] 72.1 90.8 88.5 97.4 42.5 46.8 89.8 30.5 75.4
RD++ [31] 56.1 83.8 88.0 97.3 36.7 39.6 88.5 24.7 70.0
DesTSeg [44] 70.7 90.5 88.6 81.0 50.2 50.4 51.7 33.7 69.0
Table A61: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
dowel Aug. DRAEM [37] 41.9 77.5 88.9 34.5 00.5 02.5 17.0 01.2 37.5
SimpleNet [27] 86.2 96.5 90.0 96.6 13.7 21.9 85.2 12.3 70.0
Emb. CFA [21] 60.1 82.8 90.4 58.5 00.4 00.9 21.1 00.4 44.9
PatchCore [29] 98.2 99.6 97.5 99.3 44.5 46.7 96.3 30.5 85.1
CFLOW-AD [14] 94.9 98.8 95.2 98.9 29.9 34.1 94.8 20.6 78.1
PyramidalFlow [22] 61.3 85.5 89.3 93.5 05.6 10.7 74.2 05.7 60.0
Rec. RD [10] 99.4 99.9 98.6 99.7 51.6 51.4 98.8 34.5 85.6
DiAD [16] 69.9 90.8 92.4 78.6 04.5 09.7 48.2 05.1 56.3
ViTAD [40] 94.3 98.6 93.8 99.2 37.0 41.1 95.9 25.9 80.0
InvAD [42] 99.2 99.8 99.0 99.7 50.5 50.6 98.3 33.8 85.3
InvAD-lite [42] 96.5 99.2 96.2 99.6 50.0 50.0 97.4 33.3 84.1
MambaAD [15] 97.6 99.4 96.1 99.5 51.6 50.4 97.4 33.7 84.6
Hybrid UniAD [35] 94.0 98.6 93.0 99.1 35.6 41.2 96.1 25.9 79.7
RD++ [31] 99.6 99.9 99.0 99.7 52.2 50.6 98.7 33.9 85.7
DesTSeg [44] 82.3 95.5 89.8 84.4 39.8 45.8 63.0 29.7 71.5
foam Aug. DRAEM [37] 50.1 83.0 88.9 57.0 00.3 00.4 19.2 00.2 42.7
SimpleNet [27] 74.3 93.3 88.9 86.4 10.8 21.9 64.2 12.3 62.8
Emb. CFA [21] 61.7 88.7 88.9 44.4 06.4 16.1 32.5 08.8 48.4
PatchCore [29] 78.9 94.4 89.9 94.9 18.8 29.6 83.4 17.4 72.2
CFLOW-AD [14] 73.8 93.3 89.3 85.3 16.2 26.4 57.3 15.2 63.1
PyramidalFlow [22] 55.5 81.3 89.9 61.2 00.4 00.5 32.7 00.3 45.9
Rec. RD [10] 74.7 92.9 89.9 93.5 14.1 27.3 79.6 15.8 67.4
DiAD [16] 79.9 95.3 88.9 88.3 02.0 06.3 62.1 03.2 60.4
ViTAD [40] 79.4 94.4 90.8 94.0 14.2 25.6 80.7 14.7 68.4
InvAD [42] 87.4 96.8 91.6 93.8 18.9 30.1 81.3 17.7 71.4
InvAD-lite [42] 86.4 96.5 90.5 94.9 24.2 32.9 83.6 19.7 72.7
MambaAD [15] 84.3 95.9 89.9 95.3 25.2 33.6 84.8 20.2 72.7
Hybrid UniAD [35] 75.0 93.2 88.9 83.2 09.8 22.4 57.1 12.6 61.4
RD++ [31] 69.2 91.5 89.0 92.4 14.7 26.6 75.8 15.4 65.6
DesTSeg [44] 85.0 96.4 89.8 85.5 44.7 49.6 74.4 33.0 75.1
peach Aug. DRAEM [37] 48.5 82.0 89.1 47.4 00.6 02.3 19.3 01.2 41.3
SimpleNet [27] 65.4 87.9 89.1 94.3 05.8 12.1 80.3 06.5 62.1
Emb. CFA [21] 61.2 87.5 89.4 43.3 01.7 07.6 24.4 04.0 45.0
PatchCore [29] 75.2 92.7 91.2 98.2 26.4 32.4 93.3 19.3 74.5
CFLOW-AD [14] 74.4 91.8 90.5 97.9 16.2 18.4 92.3 10.2 68.8
PyramidalFlow [22] 68.7 90.0 89.6 95.2 06.1 11.7 83.8 06.2 63.6
Rec. RD [10] 64.3 86.3 91.7 98.3 12.9 20.5 93.2 11.4 66.7
DiAD [16] 92.2 94.1 97.7 95.6 04.3 08.7 84.7 04.5 68.2
ViTAD [40] 61.2 87.2 89.1 98.0 18.5 24.8 92.0 14.1 67.3
InvAD [42] 82.5 94.2 93.3 99.2 36.9 39.5 96.8 24.6 77.5
InvAD-lite [42] 85.6 96.1 90.6 99.0 33.2 36.9 95.6 22.6 76.7
MambaAD [15] 89.5 96.9 94.2 99.4 44.0 44.3 97.2 28.5 80.8
Hybrid UniAD [35] 70.0 89.8 91.2 98.0 18.5 23.5 92.6 13.3 69.1
RD++ [31] 73.0 89.2 92.6 98.8 22.4 29.9 94.7 17.6 71.5
DesTSeg [44] 80.3 94.9 89.8 84.9 46.6 46.6 70.6 30.4 73.4
potato Aug. DRAEM [37] 49.0 82.9 89.3 58.8 00.9 03.3 36.1 01.7 45.8
SimpleNet [27] 61.1 86.8 89.3 95.4 03.2 08.0 85.0 04.2 61.3
Emb. CFA [21] 41.6 75.9 89.3 14.3 00.5 01.4 06.9 00.7 32.8
PatchCore [29] 56.8 83.9 90.2 98.9 09.8 17.0 96.0 09.3 66.2
CFLOW-AD [14] 63.9 86.8 89.3 98.5 10.3 13.8 95.0 07.4 65.4
PyramidalFlow [22] 62.3 87.3 90.5 97.0 08.2 15.1 90.5 08.2 64.4
Rec. RD [10] 62.9 87.4 91.1 99.1 14.6 21.7 96.4 12.2 67.6
DiAD [16] 87.9 97.0 93.0 98.5 08.4 14.1 94.7 07.6 70.5
ViTAD [40] 57.9 86.4 89.8 98.8 13.0 18.2 95.6 10.0 65.7
InvAD [42] 67.0 90.2 90.6 99.1 15.7 23.6 96.4 13.4 68.9
InvAD-lite [42] 63.3 87.8 90.4 98.9 11.6 17.5 95.2 09.6 66.4
MambaAD [15] 58.6 85.3 90.2 99.0 16.4 22.8 95.5 12.9 66.8
Hybrid UniAD [35] 51.6 82.7 89.3 97.9 07.2 11.9 92.6 06.3 61.9
RD++ [31] 62.3 88.0 90.1 99.1 15.2 22.1 96.5 12.4 67.6
DesTSeg [44] 64.7 89.2 90.2 54.3 12.8 10.9 21.7 05.7 49.1
Table A62: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
rope Aug. DRAEM [37] 54.8 67.5 82.4 41.6 00.8 01.0 18.6 00.5 38.1
SimpleNet [27] 95.5 98.2 92.9 98.9 41.7 46.1 90.4 30.0 80.5
Emb. CFA [21] 93.0 97.2 92.4 89.8 38.0 45.6 70.3 29.6 75.2
PatchCore [29] 96.7 98.7 95.5 99.6 55.6 56.9 97.4 39.7 87.2
CFLOW-AD [14] 93.8 97.6 92.2 99.3 43.6 47.2 95.4 30.9 81.3
PyramidalFlow [22] 15.9 51.0 81.2 79.2 01.1 01.9 43.2 01.0 39.1
Rec. RD [10] 96.0 98.4 93.1 99.6 48.2 53.8 97.7 36.8 83.8
DiAD [16] 76.9 86.1 85.1 94.5 06.8 13.8 74.5 07.4 62.5
ViTAD [40] 97.0 98.9 95.6 98.5 41.3 42.7 89.0 27.1 80.4
InvAD [42] 96.0 98.5 95.5 99.6 52.4 56.0 96.5 38.9 84.9
InvAD-lite [42] 95.4 98.2 94.7 99.2 45.1 46.8 95.4 30.5 82.1
MambaAD [15] 95.4 98.0 93.1 99.3 49.2 48.9 95.8 32.4 82.8
Hybrid UniAD [35] 97.4 99.0 96.2 99.3 39.4 45.5 96.5 29.5 81.9
RD++ [31] 95.4 98.2 94.2 99.6 49.2 54.1 97.6 37.1 84.0
DesTSeg [44] 94.1 97.9 93.3 96.9 55.6 56.8 92.3 39.6 83.8
tire Aug. DRAEM [37] 77.4 89.3 93.5 55.1 00.3 00.9 22.9 00.5 48.5
SimpleNet [27] 68.4 87.4 89.0 86.0 05.0 10.0 59.0 05.3 57.8
Emb. CFA [21] 75.9 92.9 87.4 48.8 00.7 01.7 17.8 00.8 46.5
PatchCore [29] 80.1 93.6 87.9 99.1 22.6 31.1 95.8 18.4 74.7
CFLOW-AD [14] 58.0 83.3 87.9 97.6 07.7 14.8 91.0 08.0 62.9
PyramidalFlow [22] 56.6 83.6 88.3 94.7 07.2 13.9 82.5 07.4 61.0
Rec. RD [10] 76.8 91.6 89.1 99.1 18.9 28.6 96.0 16.7 71.4
DiAD [16] 57.7 78.0 87.4 93.2 04.9 09.6 80.6 05.0 58.7
ViTAD [40] 71.0 89.2 88.7 97.6 17.0 28.2 89.1 16.4 68.7
InvAD [42] 89.7 96.8 91.0 99.5 31.3 39.6 97.7 24.7 77.9
InvAD-lite [42] 91.5 97.6 93.2 99.4 42.9 46.5 97.0 30.3 81.2
MambaAD [15] 88.2 96.3 92.3 99.5 42.3 47.2 97.1 30.9 80.4
Hybrid UniAD [35] 72.8 91.3 88.8 98.1 14.0 24.6 90.9 14.0 68.6
RD++ [31] 76.5 92.1 88.0 99.2 20.3 27.7 96.4 16.1 71.5
DesTSeg [44] 85.6 95.9 90.1 84.3 40.9 42.5 65.1 27.0 72.1
Avg Aug. DRAEM [37] 49.2 79.4 89.0 50.7 00.7 02.2 23.4 01.1 42.1
SimpleNet [27] 74.9 91.8 90.2 94.0 14.4 21.4 79.2 12.4 66.6
Emb. CFA [21] 60.1 86.3 89.5 52.0 07.0 11.9 28.7 06.9 47.9
PatchCore [29] 84.1 95.1 92.5 98.6 33.7 38.5 94.4 24.5 78.4
CFLOW-AD [14] 74.8 91.4 90.5 96.9 21.6 26.8 89.2 15.9 70.2
PyramidalFlow [22] 52.5 80.8 88.6 87.2 05.9 10.4 68.0 05.6 56.2
Rec. RD [10] 79.3 93.0 92.3 98.3 28.6 35.4 93.5 22.1 74.3
DiAD [16] 78.8 91.3 91.7 91.8 05.2 10.5 75.1 05.6 63.5
ViTAD [40] 78.7 93.3 91.5 98.0 27.0 33.0 91.3 20.2 73.3
InvAD [42] 86.3 95.9 93.4 98.7 36.2 41.1 94.9 26.4 78.1
InvAD-lite [42] 85.1 95.6 92.3 98.4 34.4 39.0 93.4 24.6 76.9
MambaAD [15] 85.8 95.5 92.6 98.6 37.1 40.8 94.1 26.0 77.8
Hybrid UniAD [35] 77.0 92.6 91.0 96.8 23.7 30.6 89.4 18.6 71.6
RD++ [31] 80.1 93.5 92.2 98.4 31.5 37.2 93.6 23.4 75.2
DesTSeg [44] 82.5 95.1 90.9 82.1 43.8 45.2 65.1 30.2 72.1
Appendix Q Detailed Quantitative Results on MVTec LOCO Dataset Under 300 epochs
Table A63: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
breakfast_box Aug. DRAEM [37] 48.8 68.9 77.2 50.1 06.5 12.5 14.3 06.7 39.8
SimpleNet [27] 89.4 94.5 85.9 89.6 58.6 56.1 75.4 39.0 78.5
Emb. CFA [21] 72.0 83.2 78.7 75.6 34.3 32.6 39.1 19.5 59.4
PatchCore [29] 77.1 88.4 78.0 89.4 55.4 49.3 76.1 32.8 74.4
CFLOW-AD [14] 75.8 86.7 77.8 88.8 50.1 48.1 71.7 31.7 71.3
PyramidalFlow [22] 61.2 78.0 77.2 70.6 34.9 32.5 34.4 19.4 55.5
Rec. RD [10] 63.4 79.2 77.2 83.7 26.6 35.6 63.9 21.6 61.4
DiAD [16] 56.8 74.3 77.2 81.6 22.9 34.2 35.5 20.6 54.6
ViTAD [40] 75.0 86.2 78.7 88.1 52.4 50.8 65.7 34.1 71.0
InvAD [42] 78.8 88.5 79.2 89.7 58.9 56.2 75.7 39.1 75.3
InvAD-lite [42] 71.7 83.9 78.0 88.3 47.6 47.7 73.6 31.3 70.1
MambaAD [15] 69.3 82.9 77.2 87.4 37.0 43.6 70.7 27.9 66.9
Hybrid UniAD [35] 77.5 88.4 77.4 88.0 51.8 50.1 72.8 33.4 72.3
RD++ [31] 68.2 81.7 77.2 86.6 39.7 43.9 67.0 28.2 66.3
DesTSeg [44] 84.0 91.1 82.2 77.5 64.0 57.9 68.0 40.8 75.0
juice_bottle Aug. DRAEM [37] 52.9 75.6 83.4 40.0 05.3 10.2 14.7 05.4 40.3
SimpleNet [27] 96.0 98.6 93.6 86.4 47.5 55.1 84.8 38.0 80.3
Emb. CFA [21] 73.0 88.2 83.4 85.1 26.9 29.1 57.4 17.1 63.3
PatchCore [29] 95.2 98.3 94.1 89.9 56.2 57.8 87.1 40.6 84.3
CFLOW-AD [14] 96.2 98.6 93.9 91.7 44.7 47.8 85.4 31.4 79.8
PyramidalFlow [22] 77.3 91.1 83.7 51.9 36.9 10.1 17.6 05.3 52.7
Rec. RD [10] 89.5 96.4 90.3 89.4 47.3 48.2 84.2 31.8 77.9
DiAD [16] 96.0 98.3 94.1 81.1 41.0 45.2 76.0 29.2 76.0
ViTAD [40] 90.5 96.4 88.6 88.6 48.1 49.2 83.4 32.6 77.8
InvAD [42] 96.1 98.6 94.8 90.9 56.1 58.7 88.2 41.5 83.3
InvAD-lite [42] 94.7 98.1 93.7 90.8 52.9 56.8 87.6 39.7 82.1
MambaAD [15] 93.2 97.6 91.7 89.8 53.9 55.8 86.5 38.7 81.2
Hybrid UniAD [35] 95.4 98.3 93.2 90.7 55.5 58.3 89.3 41.2 83.0
RD++ [31] 90.7 96.7 89.6 90.1 49.7 52.4 84.7 35.5 79.1
DesTSeg [44] 99.5 99.8 97.9 79.5 59.9 59.1 79.6 41.9 82.2
pushpins Aug. DRAEM [37] 62.2 66.8 72.2 53.9 00.8 01.9 11.5 01.0 38.5
SimpleNet [27] 76.5 83.6 74.9 55.8 02.7 07.4 59.6 03.8 51.5
Emb. CFA [21] 47.2 57.8 71.4 51.1 00.9 01.7 22.7 00.9 36.1
PatchCore [29] 74.3 82.8 72.7 56.1 03.1 07.5 63.6 03.9 54.6
CFLOW-AD [14] 63.3 71.4 71.4 39.5 01.2 03.4 52.8 01.7 43.3
PyramidalFlow [22] 45.7 52.1 71.5 55.2 00.8 01.7 18.8 00.9 35.1
Rec. RD [10] 71.5 79.6 72.3 51.1 02.2 07.5 68.2 03.9 50.3
DiAD [16] 69.0 73.8 72.9 61.4 01.6 04.6 49.7 02.4 47.6
ViTAD [40] 68.3 74.6 73.6 46.5 02.3 05.7 57.4 02.9 46.9
InvAD [42] 82.6 88.3 77.9 72.2 04.2 09.3 77.4 04.9 58.8
InvAD-lite [42] 78.2 84.1 76.4 64.0 03.2 08.5 70.3 04.5 55.0
MambaAD [15] 78.8 85.1 76.9 61.1 02.8 08.7 70.5 04.5 54.8
Hybrid UniAD [35] 71.3 79.9 72.2 46.3 02.3 08.1 62.0 04.2 48.9
RD++ [31] 75.7 82.7 73.4 54.1 02.7 07.7 67.4 04.0 52.0
DesTSeg [44] 72.2 81.3 71.4 55.1 04.3 10.1 64.0 05.3 51.2
screw_bag Aug. DRAEM [37] 49.1 66.4 78.4 45.6 06.0 13.8 10.6 07.4 38.6
SimpleNet [27] 65.3 78.1 78.3 55.5 09.2 14.8 48.8 08.0 50.0
Emb. CFA [21] 48.9 65.2 78.2 51.1 08.0 13.6 26.4 07.3 41.6
PatchCore [29] 70.1 82.8 78.2 72.1 13.1 22.0 62.1 12.4 59.7
CFLOW-AD [14] 64.8 78.7 78.4 68.9 11.4 20.6 66.8 11.5 55.7
PyramidalFlow [22] 54.2 66.8 78.2 49.9 05.8 13.6 15.9 07.3 40.6
Rec. RD [10] 63.7 78.1 78.2 71.6 12.6 21.8 52.0 12.3 54.0
DiAD [16] 81.7 91.1 80.6 71.4 12.0 22.2 58.9 12.5 59.7
ViTAD [40] 64.1 76.9 79.6 72.4 13.1 21.6 40.9 12.1 52.7
InvAD [42] 70.3 82.4 78.2 73.4 13.7 22.8 52.8 12.9 56.2
InvAD-lite [42] 63.7 77.4 78.5 72.7 13.2 22.4 48.1 12.6 53.7
MambaAD [15] 62.6 76.4 78.2 73.2 13.4 22.7 49.0 12.8 53.6
Hybrid UniAD [35] 65.0 77.1 78.5 67.5 11.3 19.3 62.8 10.7 54.5
RD++ [31] 64.9 79.3 78.2 69.8 12.2 20.5 51.3 11.4 53.7
DesTSeg [44] 70.0 81.1 79.0 50.3 08.6 13.6 46.4 07.3 49.9
Table A64: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
splicing_connectors Aug. DRAEM [37] 44.0 56.7 76.4 50.5 07.4 14.1 17.5 07.6 38.1
SimpleNet [27] 81.7 90.9 80.0 67.0 23.5 27.2 37.6 15.8 58.3
Emb. CFA [21] 61.6 74.3 76.5 60.3 12.4 16.6 36.1 09.1 48.3
PatchCore [29] 85.6 92.9 84.4 68.1 21.5 22.3 60.8 12.6 65.4
CFLOW-AD [14] 80.0 89.7 78.2 65.4 20.4 21.9 73.5 12.3 61.3
PyramidalFlow [22] 52.3 64.1 76.4 49.8 05.9 14.0 15.1 07.5 39.7
Rec. RD [10] 80.5 89.6 78.9 58.0 16.4 18.0 69.0 09.9 58.6
DiAD [16] 82.3 85.5 82.7 65.2 15.6 20.2 51.7 11.2 57.6
ViTAD [40] 82.8 91.0 79.8 70.6 23.6 25.4 62.5 14.6 62.2
InvAD [42] 83.8 91.6 81.7 62.4 21.4 23.7 70.1 13.4 62.1
InvAD-lite [42] 81.5 90.3 81.4 69.0 23.1 24.4 67.5 13.9 62.5
MambaAD [15] 87.3 93.7 84.8 76.0 32.7 33.7 66.1 20.3 67.8
Hybrid UniAD [35] 84.2 91.6 83.0 80.5 42.2 43.2 66.6 27.5 70.2
RD++ [31] 78.7 89.0 78.1 60.0 16.6 18.0 69.3 09.9 58.5
DesTSeg [44] 80.1 89.8 78.8 56.3 22.1 18.5 37.9 10.2 54.8
Avg Aug. DRAEM [37] 51.4 66.9 77.5 48.0 05.2 10.5 13.7 05.6 39.0
SimpleNet [27] 81.8 89.1 82.5 70.9 28.3 32.1 61.2 20.9 63.7
Emb. CFA [21] 60.6 73.7 77.6 64.7 16.5 18.8 36.4 10.8 49.8
PatchCore [29] 80.5 89.0 81.5 75.1 29.9 31.8 69.9 20.4 67.7
CFLOW-AD [14] 76.0 85.0 79.9 70.9 25.6 28.4 70.0 17.7 62.3
PyramidalFlow [22] 58.1 70.4 77.4 55.5 16.8 14.4 20.4 08.1 44.7
Rec. RD [10] 73.7 84.6 79.4 70.7 21.0 26.2 67.5 15.9 60.4
DiAD [16] 77.2 84.6 81.5 72.1 18.6 25.3 54.4 15.2 59.1
ViTAD [40] 76.2 85.0 80.1 73.2 27.9 30.6 62.0 19.3 62.1
InvAD [42] 82.3 89.9 82.4 77.7 30.9 34.1 72.8 22.4 67.2
InvAD-lite [42] 78.0 86.8 81.6 77.0 28.0 32.0 69.4 20.4 64.7
MambaAD [15] 78.2 87.1 81.8 77.5 28.0 32.9 68.6 20.8 64.9
Hybrid UniAD [35] 78.7 87.1 80.9 74.6 32.6 35.8 70.7 23.4 65.8
RD++ [31] 75.7 85.9 79.3 72.1 24.2 28.5 67.9 17.8 61.9
DesTSeg [44] 81.2 88.6 81.8 63.7 31.8 31.9 59.2 21.1 62.6
Appendix R Detailed Quantitative Results on VisA Dataset Under 300 epochs
Table A65: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
candle Aug. DRAEM [37] 58.1 59.6 67.6 45.7 00.1 00.6 16.7 00.3 35.5
SimpleNet [27] 83.3 85.8 77.7 91.3 07.6 17.6 80.4 09.7 63.4
RealNet [45] 53.6 63.1 66.7 53.5 08.8 09.0 53.6 04.7 44.0
Emb. CFA [21] 86.6 85.0 80.0 79.4 02.6 09.0 58.0 04.7 57.2
CFLOW-AD [14] 92.4 93.2 85.4 98.8 14.7 24.0 93.7 13.6 71.7
PyramidalFlow [22] 80.6 75.9 77.0 84.5 02.2 07.7 79.7 04.0 58.2
Rec. RD [10] 91.8 92.7 83.9 99.0 19.1 30.5 95.2 18.0 73.2
DiAD [16] 99.3 99.3 96.0 80.1 01.0 02.6 41.4 01.3 60.0
ViTAD [40] 90.0 91.5 82.4 96.1 17.1 27.6 86.5 16.0 70.2
InvAD [42] 95.2 94.9 92.2 99.3 23.3 33.5 95.7 20.1 76.3
InvAD-lite [42] 97.2 97.5 92.9 99.1 20.5 30.9 96.2 18.3 76.3
MambaAD [15] 95.5 95.8 89.6 99.1 22.4 32.4 96.1 19.3 75.8
Hybrid UniAD [35] 97.3 97.4 91.6 99.3 24.4 33.9 94.8 20.4 77.0
RD++ [31] 92.7 92.8 87.6 99.3 25.5 36.3 95.4 22.2 75.7
DesTSeg [44] 89.8 89.8 83.0 74.5 39.5 41.4 49.9 26.1 66.8
capsules Aug. DRAEM [37] 71.9 80.7 77.5 39.1 00.3 01.0 10.7 00.5 40.2
SimpleNet [27] 66.9 79.4 77.2 90.0 21.2 28.3 57.5 16.5 60.1
RealNet [45] 81.6 90.5 79.8 74.6 34.2 41.1 39.6 25.9 63.1
Emb. CFA [21] 59.6 70.8 78.7 60.6 00.5 01.0 24.3 00.5 42.2
CFLOW-AD [14] 57.5 73.8 77.5 94.5 27.9 27.8 65.9 16.1 60.7
PyramidalFlow [22] 60.5 72.8 77.8 81.9 02.3 06.5 61.3 03.3 51.9
Rec. RD [10] 81.0 89.7 81.6 99.3 57.3 59.4 91.9 42.3 80.0
DiAD [16] 66.7 80.5 79.7 94.3 08.0 16.4 55.3 08.9 57.3
ViTAD [40] 80.2 88.2 80.2 98.3 31.1 41.3 76.0 26.0 70.8
InvAD [42] 87.2 92.8 84.5 99.6 67.1 61.9 94.9 44.8 84.0
InvAD-lite [42] 88.8 92.5 87.4 99.3 60.7 59.1 94.9 42.0 83.2
MambaAD [15] 83.4 88.4 83.8 98.8 32.8 41.6 87.6 26.2 73.8
Hybrid UniAD [35] 73.5 85.6 79.1 98.4 52.1 53.6 80.8 36.6 74.7
RD++ [31] 79.5 88.6 79.8 99.1 56.8 59.7 91.4 42.6 79.3
DesTSeg [44] 86.5 92.9 85.6 95.7 38.7 42.8 77.8 27.2 74.3
cashew Aug. DRAEM [37] 19.8 54.8 80.0 53.0 01.5 03.4 40.8 01.7 36.2
SimpleNet [27] 87.1 93.3 87.2 98.5 68.0 65.2 85.0 48.3 83.5
RealNet [45] 89.5 95.4 86.8 53.5 25.4 10.1 29.0 05.3 55.7
Emb. CFA [21] 78.3 87.5 83.8 82.3 23.2 29.5 28.8 17.3 59.1
CFLOW-AD [14] 95.6 98.1 92.7 99.0 52.5 54.4 94.6 37.4 83.8
PyramidalFlow [22] 89.0 93.8 90.6 49.8 00.6 02.2 15.2 01.1 48.7
Rec. RD [10] 93.3 96.5 91.5 92.3 44.9 50.1 84.5 33.4 79.0
DiAD [16] 99.8 99.9 99.0 89.1 45.6 50.1 21.5 33.4 72.1
ViTAD [40] 87.6 94.0 87.5 98.3 62.6 61.4 79.3 44.3 81.5
InvAD [42] 96.0 98.3 93.9 97.6 65.5 65.5 89.7 48.7 86.6
InvAD-lite [42] 92.7 96.4 88.8 94.6 55.4 57.9 88.9 40.7 82.1
MambaAD [15] 92.5 96.3 90.3 94.7 50.8 55.0 86.2 37.9 80.8
Hybrid UniAD [35] 90.4 95.2 87.7 97.9 44.5 48.9 89.0 32.4 79.1
RD++ [31] 93.5 96.4 92.7 95.9 56.3 56.7 85.1 39.6 82.4
DesTSeg [44] 86.6 93.1 86.5 84.4 48.2 52.0 55.8 35.2 72.4
chewinggum Aug. DRAEM [37] 33.5 58.0 79.0 32.3 00.3 01.1 04.2 00.6 29.8
SimpleNet [27] 94.6 97.6 91.3 97.9 39.7 45.3 74.7 29.3 77.3
RealNet [45] 89.0 95.3 87.0 82.2 63.8 61.5 42.3 44.5 74.4
Emb. CFA [21] 75.9 87.8 82.7 84.4 16.6 26.5 43.1 15.3 59.6
CFLOW-AD [14] 97.2 98.8 95.2 99.2 61.9 60.3 87.8 43.1 85.8
PyramidalFlow [22] 74.5 84.3 83.3 50.1 00.7 01.1 15.1 00.6 44.2
Rec. RD [10] 95.5 98.0 93.8 98.7 58.4 61.0 78.1 43.9 83.4
DiAD [16] 97.6 98.4 96.6 68.5 15.6 30.6 18.6 18.0 60.8
ViTAD [40] 95.2 97.8 92.3 97.7 60.7 58.5 71.4 41.3 81.9
InvAD [42] 97.7 99.0 95.8 98.6 57.2 59.0 80.0 41.9 83.9
InvAD-lite [42] 97.6 99.0 95.3 98.3 56.7 56.4 78.8 39.3 83.2
MambaAD [15] 97.3 98.8 94.7 97.3 27.4 38.0 78.2 23.4 76.0
Hybrid UniAD [35] 99.1 99.5 96.8 99.3 62.4 59.6 88.2 42.5 86.4
RD++ [31] 96.2 98.3 93.0 98.8 58.8 61.8 80.1 44.7 83.9
DesTSeg [44] 96.8 98.7 95.3 97.5 78.3 70.2 68.4 54.1 86.5
Table A66: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
fryum Aug. DRAEM [37] 46.2 65.7 80.0 51.2 03.5 08.7 25.5 04.6 40.1
SimpleNet [27] 82.0 91.5 84.0 93.4 40.6 41.7 78.5 26.4 73.1
RealNet [45] 80.1 90.6 81.0 57.0 32.4 17.8 22.8 09.8 54.5
Emb. CFA [21] 65.4 81.3 80.6 76.7 16.2 22.1 41.3 12.4 54.8
CFLOW-AD [14] 91.2 96.0 90.3 97.6 53.8 55.8 87.4 38.7 81.7
PyramidalFlow [22] 80.2 90.0 82.2 89.3 19.2 18.3 74.5 10.1 64.8
Rec. RD [10] 95.8 98.1 92.4 97.0 48.1 51.8 93.1 35.0 82.3
DiAD [16] 97.0 98.3 95.6 97.5 59.0 61.8 68.9 44.7 82.6
ViTAD [40] 94.3 97.5 91.1 97.5 47.2 52.0 88.9 35.1 81.2
InvAD [42] 97.0 98.6 93.1 97.5 52.5 56.2 92.1 39.1 83.9
InvAD-lite [42] 96.0 98.1 92.8 97.3 50.6 53.2 93.2 36.3 83.0
MambaAD [15] 96.3 98.2 93.1 97.1 49.7 52.9 92.4 36.0 82.8
Hybrid UniAD [35] 91.0 96.0 87.9 97.3 45.3 52.4 85.5 35.5 79.3
RD++ [31] 95.7 98.1 92.2 97.2 49.9 53.6 92.0 36.6 82.7
DesTSeg [44] 92.3 96.6 88.5 67.0 47.6 41.3 35.4 26.0 67.0
macaroni1 Aug. DRAEM [37] 89.4 86.6 81.5 38.7 00.0 00.1 06.2 00.0 43.2
SimpleNet [27] 65.8 60.9 65.6 91.4 01.5 04.6 66.5 02.4 50.9
RealNet [45] 69.9 75.2 66.7 54.7 14.6 12.4 19.7 06.6 44.7
Emb. CFA [21] 56.0 57.4 66.9 87.7 00.4 02.4 56.0 01.2 46.7
CFLOW-AD [14] 81.8 78.7 75.7 99.1 07.9 13.2 95.2 07.1 64.5
PyramidalFlow [22] 57.6 53.5 70.0 91.0 00.5 02.0 60.0 01.0 47.8
Rec. RD [10] 95.2 93.7 88.8 99.8 23.6 31.1 97.2 18.4 75.6
DiAD [16] 96.3 94.9 92.4 55.2 00.0 00.1 03.6 00.0 48.9
ViTAD [40] 85.3 83.5 76.0 98.6 07.1 15.4 90.7 08.3 65.2
InvAD [42] 95.2 93.4 88.7 99.7 21.8 30.5 97.3 18.0 75.2
InvAD-lite [42] 94.2 92.5 86.2 99.8 19.3 27.7 97.7 16.1 73.9
MambaAD [15] 91.9 89.4 83.2 99.3 13.4 21.9 95.4 12.3 70.6
Hybrid UniAD [35] 88.1 86.2 80.0 99.4 10.8 20.1 95.6 11.2 68.6
RD++ [31] 92.8 91.2 86.5 99.7 24.0 32.4 96.7 19.3 74.8
DesTSeg [44] 93.5 89.9 90.6 73.3 15.2 25.7 36.6 14.7 60.7
macaroni2 Aug. DRAEM [37] 57.7 53.4 63.4 58.9 00.0 00.1 25.1 00.1 36.9
SimpleNet [27] 57.9 53.2 63.3 89.8 00.4 01.7 66.4 00.9 47.5
RealNet [45] 59.0 60.2 67.3 53.5 06.8 08.3 20.1 04.3 39.3
Emb. CFA [21] 58.6 56.7 66.9 82.1 00.1 00.8 48.5 00.4 44.8
CFLOW-AD [14] 67.1 65.5 65.6 97.4 01.9 06.3 89.9 03.3 56.2
PyramidalFlow [22] 62.5 59.1 71.6 83.2 00.1 00.2 35.0 00.1 44.5
Rec. RD [10] 85.6 79.3 81.0 99.6 12.3 20.2 97.0 11.3 67.9
DiAD [16] 56.0 46.8 68.2 56.8 00.0 00.1 15.8 00.0 34.8
ViTAD [40] 77.2 70.3 74.2 98.1 03.5 09.5 88.1 05.0 60.1
InvAD [42] 88.3 83.9 83.2 99.6 11.4 19.1 97.8 10.6 69.0
InvAD-lite [42] 88.1 84.6 80.4 99.7 11.7 19.8 98.2 11.0 68.9
MambaAD [15] 80.5 74.3 75.7 96.5 02.0 07.3 87.1 03.8 60.5
Hybrid UniAD [35] 81.3 81.1 71.9 98.3 05.6 13.2 92.8 07.0 63.5
RD++ [31] 82.8 76.9 80.2 99.5 09.6 18.3 96.2 10.1 66.2
DesTSeg [44] 72.2 70.4 73.1 62.8 10.3 17.2 36.0 09.4 48.9
pcb1 Aug. DRAEM [37] 80.6 78.0 78.2 53.2 00.6 00.9 21.4 00.5 44.7
SimpleNet [27] 85.2 87.0 78.1 98.6 73.1 69.7 79.0 53.5 81.5
RealNet [45] 82.7 83.6 77.5 73.6 40.1 43.6 27.3 27.9 61.2
Emb. CFA [21] 58.0 63.9 66.7 88.4 19.7 29.9 39.1 17.5 52.2
CFLOW-AD [14] 94.9 95.2 88.2 99.1 73.0 68.8 88.3 52.4 86.8
PyramidalFlow [22] 67.7 64.6 71.9 53.4 34.1 11.3 15.3 06.0 45.5
Rec. RD [10] 96.4 96.4 91.8 99.5 66.8 62.2 96.1 45.1 87.0
DiAD [16] 91.7 94.1 89.5 95.7 27.7 39.5 56.5 24.6 70.7
ViTAD [40] 95.6 94.1 91.3 99.4 65.6 61.9 89.2 44.8 85.3
InvAD [42] 96.8 96.2 92.9 99.8 80.3 75.4 96.7 60.5 91.2
InvAD-lite [42] 97.7 97.2 94.2 99.8 79.3 73.5 97.1 58.1 91.3
MambaAD [15] 96.0 94.7 94.1 99.8 79.8 72.8 94.9 57.2 90.3
Hybrid UniAD [35] 96.3 95.5 92.8 99.4 68.1 63.3 90.5 46.3 86.6
RD++ [31] 96.3 96.0 90.9 99.4 64.9 59.1 95.5 42.0 86.0
DesTSeg [44] 78.6 77.3 74.5 96.0 66.5 64.3 68.6 47.3 75.1
Table A67: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pcb2 Aug. DRAEM [37] 35.0 44.1 66.7 44.9 00.2 00.4 10.0 00.2 28.8
SimpleNet [27] 82.5 83.8 75.6 94.9 07.2 15.1 76.0 08.2 62.2
RealNet [45] 87.0 89.3 79.2 72.2 13.9 23.0 40.3 13.0 57.8
Emb. CFA [21] 57.2 65.9 66.7 80.0 02.7 07.5 47.7 03.9 46.8
CFLOW-AD [14] 92.1 93.0 85.6 96.5 12.6 18.9 82.5 10.5 68.7
PyramidalFlow [22] 66.8 71.1 68.7 92.7 07.4 08.6 74.9 04.5 55.7
Rec. RD [10] 97.3 96.9 94.2 97.7 21.1 30.1 91.7 17.7 75.6
DiAD [16] 89.9 91.7 82.9 82.2 04.4 12.2 48.1 06.5 58.8
ViTAD [40] 90.0 89.0 84.2 98.0 12.5 20.4 82.6 11.4 68.1
InvAD [42] 97.0 97.3 93.9 99.1 15.7 23.4 92.8 13.2 74.2
InvAD-lite [42] 96.2 96.4 90.1 99.0 11.7 23.4 92.5 13.3 72.8
MambaAD [15] 96.1 96.3 91.0 99.0 14.2 24.1 91.2 13.7 73.1
Hybrid UniAD [35] 93.5 94.1 87.2 98.3 13.0 19.5 86.0 10.8 70.2
RD++ [31] 97.9 98.3 96.0 98.0 25.5 33.4 91.7 20.0 77.3
DesTSeg [44] 82.8 83.2 78.1 94.4 19.1 24.0 69.8 13.6 64.5
pcb3 Aug. DRAEM [37] 45.6 50.3 66.7 41.2 00.2 00.5 11.0 00.3 30.8
SimpleNet [27] 84.2 86.2 76.6 97.3 12.9 22.4 82.7 12.6 66.0
RealNet [45] 87.0 90.8 82.9 76.9 24.6 32.2 40.9 19.2 62.2
Emb. CFA [21] 56.3 60.8 67.1 85.6 13.8 22.5 49.6 12.7 50.8
CFLOW-AD [14] 80.8 84.1 74.2 96.4 20.6 26.0 81.1 14.9 66.2
PyramidalFlow [22] 66.3 68.6 67.6 90.7 05.0 12.3 58.0 06.5 52.6
Rec. RD [10] 96.8 96.7 90.6 98.0 23.8 34.6 93.6 20.9 76.3
DiAD [16] 94.0 94.4 88.1 87.2 04.0 10.7 41.8 05.7 60.0
ViTAD [40] 91.2 91.1 86.1 98.3 22.6 28.7 88.0 16.8 72.3
InvAD [42] 97.7 97.9 93.3 99.2 15.8 28.0 93.2 16.2 75.0
InvAD-lite [42] 96.4 96.5 90.6 99.2 17.8 28.0 93.6 16.3 74.6
MambaAD [15] 95.2 95.5 88.7 99.1 18.9 27.4 92.5 15.8 73.9
Hybrid UniAD [35] 89.8 90.6 83.3 98.5 13.4 23.9 86.1 13.5 69.4
RD++ [31] 95.7 95.6 90.8 98.0 30.9 34.1 93.3 20.5 76.9
DesTSeg [44] 95.8 96.3 89.8 90.6 27.5 31.6 61.2 18.7 70.4
pcb4 Aug. DRAEM [37] 73.1 66.8 77.3 36.1 00.4 01.3 06.1 00.7 37.3
SimpleNet [27] 95.4 95.1 91.1 92.2 13.9 23.4 67.1 13.2 68.3
RealNet [45] 92.1 92.1 84.4 68.6 37.2 38.7 31.4 24.0 63.5
Emb. CFA [21] 89.5 87.4 83.7 91.4 07.6 14.1 70.0 07.6 63.4
CFLOW-AD [14] 98.9 98.8 95.7 96.8 22.2 29.7 85.6 17.4 75.4
PyramidalFlow [22] 62.6 64.8 69.5 92.7 06.9 12.4 78.4 06.6 55.3
Rec. RD [10] 99.9 99.9 98.5 97.7 30.0 35.4 88.4 21.5 78.5
DiAD [16] 99.8 99.7 99.0 95.4 09.9 18.4 82.4 10.1 72.1
ViTAD [40] 98.9 98.5 97.0 99.0 42.2 47.7 93.5 31.3 82.4
InvAD [42] 99.9 99.9 98.5 98.5 44.1 43.4 90.6 27.7 82.1
InvAD-lite [42] 99.7 99.7 97.5 98.8 49.7 48.7 92.0 32.2 83.7
MambaAD [15] 99.7 99.7 96.4 98.6 44.0 44.9 90.6 28.9 82.0
Hybrid UniAD [35] 99.4 99.4 96.2 97.6 34.3 39.0 85.2 24.2 78.7
RD++ [31] 99.9 99.9 99.0 97.1 27.4 35.9 85.8 21.9 77.9
DesTSeg [44] 98.0 98.0 93.0 94.4 53.8 52.1 74.9 35.3 80.6
pipe_fryum Aug. DRAEM [37] 63.2 76.9 80.8 46.0 01.3 03.1 14.3 01.6 40.8
SimpleNet [27] 83.7 92.1 84.2 97.9 64.1 61.6 76.6 44.5 80.0
RealNet [45] 79.1 91.2 80.3 64.1 48.5 37.7 39.6 23.2 62.9
Emb. CFA [21] 63.7 81.4 80.0 96.8 61.0 59.4 77.8 42.3 74.3
CFLOW-AD [14] 97.6 98.8 95.0 99.2 61.6 61.4 94.9 44.3 86.9
PyramidalFlow [22] 59.7 76.7 80.0 89.8 15.7 21.5 64.2 12.1 58.2
Rec. RD [10] 98.2 99.1 96.6 99.1 55.1 58.1 95.9 41.0 86.0
DiAD [16] 97.7 98.7 98.0 98.6 55.1 58.1 77.9 41.0 83.5
ViTAD [40] 97.6 98.9 94.6 99.5 65.1 66.3 94.9 49.6 88.1
InvAD [42] 99.5 99.7 98.0 99.5 69.7 67.1 95.1 50.5 89.8
InvAD-lite [42] 98.3 99.2 95.6 99.2 61.5 60.5 95.1 43.4 87.1
MambaAD [15] 98.8 99.3 97.6 98.8 52.4 53.0 94.2 36.0 84.9
Hybrid UniAD [35] 97.0 98.5 95.1 98.9 50.1 55.0 93.9 37.9 84.1
RD++ [31] 94.1 96.9 91.9 98.9 54.9 57.0 94.2 39.9 84.0
DesTSeg [44] 95.0 97.5 93.1 86.8 74.3 69.7 55.9 53.5 81.8
Table A68: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 56.2 64.6 74.9 45.0 00.7 01.8 16.0 00.9 37.0
SimpleNet [27] 80.7 83.8 79.3 94.4 29.2 33.1 74.2 22.1 67.8
RealNet [45] 79.2 84.8 78.3 65.4 29.2 27.9 33.9 17.4 57.0
Emb. CFA [21] 67.1 73.8 75.3 83.0 13.7 18.7 48.7 11.3 54.3
CFLOW-AD [14] 87.2 89.5 85.1 97.8 34.2 37.2 87.3 24.9 74.0
PyramidalFlow [22] 69.0 72.9 75.8 79.1 07.9 08.7 52.6 04.7 52.3
Rec. RD [10] 93.9 94.8 90.4 98.1 38.4 43.7 91.9 29.0 78.7
DiAD [16] 90.5 91.4 90.4 83.4 19.2 25.0 44.3 16.2 63.5
ViTAD [40] 90.3 91.2 86.4 98.2 36.4 40.9 85.8 27.5 75.6
InvAD [42] 95.6 96.0 92.3 99.0 43.7 46.9 93.0 32.6 80.9
InvAD-lite [42] 95.3 95.8 91.0 98.7 41.2 44.9 93.2 30.6 80.0
MambaAD [15] 93.6 93.9 89.8 98.2 34.0 39.3 90.5 25.9 77.0
Hybrid UniAD [35] 91.4 93.3 87.5 98.5 35.3 40.2 89.0 26.5 76.5
RD++ [31] 93.1 94.1 90.0 98.4 40.4 44.8 91.4 29.9 78.9
DesTSeg [44] 89.0 90.3 85.9 84.8 43.3 44.4 57.5 30.1 70.7
Appendix S Detailed Quantitative Results on BTAD Dataset Under 300 epochs
Table A69: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
1.0 Aug. DRAEM [37] 75.0 88.7 86.7 44.8 02.8 06.3 13.0 03.3 45.3
SimpleNet [27] 99.2 99.7 98.0 94.6 38.0 42.6 64.2 27.1 76.6
RealNet [45] 100. 100. 100. 94.2 63.8 63.6 84.3 46.6 86.6
Emb. CFA [21] 98.2 99.4 99.0 94.6 37.3 43.0 64.5 27.4 76.6
PatchCore [29] 99.6 99.8 99.0 97.1 55.5 56.6 75.8 39.5 85.4
CFLOW-AD [14] 95.5 98.4 93.6 94.8 36.2 45.6 66.4 29.5 75.8
PyramidalFlow [22] 90.3 94.1 94.2 51.8 14.7 06.3 17.1 03.3 52.6
Rec. RD [10] 99.9 100. 99.0 97.6 55.3 57.2 82.6 40.1 84.5
DiAD [16] 99.5 99.8 99.0 90.1 24.3 30.8 72.6 18.2 73.7
ViTAD [40] 98.3 99.3 96.9 96.5 55.0 55.2 69.2 38.2 81.5
InvAD [42] 99.9 100. 99.0 97.5 57.4 58.8 80.7 41.7 84.8
InvAD-lite [42] 98.2 99.4 96.9 97.1 56.7 59.5 75.4 42.3 83.3
MambaAD [15] 97.7 99.2 96.0 97.1 55.3 59.0 75.9 41.8 82.9
Hybrid UniAD [35] 99.8 99.9 99.0 97.4 55.7 57.7 80.1 40.6 84.2
RD++ [31] 99.4 99.8 97.9 97.5 53.5 56.9 80.6 39.8 83.7
DesTSeg [44] 98.3 99.4 96.0 91.9 29.9 43.9 69.9 28.1 75.6
2.0 Aug. DRAEM [37] 63.6 92.6 92.8 47.9 06.6 10.4 16.2 05.5 47.2
SimpleNet [27] 83.2 97.1 93.2 94.6 49.2 50.3 47.2 33.6 73.5
RealNet [45] 80.0 96.7 93.0 90.4 31.8 51.8 44.0 35.0 69.7
Emb. CFA [21] 80.3 96.5 93.2 94.4 47.5 51.2 47.6 34.5 73.0
PatchCore [29] 83.8 97.3 92.8 96.0 59.3 59.8 54.1 42.7 79.3
CFLOW-AD [14] 79.9 96.7 92.8 96.3 59.3 62.8 54.3 45.7 77.4
PyramidalFlow [22] 73.7 95.2 92.8 81.9 63.2 52.9 32.5 35.9 70.3
Rec. RD [10] 82.9 97.2 92.8 96.7 63.3 64.2 58.3 47.2 79.3
DiAD [16] 72.9 95.2 92.8 87.8 29.4 37.1 45.1 22.8 65.8
ViTAD [40] 84.0 97.3 93.2 96.7 72.9 67.2 51.2 50.6 80.4
InvAD [42] 86.4 97.7 93.8 96.4 60.2 61.1 57.1 44.0 79.0
InvAD-lite [42] 84.6 97.5 92.8 96.4 59.8 60.7 58.5 43.5 78.6
MambaAD [15] 81.6 97.0 92.8 96.1 54.2 59.2 57.6 42.0 76.9
Hybrid UniAD [35] 84.0 97.3 92.8 95.1 43.5 53.6 57.9 36.6 74.9
RD++ [31] 84.3 97.4 93.2 96.6 64.1 63.4 57.3 46.4 79.5
DesTSeg [44] 82.7 97.0 92.8 93.6 58.6 51.0 52.7 34.2 75.5
3.0 Aug. DRAEM [37] 69.2 49.0 59.1 50.4 00.3 01.7 19.2 00.9 35.6
SimpleNet [27] 99.7 96.9 90.6 99.4 35.6 38.1 97.2 23.5 79.6
RealNet [45] 99.3 92.2 84.5 76.9 48.6 51.1 45.3 34.4 71.1
Emb. CFA [21] 99.8 97.5 91.5 99.5 55.4 54.9 95.0 37.8 84.8
PatchCore [29] 99.8 97.3 92.1 99.5 50.1 48.2 98.0 31.8 85.2
CFLOW-AD [14] 99.3 91.8 82.7 99.4 42.5 38.5 97.5 23.9 78.8
PyramidalFlow [22] 92.0 55.9 50.0 94.8 05.2 11.2 79.7 05.9 55.5
Rec. RD [10] 99.6 93.3 89.6 99.7 52.7 52.6 98.9 35.7 83.8
DiAD [16] 98.3 69.9 86.1 97.9 07.7 13.1 93.3 07.0 66.6
ViTAD [40] 99.6 94.3 90.9 99.6 47.0 47.1 98.1 30.8 82.4
InvAD [42] 99.6 94.1 88.2 99.7 58.6 56.4 98.7 39.3 85.0
InvAD-lite [42] 99.8 96.8 93.5 99.6 50.7 51.0 98.2 34.2 84.2
MambaAD [15] 99.5 92.3 90.3 99.6 44.2 47.1 98.3 30.8 81.6
Hybrid UniAD [35] 99.8 97.8 93.1 99.7 57.8 55.1 98.7 38.0 86.0
RD++ [31] 99.7 95.6 91.8 99.8 55.4 54.3 98.8 37.3 85.1
DesTSeg [44] 99.6 93.9 92.5 98.9 28.8 20.6 96.1 11.5 75.8
Avg Aug. DRAEM [37] 69.3 76.8 79.5 47.7 03.2 06.1 16.1 03.2 42.7
SimpleNet [27] 94.0 97.9 93.9 96.2 41.0 43.7 69.6 28.1 76.6
RealNet [45] 93.1 96.3 92.5 87.2 48.0 55.5 57.9 38.7 75.8
Emb. CFA [21] 92.7 97.8 94.6 96.2 46.7 49.7 69.0 33.2 78.1
PatchCore [29] 94.4 98.2 94.6 97.5 55.0 54.9 76.0 38.0 83.3
CFLOW-AD [14] 91.6 95.6 89.7 96.9 46.0 49.0 72.7 33.0 77.4
PyramidalFlow [22] 85.3 81.7 79.0 76.1 27.7 23.4 43.1 15.0 59.5
Rec. RD [10] 94.1 96.8 93.8 98.0 57.1 58.0 79.9 41.0 82.5
DiAD [16] 90.2 88.3 92.6 91.9 20.5 27.0 70.3 16.0 68.7
ViTAD [40] 94.0 97.0 93.7 97.6 58.3 56.5 72.8 39.9 81.4
InvAD [42] 95.3 97.3 93.7 97.9 58.7 58.8 78.8 41.7 82.9
InvAD-lite [42] 94.2 97.9 94.4 97.7 55.7 57.0 77.4 40.0 82.0
MambaAD [15] 92.9 96.2 93.0 97.6 51.2 55.1 77.3 38.2 80.5
Hybrid UniAD [35] 94.5 98.4 94.9 97.4 52.4 55.5 78.9 38.4 81.7
RD++ [31] 94.5 97.6 94.3 98.0 57.7 58.2 78.9 41.2 82.7
DesTSeg [44] 93.5 96.7 93.8 94.8 39.1 38.5 72.9 24.6 75.6
Appendix T Detailed Quantitative Results on MPDD Dataset Under 300 epochs
Table A70: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bracket_black Aug. DRAEM [37] 45.3 54.3 77.0 52.8 00.2 00.5 24.9 00.3 36.4
SimpleNet [27] 78.7 85.3 80.4 94.9 03.6 09.1 89.3 04.7 63.0
RealNet [45] 72.0 84.5 76.4 64.4 09.5 16.7 45.9 09.1 52.8
Emb. CFA [21] 87.9 92.7 84.1 93.8 02.9 08.8 83.8 04.6 64.9
PatchCore [29] 88.4 92.7 83.9 98.5 11.0 21.0 96.7 11.8 72.6
CFLOW-AD [14] 64.8 74.6 76.7 95.8 02.5 05.4 86.5 02.8 58.0
PyramidalFlow [22] 75.3 81.8 79.3 89.5 00.4 01.0 66.2 00.5 56.2
Rec. RD [10] 83.7 87.0 85.7 97.6 05.9 12.3 93.3 06.5 66.5
DiAD [16] 100. 100. 100. 90.8 00.5 01.4 63.8 00.7 65.2
ViTAD [40] 78.7 86.1 81.5 96.0 02.7 07.8 89.8 04.1 63.2
InvAD [42] 88.2 93.2 83.8 98.1 08.2 15.0 95.5 08.1 68.9
InvAD-lite [42] 81.2 87.5 81.7 95.8 07.4 16.2 90.6 08.8 65.8
MambaAD [15] 82.8 88.6 80.7 94.8 04.6 13.1 89.1 07.0 64.8
Hybrid UniAD [35] 96.3 96.6 96.9 94.8 00.9 02.2 84.5 01.1 67.5
RD++ [31] 82.3 86.2 81.7 97.6 04.4 09.6 93.8 05.0 65.1
DesTSeg [44] 92.9 95.0 91.3 91.9 06.0 14.2 87.6 07.6 68.4
bracket_brown Aug. DRAEM [37] 53.9 69.1 80.3 54.4 00.7 01.3 28.8 00.6 41.2
SimpleNet [27] 92.4 95.3 93.5 94.9 08.8 17.0 87.4 09.3 69.9
RealNet [45] 87.4 90.4 92.6 51.4 10.1 03.5 17.8 01.8 50.5
Emb. CFA [21] 93.5 95.8 96.2 91.5 07.4 17.7 69.6 09.7 67.4
PatchCore [29] 95.6 97.2 95.3 98.6 26.2 32.0 95.4 19.0 79.5
CFLOW-AD [14] 77.8 81.6 87.7 96.4 07.2 14.0 92.5 07.5 65.3
PyramidalFlow [22] 70.0 78.6 84.4 93.9 05.5 12.6 75.4 06.8 60.1
Rec. RD [10] 83.1 87.0 91.7 96.0 14.2 23.2 94.0 13.1 69.9
DiAD [16] 88.6 94.2 89.5 94.8 22.5 29.9 69.8 17.6 69.9
ViTAD [40] 86.0 87.8 91.9 98.3 16.2 26.2 95.0 15.1 71.6
InvAD [42] 96.8 98.0 98.1 96.3 18.7 29.4 91.4 17.2 75.5
InvAD-lite [42] 93.4 96.2 92.7 98.5 21.4 31.6 95.1 18.8 75.6
MambaAD [15] 97.2 98.5 97.1 97.8 19.6 27.3 92.8 15.8 75.8
Hybrid UniAD [35] 93.1 95.8 93.5 98.6 39.3 47.0 90.7 30.8 79.7
RD++ [31] 85.7 90.5 92.7 96.8 19.8 27.7 95.1 16.1 72.6
DesTSeg [44] 91.4 92.4 96.2 78.9 05.1 08.8 66.9 04.6 62.8
bracket_white Aug. DRAEM [37] 69.4 70.9 74.0 66.1 00.1 00.2 23.6 00.1 43.5
SimpleNet [27] 88.4 91.4 83.0 97.8 02.2 05.6 86.4 02.9 65.0
RealNet [45] 72.0 82.3 78.4 81.4 30.1 39.6 40.1 24.7 60.6
Emb. CFA [21] 89.2 91.7 81.5 94.9 00.8 02.9 72.4 01.4 61.9
PatchCore [29] 93.3 94.8 88.9 98.8 10.0 19.3 93.9 10.7 73.9
CFLOW-AD [14] 75.3 76.4 73.0 98.6 02.1 04.6 93.8 02.4 60.5
PyramidalFlow [22] 76.3 78.5 76.9 99.0 06.3 14.2 90.1 07.7 63.0
Rec. RD [10] 85.0 86.9 80.6 98.9 01.7 05.0 94.4 02.5 64.6
DiAD [16] 81.0 82.5 75.4 94.2 00.2 00.6 80.0 00.3 59.1
ViTAD [40] 79.4 85.4 77.2 95.4 00.8 02.6 86.6 01.3 61.1
InvAD [42] 95.8 96.9 93.1 99.3 10.7 20.9 94.7 11.7 73.1
InvAD-lite [42] 93.6 95.1 87.3 99.3 06.8 10.7 94.9 05.6 69.7
MambaAD [15] 93.6 95.1 89.7 99.3 08.9 18.9 95.0 10.5 71.5
Hybrid UniAD [35] 82.4 80.5 79.3 95.9 00.8 02.8 82.0 01.4 60.5
RD++ [31] 83.4 86.4 76.2 98.7 01.6 04.3 93.4 02.2 63.4
DesTSeg [44] 96.3 93.5 94.9 91.6 00.9 02.2 52.6 01.1 61.7
connector Aug. DRAEM [37] 63.3 47.7 53.1 30.3 00.5 01.5 12.2 00.8 29.8
SimpleNet [27] 99.3 98.5 96.6 98.9 56.0 55.2 96.5 38.2 85.9
RealNet [45] 93.3 86.8 85.7 70.7 47.9 42.9 48.0 27.3 67.9
Emb. CFA [21] 95.7 86.4 93.3 97.1 31.9 39.9 90.8 25.0 76.4
PatchCore [29] 100. 100. 100. 99.5 71.2 63.3 98.4 46.3 91.5
CFLOW-AD [14] 96.7 92.9 89.7 97.4 33.1 33.2 91.5 19.9 76.4
PyramidalFlow [22] 58.3 38.3 54.5 91.3 04.4 08.5 73.7 04.4 47.0
Rec. RD [10] 99.8 99.5 96.6 99.1 58.3 54.5 97.2 37.5 86.4
DiAD [16] 88.3 81.5 74.3 96.2 11.6 22.0 87.7 12.4 66.0
ViTAD [40] 96.0 91.7 86.7 99.3 39.8 46.9 97.6 30.6 79.7
InvAD [42] 98.3 95.7 96.6 99.2 62.3 57.6 97.3 40.5 86.7
InvAD-lite [42] 96.7 83.4 96.6 99.4 61.5 57.4 97.9 40.2 84.7
MambaAD [15] 98.8 97.9 92.9 99.2 55.4 53.4 97.5 36.4 85.0
Hybrid UniAD [35] 82.9 65.0 72.7 97.6 14.0 24.0 92.1 13.7 64.0
RD++ [31] 100. 100. 100. 99.3 66.3 59.8 97.7 42.7 89.0
DesTSeg [44] 90.7 77.3 87.5 89.1 31.4 38.3 81.2 23.7 70.8
Table A71: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
metal_plate Aug. DRAEM [37] 71.7 89.7 84.5 57.5 30.0 21.2 25.7 11.8 54.3
SimpleNet [27] 100. 100. 100. 98.2 88.9 81.3 88.1 68.5 93.8
RealNet [45] 100. 100. 100. 90.2 84.6 80.4 80.7 67.2 90.8
Emb. CFA [21] 100. 100. 100. 98.8 92.6 85.3 90.8 74.4 95.4
PatchCore [29] 100. 100. 100. 99.0 93.5 86.1 94.1 75.7 96.6
CFLOW-AD [14] 100. 100. 100. 98.6 89.7 83.7 92.9 72.0 95.0
PyramidalFlow [22] 78.3 92.4 87.5 68.0 49.7 40.2 29.1 25.1 63.6
Rec. RD [10] 100. 100. 100. 99.1 92.6 87.4 96.6 77.6 96.5
DiAD [16] 98.3 99.5 98.6 87.2 54.3 55.1 45.0 38.0 76.9
ViTAD [40] 100. 100. 100. 99.3 95.5 88.2 94.8 78.9 96.8
InvAD [42] 100. 100. 100. 99.0 91.6 86.6 95.7 76.3 96.1
InvAD-lite [42] 100. 100. 100. 98.6 88.9 84.5 94.5 73.2 95.2
MambaAD [15] 99.8 99.9 99.3 98.5 86.2 84.3 94.7 72.8 94.7
Hybrid UniAD [35] 58.8 79.6 85.0 93.3 50.4 62.8 80.8 45.8 73.0
RD++ [31] 99.9 100. 99.3 99.0 91.1 87.0 95.5 77.0 96.0
DesTSeg [44] 100. 100. 100. 96.9 93.8 87.2 91.4 77.2 95.6
tubes Aug. DRAEM [37] 57.4 73.4 81.7 40.9 01.0 03.5 15.6 01.8 39.1
SimpleNet [27] 84.6 94.1 84.6 97.9 42.4 46.1 92.6 29.9 77.5
RealNet [45] 91.3 95.9 90.6 90.3 53.0 55.2 81.2 38.1 79.6
Emb. CFA [21] 85.6 94.4 85.9 97.8 38.1 43.7 91.7 27.9 76.7
PatchCore [29] 90.4 96.6 92.4 98.8 64.5 63.8 95.5 46.9 86.9
CFLOW-AD [14] 68.0 85.0 81.2 96.9 35.7 39.6 88.2 24.7 70.7
PyramidalFlow [22] 76.8 87.4 83.2 92.5 08.4 18.0 76.8 09.9 63.3
Rec. RD [10] 90.4 96.2 88.5 98.9 64.7 61.2 95.7 44.1 85.1
DiAD [16] 58.7 77.3 81.2 85.6 02.8 05.9 50.3 03.0 51.7
ViTAD [40] 84.3 93.6 84.5 98.5 52.5 55.0 93.6 38.0 80.3
InvAD [42] 96.6 98.7 95.5 99.3 73.1 69.0 97.4 52.7 89.9
InvAD-lite [42] 91.6 96.8 90.1 98.4 62.6 63.0 94.3 45.9 85.3
MambaAD [15] 62.9 78.6 81.9 96.5 26.7 34.5 87.4 20.8 66.9
Hybrid UniAD [35] 67.3 81.6 83.3 92.1 08.3 14.7 72.7 07.9 60.0
RD++ [31] 92.1 96.9 90.0 99.2 73.6 69.7 96.6 53.5 88.3
DesTSeg [44] 84.0 92.3 87.2 96.5 46.3 46.7 89.9 30.4 77.6
Avg Aug. DRAEM [37] 60.2 67.5 75.1 50.4 05.4 04.7 21.8 02.6 40.7
SimpleNet [27] 90.6 94.1 89.7 97.1 33.6 35.7 90.0 25.6 75.8
RealNet [45] 86.0 90.0 87.3 74.7 39.2 39.7 52.3 28.0 67.0
Emb. CFA [21] 92.0 93.5 90.2 95.7 29.0 33.0 83.2 23.8 73.8
PatchCore [29] 94.6 96.9 93.4 98.9 46.1 47.6 95.7 35.0 83.5
CFLOW-AD [14] 80.4 85.1 84.7 97.3 28.4 30.1 90.9 21.5 71.0
PyramidalFlow [22] 72.5 76.2 77.6 89.0 12.4 15.8 68.5 09.1 58.9
Rec. RD [10] 90.3 92.8 90.5 98.3 39.6 40.6 95.2 30.2 78.2
DiAD [16] 85.8 89.2 86.5 91.4 15.3 19.2 66.1 12.0 64.8
ViTAD [40] 87.4 90.8 87.0 97.8 34.6 37.8 92.9 28.0 75.5
InvAD [42] 96.0 97.1 94.5 98.5 44.1 46.4 95.3 34.4 81.7
InvAD-lite [42] 92.8 93.2 91.4 98.3 41.4 43.9 94.6 32.1 79.4
MambaAD [15] 89.2 93.1 90.3 97.7 33.5 38.6 92.8 27.2 76.5
Hybrid UniAD [35] 80.1 83.2 85.1 95.4 19.0 25.6 83.8 16.8 67.5
RD++ [31] 90.6 93.3 90.0 98.4 42.8 43.0 95.3 32.7 79.1
DesTSeg [44] 92.6 91.8 92.8 90.8 30.6 32.9 78.3 24.1 72.8
Appendix U Detailed Quantitative Results on MAD_Real Dataset Under 300 epochs
Table A72: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Bear Aug. DRAEM [37] 62.5 91.6 90.6 47.4 00.3 01.8 25.3 00.9 45.6
SimpleNet [27] 17.5 69.8 90.6 74.6 01.2 04.9 56.3 02.5 45.0
Emb. CFA [21] 13.3 68.8 90.6 99.0 03.9 09.6 96.2 05.1 54.5
PatchCore [29] 84.2 96.8 91.7 99.6 17.6 29.5 98.6 17.3 76.1
CFLOW-AD [14] 77.5 94.6 92.3 99.4 06.3 12.2 97.7 06.5 68.6
PyramidalFlow [22] 07.5 67.7 90.6 98.5 02.8 05.3 94.9 02.7 52.5
Rec. RD [10] 86.7 96.4 96.0 97.7 06.9 10.5 92.9 05.5 69.6
DiAD [16] 85.0 97.0 90.6 97.5 01.8 05.4 92.0 02.8 67.0
ViTAD [40] 42.5 79.5 90.6 96.3 03.1 04.4 88.0 02.2 57.8
InvAD [42] 80.8 96.0 92.0 99.6 15.9 25.0 98.5 14.3 72.5
InvAD-lite [42] 47.5 82.3 90.6 97.8 04.2 05.7 93.3 02.9 60.2
MambaAD [15] 58.3 87.2 90.6 99.2 11.2 18.8 97.1 10.4 66.1
Hybrid UniAD [35] 78.3 95.5 92.3 99.0 05.3 12.8 96.4 06.8 68.5
RD++ [31] 74.2 93.9 92.3 99.0 08.1 15.0 96.6 08.1 68.4
DesTSeg [44] 22.5 75.9 90.6 91.4 00.5 01.2 71.4 00.6 50.5
Bird Aug. DRAEM [37] 23.6 74.8 89.8 58.7 00.7 01.4 20.0 00.7 38.4
SimpleNet [27] 60.0 86.9 89.8 58.3 00.6 01.4 20.8 00.7 45.4
Emb. CFA [21] 54.5 83.0 89.8 86.7 04.7 10.9 57.1 05.8 55.2
PatchCore [29] 71.8 92.6 93.6 96.4 32.2 39.7 84.2 24.8 74.6
CFLOW-AD [14] 38.2 80.2 89.8 89.5 02.9 06.3 64.2 03.2 53.0
PyramidalFlow [22] 15.5 68.6 89.8 90.7 02.7 06.4 66.5 03.3 48.6
Rec. RD [10] 23.6 72.4 89.8 87.0 06.8 16.3 56.7 08.9 50.4
DiAD [16] 90.9 98.3 95.2 94.5 19.5 28.0 77.3 16.3 72.0
ViTAD [40] 62.7 87.5 91.7 89.9 08.3 16.9 61.9 09.2 59.8
InvAD [42] 89.1 97.6 91.7 93.4 25.0 37.3 75.4 22.9 72.8
InvAD-lite [42] 84.5 96.4 91.7 96.2 23.1 29.3 83.5 17.2 72.1
MambaAD [15] 77.3 93.1 95.7 95.0 11.3 21.4 78.7 12.0 67.5
Hybrid UniAD [35] 41.8 78.2 89.8 86.9 02.6 06.3 65.7 03.2 53.0
RD++ [31] 50.9 86.6 89.8 90.0 08.2 17.8 64.4 09.8 58.2
DesTSeg [44] 72.7 93.3 91.7 86.2 10.5 15.2 59.3 08.2 61.3
Elephant Aug. DRAEM [37] 62.2 87.9 87.8 47.6 00.3 00.7 13.5 00.4 42.9
SimpleNet [27] 05.6 61.1 87.8 64.1 00.8 02.8 24.7 01.4 35.3
Emb. CFA [21] 33.3 70.9 87.8 87.5 06.0 11.0 59.6 05.8 50.9
PatchCore [29] 27.8 69.5 87.8 93.7 11.1 18.3 78.8 10.1 56.1
CFLOW-AD [14] 30.0 74.7 87.8 90.9 07.0 11.4 67.3 06.1 52.7
PyramidalFlow [22] 60.0 86.0 87.8 92.5 02.7 05.4 75.1 02.8 58.5
Rec. RD [10] 48.9 84.7 87.8 88.4 02.1 05.2 63.2 02.7 54.3
DiAD [16] 20.0 64.3 90.0 91.8 02.0 04.4 73.1 02.2 49.4
ViTAD [40] 43.3 79.9 87.8 95.5 09.8 15.9 83.6 08.7 59.4
InvAD [42] 35.6 72.2 87.8 92.9 05.0 11.3 76.6 06.0 54.5
InvAD-lite [42] 35.6 78.4 87.8 91.6 04.9 11.3 73.7 06.0 54.8
MambaAD [15] 24.4 67.4 87.8 89.2 02.3 05.6 70.6 02.9 49.6
Hybrid UniAD [35] 13.3 63.1 87.8 78.4 01.6 05.6 38.8 02.9 41.2
RD++ [31] 40.0 77.1 87.8 87.5 02.1 05.0 63.9 02.6 51.9
DesTSeg [44] 12.2 64.6 87.8 68.7 00.8 02.3 36.5 01.2 39.0
Parrot Aug. DRAEM [37] 30.4 76.4 92.0 49.3 00.6 01.3 13.5 00.6 37.6
SimpleNet [27] 24.3 73.9 90.2 58.3 00.8 01.8 35.1 00.9 40.6
Emb. CFA [21] 05.2 66.2 90.2 42.0 00.5 01.3 16.9 00.7 31.8
PatchCore [29] 05.2 66.2 90.2 81.7 02.1 05.1 64.0 02.6 46.0
CFLOW-AD [14] 15.7 67.9 90.2 66.7 01.0 02.5 41.7 01.2 40.8
PyramidalFlow [22] 09.6 69.4 90.2 67.7 01.0 02.2 32.4 01.1 38.9
Rec. RD [10] 05.2 65.5 90.2 64.5 00.9 02.1 37.0 01.1 37.9
DiAD [16] 00.0 64.3 90.2 56.3 00.7 01.7 37.9 00.9 35.9
ViTAD [40] 10.4 67.8 90.2 70.3 01.1 02.6 40.7 01.3 40.4
InvAD [42] 05.2 65.7 90.2 68.7 01.1 02.4 44.6 01.2 39.7
InvAD-lite [42] 03.5 65.1 90.2 71.4 01.1 02.5 42.2 01.3 39.4
MambaAD [15] 04.3 65.4 90.2 72.5 01.1 02.5 46.1 01.3 40.3
Hybrid UniAD [35] 10.4 67.2 90.2 61.5 00.8 01.9 38.0 00.9 38.6
RD++ [31] 02.6 64.8 90.2 64.2 00.9 02.0 35.7 01.0 37.2
DesTSeg [44] 07.8 67.0 90.2 66.4 01.1 02.5 30.5 01.3 37.9
Table A73: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Pig Aug. DRAEM [37] 43.5 72.6 89.5 61.9 00.6 00.8 26.9 00.4 42.3
SimpleNet [27] 14.1 63.4 87.2 76.6 01.7 06.1 37.0 03.2 40.9
Emb. CFA [21] 56.5 78.2 89.5 93.0 05.0 11.3 72.6 06.0 58.0
PatchCore [29] 56.5 78.8 89.5 98.5 19.9 28.7 89.7 16.7 67.1
CFLOW-AD [14] 57.6 78.4 87.2 96.8 07.1 12.7 86.1 06.8 60.8
PyramidalFlow [22] 50.6 80.7 89.5 96.4 06.8 13.4 83.4 07.2 60.1
Rec. RD [10] 60.0 80.1 87.2 98.0 20.0 29.6 90.2 17.4 66.4
DiAD [16] 57.6 79.8 87.2 96.4 09.0 18.9 81.6 10.4 61.5
ViTAD [40] 44.7 75.3 87.2 97.6 09.2 19.4 88.0 10.7 60.2
InvAD [42] 64.7 81.2 88.9 98.0 11.2 24.2 88.1 13.8 65.2
InvAD-lite [42] 57.6 78.1 88.2 98.6 17.5 31.3 93.2 18.5 66.4
MambaAD [15] 42.4 74.5 87.2 97.3 08.5 18.6 87.3 10.2 59.4
Hybrid UniAD [35] 48.2 76.9 89.5 89.5 02.1 04.9 61.6 02.5 53.2
RD++ [31] 56.5 79.2 89.5 98.0 13.8 23.6 91.2 13.3 64.5
DesTSeg [44] 45.9 75.2 87.2 84.8 01.4 03.3 50.5 01.7 49.8
Puppy Aug. DRAEM [37] 57.0 87.0 88.9 31.2 00.2 00.7 13.7 00.3 39.8
SimpleNet [27] 52.0 82.2 88.9 90.8 01.7 04.4 65.1 02.3 55.0
Emb. CFA [21] 82.0 94.8 93.0 98.5 06.7 14.0 92.0 07.5 68.7
PatchCore [29] 100. 100. 100. 99.2 13.4 25.1 92.1 14.3 78.7
CFLOW-AD [14] 70.0 91.3 93.0 99.2 31.8 34.5 93.1 20.9 73.3
PyramidalFlow [22] 68.0 91.2 88.9 95.9 11.7 20.7 77.6 11.5 64.9
Rec. RD [10] 82.0 95.6 88.9 95.7 05.5 08.5 83.8 04.4 65.7
DiAD [16] 91.0 97.7 95.2 95.8 03.4 09.1 80.8 04.8 67.6
ViTAD [40] 65.0 90.2 88.9 90.5 01.0 02.6 69.6 01.3 58.3
InvAD [42] 100. 100. 100. 99.4 28.6 31.3 93.7 18.5 79.0
InvAD-lite [42] 97.0 99.3 97.4 95.9 03.4 09.2 82.8 04.8 69.3
MambaAD [15] 92.0 97.8 95.2 98.7 26.1 36.1 94.2 22.0 77.2
Hybrid UniAD [35] 78.0 94.7 88.9 98.7 12.0 20.2 92.8 11.2 69.3
RD++ [31] 65.0 90.3 90.9 97.3 04.0 07.5 90.5 03.9 63.6
DesTSeg [44] 54.0 82.7 88.9 93.1 01.5 03.9 66.6 02.0 55.8
Scorpion Aug. DRAEM [37] 46.1 83.9 90.2 41.4 01.2 02.9 10.8 01.5 39.5
SimpleNet [27] 47.0 84.1 90.2 83.5 09.0 15.8 50.3 08.6 54.3
Emb. CFA [21] 27.8 80.3 90.2 61.0 05.4 13.1 26.3 07.0 43.4
PatchCore [29] 13.0 70.7 90.2 88.2 08.8 17.0 61.6 09.3 50.9
CFLOW-AD [14] 54.8 87.2 90.2 77.9 04.2 07.5 34.0 03.9 50.8
PyramidalFlow [22] 47.0 84.3 90.2 65.5 02.6 05.8 20.3 03.0 45.1
Rec. RD [10] 55.7 87.9 90.2 66.3 02.1 04.9 21.8 02.5 47.0
DiAD [16] 11.3 71.8 90.2 71.9 03.1 06.8 30.7 03.5 40.8
ViTAD [40] 29.6 79.5 90.2 62.7 01.9 04.2 16.8 02.1 40.7
InvAD [42] 31.3 80.4 90.2 74.4 03.1 06.5 30.8 03.4 45.2
InvAD-lite [42] 32.2 79.2 90.2 74.6 03.0 06.0 28.1 03.1 44.8
MambaAD [15] 42.6 80.8 90.2 75.0 02.9 06.3 27.1 03.2 46.4
Hybrid UniAD [35] 27.0 73.9 90.2 78.6 04.2 09.1 41.2 04.8 46.3
RD++ [31] 25.2 75.8 90.2 70.7 02.5 05.8 24.6 03.0 42.1
DesTSeg [44] 27.8 77.9 90.2 68.1 03.2 08.0 25.3 04.2 42.9
Turtle Aug. DRAEM [37] 58.1 85.3 91.3 24.4 00.3 01.5 09.2 00.8 38.6
SimpleNet [27] 63.8 87.9 89.4 94.6 07.0 15.1 81.8 08.1 62.8
Emb. CFA [21] 100. 100. 100. 99.3 41.5 42.7 96.8 27.2 82.9
PatchCore [29] 100. 100. 100. 99.5 43.1 48.4 97.9 31.9 86.1
CFLOW-AD [14] 99.0 99.8 97.7 99.1 28.4 38.0 96.5 23.5 79.8
PyramidalFlow [22] 93.3 98.5 93.3 98.4 23.9 33.3 95.3 20.0 76.6
Rec. RD [10] 95.2 99.0 95.2 98.5 23.2 30.4 94.9 18.0 76.6
DiAD [16] 74.3 94.2 89.4 93.2 07.5 13.5 83.5 07.2 65.1
ViTAD [40] 83.8 96.1 93.3 95.6 06.2 13.2 85.6 07.0 67.7
InvAD [42] 100. 100. 100. 99.7 51.9 50.0 98.4 33.3 85.7
InvAD-lite [42] 99.0 99.8 97.7 98.0 15.5 21.6 94.5 12.1 75.2
MambaAD [15] 100. 100. 100. 99.3 32.8 39.2 97.2 24.4 81.2
Hybrid UniAD [35] 98.1 99.6 97.6 98.2 22.8 31.5 91.0 18.7 77.0
RD++ [31] 100. 100. 100. 99.2 25.1 35.8 97.0 21.8 79.6
DesTSeg [44] 54.3 86.7 89.4 95.2 09.8 16.1 77.0 08.8 61.2
Table A74: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Unicorn Aug. DRAEM [37] 36.2 79.8 89.4 43.0 00.2 00.5 09.9 00.2 37.0
SimpleNet [27] 19.0 68.8 89.4 81.8 00.9 02.2 49.5 01.1 44.5
Emb. CFA [21] 31.4 73.8 89.4 67.8 00.8 03.6 33.2 01.9 42.9
PatchCore [29] 43.8 74.1 93.3 97.3 06.8 14.1 90.3 07.6 61.3
CFLOW-AD [14] 63.8 90.9 91.3 90.1 01.5 03.3 65.7 01.7 58.1
PyramidalFlow [22] 65.7 84.4 89.4 82.2 00.9 02.2 51.8 01.1 53.8
Rec. RD [10] 21.0 71.0 89.4 73.5 00.5 01.3 33.8 00.6 41.5
DiAD [16] 62.9 91.4 89.4 86.0 01.0 02.4 60.4 01.2 56.2
ViTAD [40] 20.0 72.2 89.4 73.9 00.5 01.1 31.7 00.6 41.3
InvAD [42] 31.4 72.8 89.4 89.9 01.6 04.3 67.3 02.2 51.0
InvAD-lite [42] 32.4 72.3 89.4 84.4 01.4 04.3 57.0 02.2 48.7
MambaAD [15] 18.1 67.6 89.4 81.7 00.8 02.1 50.0 01.1 44.2
Hybrid UniAD [35] 16.2 66.6 89.4 89.3 01.4 03.7 66.1 01.9 47.5
RD++ [31] 29.5 75.7 89.4 81.4 00.7 01.5 43.7 00.8 46.0
DesTSeg [44] 21.0 69.4 89.4 76.4 00.5 01.3 31.0 00.6 41.3
Whale Aug. DRAEM [37] 43.1 85.6 92.8 20.2 00.1 00.4 06.4 00.2 35.5
SimpleNet [27] 58.1 91.5 92.8 89.1 02.5 08.2 68.5 04.3 58.7
Emb. CFA [21] 80.6 96.2 93.5 98.6 29.3 32.2 90.5 19.2 74.4
PatchCore [29] 95.0 99.1 98.5 99.7 46.8 47.4 97.8 31.0 85.2
CFLOW-AD [14] 67.5 94.4 92.8 99.1 17.5 24.7 95.1 14.1 70.2
PyramidalFlow [22] 89.4 98.2 95.5 98.0 20.2 25.6 90.6 14.7 73.9
Rec. RD [10] 81.2 96.7 92.8 98.9 19.2 23.5 94.9 13.3 72.5
DiAD [16] 86.9 98.0 94.1 94.6 02.9 05.4 77.2 02.8 65.6
ViTAD [40] 42.5 85.4 92.8 96.3 03.6 08.5 83.6 04.5 59.0
InvAD [42] 96.9 99.5 97.0 99.8 45.3 50.1 98.7 33.4 83.9
InvAD-lite [42] 83.8 96.6 95.5 99.5 39.4 41.9 98.2 26.5 79.3
MambaAD [15] 87.5 97.6 95.5 99.7 33.8 41.7 98.6 26.3 79.2
Hybrid UniAD [35] 42.5 87.3 92.8 98.1 11.1 19.6 91.1 10.9 63.2
RD++ [31] 79.4 95.9 92.8 99.2 27.9 29.8 96.1 17.5 74.4
DesTSeg [44] 42.5 86.4 92.8 95.1 02.6 06.0 79.3 03.1 57.8
Avg Aug. DRAEM [37] 46.3 82.5 90.2 42.5 00.5 01.2 14.9 00.6 39.7
SimpleNet [27] 36.1 77.0 89.6 77.2 02.6 06.3 48.9 03.3 48.2
Emb. CFA [21] 48.5 81.2 91.4 83.3 10.4 15.0 64.1 08.6 56.3
PatchCore [29] 59.7 84.8 93.5 95.4 20.2 27.3 85.5 16.6 68.2
CFLOW-AD [14] 57.4 85.9 91.2 90.9 10.8 15.3 74.1 08.8 60.8
PyramidalFlow [22] 50.6 82.9 90.5 88.6 07.5 12.0 68.8 06.7 57.3
Rec. RD [10] 56.0 84.9 90.7 86.9 08.7 13.2 66.9 07.4 58.2
DiAD [16] 58.0 85.7 91.1 87.8 05.1 09.6 69.4 05.2 58.1
ViTAD [40] 44.5 81.3 90.2 86.9 04.5 08.9 65.0 04.8 54.5
InvAD [42] 63.5 86.5 92.7 91.6 18.9 24.2 77.2 14.9 64.9
InvAD-lite [42] 57.3 84.8 91.9 90.8 11.4 16.3 74.7 09.5 61.0
MambaAD [15] 54.7 83.1 92.2 90.8 13.1 19.2 74.7 11.4 61.1
Hybrid UniAD [35] 45.4 80.3 90.8 87.8 06.4 11.6 68.3 06.4 55.8
RD++ [31] 52.3 84.0 91.3 88.7 09.3 14.4 70.4 08.2 58.6
DesTSeg [44] 36.1 77.9 89.8 82.5 03.2 06.0 52.7 03.2 49.7
Appendix V Detailed Quantitative Results on MAD_Sim Dataset Under 300 epochs
Table A75: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
01Gorilla Aug. DRAEM [37] 54.4 91.6 95.1 45.4 00.7 02.0 19.4 01.0 44.1
SimpleNet [27] 53.0 91.5 94.9 90.9 02.7 05.7 65.5 02.9 57.7
Emb. CFA [21] 55.9 93.6 94.9 79.0 07.2 15.4 55.8 08.3 57.4
CFLOW-AD [14] 54.4 92.8 94.9 90.8 04.8 05.1 65.4 02.6 58.3
PyramidalFlow [22] 46.5 90.4 94.9 91.5 03.7 06.0 75.1 03.1 58.3
Rec. RD [10] 62.7 93.3 95.1 93.1 04.5 09.0 76.4 04.7 62.0
DiAD [16] 64.4 94.6 94.9 91.6 03.6 08.3 64.6 04.3 60.3
ViTAD [40] 66.1 93.9 95.1 93.7 05.9 10.6 76.3 05.6 63.1
InvAD [42] 67.4 95.7 94.9 95.2 17.1 24.1 82.0 13.7 68.1
InvAD-lite [42] 70.6 95.5 94.9 94.1 08.0 15.1 78.3 08.2 65.2
MambaAD [15] 65.2 93.8 94.9 92.8 05.0 10.7 75.6 05.7 62.6
Hybrid UniAD [35] 52.6 92.3 94.9 92.7 06.3 11.0 79.7 05.8 61.4
RD++ [31] 53.6 91.1 94.9 92.7 04.0 08.3 74.6 04.3 59.9
DesTSeg [44] 46.5 90.3 94.9 58.2 05.0 09.0 34.4 04.7 48.3
02Unicorn Aug. DRAEM [37] 41.9 90.6 95.2 47.3 00.4 00.8 14.1 00.4 41.5
SimpleNet [27] 55.0 91.6 95.2 83.2 01.3 02.6 61.3 01.3 55.7
Emb. CFA [21] 63.8 94.8 95.5 67.8 01.4 03.9 32.1 02.0 51.3
CFLOW-AD [14] 59.0 93.5 95.2 84.4 02.3 03.4 55.4 01.7 56.2
PyramidalFlow [22] 37.0 87.7 95.5 80.3 01.1 02.3 52.2 01.2 50.9
Rec. RD [10] 61.4 94.1 95.2 85.3 02.0 04.1 70.7 02.1 59.0
DiAD [16] 79.3 97.2 95.4 80.0 01.6 04.5 48.0 02.3 58.0
ViTAD [40] 63.9 94.7 95.2 86.5 02.4 05.8 67.8 03.0 59.5
InvAD [42] 79.7 97.6 95.4 87.3 03.8 09.8 74.6 05.2 64.0
InvAD-lite [42] 71.3 96.0 95.2 86.6 02.9 06.4 73.2 03.3 61.7
MambaAD [15] 65.7 95.2 95.4 84.3 02.1 04.7 68.7 02.4 59.4
Hybrid UniAD [35] 71.4 96.2 95.4 87.1 02.5 06.3 64.5 03.3 60.5
RD++ [31] 60.6 93.8 95.2 85.1 02.3 04.6 69.2 02.3 58.7
DesTSeg [44] 72.0 96.1 95.5 55.9 06.4 10.5 40.0 05.6 53.8
03Mallard Aug. DRAEM [37] 36.6 88.0 95.4 46.6 01.3 01.5 15.3 00.8 40.7
SimpleNet [27] 55.2 91.6 95.4 85.5 02.6 05.6 58.6 02.9 56.4
Emb. CFA [21] 31.3 87.1 95.4 63.8 02.9 10.0 36.7 05.3 46.7
CFLOW-AD [14] 40.1 90.5 95.4 84.8 03.8 05.1 53.1 02.6 53.3
PyramidalFlow [22] 44.2 89.9 95.4 83.9 02.5 05.0 56.3 02.6 53.9
Rec. RD [10] 44.4 89.1 95.4 83.5 03.4 07.9 56.7 04.1 54.3
DiAD [16] 73.9 96.0 95.4 83.0 02.4 04.6 48.7 02.4 57.7
ViTAD [40] 46.6 91.4 95.4 85.2 05.3 11.2 60.4 05.9 56.5
InvAD [42] 46.5 91.3 95.4 85.9 07.8 13.4 61.5 07.2 57.4
InvAD-lite [42] 40.9 90.2 95.4 85.2 06.0 11.1 60.2 05.9 55.6
MambaAD [15] 55.2 93.6 95.4 84.2 04.9 10.5 57.9 05.6 57.4
Hybrid UniAD [35] 33.9 87.0 95.4 86.1 03.8 08.9 61.8 04.7 53.8
RD++ [31] 37.1 87.9 95.4 84.0 03.4 07.8 57.1 04.1 53.2
DesTSeg [44] 33.2 88.1 95.4 58.7 03.4 08.0 39.0 04.1 46.5
04Turtle Aug. DRAEM [37] 45.6 87.8 94.9 47.2 00.6 01.3 12.0 00.7 41.3
SimpleNet [27] 48.7 90.2 94.9 85.7 02.3 04.7 60.9 02.4 55.3
Emb. CFA [21] 40.8 90.8 94.9 71.1 02.4 03.6 38.6 01.8 48.9
CFLOW-AD [14] 63.1 94.7 94.9 86.3 04.1 06.3 57.1 03.2 58.1
PyramidalFlow [22] 16.8 81.8 94.9 76.6 01.2 03.1 39.7 01.6 44.9
Rec. RD [10] 64.1 94.1 94.9 91.1 09.3 16.8 72.6 09.2 63.3
DiAD [16] 85.5 98.3 96.3 91.8 07.1 14.3 63.1 07.7 65.2
ViTAD [40] 59.0 93.5 94.9 91.0 07.9 15.4 70.5 08.3 61.7
InvAD [42] 83.0 97.7 95.1 93.4 15.5 25.9 78.8 14.9 69.9
InvAD-lite [42] 76.8 96.9 95.6 92.3 12.4 21.3 75.4 11.9 67.2
MambaAD [15] 66.6 95.2 94.9 91.0 09.3 17.7 70.8 09.7 63.6
Hybrid UniAD [35] 55.7 92.0 94.9 88.9 04.3 07.6 70.9 04.0 59.2
RD++ [31] 60.1 94.2 94.9 90.2 07.0 13.2 69.8 07.1 61.3
DesTSeg [44] 34.1 88.5 94.9 55.0 01.6 03.6 38.8 01.8 45.2
Table A76: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
05Whale Aug. DRAEM [37] 50.5 90.3 94.4 47.6 00.6 01.5 20.6 00.8 43.6
SimpleNet [27] 50.4 89.2 94.4 85.1 01.8 03.7 62.4 01.9 55.3
Emb. CFA [21] 51.6 91.5 94.4 71.3 02.2 04.1 41.0 02.1 50.9
CFLOW-AD [14] 54.0 91.1 94.4 85.3 02.8 04.4 63.2 02.2 56.5
PyramidalFlow [22] 27.7 82.7 94.4 78.9 01.2 02.8 52.5 01.4 48.6
Rec. RD [10] 52.4 91.0 94.4 86.8 03.3 07.1 74.0 03.7 58.4
DiAD [16] 51.0 90.8 94.4 81.1 02.6 07.5 51.9 03.9 54.2
ViTAD [40] 58.5 92.7 94.4 88.3 04.1 08.3 76.5 04.3 60.4
InvAD [42] 74.0 96.3 94.4 89.9 08.8 16.9 81.1 09.2 65.9
InvAD-lite [42] 68.1 95.2 94.4 89.3 06.2 11.2 78.5 05.9 63.3
MambaAD [15] 60.4 93.5 94.4 87.3 04.4 08.9 74.1 04.6 60.4
Hybrid UniAD [35] 79.4 97.1 94.4 90.8 05.2 11.6 80.1 06.2 65.5
RD++ [31] 54.1 91.8 94.4 86.5 03.5 07.0 73.0 03.6 58.6
DesTSeg [44] 57.1 92.1 94.4 61.4 03.2 08.3 55.5 04.3 53.1
06Bird Aug. DRAEM [37] 44.1 87.6 94.5 48.6 01.0 02.3 27.2 01.1 43.6
SimpleNet [27] 52.9 91.4 94.4 91.8 05.1 09.4 75.9 04.9 60.1
Emb. CFA [21] 65.1 92.8 94.5 73.3 04.0 09.7 47.2 05.1 55.2
CFLOW-AD [14] 63.8 91.8 94.4 90.4 03.8 07.1 70.8 03.7 60.3
PyramidalFlow [22] 77.2 96.3 94.4 92.7 05.9 11.1 77.8 05.9 65.1
Rec. RD [10] 60.4 92.0 94.4 90.6 04.9 08.4 75.2 04.4 60.8
DiAD [16] 77.3 97.0 94.4 93.6 06.5 12.9 77.3 06.9 65.6
ViTAD [40] 65.7 93.5 94.5 91.2 04.9 09.8 76.1 05.2 62.2
InvAD [42] 82.8 97.4 94.9 91.9 06.3 12.8 78.0 06.8 66.3
InvAD-lite [42] 72.0 95.5 94.5 91.6 05.8 10.0 77.0 05.3 63.8
MambaAD [15] 66.9 94.7 94.7 91.1 04.7 08.6 76.1 04.5 62.4
Hybrid UniAD [35] 67.0 93.2 94.4 90.8 04.8 09.1 74.8 04.8 62.0
RD++ [31] 60.0 91.4 94.5 90.8 04.8 08.3 75.4 04.3 60.7
DesTSeg [44] 68.2 93.5 94.5 60.3 04.1 08.8 40.5 04.6 52.8
07Owl Aug. DRAEM [37] 56.3 87.8 92.8 45.4 00.5 01.2 16.4 00.6 42.9
SimpleNet [27] 55.0 87.5 92.8 87.7 01.9 04.1 64.6 02.1 56.2
Emb. CFA [21] 60.8 91.7 93.0 77.1 03.2 07.0 47.4 03.6 54.3
CFLOW-AD [14] 66.9 92.8 92.8 89.9 03.8 05.0 66.2 02.6 59.6
PyramidalFlow [22] 48.4 85.1 93.0 90.0 03.5 07.8 71.1 04.1 57.0
Rec. RD [10] 62.2 91.8 93.0 91.9 05.3 09.4 77.7 04.9 61.6
DiAD [16] 62.0 91.4 92.8 94.4 06.1 13.1 72.8 07.0 61.8
ViTAD [40] 65.6 91.9 92.8 92.7 05.2 10.9 77.8 05.7 62.4
InvAD [42] 72.6 94.3 93.0 94.6 12.5 22.7 83.6 12.8 67.6
InvAD-lite [42] 66.6 93.1 93.2 93.3 07.6 13.7 80.1 07.3 63.9
MambaAD [15] 67.5 93.5 92.8 91.5 04.8 09.3 76.0 04.9 62.2
Hybrid UniAD [35] 63.7 92.1 93.0 92.5 04.9 09.8 77.6 05.2 61.9
RD++ [31] 66.1 92.5 92.8 92.2 05.2 09.0 78.4 04.7 62.3
DesTSeg [44] 64.8 92.4 93.2 54.3 11.1 10.0 39.1 05.3 52.1
08Sabertooth Aug. DRAEM [37] 58.0 88.3 93.0 47.1 00.5 01.1 18.6 00.6 43.8
SimpleNet [27] 55.1 88.3 93.2 87.3 01.9 03.9 65.6 02.0 56.5
Emb. CFA [21] 53.8 90.6 93.0 72.0 02.2 03.8 36.8 01.9 50.3
CFLOW-AD [14] 60.8 90.8 93.2 86.6 01.7 03.4 58.3 01.7 56.4
PyramidalFlow [22] 65.2 93.2 93.0 82.5 01.2 02.6 58.6 01.3 56.6
Rec. RD [10] 66.0 93.2 93.4 88.0 02.0 03.6 67.2 01.9 59.1
DiAD [16] 69.2 94.0 94.5 86.1 01.9 03.4 51.3 01.7 57.2
ViTAD [40] 61.6 91.3 93.2 89.0 02.2 04.0 68.0 02.0 58.5
InvAD [42] 78.0 95.8 93.4 91.3 03.9 06.2 74.7 03.2 63.3
InvAD-lite [42] 73.7 94.6 93.4 89.4 03.1 04.6 69.4 02.4 61.2
MambaAD [15] 64.4 92.6 93.4 87.1 02.2 03.5 63.9 01.8 58.2
Hybrid UniAD [35] 62.5 91.0 93.2 89.5 02.5 04.2 74.5 02.1 59.6
RD++ [31] 63.2 91.9 93.0 88.1 02.3 03.7 66.6 01.9 58.4
DesTSeg [44] 70.0 93.6 93.2 52.0 02.3 04.0 33.9 02.0 49.9
Table A77: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
09Swan Aug. DRAEM [37] 46.6 86.6 93.1 47.7 00.6 01.4 19.9 00.7 42.3
SimpleNet [27] 49.8 85.9 93.1 86.1 01.9 03.9 62.4 02.0 54.7
Emb. CFA [21] 54.9 88.2 93.1 72.0 02.0 06.7 37.8 03.5 50.7
CFLOW-AD [14] 47.0 86.9 93.1 85.7 02.2 03.9 60.1 02.0 54.1
PyramidalFlow [22] 43.7 85.4 93.1 80.2 01.3 02.9 53.9 01.5 51.5
Rec. RD [10] 55.8 89.3 93.1 87.7 02.4 04.6 71.0 02.3 57.7
DiAD [16] 78.8 95.5 94.3 87.8 09.2 16.7 57.8 09.1 62.9
ViTAD [40] 54.1 88.1 93.2 90.1 03.1 06.2 73.7 03.2 58.4
InvAD [42] 65.8 92.6 93.2 91.8 06.2 11.6 80.0 06.2 63.0
InvAD-lite [42] 60.1 90.5 93.1 90.1 04.1 08.3 76.8 04.3 60.4
MambaAD [15] 56.3 89.5 93.1 87.3 02.5 04.4 69.4 02.2 57.5
Hybrid UniAD [35] 58.4 89.7 93.1 89.5 03.2 06.8 74.7 03.5 59.3
RD++ [31] 53.9 89.4 93.2 87.6 02.4 04.4 69.8 02.3 57.2
DesTSeg [44] 53.3 89.8 93.1 51.5 02.2 03.2 36.0 01.6 47.0
10Sheep Aug. DRAEM [37] 36.8 81.6 92.7 47.2 00.6 01.3 20.8 00.7 40.1
SimpleNet [27] 44.9 85.1 92.7 91.3 02.8 05.2 76.4 02.7 56.9
Emb. CFA [21] 66.4 92.3 92.7 71.9 02.9 07.4 50.2 03.8 54.8
CFLOW-AD [14] 63.0 90.1 92.9 91.9 03.9 06.8 79.2 03.5 61.1
PyramidalFlow [22] 58.6 90.6 92.7 90.6 02.2 05.0 75.9 02.6 59.4
Rec. RD [10] 62.6 91.8 92.7 92.5 05.1 08.8 84.1 04.6 62.5
DiAD [16] 87.4 97.3 94.3 93.6 04.5 08.8 76.6 04.6 66.1
ViTAD [40] 65.0 91.1 92.7 93.4 04.4 08.7 84.7 04.5 62.9
InvAD [42] 84.5 96.8 93.5 94.3 08.1 13.9 87.0 07.5 68.3
InvAD-lite [42] 80.5 96.2 93.9 93.7 07.0 10.7 85.7 05.7 66.8
MambaAD [15] 66.0 93.0 92.7 92.6 05.2 09.4 83.9 05.0 63.3
Hybrid UniAD [35] 65.9 92.4 92.9 93.6 05.6 09.9 84.2 05.2 63.5
RD++ [31] 70.4 94.2 92.7 92.6 06.5 10.6 84.1 05.6 64.4
DesTSeg [44] 71.1 94.1 92.9 54.8 05.9 08.2 43.4 04.3 52.9
11Pig Aug. DRAEM [37] 52.4 89.4 94.0 47.2 00.5 01.1 22.2 00.5 43.8
SimpleNet [27] 50.2 88.7 94.0 89.8 02.0 04.2 73.3 02.1 57.5
Emb. CFA [21] 57.1 91.4 94.0 84.7 09.7 17.7 56.1 09.7 58.7
CFLOW-AD [14] 52.6 90.1 94.0 92.4 06.5 05.6 74.5 02.9 59.4
PyramidalFlow [22] 50.9 89.4 94.0 91.3 02.6 05.7 71.5 02.9 57.9
Rec. RD [10] 59.3 92.0 94.0 94.1 06.0 11.7 83.2 06.2 62.9
DiAD [16] 53.1 92.4 94.0 93.9 09.2 16.3 71.5 08.9 61.5
ViTAD [40] 58.8 91.6 94.5 95.7 08.1 15.6 83.0 08.5 63.9
InvAD [42] 72.2 94.9 94.6 96.8 17.6 25.1 87.6 14.3 69.8
InvAD-lite [42] 66.3 93.6 94.2 96.1 13.6 21.2 85.8 11.8 67.3
MambaAD [15] 63.5 92.9 94.0 95.1 08.8 15.5 83.8 08.4 64.8
Hybrid UniAD [35] 52.6 88.9 94.0 94.5 05.1 09.6 81.2 05.0 60.8
RD++ [31] 60.8 92.4 94.0 94.5 06.5 12.7 83.6 06.8 63.5
DesTSeg [44] 65.5 93.0 94.0 62.3 08.1 13.1 48.0 07.0 54.9
12Zalika Aug. DRAEM [37] 50.4 87.0 92.9 40.9 00.9 02.2 17.0 01.1 41.6
SimpleNet [27] 50.4 86.8 92.9 84.1 02.5 05.0 57.9 02.5 54.2
Emb. CFA [21] 56.2 90.5 92.9 70.8 04.3 08.7 47.5 04.5 53.0
CFLOW-AD [14] 53.0 88.4 92.9 86.0 04.4 07.0 55.4 03.6 55.3
PyramidalFlow [22] 42.1 85.2 92.9 83.4 02.3 04.7 63.7 02.4 53.5
Rec. RD [10] 56.8 91.0 92.9 87.3 05.0 08.3 69.2 04.3 58.6
DiAD [16] 34.5 82.3 92.9 85.4 03.5 06.1 43.2 03.2 49.7
ViTAD [40] 56.4 91.0 92.9 87.4 04.8 08.6 68.2 04.5 58.5
InvAD [42] 61.2 92.2 92.9 90.2 09.4 15.5 75.4 08.4 62.4
InvAD-lite [42] 58.6 91.2 92.9 88.9 07.1 11.7 72.6 06.2 60.4
MambaAD [15] 58.6 91.5 92.9 86.8 04.7 08.0 68.0 04.2 58.6
Hybrid UniAD [35] 50.8 88.9 92.9 86.8 03.8 06.7 67.9 03.4 56.8
RD++ [31] 52.7 90.3 92.9 87.3 05.5 08.8 69.1 04.6 58.1
DesTSeg [44] 50.7 88.9 92.9 52.8 04.7 06.0 38.8 03.1 47.8
Table A78: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
13Pheonix Aug. DRAEM [37] 48.9 89.6 94.3 51.2 00.6 01.6 27.8 00.8 44.9
SimpleNet [27] 49.2 89.3 94.3 82.2 01.3 02.8 52.3 01.4 53.1
Emb. CFA [21] 56.9 91.0 94.4 66.2 01.6 04.6 50.0 02.4 52.1
CFLOW-AD [14] 52.5 90.2 94.4 82.3 01.5 02.9 57.6 01.5 54.5
PyramidalFlow [22] 44.6 86.1 94.3 78.0 01.1 02.5 61.6 01.2 52.6
Rec. RD [10] 52.1 90.1 94.4 83.5 02.2 03.3 69.3 01.7 56.4
DiAD [16] 39.7 87.8 94.3 83.6 01.5 03.0 49.9 01.5 51.4
ViTAD [40] 53.3 90.5 94.3 84.6 01.7 03.4 68.0 01.8 56.5
InvAD [42] 64.2 92.8 94.6 87.5 03.7 08.2 77.4 04.3 61.2
InvAD-lite [42] 56.3 91.9 94.5 86.0 03.1 04.6 74.2 02.3 58.7
MambaAD [15] 57.5 92.1 94.4 83.4 02.7 03.4 69.8 01.7 57.6
Hybrid UniAD [35] 55.3 90.6 94.3 85.2 01.8 03.6 68.1 01.8 57.0
RD++ [31] 56.6 91.9 94.4 84.1 01.9 03.7 70.3 01.9 57.6
DesTSeg [44] 56.9 92.0 94.3 52.9 02.3 04.8 35.7 02.5 48.4
14Elephant Aug. DRAEM [37] 44.6 87.2 93.9 50.3 01.0 01.9 19.3 00.9 42.6
SimpleNet [27] 48.0 87.8 93.9 71.8 01.8 04.0 49.8 02.0 51.0
Emb. CFA [21] 57.4 91.5 93.9 54.1 01.8 04.1 39.0 02.1 48.8
CFLOW-AD [14] 55.8 91.3 93.9 68.1 02.2 03.5 49.6 01.8 52.1
PyramidalFlow [22] 47.0 88.0 93.9 63.8 01.4 03.2 51.1 01.6 49.8
Rec. RD [10] 58.9 91.7 93.9 69.8 02.8 06.1 64.1 03.2 55.3
DiAD [16] 65.8 94.4 93.9 74.3 02.0 03.9 45.9 02.0 54.3
ViTAD [40] 57.0 91.9 93.9 70.7 03.1 07.2 64.2 03.7 55.4
InvAD [42] 66.7 93.9 94.0 72.5 04.6 11.5 68.5 06.1 58.8
InvAD-lite [42] 65.9 93.9 93.9 70.8 03.6 08.2 65.8 04.3 57.4
MambaAD [15] 63.3 93.4 93.9 69.6 02.8 06.3 62.8 03.2 56.0
Hybrid UniAD [35] 58.2 91.5 93.9 69.5 02.2 05.1 62.2 02.6 54.7
RD++ [31] 58.7 92.3 94.0 70.2 02.9 06.5 63.3 03.3 55.4
DesTSeg [44] 63.7 92.9 94.0 54.5 04.9 08.0 50.1 04.2 52.6
15Parrot Aug. DRAEM [37] 52.1 86.0 91.8 48.6 00.8 01.8 18.7 00.9 42.8
SimpleNet [27] 49.4 82.6 91.8 83.3 02.4 04.8 62.1 02.5 53.8
Emb. CFA [21] 57.1 89.2 91.8 65.9 02.3 04.8 48.2 02.4 51.3
CFLOW-AD [14] 53.0 87.0 92.0 83.4 03.0 05.3 60.2 02.7 54.8
PyramidalFlow [22] 59.2 86.3 92.2 80.0 02.2 04.7 59.7 02.4 54.9
Rec. RD [10] 50.7 86.7 92.0 84.0 03.2 06.9 67.1 03.6 55.8
DiAD [16] 53.1 88.9 91.8 83.4 02.5 05.8 63.9 03.0 55.6
ViTAD [40] 49.8 86.9 91.8 83.4 03.1 06.7 66.4 03.5 55.4
InvAD [42] 53.9 88.6 92.0 86.6 06.6 14.5 74.0 07.8 59.5
InvAD-lite [42] 50.0 87.3 91.8 85.4 04.6 09.6 69.8 05.0 56.9
MambaAD [15] 49.9 86.8 91.8 83.9 03.2 06.3 65.1 03.3 55.3
Hybrid UniAD [35] 46.9 86.0 91.8 85.5 03.5 06.5 69.9 03.4 55.7
RD++ [31] 53.9 88.1 92.0 84.1 03.1 06.7 66.5 03.5 56.3
DesTSeg [44] 54.4 88.4 91.8 52.7 04.9 06.0 25.8 03.1 46.3
16Cat Aug. DRAEM [37] 53.8 90.8 94.3 51.5 00.5 01.6 27.3 00.8 45.7
SimpleNet [27] 54.2 90.1 94.3 93.2 02.5 05.0 79.3 02.6 59.8
Emb. CFA [21] 53.5 90.3 94.4 71.8 01.4 04.9 42.5 02.5 51.3
CFLOW-AD [14] 54.5 90.6 94.3 93.1 02.4 04.8 80.3 02.5 60.0
PyramidalFlow [22] 46.7 87.7 94.4 91.7 01.7 04.1 74.2 02.1 57.2
Rec. RD [10] 49.3 88.8 94.3 93.6 03.0 06.1 84.3 03.1 59.9
DiAD [16] 63.5 93.3 94.3 93.9 02.9 05.6 82.0 02.9 62.2
ViTAD [40] 57.2 91.7 94.3 93.9 03.2 06.4 84.6 03.3 61.6
InvAD [42] 57.8 92.4 94.3 94.4 05.1 10.7 86.2 05.6 63.0
InvAD-lite [42] 56.8 91.4 94.3 94.3 03.9 08.0 85.6 04.2 62.0
MambaAD [15] 53.7 90.3 94.3 93.9 03.2 06.5 84.5 03.4 60.9
Hybrid UniAD [35] 52.8 89.8 94.3 94.1 03.0 06.4 83.3 03.3 60.5
RD++ [31] 54.2 91.3 94.3 93.5 03.3 06.1 84.6 03.1 61.0
DesTSeg [44] 54.8 90.7 94.3 57.7 01.6 04.5 45.0 02.3 49.8
Table A79: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
17Scorpion Aug. DRAEM [37] 48.2 86.3 92.9 50.6 00.5 01.2 21.5 00.6 43.0
SimpleNet [27] 52.6 88.2 92.9 83.8 01.5 03.0 55.5 01.5 53.9
Emb. CFA [21] 63.8 92.1 93.1 76.6 03.2 06.8 46.9 03.5 54.6
CFLOW-AD [14] 62.4 91.7 92.9 89.5 03.3 04.8 67.4 02.5 58.9
PyramidalFlow [22] 55.3 86.5 92.9 86.4 02.0 05.1 58.1 02.6 55.2
Rec. RD [10] 62.9 91.2 92.9 91.4 04.7 07.4 81.9 03.8 61.8
DiAD [16] 70.9 94.4 93.9 87.4 02.1 04.1 56.6 02.1 58.5
ViTAD [40] 59.1 89.8 93.1 91.1 03.3 07.2 77.2 03.8 60.1
InvAD [42] 80.9 96.6 93.4 93.7 07.2 10.2 85.4 05.4 66.8
InvAD-lite [42] 74.5 95.0 93.0 93.4 06.1 08.5 84.6 04.4 65.0
MambaAD [15] 65.1 92.4 92.9 92.7 04.8 07.3 82.3 03.8 62.5
Hybrid UniAD [35] 70.1 93.4 92.9 92.2 05.7 09.1 81.1 04.8 63.5
RD++ [31] 64.2 92.4 92.9 91.6 04.8 06.9 81.6 03.6 62.1
DesTSeg [44] 70.0 93.8 93.1 62.6 04.3 08.0 48.8 04.2 54.4
18Obesobeso Aug. DRAEM [37] 50.7 89.7 94.1 51.4 00.9 01.9 24.9 01.0 44.8
SimpleNet [27] 49.3 89.8 94.1 90.9 03.5 06.5 72.7 03.3 58.1
Emb. CFA [21] 61.4 92.8 94.1 78.8 12.4 18.1 60.4 10.0 59.7
CFLOW-AD [14] 62.0 91.9 94.1 92.6 06.0 07.8 79.6 04.0 62.0
PyramidalFlow [22] 62.9 93.1 94.1 93.5 05.5 10.6 79.8 05.6 62.8
Rec. RD [10] 63.0 93.4 94.1 93.4 10.7 19.2 87.0 10.6 65.8
DiAD [16] 92.2 99.0 95.2 95.8 13.2 21.9 84.5 12.3 71.7
ViTAD [40] 66.9 94.4 94.1 94.2 16.5 24.1 88.1 13.7 68.3
InvAD [42] 83.7 97.5 95.1 94.9 16.6 25.8 90.4 14.8 72.0
InvAD-lite [42] 73.4 95.7 94.2 93.7 12.2 20.2 87.9 11.2 68.2
MambaAD [15] 69.6 95.1 94.2 93.1 11.2 21.6 86.5 12.1 67.3
Hybrid UniAD [35] 71.7 94.8 94.2 93.1 09.5 15.8 84.8 08.6 66.3
RD++ [31] 67.0 93.2 94.2 93.1 07.9 15.6 86.0 08.4 65.3
DesTSeg [44] 71.3 95.4 94.1 63.7 18.5 23.7 53.0 13.4 60.0
19Bear Aug. DRAEM [37] 52.3 88.6 93.3 48.7 00.6 01.3 19.1 00.6 43.4
SimpleNet [27] 50.9 87.4 93.3 85.0 02.0 03.9 65.0 02.0 55.4
Emb. CFA [21] 56.1 90.7 93.5 78.6 04.2 08.8 50.1 04.6 54.6
CFLOW-AD [14] 61.1 91.7 93.3 88.9 04.0 06.3 64.4 03.3 58.5
PyramidalFlow [22] 53.6 89.5 93.3 87.3 03.1 05.7 61.6 02.9 56.3
Rec. RD [10] 61.5 92.1 93.6 90.8 05.6 10.0 76.1 05.3 61.4
DiAD [16] 64.3 93.7 93.3 90.1 05.1 13.0 65.0 07.0 60.6
ViTAD [40] 60.1 90.8 93.5 91.1 05.3 11.1 75.0 05.9 61.0
InvAD [42] 71.0 94.5 93.8 92.5 11.7 18.3 81.2 10.1 66.1
InvAD-lite [42] 67.9 93.4 93.4 92.2 11.2 17.0 78.8 09.3 64.8
MambaAD [15] 63.5 92.3 93.5 91.2 07.2 12.7 75.9 06.8 62.3
Hybrid UniAD [35] 61.7 91.9 93.6 90.6 05.0 09.4 79.4 04.9 61.7
RD++ [31] 64.9 92.9 93.5 90.9 06.5 11.2 76.6 05.9 62.4
DesTSeg [44] 68.3 93.5 93.5 60.4 06.8 11.7 52.0 06.2 55.2
20Puppy Aug. DRAEM [37] 51.7 86.3 92.7 47.5 00.5 01.0 17.2 00.5 42.4
SimpleNet [27] 55.6 88.5 92.7 83.1 01.3 02.6 55.6 01.3 54.2
Emb. CFA [21] 50.2 88.6 92.9 68.5 01.9 05.0 42.2 02.6 49.9
CFLOW-AD [14] 59.0 90.3 92.7 84.9 02.0 04.4 56.5 02.2 55.7
PyramidalFlow [22] 52.4 87.6 92.9 84.8 01.9 04.9 59.9 02.5 54.9
Rec. RD [10] 50.9 89.0 92.7 87.0 03.5 05.6 66.1 02.9 56.4
DiAD [16] 46.2 88.9 92.7 83.4 01.4 02.8 52.7 01.4 52.6
ViTAD [40] 55.4 88.2 93.0 87.5 02.4 04.8 66.3 02.4 56.8
InvAD [42] 61.9 92.0 92.9 89.7 05.7 09.7 73.8 05.1 60.8
InvAD-lite [42] 55.2 90.0 92.7 87.9 03.7 05.9 69.1 03.0 57.8
MambaAD [15] 48.8 87.3 92.7 86.6 02.3 04.9 64.6 02.5 55.3
Hybrid UniAD [35] 56.7 90.2 92.7 87.2 04.5 10.8 70.9 05.7 59.0
RD++ [31] 59.3 91.4 93.0 86.9 03.5 06.9 65.4 03.6 58.1
DesTSeg [44] 57.5 91.4 92.9 55.2 11.9 11.7 37.9 06.2 51.2
Appendix W Detailed Quantitative Results on Uni-Medical Dataset Under 300 epochs
Table A80: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
brain Aug. DRAEM [37] 53.0 84.2 90.6 28.7 02.0 06.2 05.7 03.2 38.6
SimpleNet [27] 77.7 94.1 90.7 93.7 36.4 39.3 71.1 24.4 71.9
Emb. CFA [21] 67.0 90.8 90.6 45.2 06.2 07.4 07.5 03.8 45.0
CFLOW-AD [14] 85.8 96.2 92.1 94.5 55.8 17.8 73.6 09.8 73.7
PyramidalFlow [22] 66.6 89.6 90.6 87.0 12.4 18.9 54.6 10.4 60.0
Rec. RD [10] 85.0 95.4 92.1 96.4 45.6 49.0 82.3 32.5 78.0
DiAD [16] 93.6 98.2 95.2 95.3 37.3 42.3 80.4 26.9 77.5
ViTAD [40] 89.0 97.3 92.9 97.6 60.2 58.0 83.6 40.9 82.7
InvAD [42] 89.5 97.2 92.5 97.6 56.5 57.4 85.6 40.3 82.3
InvAD-lite [42] 84.4 95.5 91.7 96.3 44.8 48.1 81.9 31.7 77.5
MambaAD [15] 91.9 98.0 93.5 97.6 56.6 58.0 86.4 40.8 83.1
Hybrid UniAD [35] 90.2 97.6 92.9 97.6 58.8 57.2 83.3 40.0 82.5
RD++ [31] 85.8 96.0 91.6 96.9 48.3 52.2 83.8 35.3 79.2
DesTSeg [44] 80.4 93.8 91.2 82.1 49.3 51.0 55.4 34.2 71.9
liver Aug. DRAEM [37] 56.2 47.3 64.0 26.2 00.5 02.4 09.8 01.2 29.5
SimpleNet [27] 52.5 44.7 62.3 96.8 08.5 13.7 86.9 07.4 52.2
Emb. CFA [21] 37.1 35.0 61.2 14.0 00.2 00.8 00.7 00.4 21.3
CFLOW-AD [14] 63.8 54.0 63.5 96.0 02.9 05.7 87.8 02.9 53.4
PyramidalFlow [22] 65.7 56.1 65.7 82.0 01.7 02.3 41.4 01.2 45.0
Rec. RD [10] 55.4 45.4 64.1 96.5 05.3 10.0 90.6 05.3 52.5
DiAD [16] 65.8 60.1 63.9 96.9 07.6 13.9 91.5 07.5 57.1
ViTAD [40] 63.5 54.5 65.4 97.8 11.1 18.9 92.5 10.4 57.7
InvAD [42] 62.5 52.1 65.2 96.8 06.8 12.6 92.3 06.7 55.5
InvAD-lite [42] 56.4 46.1 63.5 96.1 04.8 09.3 89.7 04.9 52.3
MambaAD [15] 61.1 50.9 63.8 96.8 08.3 15.5 91.7 08.4 55.4
Hybrid UniAD [35] 64.7 51.5 64.9 97.1 07.9 14.1 92.8 07.6 56.1
RD++ [31] 55.5 44.6 64.1 96.4 05.3 09.9 90.9 05.2 52.4
DesTSeg [44] 74.8 72.4 66.0 87.6 38.6 41.8 84.5 26.4 66.5
retinal Aug. DRAEM [37] 66.1 57.9 60.4 42.5 03.3 09.6 11.9 05.1 36.0
SimpleNet [27] 87.3 87.1 77.1 94.2 62.8 58.0 79.6 40.9 78.0
Emb. CFA [21] 64.6 62.6 59.5 55.2 08.1 11.5 18.6 06.1 40.0
CFLOW-AD [14] 87.4 85.5 76.3 95.6 60.0 55.5 83.7 38.4 77.7
PyramidalFlow [22] 47.7 42.8 59.5 77.7 12.2 23.4 44.7 13.3 44.0
Rec. RD [10] 88.8 86.7 77.7 96.4 65.7 60.6 86.6 43.4 80.4
DiAD [16] 89.2 87.5 79.0 95.7 62.7 58.3 84.3 41.2 79.5
ViTAD [40] 92.0 90.0 82.0 95.5 69.3 63.8 83.3 46.8 82.3
InvAD [42] 92.1 90.6 82.1 97.2 72.9 66.5 89.0 49.8 84.3
InvAD-lite [42] 92.1 90.2 82.8 96.1 66.5 61.8 85.7 44.7 82.2
MambaAD [15] 93.8 87.9 86.6 95.9 64.2 64.7 85.5 47.8 82.7
Hybrid UniAD [35] 86.2 80.7 74.8 94.9 50.4 52.7 81.2 35.8 74.4
RD++ [31] 90.3 87.7 79.9 96.7 69.5 63.8 87.5 46.9 82.2
DesTSeg [44] 90.9 88.9 81.9 80.0 64.0 58.7 58.5 41.6 74.7
Avg Aug. DRAEM [37] 58.5 63.2 71.7 32.5 02.0 06.1 09.1 03.2 34.7
SimpleNet [27] 72.5 75.3 76.7 94.9 35.9 37.0 79.2 24.2 67.4
Emb. CFA [21] 56.3 62.8 70.4 38.1 04.8 06.6 08.9 03.4 35.4
CFLOW-AD [14] 79.0 78.5 77.3 95.4 39.6 26.3 81.7 17.0 68.3
PyramidalFlow [22] 60.0 62.8 71.9 82.2 08.8 14.9 46.9 08.3 49.6
Rec. RD [10] 76.4 75.8 77.9 96.4 38.9 39.8 86.5 27.0 70.2
DiAD [16] 82.9 81.9 79.4 96.0 35.9 38.2 85.4 25.2 71.4
ViTAD [40] 81.5 80.6 80.1 97.0 46.8 46.9 86.5 32.7 74.2
InvAD [42] 81.3 80.0 79.9 97.2 45.4 45.5 89.0 32.3 74.0
InvAD-lite [42] 77.6 77.2 79.3 96.2 38.7 39.7 85.7 27.1 70.6
MambaAD [15] 82.3 78.9 81.3 96.7 43.0 46.1 87.9 32.4 73.7
Hybrid UniAD [35] 80.4 76.6 77.5 96.5 39.0 41.3 85.8 27.8 71.0
RD++ [31] 77.2 76.1 78.5 96.7 41.0 41.9 87.4 29.1 71.3
DesTSeg [44] 82.0 85.0 79.7 83.2 50.6 50.5 66.1 34.1 71.0